### v1

# GEORCE: A Fast New Control Algorithm for Computing Geodesics

Abstract
Computing geodesics for Riemannian manifolds is a difficult task that often relies on numerical approximations. However, these approximations tend to be either numerically unstable, have slow convergence, or scale poorly with manifold dimension and number of grid points. We introduce a new algorithm called GEORCE that computes geodesics via a transformation into a discrete control problem. We show that GEORCE has global convergence and quadratic local convergence. In addition, we show that it extends to Finsler manifolds. For both Finslerian and Riemannian manifolds, we thoroughly benchmark GEORCE against several alternative optimization algorithms and show empirically that it has a much faster and more accurate performance for a variety of manifolds, including key manifolds from information theory and manifolds that are learned using generative models.
keywords:
Riemannian Manifolds, Finsler Manifolds, Geodesics, Optimization, Control Problem
Mathematics Subject Classification 53C22, 49Q99, 65D15
1 Introduction
Reliably computing geodesics that connect point pairs is a constant source of frustration to practitioners working with non-trivial Riemannian manifolds. Such manifolds see common use across many fields, including medical image analysis [1, 2, 3], protein modeling [4, 5, 6], robotics [7, 8, 9, 10], generative modeling [11, 12, 13, 14, 15, 16] and information geometry [17, 18, 19, 20, 21].
Geodesics, or locally shortest paths, are the key computational tools for engaging with data residing on a Riemannian manifold. The lengths of geodesic curves inform us about manifold distances. Likewise, the exponential and logarithm maps, which generalize addition and subtraction, are rooted in geodesics [22]. Unfortunately, the solution to the boundary-value problem for geodesics are only available in closed form for the simplest manifolds, and numerical approximations are therefore required in practice.
Such approximations are typically achieved by minimizing the energy functional over a suitable set of curves connecting the given boundary points and parametrized by $\gamma:[0,1]\rightarrow\mathcal{M}$ [23]
$\mathcal{E}(\gamma)=\frac{1}{2}\int_{0}^{1}\dot{\gamma}(t)^{\top}G\left(\gamma(t)\right)\dot{\gamma}(t)\,\mathrm{d}t,$ (1)
where $G$ denotes the metric matrix function of the manifold in local coordinates. This energy is commonly minimized with general-purpose gradient-based optimization methods [24, 25, 12] or, in particular when considering the initial value problem and using shooting methods, by solving the system of ordinary differential equations that arise by applying the Euler-Lagrange equations [26, 27]. Unfortunately, both approaches are often numerically brittle, exhibit slow convergence, or scale poorly with the number of grid points and manifold dimensions.
Figure 1: Comparison between GEORCE and baseline methods for computing locally length minimizing geodesics between given point pairs on four different manifolds. All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional (5) is less than $10^{-4}$, or the number of iterations exceeds $100$. Note that sometimes the respective solutions overlap, making it difficult to distinguish them. Also note, that in the case of the torus the initial curve determines the direction around the hole so that the obtained geodesics follow this direction and are therefore clearly not the globally shortest geodesics between the given start and end point. Experimental details and benchmark data are in Appendix E.
In this paper, we introduce the GEORCE algorithm (GEodesic Optimization algoRithm using Control tEchniques) that computes numerical solutions to the geodesic boundary value problem via an optimal control strategy. GEORCE applies to general Riemannian as well as general Finslerian manifolds. We show that the algorithm has global convergence similar to gradient descent and that it has quadratic local convergence similar to the Newton method. Empirically, we compare GEORCE with alternative optimization algorithms on the discretized energy functional and see that GEORCE is much faster and attains lower Riemannian length as illustrated in Fig. 1.
2 Background and related work
A Riemannian manifold, $(\mathcal{M},g)$, can indirectly be considered as a smooth $n$-dimensional surface in a Euclidean space of sufficiently high dimension, see [28, 29]. In this way $\mathcal{M}$ inherits its metric $g$ from the ambient space. All the intrinsic geometry of $(\mathcal{M},g)$, including curvature and geodesic distances, are encoded in the metric $g$. The metric can be – and often is – constructed independently of the formal embedding alluded to above, namely directly as a symmetric positive definite matrix valued function on the domain of a parametrization of the manifold in question. The metric, $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$, defines an inner product that smoothly varies over the tangent spaces of $\mathcal{M}$ such that in a local chart $g(v,w)=v^{\top}G(x)w$, where $G$ is the metric matrix function evaluated in $x\in\mathcal{M}$ for $v,w\in T_{x}\mathcal{M}$. The tangent space, $T_{x}\mathcal{M}$, consists of tangents to all curves at $x\in\mathcal{M}$ [23].
A geodesic can then be defined using the inner product in each tangent space. The length of a curve, $\gamma:U\subseteq[0,1]\rightarrow\mathcal{M}$ parameterized by $t\in[0,1]$ is given by
$\mathcal{L}(\gamma)=\int_{0}^{1}\sqrt{\dot{\gamma}(t)^{\top}G(\gamma(t))\dot{\gamma}(t)}\,\mathrm{d}t,$ (2)
with the corresponding energy given by eq. 1. A geodesic is a smooth curve that locally minimizes the length between two points $x,y\in\mathcal{M}$ and gives rise to the Riemannian distance, $\mathrm{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}_{+}$, which is defined as the length of the corresponding shortest geodesic between the given two points on the manifold [23]. It can be shown that if a curve $\gamma$ locally minimizes the energy (1), then the curve has constant velocity and locally minimizes the length [30]. For general Riemannian manifolds, existence and uniqueness are only guaranteed in a star-shaped neighborhood around any point $x\in\mathcal{M}$ [23]. All points, where the geodesics from a point $x\in\mathcal{M}$ are no longer length minimizing, are denoted the cut locus of $x$ [23]. In this paper, we assume that the manifold is geodesically complete such that at least one length-minimizing geodesic exists between any two points $x,y\in\mathcal{M}$.
A Riemannian manifold can be equipped with a unique metrically compatible connection, $\nabla$, known as the Levi-Civita Connection [23], where uniqueness and existence are guaranteed by completeness of the metric [23]. An equivalent way to define geodesics is to define these as curves, $\gamma$, with zero acceleration using the Levi-Civita connection as the covariant derivative, i.e., $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$. In a local chart, this gives rise to the following (usually non-linear) ode-system [23]
$\frac{\mathrm{d}^{2}\gamma^{k}}{\mathrm{d}t^{2}}+\Gamma_{ij}^{k}\frac{\mathrm{d}\gamma^{i}}{\mathrm{d}t}\frac{\mathrm{d}\gamma^{j}}{\mathrm{d}t}=0,\quad\gamma(0)=a,\gamma(1)=b,$ (3)
written in Einstein notation [^1] In Einstein notation, upper and lower indices constitute a sum, i.e. $a^{i}b_{i}=\sum_{i}a_{i}b_{i}$, where $\left\{\Gamma_{ij}^{k}\right\}_{i,j,k}$ denote the Christoffel symbols. The exponential map $\mathrm{Exp}_{x}(v):T_{x}\mathcal{M}\rightarrow\mathcal{M}$ is then defined as $\mathrm{Exp}_{x}(v)=\gamma(1)$, where $\gamma$ is a geodesic with $\gamma(0)=x\in\mathcal{M}$ and initial velocity vector, $v\in T_{x}\mathcal{M}$, i.e. $\dot{\gamma}(0)=v$. Within the cut locus of $x\in\mathcal{M}$ the exponential map, $\mathrm{Exp}_{x}$, is a diffeomorphism and its inverse is the logarithmic map, $\mathrm{Log}_{x}:\mathcal{M}\rightarrow T_{x}\mathcal{M}$.
Examples of Riemannian manifolds include information geometric spaces, i.e., statistical manifolds, where “points” in the manifold correspond to statistical distributions. These can be equipped with the so-called Fisher-Rao metric, see [31, 18]:
$G_{ij}(\theta)=\int_{\mathcal{X}}p(x|\theta)\left(\frac{\partial}{\partial\theta^{i}}\log p(x|\theta)\right)\left(\frac{\partial}{\partial\theta^{j}}\log p(x|\theta)\right)\,\mathrm{d}\mu(x),$ (4)
where $\theta$ corresponds to the parameters of the probability density function $p$ under a measure $\mu$. Informally, the Fisher-Rao metric can be seen as a measure of the amount of information in the data around a parameter $\theta$ [32].
The energy functional in eq. 1 can be discretized and minimized using standard optimization methods. Consider fixed start and end points $x_{0}=a\in\mathcal{M}$ and $x_{T}=b\in\mathcal{M}$, and let $x_{0:T}:=\{x_{t}\}_{t=0}^{T}$ denote a discretized version of the candidate curve $\gamma$ in $T+1$ points. Minimizing the energy functional can be re-written as a discrete constrained optimization problem, i.e., subdivision of $[0,1]$ into $T$ subintervals, in the chosen local coordinate system [25, 12]
$\begin{split}E(x_{0:T})=\quad\min_{x_{0:T}}\quad&\sum_{t=0}^{T-1}(x_{t+1}-x_{t})^{\top}G(x_{t})(x_{t+1}-x_{t})\\ \text{s.t.}\quad&x_{0}=a,x_{T}=b,\end{split}$ (5)
where $G\left(\cdot\right)$ denotes the metric matrix function in local coordinates, and the tangent vectors of the coordinate expression for $\gamma$ are discretized as $\dot{\gamma}(t)\approx\left(x_{t+1}-x_{t}\right)T$. Note that the constant scaling by $\nicefrac{{1}}{{2}}$ is omitted since it does not affect the optimal solution. With this formulation, the curve is approximated as a piece-wise linear function. This is not restrictive since $\gamma$ can be approximated into a curve for any given precision $T$ or converted to a smooth function, e.g., using splines [24].
The optimization problem in eq. 5 can be solved using standard solvers in the variables $x_{1:(T-1)}$. Gradient descent converges to a local minimum (global convergence), but will have local linear convergence, while a Newton method or quasi-Newton method will have local quadratic and super-linear convergence, respectively, but will involve solving a linear equation system and will, in general, not have global convergence [33].
Limitations of existing numerical solvers. Solving the ode-system in eq. 3 as a boundary value problem (bvp) can have difficulties converging depending on the choice of initial velocity vector [26] and does not scale well in higher dimension. We illustrate this in Fig. 2, where we show the estimated length and runtime for $\mathbb{S}^{n}$ with $n=2,3,5,10,20$ for different integration methods compared to our proposed algorithm (GEORCE). Different approaches try to circumvent this by solving the bvp on subintervals [26, 34] or solving the bvp-problem as a fixed-point problem in a probabilistic manner [35, 36]. However, these methods often require access to higher-order information about the metric tensor, e.g in the form of Christoffel symbols.
Figure 2: The estimates of bvp-solvers with different integration methods compared to GEORCE on $\mathbb{S}^{n}$ for $n=2,3,5,10$. The bvp-solvers minimize the squared error to the boundary condition using BFGS [37, 38, 39, 40]. All methods are terminated if they take more than 24 hours or use more than 10 GB of memory on a CPU. Details on the experiment is found in Appendix E.1.
Directly minimizing the energy in Eq. 5 can, e.g., be done using off-the-self solvers like ADAM [41] or BFGS [37, 38, 39, 40]. However, these methods tend to have slow convergence and scale poorly in the manifold dimension and number of grid points (see Sec. 5). Alternatively, the geodesic can be estimated by discretizing the smooth manifold by simplicial complexes [42], for which there exist different geodesic solvers [43, 44, 45, 46, 47]. Although discretized manifolds appear naturally in computational geometry, the approach is not scalable to discretize high-dimensional smooth manifolds.
3 GEORCE
To develop an efficient primal-dual algorithm for computing geodesics, we will first reformulate the discretized energy functional (5) as a discrete control problem. We then derive the necessary conditions for a minimum for the control problem using techniques from optimal control to decompose the discrete control problem into strictly convex subproblems; from this, we deduce an iterative scheme to compute geodesics. Finally, we prove that our proposed algorithm exhibits global convergence and local quadratic convergence.
As a starting point, we reformulate the problem of finding a minimum of the discrete energy functional (5) to a control problem
$\begin{split}\min_{(x_{1:T-1},u_{0:T-1})}\quad&\left\{\sum_{t=0}^{T-1}u_{t}^{\top}G(x_{t})u_{t}\right\}\\ \text{s.t.}\quad&x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a,x_{T}=b.\end{split}$ (6)
Here, $x_{0:T}$ denotes the state variables, while $u_{0:(T-1)}$ corresponds to the control variables, i.e., the velocity vectors along the curve (modulo scaling). A key insight into our approach is that rephrasing into an optimal control problem leads to a set of convex subproblems. Using this, we derive the following proposition.
Proposition 1.
The necessary conditions for a minimum in Eq. 6 is
$\begin{split}&2G(x_{t})u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &{\left.\kern-1.2pt\nabla_{y}\left[u_{t}^{\top}G(y)u_{t}\right]\vphantom{\big{|}}\right|_{y=x_{t}}}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1.\\ &x_{0}=a,x_{T}=b,\end{split}$ (7)
where $\mu_{t}\in\mathbb{R}^{d}$ for $t=0,\dots,T-1$.
Proof.
See Appendix A.2. ∎
Proposition 1 decomposes the minimization problem in Eq. 6 into a system of equations for each time step $t$. Note that these equations are necessary conditions for the optimum for eq. 6 and are sufficient in the case where the objective function in eq. 6 is convex in both $x_{t}$ and $u_{t}$. However, this is generally not the case.
The problem with the necessary conditions in Proposition 1 is that the equations cannot be solved in closed form for a general metric matrix function, $G$. To circumvent this, we will derive an iterative scheme. Consider the variables $x_{0:T}^{(i)}$ and $u_{0:T}^{(i)}$ in iteration $i$. We apply the “trick” of fixing the metric matrix function, and the derivative of the inner product in iteration $i$.
$\begin{split}&\nu_{t}:={\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{\top}G(y)u_{t}\right)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)},u_{t}=u_{t}^{(i)}}},\quad t=1,\dots,T-1,\\ &G_{t}:=G\left(x_{t}^{(i)}\right),\quad t=0,\dots,T-1,\end{split}$ (8)
Fixing these in iteration $i$, the system of equations in Eq. 7 reduces to
$\begin{split}&2G_{t}u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &\nu_{t}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1,\\ &\sum_{t=0}^{T-1}u_{t}=b-a,\\ \end{split}$ (9)
where $\nu_{1:(T-1)}$ and $G_{0:(T-1)}$ are fixed (8). Since $G_{0:(T-1)}$ are positive definite, the modified system of equations in Eq. 9 can be explicitly solved in iteration $i$ for $u_{0:(T-1)}$ and $\mu_{0:(T-1)}$ as shown in the following proposition.
Proposition 2.
The update scheme for $u_{t},\mu_{t}$ and $x_{t}$ is
$\begin{split}&\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$ (10)
Proof.
See Appendix A.3. ∎
Algorithm 1 GEORCE for Riemannian manifolds
1:Input: $\mathrm{tol}$, $T$
2:Output: Geodesic estimate $x_{0:T}$
3:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$, $u_{t}^{(0)}\leftarrow\frac{b-a}{T}$ for $t=0.,\dots,T$ and $i\leftarrow 0$
4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
5:     $G_{t}\leftarrow G\left(x_{t}^{(i)}\right)$ for $t=0,\dots,T-1$
6:     $\nu_{t}\leftarrow{\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{(i)}G\left(y\right)u_{t}^{(i)}\right)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}$ for $t=,1\dots,T-1$
7:     $\mu_{T-1}\leftarrow\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right)$
8:     $u_{t}\leftarrow-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right)$ for $t=0,\dots,T-1$
9:     $x_{t+1}\leftarrow x_{t}+u_{t}$ for $t=0,\dots,T-1$
10:     Using line search find $\alpha^{*}$ for the following optimization problem with the discrete energy functional $E$
$\begin{split}\alpha^{*}=\operatorname*{arg\,min}_{\alpha}\quad&E\left(x_{0:T}\right)\quad\text{(exact line search)}\\ \text{s.t.}\quad&x_{t+1}=x_{t}+\alpha u_{t}+(1-\alpha)u_{t}^{(i)},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
11:     Set $u_{t}^{(i+1)}\leftarrow\alpha^{*}u_{t}+(1-\alpha^{*})u_{t}^{(i)}$ for $t=0,\dots,T-1$
12:     Set $x_{t+1}^{(i+1)}\leftarrow x_{t}^{(i+1)}+u_{t}^{(i+1)}$ for $t=0,\dots,T-1$
13:     $i\leftarrow i+1$
14:end while
15:return $x_{t}$ for $t=0,\dots,T-1$
Proposition 2 immediately leads to an iterative updating of $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ to estimate the geodesic, which we denote the GEORCE-algorithm. GEORCE is described in pseudo-code in Algorithm 1.
Line search and stopping criteria
We find that an appropriate step size is necessary to ensure global convergence for GEORCE (See Proposition 3), and therefore we update $u_{t}^{(i+1)}\leftarrow\alpha u_{t}^{(i+1)}+(1-\alpha)u_{t}^{(i)}$ in iteration $i$. Here, $\alpha$ denotes the step size, which we estimate with line search. One approach is to estimate the optimal step size in each iteration, but, in our experience, this is unnecessarily expensive. We instead use soft line search, which merely ensures that the objective function decreases in each iteration [33]. We implement this using backtracking, where an initial step size is multiplied by a decay rate, $10<\rho<$ 1, until some condition is met, e.g., the Wolfe condition [48], the Armijo condition [49], or the curvature condition [48]. In practice, we find that the Armijo condition [49] with $\rho=0.5$ is sufficient for most manifolds, although the decay rate can be fine-tuned to increase the performance of GEORCE.
We threshold the $\ell^{2}$-norm of the gradient of the discrete energy functional as a stopping criteria for GEORCE. In all empirical comparisons, identical stopping criteria are applied to all optimization algorithms.
Convergence results
In the following, we show that GEORCE is guaranteed to converge to a local minimum, i.e., it exhibits global convergence.
Proposition 3 (Global convergence).
Let $E^{(i)}$ be the value of the discretized energy functional for the solution after iteration $i$ (with line search) in GEORCE. If the starting point $\left(x^{(0)},u^{(0)}\right)$ is feasible in the sense that the geodesic candidate has the correct start and end points, then the series $\left\{E^{(i)}\right\}_{i>0}$ will converge to a (local) minimum.
Proof.
See Appendix A.5 for the proof. ∎
In theory, GEORCE could jump between two local minima with the same objective value unless the local minimum is strongly unique, as described in Appendix A.1, and seen in conjugate points [23]. Proposition 3 ensures that GEORCE converges to a local minimum (global convergence) similar to gradient descent. In contrast, gradient descent has only linear local convergence, but we show in the following that under sufficient regularity of the discretized energy functional (see Appendix A.1), GEORCE has quadratic local convergence.
Figure 3: The figure shows the $\ell^{2}$-norm of the gradient of eq. 5 for each iteration for GEORCE and alternative methods applied to construct geodesics for $100$ iterations for four different manifolds corresponding to Fig. 1. We see that the $\ell^{2}$-norm of the gradient of eq. 5 for GEORCE converges considerably faster than for alternative algorithms.
Proposition 4 (Local quadratic convergence).
Under sufficient regularity of the discretized energy functional (see Appendix A.1) it holds that if the discretized energy functional has a strongly unique (local) minimum point $z^{*}$ (see Appendix A.1) and locally the step-size $\alpha^{*}=1$, i.e., no line search, then GEORCE has locally quadratic convergence, i.e.,
$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*})\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
where $z^{(i)}=(x^{(i)},u^{(i)})$ is the solution from GEORCE in iteration $i$, and $B_{\epsilon}\left(z^{*}\right)$ is the ball with radius $\epsilon>0$ and center $z^{*}$.
Proof.
See Appendix A.6 for the proof and Appendix A.1 for the assumptions. ∎
Fig. 3 demonstrates the convergence by plotting the $\ell^{2}$-norm of the gradient of the discretized energy functional in each iteration for four different manifolds. Note that by formulating the geodesic boundary problem as a control problem, we estimate the discretized tangent vectors along the curve $\{u_{t}\}_{t=0}^{T}$. Modulo scaling $u_{0}$ will in the limit of a sufficiently fine grid correspond to the logarithmic map, $\mathrm{Log}_{a}b$.
Algorithm Global convergence Local convergence Gradient Hessian Complexity
Gradient descent Linear $\mathcal{O}\left(Td^{2}\right)$
Quasi-Newton Super linear $\mathcal{O}\left(T^{2}d^{2}\right)$
Newton method Quadratic $\mathcal{O}\left(Td^{3}\right)$
GEORCE Quadratic $\mathcal{O}\left(Td^{3}\right)$
Table 1: Convergence, complexity, and use of higher order derivatives for different optimization algorithms in each iteration, where $T$ is the number of grid points and $d$ is the manifold dimension.
Comparison with other optimization methods
GEORCE has complexity $\mathcal{O}\left(Td^{3}\right)$ due to matrix inversions along the discretized curve but scales only linearly in $T$, unlike e.g. quasi-Newton methods (Table 1). Although the computational complexity in the manifold dimension is of a lower order for quasi-Newton methods and gradient descent, GEORCE exhibits faster convergence. Since the Hessian of the discretized energy functional in Eq. 5 is sparse, the update step in the Newton method can be simplified in complexity compared to the standard Newton method that has complexity $\mathcal{O}\left(T^{3}d^{3}\right)$ (see Appendix B.2). However, the sparse Newton step requires second-order derivatives, unlike GEORCE, which only requires first-order derivatives of the discretized energy.
4 Extension to Finslerian geometry
A Finsler geometry can be seen as a generalization of Riemannian geometry, where the tangent spaces are not equipped with an quadratic form metric but with a Finsler metric instead. A Finsler metric, $F:\mathcal{M}\times T\mathcal{M}\rightarrow\mathbb{R}_{+}$, where $\mathcal{M}$ is a differentiable manifold, is a Minkowski norm in the sense that $F$ is $C^{\infty}$ on $T\mathcal{M}\setminus\{0\}$, $F(x,cv)=cF(x,v)$ for all $v\in T\mathcal{M}$, $x\in\mathcal{M}$ and $c>0$ and for any $v\in T\mathcal{M}$, then $\frac{\partial^{2}F}{\partial v^{i}\partial v^{j}}(x,v)$ is positive definite [50]. Geodesics on a Finsler manifold can be found equivalently to the Riemannian case by minimizing the energy functional, $\int_{0}^{1}F^{2}\left(\gamma(t),\dot{\gamma}(t)\right)\,\mathrm{d}t$, on a suitable set of curves connecting the two points thereby approximating a geodesic $\gamma$, [50]. GEORCE is similarly applicable to Finsler geometry. First, we note that the energy functional can be written as [50]
$\frac{1}{2}\int_{0}^{1}F^{2}\left(\gamma(t),\dot{\gamma}(t)\right)\,\mathrm{d}t=\int_{0}^{1}\dot{\gamma}(t)^{\top}G\left(\gamma(t),\dot{\gamma}(t)\right)\dot{\gamma}(t)\,\mathrm{d}t,$ (11)
where $G(x,v)=\frac{1}{2}\frac{\partial^{2}F^{2}}{\partial v^{i}\partial v^{j}}\left(x,v\right)$ denotes the fundamental tensor in local coordinates [^2] In literature, the fundamental tensor is often denoted with a lower case $g$. To have the notation consistent with the Riemannian case, we opt to use an upper case $G$ instead., which is positive definite [50]. Unlike the Riemannian case, the fundamental tensor, $G$, is now a function of both the position and velocity, which implies the update formulas in Proposition 2. Similar to Riemannian case, we can formulate the discrete energy functional for the Finsler case as the following discretized energy functional in Eq. 11 becomes
$\begin{split}E_{F}(x_{0:T}):=\quad\min_{x_{0:T}}\quad&\sum_{t=0}^{T-1}(x_{t+1}-x_{t})^{\top}G(x_{t},x_{t+1}-x_{t})(x_{t+1}-x_{t})\\ \text{s.t.}\quad&x_{0}=a,x_{T}=b,\end{split}$ (12)
where the corresponding discrete control problem becomes
$\begin{split}\min_{(x_{1:T-1},u_{0:T-1})}\quad&\sum_{t=0}^{T-1}u_{t}^{\top}G\left(x_{t},u_{t}\right)u_{t},\quad t=0,\dots,T-1,\\ \text{s.t.}\quad&x_{t+1}=x_{t}+u_{t}\\ &x_{0}=a,x_{T}=b,\end{split}$ (13)
Since, the fundamental tensor, $G$, depends on both the state, $x_{0:T}$, and control variable $u_{0:T}$, the necessary conditions in Proposition 1 have to be modified slightly to take this into account. Taking this into account and following the same approach as in the Riemannian case, we get the following update scheme in the Finslerian case (see Appendix A.4 for details)
$\begin{split}&\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\left(\zeta_{t}+\sum_{t>j}^{T-1}\nu_{j}\right)\right),\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\zeta_{t}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a,\end{split}$ (14)
where in iteration $i$ we fix
$\begin{split}&\nu_{t}:={\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{\top}G\left(y,u_{t}\right)u_{t}\right)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)},u_{t}=u_{t}^{(i)}}},\quad t=1,\dots,T-1,\\ &\zeta_{t}:={\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{\top}G\left(x_{t},y\right)u_{t}\right)\vphantom{\big{|}}\right|_{x_{t}=x_{t}^{(i)},u_{t}=u_{t}^{(i)},y=u_{t}^{(i)}}},\quad t=1,\dots,T-1,\\ &G_{t}:=G\left(x_{t}^{(i)},u_{t}^{(i)}\right),\quad t=0,\dots,T-1.\end{split}$
We see that the only extra term in GEORCE for Finsler manifolds is $\{\zeta_{t}\}_{t=0}^{T-1}$. Using the modified update scheme in Eq. 14, we estimate geodesics for Finsler manifolds similar to Algorithm 1. For completeness, we show the algorithm in pseudo-code for the Finsler case in Appendix B.1.
Using the same technique as in the Riemannian case, we show that GEORCE for Finsler manifolds has global convergence and local quadratic convergence. We refer to the details in Appendix A.7.
5 Experiments
Figure 4: The first two figures display the length and runtime of geodesics for $\mathbb{S}^{n}$ estimated using different algorithms from Table 2 and other methods in Appendix E.3. The latter row shows the length and runtime, respectively for the $n$-Ellipsoid with half axes of $n$ equally spaced points between $0.5$ and $1.0$. The optimization solvers are described in Appendix E.1.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2742$ $57.4164\pm 0.0616$ $2.2772$ $0.2335\pm 0.0005$ $\boldsymbol{2.2741}$ $\boldsymbol{0.0078}\pm\boldsymbol{0.0000}$
$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $67.4657\pm 0.4874$ $2.4377$ $0.3475\pm 0.0010$ $2.4342$ $\boldsymbol{0.0635}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{5}$ $2.3757$ $77.6975\pm 0.0741$ $2.4222$ $1.9894\pm 0.0188$ $\boldsymbol{2.3755}$ $\boldsymbol{0.2219}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{10}$ $1.9529$ $150.0621\pm 0.8919$ $2.1016$ $4.4208\pm 0.0308$ $\boldsymbol{1.9523}$ $\boldsymbol{0.3184}\pm\boldsymbol{0.0002}$
$\mathbb{S}^{20}$ $1.4810$ $1068.8645\pm 17.2001$ $1.6839$ $8.7097\pm 0.0352$ $\boldsymbol{1.4808}$ $\boldsymbol{0.6732}\pm\boldsymbol{0.0204}$
$\mathbb{S}^{50}$ $1.0464$ $7350.6001\pm 244.3737$ $1.1463$ $61.8926\pm 0.3034$ $\boldsymbol{0.9749}$ $\boldsymbol{4.1016}\pm\boldsymbol{0.0201}$
$\mathbb{S}^{100}$- -- $0.8466$ $277.4217\pm 0.2280$ $\boldsymbol{0.4187}$ $\boldsymbol{10.9208}\pm\boldsymbol{0.0071}$
$\mathbb{S}^{250}$- -- $0.6244$ $1424.9031\pm 22.1592$ $\boldsymbol{0.3192}$ $\boldsymbol{78.2937}\pm\boldsymbol{0.0991}$
$\mathbb{S}^{500}$- -- $0.5717$ $2242.4697\pm 17.9769$ $\boldsymbol{0.2198}$ $\boldsymbol{224.2761}\pm\boldsymbol{0.4894}$
$\mathbb{S}^{1000}$- -- $0.5279$ $4329.6621\pm 38.0163$ $\boldsymbol{0.1573}$ $\boldsymbol{893.4682}\pm\boldsymbol{7.3130}$
$\mathrm{E}\left(2\right)$ $1.9595$ $56.9650\pm 0.4106$ $1.9596$ $0.2328\pm 0.0001$ $\boldsymbol{1.9595}$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0499}$ $94.1769\pm 0.7044$ $2.0525$ $0.3392\pm 0.0024$ $2.0501$ $\boldsymbol{0.0119}\pm\boldsymbol{0.0000}$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8403}$ $194.5403\pm 1.3293$ $1.9521$ $2.2227\pm 0.0186$ $1.8410$ $\boldsymbol{0.4220}\pm\boldsymbol{0.0038}$
$\mathrm{E}\left(10\right)$ $\boldsymbol{1.5114}$ $427.6278\pm 4.1044$ $1.6384$ $4.5697\pm 0.0541$ $1.5139$ $\boldsymbol{0.4566}\pm\boldsymbol{0.0016}$
$\mathrm{E}\left(20\right)$ $1.1795$ $1669.2692\pm 10.2877$ $1.3049$ $9.4335\pm 0.0507$ $\boldsymbol{0.6548}$ $\boldsymbol{1.9786}\pm\boldsymbol{0.0031}$
$\mathrm{E}\left(50\right)$ $0.8338$ $10685.6133\pm 321.6447$ $0.9244$ $73.5682\pm 0.3306$ $\boldsymbol{0.4642}$ $\boldsymbol{2.4208}\pm\boldsymbol{0.0222}$
$\mathrm{E}\left(100\right)$- -- $0.6872$ $289.4196\pm 2.5224$ $\boldsymbol{0.3249}$ $\boldsymbol{7.7128}\pm\boldsymbol{0.0714}$
$\mathrm{E}\left(250\right)$- -- $0.4978$ $1620.6428\pm 16.9546$ $\boldsymbol{0.2178}$ $\boldsymbol{56.8527}\pm\boldsymbol{0.1526}$
$\mathrm{E}\left(500\right)$- -- $0.4559$ $2135.6899\pm 30.5699$ $\boldsymbol{0.1467}$ $\boldsymbol{220.1656}\pm\boldsymbol{0.9902}$
$\mathrm{E}\left(1000\right)$- -- $0.4232$ $4398.9453\pm 25.8738$ $\boldsymbol{0.1131}$ $\boldsymbol{641.7199}\pm\boldsymbol{20.9901}$
$\mathbb{T}^{2}$ $9.9851$ $92.7400\pm 3.1267$ $9.9855$ $0.1060\pm 0.0002$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0182}\pm\boldsymbol{0.0003}$
$\mathbb{H}^{2}$ $1.2499$ $75.0079\pm 0.1710$ $1.2499$ $0.0988\pm 0.0001$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0025}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0669}$ $88.2739\pm 0.4152$ $1.0879$ $0.2611\pm 0.0004$ $1.0669$ $\boldsymbol{0.0141}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2639}$ $168.5082\pm 0.3039$ $2.2896$ $2.3780\pm 0.0046$ $2.2639$ $\boldsymbol{0.1442}\pm\boldsymbol{0.0002}$
Gaussian Distribution $\boldsymbol{2.5952}$ $47.6288\pm 0.1359$ $2.5956$ $0.0167\pm 0.0001$ $2.5952$ $\boldsymbol{0.0015}\pm\boldsymbol{0.0000}$
Fréchet Distribution $\boldsymbol{1.1538}$ $51.3063\pm 0.0566$ $1.1542$ $0.0185\pm 0.0001$ $1.1539$ $\boldsymbol{0.0009}\pm\boldsymbol{0.0000}$
Cauchy Distribution $1.6454$ $36.5031\pm 0.0437$ $1.6466$ $0.0177\pm 0.0001$ $\boldsymbol{1.6452}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0000}$
Pareto Distribution $0.8297$ $39.0465\pm 0.0471$ $0.8336$ $0.0168\pm 0.0000$ $\boldsymbol{0.8297}$ $\boldsymbol{0.0007}\pm\boldsymbol{0.0000}$
Table 2: The length of the estimated geodesics for the BFGS-algorithm, ADAM and GEORCE on a CPU. The methods were terminated if the $\ell^{2}$-norm of the gradient was less than $10^{-4}$ or after $1,000$ iterations. $\mathrm{E}(n)$ denotes an Ellipsoid of dimension $n$ with half axes of $n$ equally spaced points between $0.5$ and $1.0$, while $\mathcal{P}(n)$ denotes the space of $n\times n$ symmetric positive definite matrices. When the computational time was longer than $24$ hours, the value is set to --.
Riemannian manifolds
Figure 5: The geodesics for an “egg-tray” parameterized by $(x,y,2\cos x\cos y)$, where there are no unique geodesic connecting the points. The BFGS estimate, ADAM estimate and the initial curve overlap completely.
To compare GEORCE with other optimization algorithms, we show the runtime and length of geodesics computed for different manifolds with varying dimensions in Table 2. Here we compare GEORCE with BFGS [37, 38, 39, 40] and ADAM [41]. We terminate the algorithm if the $\ell^{2}$-norm of the gradient of the discretized energy functional in Eq. 5 is less than $10^{-4}$, or if the number of iterations exceeds $1,000$. A full description of the manifolds and experimental details can be found in Appendix C and Appendix E, respectively.
We show the estimated geodesics with $T=100$ grid points and summarize the results for the corresponding runtime and length of the computed geodesics for the $n$-sphere and $n$-ellipsoid in Fig. 4. In Appendix E.3, we show similar run-time experiments for varying grid points as well as other optimization algorithms. We also show the corresponding results on a GPU with other CUDA compatible algorithms. In general, we see that GEORCE is considerably faster with lower length compared to alternative methods.
To test GEORCE on manifolds with multiple length minimizing geodesics, we consider the geodesic between $(-5,-5)$ and $(5,5)$ on an “egg-tray” parameterized by $(x,y,2\cos x\cos y)$ with $T=1,000$ grid points. We show the estimates in Fig. 5, where we see that GEORCE obtains the a signficiantly smaller length, where there is also a clear symmetric to solution to GEORCE along the other side of the “egg-tops”.
Figure 6: The connecting geodesic between four different distributions equipped with the Fisher-Rao metric using GEORCE. From left to right the figures show the connecting geodesic (blue) between two fixed distributions (black) for the Gaussian distribution, Fréchet distribution, Cauchy distribution and the Pareto distribution.
Figure 7: The geodesics computed under the Hessian of the log-partition function of a Gaussian distribution. The background color illustrates the log-partition function.
To illustrate the application of GEORCE to more abstract manifolds, we compute geodesics on the statistical manifolds, i.e., between distributions [17, 18]. Fig. 6 shows the estimated geodesic between a Gaussian distribution, a Fréchet distribution, a Cauchy distribution, and a Pareto distribution using GEORCE all equipped with the Fisher-Rao metric in eq. 4. The metric matrix function for the corresponding statistical manifold is described in [31]. We see that GEORCE converges after only a few iterations. The computed geodesics using alternative methods are summarized in Table 2, where we see that GEORCE obtains significantly shorter length much faster, also in high-dimension, compared to alternative methods.
We further illustrate the application of GEORCE to information geometry by considering potential functions encountered in Bregman geometry [18]. We consider the log-partition function for the Gaussian distribution, which is convex and smooth
$A(\theta_{1},\theta_{2})=\frac{\theta_{1}^{2}}{2\theta_{2}}+\frac{1}{2}\log\left(\frac{2\pi}{\theta_{2}}\right),$
where $\theta_{1}=\frac{\mu}{\sigma^{2}}$ and $\theta_{2}=\frac{1}{\sigma^{2}}$ for a Gaussian distribution with parameters $(\mu,\sigma)$ such that the density takes the form of an exponential family $p(x)=h(x)\exp\left(\theta_{1}x-A(\theta_{1},\theta_{2})\right)$ for a suitable $h$. We consider the corresponding Hessian of the log-partition function, which defines a Riemannian metric. Fig. 7 shows geodesics computed with both GEORCE and alternative methods using at most $100$ iterations. We see that GEORCE uses less iterations and obtains the smallest length compared to the alternative methods.
Figure 8: Comparison between GEORCE and baseline methods for computing geodesics on four different Finsler manifolds. The Finsler metrics stem from a Riemannian background metric with the force field in eq. 15 on them. The length of the force field is displayed in color, while the arrows show the direction of the force field. The algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional is less than $10^{-4}$, or if the number of iterations exceeds $100$. We see that GEORCE uses less iterations and obtains the smallest length.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.6525$ $244.9129\pm 0.0680$ $1.6536$ $0.4967\pm 0.0007$ $\boldsymbol{1.6525}$ $\boldsymbol{0.0194}\pm\boldsymbol{0.0009}$
$\mathbb{S}^{3}$ $1.7433$ $277.8015\pm 0.5819$ $1.7456$ $0.7639\pm 0.0031$ $\boldsymbol{1.7432}$ $\boldsymbol{0.0362}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{5}$ $1.7032$ $435.6987\pm 4.6022$ $1.7377$ $4.6711\pm 0.0047$ $\boldsymbol{1.7029}$ $\boldsymbol{0.4556}\pm\boldsymbol{0.0021}$
$\mathbb{S}^{10}$ $1.4151$ $812.2809\pm 13.5968$ $1.5174$ $9.7974\pm 0.0052$ $\boldsymbol{1.4109}$ $\boldsymbol{0.7473}\pm\boldsymbol{0.0029}$
$\mathbb{S}^{20}$ $1.0951$ $1446.0956\pm 4.8038$ $1.2018$ $18.3489\pm 0.0747$ $\boldsymbol{1.0760}$ $\boldsymbol{1.2856}\pm\boldsymbol{0.0016}$
$\mathbb{S}^{50}$ $0.8216$ $5992.7261\pm 56.5973$ $0.8128$ $146.4556\pm 0.9341$ $\boldsymbol{0.7285}$ $\boldsymbol{8.4758}\pm\boldsymbol{0.0234}$
$\mathbb{S}^{100}$- -- $0.6187$ $498.9705\pm 0.4032$ $\boldsymbol{0.2570}$ $\boldsymbol{15.3786}\pm\boldsymbol{0.0092}$
$\mathbb{S}^{250}$- -- $0.5195$ $1397.6732\pm 0.7793$ $\boldsymbol{0.1622}$ $\boldsymbol{132.5889}\pm\boldsymbol{0.1106}$
$\mathbb{S}^{500}$- -- $0.4850$ $1954.3231\pm 14.1107$ $\boldsymbol{0.1404}$ $\boldsymbol{364.5326}\pm\boldsymbol{1.8222}$
$\mathbb{S}^{1000}$- -- $0.3843$ $5982.0884\pm 204.1833$ $\boldsymbol{0.1140}$ $\boldsymbol{1338.9869}\pm\boldsymbol{2.3954}$
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3616}$ $230.0809\pm 1.3967$ $1.3621$ $0.5310\pm 0.0025$ $1.3618$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.4001}$ $416.6281\pm 5.5354$ $1.4087$ $0.8183\pm 0.0026$ $1.4023$ $\boldsymbol{0.1081}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2951}$ $638.7185\pm 2.7818$ $1.3265$ $4.3802\pm 0.0039$ $1.2956$ $\boldsymbol{0.4554}\pm\boldsymbol{0.0014}$
$\mathrm{E}\left(10\right)$ $\boldsymbol{1.0733}$ $741.9274\pm 12.3382$ $1.1229$ $9.7899\pm 0.0117$ $1.0804$ $\boldsymbol{0.5815}\pm\boldsymbol{0.0022}$
$\mathrm{E}\left(20\right)$ $0.8510$ $1820.3137\pm 4.0734$ $0.8980$ $19.4614\pm 0.0039$ $\boldsymbol{0.8121}$ $\boldsymbol{1.0671}\pm\boldsymbol{0.0044}$
$\mathrm{E}\left(50\right)$ $0.5981$ $11191.9424\pm 597.7002$ $0.6325$ $146.6409\pm 0.1668$ $\boldsymbol{0.2842}$ $\boldsymbol{4.4377}\pm\boldsymbol{0.0451}$
$\mathrm{E}\left(100\right)$- -- $0.4936$ $491.8223\pm 0.2346$ $\boldsymbol{0.2009}$ $\boldsymbol{14.5367}\pm\boldsymbol{0.0092}$
$\mathrm{E}\left(250\right)$- -- $0.4159$ $1475.5049\pm 4.1430$ $\boldsymbol{0.1287}$ $\boldsymbol{125.4187}\pm\boldsymbol{0.0888}$
$\mathrm{E}\left(500\right)$- -- $0.4142$ $1984.7845\pm 10.5689$ $\boldsymbol{0.0967}$ $\boldsymbol{373.2816}\pm\boldsymbol{0.4227}$
$\mathrm{E}\left(1000\right)$- -- $0.3444$ $5929.4438\pm 167.1873$ $\boldsymbol{0.0799}$ $\boldsymbol{1211.8812}\pm\boldsymbol{0.2511}$
$\mathbb{T}^{2}$ $7.3890$ $449.4798\pm 10.3282$ $7.3585$ $0.3294\pm 0.0018$ $\boldsymbol{6.8067}$ $\boldsymbol{0.1126}\pm\boldsymbol{0.0002}$
$\mathbb{H}^{2}$ $1.3340$ $384.3149\pm 0.3044$ $1.3360$ $0.2744\pm 0.0024$ $\boldsymbol{1.3337}$ $\boldsymbol{0.0094}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7291}$ $414.8355\pm 3.0837$ $0.7338$ $0.6368\pm 0.0026$ $0.7292$ $\boldsymbol{0.0222}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $536.6575\pm 0.9323$ $1.2980$ $5.2824\pm 0.0540$ $1.2856$ $\boldsymbol{0.2094}\pm\boldsymbol{0.0004}$
Gaussian Distribution $1.6878$ $163.4399\pm 0.1859$ $1.6936$ $0.1145\pm 0.0001$ $\boldsymbol{1.6878}$ $\boldsymbol{0.0047}\pm\boldsymbol{0.0001}$
Fréchet Distribution $0.5408$ $123.8418\pm 0.0499$ $0.5406$ $0.0585\pm 0.0001$ $\boldsymbol{0.5405}$ $\boldsymbol{0.0021}\pm\boldsymbol{0.0001}$
Cauchy Distribution $1.1134$ $182.3980\pm 2.2559$ $\boldsymbol{1.1133}$ $0.1105\pm 0.0007$ $1.1133$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0002}$
Pareto Distribution $0.3884$ $84.8186\pm 0.0832$ $0.3885$ $0.0235\pm 0.0007$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0001}$
Table 3: The length of the estimated geodesics for the BFGS-algorithm, ADAM-gradient descent and GEORCE on a CPU. The algorithms have been stopped if the $\ell^{2}$-norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations have been reached. $\mathrm{E}(n)$ denotes an Ellipsoid of dimension $n$ with half axes of $n$ equally spaced points between $0.5$ and $1.0$, while $\mathcal{P}(n)$ denotes the space of $n\times n$ symmetric positive definite matrices. When the computational time was longer than $24$ hours, the value is set to --.
Finsler manifolds
To illustrate our method for different Finsler manifolds and dimensions, we consider a microswimmer on a Riemannian manifold $m$ that moves with constant self-propulsion velocity $v$ with Riemannian length, $v_{0}:=||v||_{g}$ [51] and we assume that this movement is affected by a generic force field, $f$, of the form
$f(x)=\frac{\sin x\odot\cos x}{(\cos x)^{\top}G(x)\cos x},$ (15)
where $\cos x$ and $\sin x$ denote the element-wise function for each entry of the local coordinates $x$, while $G$ denotes the corresponding Riemannian background metric such that the force field has $G$-length less than or equal to 1. If $||f(x)||_{g}<v_{0}$, then the optimal travel time of the microswimmer corresponds to a geodesic for a Finsler manifold of the form [51]
$F=\sqrt{a_{ij}\dot{r}^{i}\dot{r}^{j}}+b_{i}\dot{r}^{i},$
where $a_{ij}=g_{ij}\lambda+f_{i}f_{j}\lambda^{2}$, $b_{i}=-f_{i}\lambda$, $b_{i}=-f_{i}\lambda$, $f_{i}=g_{ij}f^{j}$ and $\lambda^{-1}=v_{0}^{2}-g_{ij}f^{i}f^{j}$ [51], while $r(t)$ denotes the position of the microswimmer at time $t$. We refer to [51] and Appendix C for details. We consider the same background metrics in Table 2 all equipped with the generic force field in eq. 15 and compare GEORCE to BFGS [37, 38, 39, 40] and ADAM [41]. Fig. 8 shows the estimated geodesics for four different Riemannian manifolds equipped with the generic force field in eq. 15. We show the estimated geodesics for $T=100$ grid points in Table 3. In Appendix E.3 we show similar runtime experiments for varying grid points as well as other optimization algorithms. We see that GEORCE is considerably faster and more accurate than the alternatives.
Applications to generative models
To illustrate the application of GEORCE to learned manifolds, we compute geodesics for generative models and compare them to similar methods. A variational autoencoder (VAE) is a generative model that learns the data distribution, $\mathcal{X}\subset\mathbb{R}^{D}$, and a latent representation of data, $\mathcal{Z}\subset\mathbb{R}^{d}$, with $d<D$ by maximizing a lower bound of the log-likelihood of the data with respect to a variational family of distributions [52]
For latent space generative models, the decoder can be seen as a smooth $d$ dimensional immersion, $f:\mathcal{Z}\rightarrow\mathcal{X}$, which defines a Riemannian geometry with the pull-back metric $G(z)=J_{f}^{\top}(z)J_{f}(z)$, where $J_{f}$ denotes the Jacobian of $z$ [11, 12].
We train a VAE for the CelebA dataset [53], which consists of $202{,}599$ facial images of size $64\times 64\times 3$, where the architecture is described in Appendix D.1. Fig. 9 illustrates the geodesics constructed between two images reconstructed using the variational autoencoder. We see that GEORCE obtains a smaller length than comparable methods for $T=100$ grid points. We refer to Appendix D.2 for runtime results for the VAE, where we see that GEORCE is significantly faster than alternative methods. Note that the tolerance for convergence is $10^{-3}$, since it was observed that it was not possible for any method to obtain an $\ell^{2}$-norm of the gradient less than $10^{-4}$. This is most likely due to potential instabilities for the learned manifold. We summarize the results in Table 11 in Appendix E.3, where we also consider the geodesic with a VAE trained on MNIST data [54] and SVHN data [55]. For the exact architecture and further experiments of the VAE, we refer to Appendix D.1 and Appendix D.2.
Figure 9: Geodesics for a variational autoencoder [52] trained on the CelebA dataset [53] using GEORCE and alternative methods. All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional in eq. 5 is less than $10^{-3}$, or if the number of iterations exceeds $100$. The algorithms are written to the left with number of iterations in parenthesis and length to the right. The first-axis shows the grid number for the image sequence. The details of the manifolds and experiments can be found Appendix D.1
6 Conclusion
In this paper, we have introduced the algorithm GEORCE for estimating geodesics between any given points on Riemannian and Finsler manifolds by formulating the discretized energy functional as a control problem. We have proved that GEORCE has global convergence and quadratic local convergence, and thus, GEORCE exhibits similar global convergence as gradient descent as well as local quadratic convergence similar to the Newton method without using second order derivatives. GEORCE scales cubicly in the manifold dimension and only linearly in the number of grid points. Empirically, we have shown that GEORCE exhibits superior convergence in terms of runtime and accuracy compared to other geodesic optimization algorithms for Riemannian and Finsler manifolds across both dimensions and manifolds.
Limitations: GEORCE exhibits both global convergence and local quadratic convergence requiring few iterations. However, GEORCE requires evaluating the inverse matrix of the metric matrix function in each iteration, which can be numerically expensive. In addition, GEORCE uses line search to determine the step size in each iteration, and if it is numerically expensive to evaluate the discretized energy functional, then this can potentially increase the runtime of GEORCE. Further research should therefore investigate efficient approximations of the inverse metric matrix function as well as possible adaptive updating schemes for the step size to increase computational efficiency for high-dimensional problems.
[…]
Supplementary information
The supplementary information in the paper consists of the appendix for proofs, details on experiments and manifolds as well as additional experiments.
[…]
Acknowledgements Many thanks to Prof. Steen Markvorsen and Dr. Andreas A. Bock for comments and corrections in the process of writing this paper. Especially, a thanks to Prof. Steen Markvorsen for hinting at the possible extension to Finsler geometry. SH is supported by a research grant (42062) from VILLUM FONDEN. This project received funding from the European Research Council (ERC) under the European Union’s Horizon research and innovation programme (grant agreement 101125993). The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606).
Declarations
  • Funding: Research grant 42062 from VILLUM FONDEN; funding from the European Research Council (ERC) under the European Union’s Horizon research and innovation programme (grant agreement 101125993); funding from the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606).
  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use): NA
  • Ethics approval and consent to participate: NA
  • Consent for publication: Currently under review.
  • Data availability: All data used are described in the appendix.
  • Materials availability: NA
  • Code availability: The code can be found at https://github.com/FrederikMR/georce
  • Author contribution: FMR derived the algorithm and implemented the method as well as benchmarks. SH supervised the project with comments and corrections. The paper was written jointly by SH and FMR. All authors reviewed the manuscript.
Appendix A Proofs and derivations
A.1 Assumptions
In the following, we will prove local quadratic convergence under the following assumptions.
Assumption 5.
We assume the following regarding the discretized energy functional.
  • A (local) minimum point $z^{*}=(x^{*},u^{*})$ to the discretized energy functional is defined as strongly unique minimum point if and only if
    $\exists\epsilon>0:\quad\exists K>0:\quad\forall z\in B_{\epsilon}(z^{*}):\quad E(z)-E(z^{*})\geq K\left\lVert z-z^{*}\right\rVert,$
    where $B_{\epsilon}(z^{*})=\left\{z|\left\lVert z-z^{*}\right\rVert<\epsilon\right\}$.
  • We assume that the discretized energy functional $E(z)$ is locally Lipschitz.
  • The first order Taylor approximation of the discretized energy functional is
    $\Delta E=\langle\nabla E(z_{0}),\Delta z\rangle+\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert,$
    i.e., we assume that $\mathcal{O}\left(\Delta z\right)$ is independent of $z_{0}$, and the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{\tilde{O}}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally (which follows from the Lipschitz condition).
Further for global convergence we will assume Lipschitz continuity of the discretized energy functional. Note that since the metric varies smoothly on the manifold, then the discretized energy functional is also smooth and hence Lipschitz. Thus, all Finsler and Riemannian manifolds satisfies the above conditions, but an even larger class of energy functionals that does not necessarily stem from a Riemannian or Finsler metric will satisfy the requirements.
A.2 Proof of necessary conditions
Proposition 6.
The necessary conditions for a minimum in Eq. 6 is
$\begin{split}&2G(x_{t})u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &{\left.\kern-1.2pt\nabla_{y}\left[u_{t}^{\top}G(y)u_{t}\right]\vphantom{\big{|}}\right|_{y=x_{t}}}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1.\\ &x_{0}=a,x_{T}=b.\end{split}$ (16)
where $\mu_{t}\in\mathbb{R}^{d}$ for $t=0,\dots,T-1$.
Proof.
The Hamiltonian of Eq. 6 is
$H_{t}(x_{t},u_{t},\mu_{t})=u_{t}^{\top}G(x_{t})u_{t}+\mu_{t}^{\top}(x_{t}+u_{t}).$ (17)
The Hamiltonian $H_{t}(x_{t},u_{t},\mu_{t})$ is strictly convex in $u_{t}$, since $G(x_{t})$ is positive definite. The time-discrete version of Pontryagins maximum principle [56] then yields the following when applied to our control problem (6)
$\begin{split}&2G(x_{t})u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &{\left.\kern-1.2pt\nabla_{y}\left[u_{t}^{\top}G(y)u_{t}\right]\vphantom{\big{|}}\right|_{y=x_{t}}}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1.\\ &x_{0}=a,x_{T}=b.\end{split}$ (18)
A.3 Proof of update scheme for Riemannian geometry
Proposition 7.
The update scheme in eq. 7 and eq. 9 for $u_{t},\mu_{t}$ and $x_{t}$ can in closed-form be written as
$\begin{split}&\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
Proof.
From the update formula $\nu_{t}+\mu_{t}=\mu_{t-1}$ for $t=1,\dots,T-1$ we deduce that $\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}$. From the update scheme $02G_{t}u_{t}+\mu_{t}=$ 0 we deduce that $u_{t}=-\frac{1}{2}G_{t}^{-1}\mu_{t}$, since $G_{t}$ is positive definite and hence has a well-defined inverse. Noting that $\sum_{t=0}^{T-1}u_{t}=b-a$ we can re-write the iterative formulas as
$\begin{split}&\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j},\quad t=0,\dots,T-1,\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &-\frac{1}{2}\sum_{t=0}^{T-1}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t+1}^{T-1}\nu_{j}\right)=b-a.\end{split}$
The last equation can be used to solve for $\mu_{T-1}$ utilizing that the sum of (inverse) positive definite matrices is positive definite and has a well-defined inverse
$\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),$
which proves the result. ∎
A.4 Proof of update scheme for Finslerian geometry
Consider the following control problem in the Finslerian case.
$\begin{split}\min_{x_{0:T}}\quad&\sum_{t=0}^{T-1}u_{t}^{\top}G\left(x_{t},u_{t}\right)u_{t},\quad t=0,\dots,T-1,\\ \text{s.t.}\quad&x_{t+1}=x_{t}+u_{t}\\ &x_{0}=a,x_{T}=b,\end{split}$ (19)
By similar computation as in Proposition. 1 with $G$ is now dependent on both $x_{t}$ and $u_{t}$, we get the following equation system to be solved in iteration $i+1$ derived from Pontryagin’s maximum principle by fixing variables
$\begin{split}&\nu_{t}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1,\\ &2G_{t}u_{t}+\zeta_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &\sum_{t=0}^{T-1}u_{t}=b-a,\\ &\nu_{t}:={\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{\top}G\left(y,u_{t}\right)u_{t}\right)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)},u_{t}=u_{t}^{(i)}}},\quad t=1,\dots,T-1,\\ &\zeta_{t}:={\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{\top}G\left(x_{t},y\right)u_{t}\right)\vphantom{\big{|}}\right|_{x_{t}=x_{t}^{(i)},u_{t}=u_{t}^{(i)},y=u_{t}^{(i)}}},\quad t=1,\dots,T-1,\\ &G_{t}:=G\left(x_{t}^{(i)},u_{t}^{(i)}\right),\quad t=0,\dots,T-1,\end{split}$ (20)
The related state variables (before line search) are found by the state equation, $x_{t+1}=x_{t}+u_{t}$ for $t=0,\dots,T-1$ with $x_{0}=a$. Using the same approach as in the proof of Proposition 2, then Eq. 20 is equivalent to
$\begin{split}&\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\left(\zeta_{t}+\sum_{t>j}^{T-1}\nu_{j}\right)\right),\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\zeta_{t}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a,\end{split}$ (21)
where it is exploited that the fundamental tensor is positive definite.
A.5 Proof of global convergence
To prove global convergence for the algorithm GEORCE we first show existence and uniqueness by proving two lemmas. Secondly, we prove that in each iteration the discretized energy functional is decreasing, unless a (local) minimum has been found. We will assume Lipschitz continuity of the metric tensor as described in 5.
[…]
Lemma 8.
Assume that $(x^{(i)},u^{(i)})$ is a feasible solution, then
  1. 1.
    There exists a unique solution $\left(x^{(i+1)},u^{(i+1)}\right)$ to the system of equations in eq. 10 based on $\left(x^{(i)},u^{(i)}\right)$.
  2. 2.
    All linear combinations $(1-\alpha)\left(x^{(i)},u^{(i)}\right)+\alpha\left(x^{(i+1)},u^{(i+1)}\right)$ for $0<\alpha\leq 1$ are feasible solutions.
Proof.
Since the matrices, $\left\{G\left(x_{t}^{(i)}\right)\right\}_{t=0}^{T}$ are positive definite, then the matrices are regular with unique inverse, and so are the sum of the inverse matrices. This means that there exists an unique solution to the update scheme in proposition 2 proving 1.
Since the solution $\left(x^{(i+1)},u^{(i+1)}\right)$ is also a feasible solution for the update scheme in proposition 2, then the linear combination of two feasible solutions will also be a feasible solution, i.e.
$\sum_{t=0}^{T-1}u_{t}^{(j)}=(b-a),\quad j=1,\dots,(i+1).$
The new solution based on the linear combination gives
$\begin{split}&(1-\alpha)\sum_{t=0}^{T-1}u_{t}^{(i)}+\alpha\sum_{t=0}^{T-1}u_{t}^{(i+1)}=(1-\alpha)(b-a)+\alpha(b-a)=(b-a)\\ &(1-\alpha)x^{(i)}+\alpha x^{(i+1)}=(1-\alpha)\left(\alpha,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b\right)+\alpha\left(a,x_{1}^{(i+1)},\dots,x_{T-1}^{(i+1)},b\right)\\ &=\left(a,(1-\alpha)x_{1}^{(i)}+\alpha x_{1}^{(i+1)},\dots,(1-\alpha)x_{T-1}^{(i)}+\alpha x_{T-1}^{(i+1)},b\right),\end{split}$
where for any $t=0,\dots,T-1$
$\begin{split}(1-\alpha)x_{t}^{(i)}+\alpha x_{t}^{(i+1)}&=(1-\alpha)\left(a+\sum_{j=0}^{t-1}u_{j}^{(i)}\right)+\alpha\left(a+\sum_{j=0}^{t-1}u_{j}^{(i+1)}\right)\\ &=a+\sum_{j=0}^{t-1}\left((1-\alpha)u_{j}^{(i)}+\alpha u_{j}^{(i+1)}\right).\end{split}$
This shows that the linear combination in the state is feasible as it produces feasible state variables in terms of start and end point. Furthermore, each state variable is feasible, if they were determined from the linear combination of the control vectors, which proves that the linear combinations of the state and control variables are also feasible solutions. ∎
Lemma 9.
Let $\left\{x_{t}^{(i)},u_{t}^{(i)}\right\}_{t=0}^{T}$ denote the feasible solution after iteration $i$ in GEORCE. If $\left\{x_{t}^{(i)},u_{t}^{(i)}\right\}_{t=0}^{T}$ is not a (local) minimum point, then the feasible solution from iteration $\left\{x_{t}^{(i+1)},u_{t}^{(i+1)}\right\}_{t=0}^{T}$ will decrease the objective function in the sense that there exists an $\eta>0$ such that for all $\alpha$ with $0<\alpha\leq\eta\leq 1$, then
$E\left(x^{(i)}+\alpha\left(x^{(i+1)}-x^{(i)}\right),u^{(i)}+\alpha\left(u^{(i+1)}-u^{(i)}\right)\right)<E\left(x^{(i)},u^{(i)}\right),$
where $x^{(i)}=\left(a,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b\right)$ and $u^{(i)}=\left(u_{0}^{(i)},u_{1}^{(i)},\dots,u_{T-2}^{(i)},u_{T-1}^{(i)}\right)$.
Proof.
$E(x,u)$ is a smooth function. The first order Taylor approximation of the energy function $E(x,u)$ in eq. 5 in $(x,u)$ is
$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{x_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta x_{t}\right\rangle+\mathcal{O}(\Delta x_{t})\left\lVert\Delta x_{t}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right),\end{split}$
where $\Delta u_{t}:=u_{t}^{(i+1)}-u_{t}^{(i)}$ and $\Delta x_{t}:=x_{t}^{(i+1)}-x_{t}^{(i)}$. By the optimally conditions in eq. 7 we have that
$\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{t-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=-\mu_{t},\quad t=0,\dots,T-1.\end{split}$ (22)
Inserting the optimality conditions in the Taylor expansion we have that
$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\langle\mu_{t-1}-\mu_{t},\Delta x_{t}\rangle+\mathcal{O}(\Delta x_{t})\left\lVert\Delta x_{t}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$
Since $\Delta x_{t}=\sum_{j=0}^{t-1}\Delta u_{j}$, then
$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\langle\mu_{t-1}-\mu_{t},\sum_{j=0}^{t-1}\Delta u_{j}\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$
Re-arranging the terms we get
$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-2}\left\langle\mu_{t},\sum_{j=0}^{t}\Delta u_{j}\right\rangle\\ &-\sum_{t=1}^{T-1}\left(\left\langle\mu_{t},\sum_{j=0}^{t-1}\Delta u_{j}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert u_{t}\right\rVert\right)\\ &=\langle\mu_{0},\Delta u_{0}\rangle+\sum_{t=1}^{T-2}\langle\mu_{t},\Delta u_{t}\rangle-\left\langle\mu_{T-1},\sum_{j=0}^{T-1}\Delta u_{j}\right\rangle+\langle\mu_{T-1},\Delta u_{T-1}\rangle\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\end{split}$
Since $\Delta u_{j}=u_{j}^{(i+1)}-u_{j}^{(i)}$ and $\Delta x_{j}=x_{j}^{(i+1)}-x_{j}^{(i)}$ for $j=0,\dots,T$, and since the solution is feasible we have that $\sum_{j=0}^{T-1}u_{j}^{(i)}=b-a$ for any $i$, which implies that $\sum_{j=0}^{T-1}\Delta u_{j}=0$. This reduces the above expression to
$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left(\langle\mu_{t},\Delta u_{t}\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\\ &=\sum_{t=0}^{T-1}\left\langle\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\sum_{t=0}^{T-1}\mathcal{O}\left(\Delta u_{t}\right)\left\lVert u_{t}\right\rVert.\end{split}$
By eq. 22 we have
$\begin{split}\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=-{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}\\ &+{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}\\ &=-2G\left(x_{t}^{(i)}\right)\left(u_{t}^{(i+1)}-u_{t}^{(i)}\right),\end{split}$ (23)
which implies that
$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$ (24)
Since $G\left(x_{t}^{(i)}\right)$ is positive definite, the first summation is negative assuming that at least one non-zero vector $\{\Delta u_{t}\}_{t=0}^{T-1}$, which is the case, since we have assumed that $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ is not a (local) minimum. Now scale all vectors, $\left\{\Delta u_{t}\right\}_{t=0}^{T-1}$ with a scalar $10<\alpha<$ 1. Note that the term, $-2G\left(x_{t}^{(i)}\right)\Delta u_{t}$, is unaffected by this, since it equals $\mu_{t}+\nabla_{u_{t}}E(x,u)\left(x_{t}^{(i)},u_{t}^{(i)}\right)$, which is independent of $\alpha$. The first order Taylor order approximation is
$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\alpha\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\alpha\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\alpha\Delta u_{j}\right\rVert+\mathcal{O}\left(\alpha\Delta u_{t}\right)\left\lVert\alpha\Delta u_{t}\right\rVert\right).\end{split}$
In the limit it now follows that there exists an $\eta>0$ such that $\frac{\Delta E(x,u)}{\alpha}<0$ for some $\alpha$ with $0<\alpha\leq\eta\leq 1$, which implies that $\Delta E(x,u)<0$. ∎
Theorem 10.
Let $E^{(i)}$ be the value of the discretized energy functional for the solution after iteration $i$ (with line search) in GEORCE. If the starting point $\left(x^{(0)},u^{(0)}\right)$ is feasible, then the series $\left\{E^{(i)}\right\}_{i>0}$ will converge to a (local) minimum.
Proof.
If the starting point $\left(x^{(0)},u^{(0)}\right)$ in GEORCE is feasible, then the points $\left(x^{(i)},u^{(i)}\right)$ after each iteration (with line-search) in GEORCE will also be feasible solutions by lemma 8.
The discretized energy functional is a positive function, i.e. a lower bound is at least 0. Assume that the series $\left\{E^{(i)}\right\}_{i>0}$ is not converging to a (local) minimum. From lemma 9 it follows that $\left\{E^{(i)}\right\}_{i>0}$ is decreasing for increasing $i$, and since $E^{(i)}$ has a lower bound, the series $\left\{E^{(i)}\right\}_{i}$ can not be non-converging. Assume now that the convergence value for $\left\{E^{(i)}\right\}_{i}$, denoted $\hat{E}$, is not a (local) minimum, and denote the convergence point $(\hat{x},\hat{u})$. According to lemma 9 GEORCE will from the solution $(\hat{x},\hat{u})$ in the following iteration produce a new point $(\tilde{x},\tilde{u})$ such that $E(\tilde{x},\tilde{u})<\hat{E}$, which is contradiction as the series $\left\{E^{(i)}\right\}_{i}$ is assumed to be converging to $\hat{E}$. The series, $\left\{E^{(i)}\right\}_{i}$, will therefore converge, and the convergence point is a (local) minimum. ∎
Under the assumption of a strongly unique minimum point in Appendix A.1 GEORCE will converge to a local minimum by 3. Thus, it is not possible that GEORCE will ”jump” between local minimas.
A.6 Proof of local quadratic convergence
In this section we show that GEORCE has locally quadratic convergence to a (local) minimum under certain assumptions satisfied by the discretized energy functional. We will assume the assumptions in 5 hold. We first prove the following result, before proving quadratic local convergence.
Lemma 11.
Assume $f(x)$ is a smooth function. Let $z^{*}$ be a strongly unique (local) minimum point for $f$. Then
$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}(z^{*})\setminus\{z^{*}\}:\quad\langle\nabla f(z),z^{*}-z\rangle<0.$
Proof.
Since $f(x)$ is a smooth function, then the first order Taylor approximation for $f(x)$ exists. Consider a point $z\in B_{\epsilon}(z^{*})\setminus\{z^{*}\}$, then
$f(z^{*})-f(z)=\langle\nabla f(z),z^{*}-z\rangle+\mathcal{O}\left(z^{*}-z\right)\left\lVert z^{*}-z\right\rVert.$
Since $z^{*}$ is a strongly unique (local) minimum point, then
$f(z^{*})-f(z)\leq-K\left\lVert z^{*}-z\right\rVert<0,\quad K>0,$
which implies that
$\frac{f(z^{*})-f(z)}{\left\lVert z^{*}-z\right\rVert}=\left\langle\nabla f(z),\frac{z^{*}-z}{\left\lVert z^{*}-z\right\rVert}\right\rangle+\mathcal{O}\left(z^{*}-z\right)\leq-K.$
Since $\lim_{z\rightarrow z^{*}}\mathcal{O}(z^{*}-z)=0$, then there exists $\epsilon>0$ such that
$\left\langle\nabla f(z),\frac{z^{*}-z}{\left\lVert z^{*}-z\right\rVert}\right\rangle<0,$
which implies that $\langle\nabla f(z),z^{*}-z\rangle<0$ for some $\epsilon>0$ for all $z\in B_{\epsilon}\left(z^{*}\right)\setminus\left\{z^{*}\right\}$. ∎
Proposition 12.
If the discretized energy functional has a strongly unique (local) minimum point $z^{*}$ and locally $\alpha^{*}=1$, i.e. no line search, then GEORCE has locally quadratic convergence, i.e.
$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
where $z^{(i)}=(x^{(i)},u^{(i)})$ is the solution from GEORCE at iteration $i$.
Proof.
Assume that $z^{*}$ is a (local) minimum point. Since GEORCE has global convergence to a (local) minimum point by theorem 3, we will assume that the algorithm has convergence point $E(z^{*})$. Consider iteration $i$ with $z^{(i)}=(x^{(i)},u^{(i)})$ from GEORCE, and assume that all following iterations in GEORCE is in the ball $B_{\epsilon_{1}}(z^{*})$, and fulfills the strongly unique assumption in 5. Note that as the series $\left\{z^{(i)}\right\}$ converge to $z^{*}$, then $\left\{z^{(i)}\right\}_{i\geq k}\subset B_{\epsilon_{1}}(z^{*})$ for some $k>0$. It then follows from the assumption of strongly unique minimizing point that for some $K>0$
$E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\geq K\left\lVert z^{(i+1)}-z^{*}\right\rVert,$
since $z^{(i+1)}\in B_{\epsilon_{1}}(z^{*})$. The left hand term above can be rearranged into
$E\left(z^{(i+1)}\right)-E\left(z^{*}\right)=\left(E\left(z^{(i+1)}\right)-E\left(z^{(i)}\right)\right)+\left(E\left(z^{(i)}\right)-E\left(z^{*}\right)\right).$
By eq. 24 in the proof of lemma 9 and utilizing $\alpha^{*}=1$, it follows that
$\begin{split}&E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\\ &=\left(E\left(z^{(i+1)}\right)-E\left(z^{(i)}\right)\right)+\left(E\left(z^{(i)}\right)-E\left(z^{*}\right)\right)\\ &=\sum_{t=0}^{T-1}\left(\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\\ &-\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\left(u_{t}^{*}-u_{t}^{(i)}\right)\right\rangle\\ &-\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert-\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\\ &=\sum_{t=0}^{T-1}\left(\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}-\left(u_{t}^{*}-u_{t}^{(i)}\right)\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert+\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\right)\\ &=\sum_{t=0}^{T-1}\left(\left\langle 2G\left(x_{t}^{(i)}\right)\Delta u_{t},u_{t}^{*}-u_{t}^{(i+1)}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert+\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\right).\end{split}$
Observe now that the term $2G\left(x_{t}^{(i)}\right)\Delta u_{t}$ is the gradient for $z_{t}^{(i)}$. Define $\epsilon_{2}$ such that the $\mathcal{O}$-terms are sufficiently close to zero for all $z\in B_{\epsilon_{2}}(z^{*})$. Thus, for $\epsilon=\min\left\{\epsilon_{1},\epsilon_{2}\right\}$ we have from lemma 11
$\sum_{t=0}^{T-1}\left\langle 2G\left(x_{t}^{(i)}\right)\Delta u_{t},u_{t}^{*}-u_{t}^{(i+1)}\right\rangle<0,$
since $z^{*}$ is assumed to be a strongly unique minimum point by assumptions 5. Combining the inequalities and utilizing that $E(z)$ is locally Lipschitz continuous by assumptions 5, then
$K\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\leq\mathcal{O}\left(\left\lVert z^{i}-z^{*}\right\rVert^{2}\right),$
which implies that there exists an $\epsilon>0$ such that
$\forall z^{(i)}\in B_{\epsilon}\left(z^{*}\right):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
for $c>0$. ∎
The assumption in the above proof that the line search parameter, $\alpha^{*}=1$ locally at the minimum point is based on that locally the approximation of the discretized energy functional becomes close to the true function.
A.7 Extension to Finslerian manifolds
The proof for global convergence and local quadratic convergence hold similar for GEORCE in the Finsler case. Let $E_{F}$ denote the corresponding discretized Finsler energy functional. For the Finsler case the optimally condition becomes
$\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E_{F}(x_{t},u_{t0})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{T-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=2G\left(x_{t}^{(i)},u_{t}^{(i)}\right)u_{t}^{(i+1)}+\zeta_{t}=-\mu_{t},\\ &\quad t=0,\dots,T-1\end{split}$ (25)
The proof of global and local quadratic convergence are completely similar for the Riemannian and Finsler case except that eq. 23 in the Finsler case becomes
$\begin{split}\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=-{\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i+1)},u_{t}^{(i+1)}\right)}}\\ &+{\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i+1)},u_{t}^{(i)}\right)}}\\ &=-\left(G\left(x_{t}^{(i)},u_{t}^{(i)}\right)u_{t}^{(i+1)}+\zeta_{t}\right)\\ &+\left(2G\left(x_{t}^{(i)},u_{t}^{(i)}\right)+\zeta_{t}\right)\\ &=-2G\left(x_{t}^{(i)},u_{t}^{(i)}\right)\left(u_{t}^{(i+1)}-u_{t}^{(i)}\right),\end{split}$ (26)
by using the properties of GEORCE in the Finsler case in eq. 25. Eq. 26 is completely similar to eq. 23 in the Riemannian case, and from that the global convergence and local quadratic convergence proofs are completely similar to the Riemannian case.
Appendix B Algorithms
B.1 GEORCE for Finsler manifolds
In Algorithm 2 we show GEORCE for Finsler manifolds in pseudo-code using the update scheme in Eq. 14.
Algorithm 2 GEORCE for Finsler Manifolds
1:Input: $\mathrm{tol}$, $T$
2:Output: Geodesic estimate $x_{0:T}$
3:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$, $u_{t}^{(0)}\leftarrow\frac{b-a}{T}$ for $t=0.,\dots,T$ and $i\leftarrow 0$
4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
5:     $G_{t}\leftarrow G\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ for $t=0,\dots,T-1$
6:     $\nu_{t}\leftarrow{\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{(i)}G\left(y,u_{t}^{(i)}\right)u_{t}^{(i)}\right)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}$ for $t=1,\dots,T-1$
7:     $\zeta_{t}\leftarrow{\left.\kern-1.2pt\nabla_{y}\left(u_{t}^{(i)}G\left(x_{t}^{(i)},y\right)u_{t}^{(i)}\right)\vphantom{\big{|}}\right|_{y=u_{t}^{(i)}}}$ for $t=1,\dots,T-1$
8:     $\mu_{T-1}\leftarrow\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\left(\zeta_{t}+\sum_{t>j}^{T-1}\nu_{j}\right)\right)$
9:     $u_{t}\leftarrow-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\zeta_{t}+\sum_{j>t}^{T-1}\nu_{j}\right)$ for $t=0,\dots,T-1$
10:     $x_{t+1}\leftarrow x_{t}+u_{t}$ for $t=0,\dots,T-1$
11:     Using line search find $\alpha^{*}$ for the following optimization problem with the discrete energy functional $E_{F}$
$\begin{split}\alpha^{*}=\operatorname*{arg\,min}_{\alpha}\quad&E_{F}\left(x_{0:T}\right)\quad\text{(exact line-search)}\\ \text{s.t.}\quad&x_{t+1}=x_{t}+\alpha u_{t}+(1-\alpha)u_{t}^{(i)},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
12:     Set $u_{t}^{(i+1)}\leftarrow\alpha^{*}u_{t}+(1-\alpha^{*})u_{t}^{(i)}$ for $t=0,\dots,T-1$
13:     Set $x_{t+1}^{(i+1)}\leftarrow x_{t}^{(i+1)}+u_{t}^{(i+1)}$ for $t=0,\dots,T-1$
14:     $i\leftarrow i+1$
15:end while
16:return $x_{t}$ for $t=0,\dots,T-1$
B.2 Sparse Newton Method
In this section we derive a sparse Newton method for minimizing eq. 5. We consider the Riemannian case, but note that it is easily generalized to the Finslerian case. The Hessian of the eq. 5 we get that
$\begin{split}\partial^{2}_{x_{t}x_{t}}E\left(x_{0:T}\right)&=\partial^{2}_{x_{t}x_{t}}{\left.\kern-1.2ptu_{t}^{\top}G(x_{t})u_{t}\vphantom{\big{|}}\right|_{u_{t}=x_{t+1}-x_{t}}}+2G(x_{t-1})-2G(x_{t})\\ &+4\nabla_{x_{t-1}}{\left.\kern-1.2ptG(x_{t-1})u_{t}\vphantom{\big{|}}\right|_{u_{t}=x_{t}-x_{t-1}}}-4\nabla_{x_{t}}{\left.\kern-1.2ptG(x_{t})u_{t}\vphantom{\big{|}}\right|_{u_{t}=x_{t+1}-x_{t}}},\quad t=1,\dots,T-1,\\ \partial^{2}_{x_{t}x_{t+1}}E\left(x_{0:T}\right)&=2\nabla_{x_{t}}{\left.\kern-1.2ptG(x_{t})u_{t}\vphantom{\big{|}}\right|_{u_{t}=x_{t+1}-x_{t}}},\quad t=1,\dots,T-2,\\ \partial^{2}_{x_{t+1}x_{t}}E\left(x_{0:T}\right)&=2{\left.\kern-1.2pt\left(G(x_{t})u_{t}\right)^{\top}\vphantom{\big{|}}\right|_{u_{t}=x_{t+1}-x_{t}}},\quad t=1,\dots,T-2.\end{split}$
where $\partial^{2}_{x_{i},x_{j}}$ denotes the second order derivative. From this we see that the Hessian can be written as
$H=\begin{pmatrix}H_{11}&H_{12}&0&0&0&0&0&\dots&0\\ H_{21}&H_{22}&H_{23}&0&0&0&0&\dots&0\\ 0&H_{23}&H_{33}&H_{34}&0&0&0&\dots&0\\ 0&0&H_{43}&H_{33}&H_{45}&0&0&\dots&0\\ 0&0&0&H_{54}&H_{55}&H_{56}&0&\dots&0\\ 0&0&0&0&H_{65}&H_{66}&H_{67}&\dots&0\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ 0&0&0&0&0&0&\dots&H_{(T-2)(T-2)}&H_{(T-1)(T-2)}\\ 0&0&0&0&0&0&\dots&H_{(T-2)(T-1)}&H_{(T-1)(T-1)}\\ \end{pmatrix},$ (27)
where we define
$H_{ij}=\partial^{2}_{x_{i}x_{j}}E(x_{0:T}),$
A Newton step has the following form at iteration $j$.
$s^{(j)}=-H^{-1}\nabla_{x_{1:(T-1)}}E(x_{0:T}),$ (28)
To simplify notation, define $s_{i}:=-\nabla_{x_{i}}E(x_{0:T})$. Using the block structure in eq. 27, we can recursively solve eq. 28 similar to the Thomas Algorithm [57].
$\begin{split}s_{i}&\leftarrow H_{ii}^{-1}s_{i},\\ P&\leftarrow H_{ii+1}^{\top},\quad i+1<T\\ H_{ii+1}&\leftarrow H_{ii}^{-1}H_{ii+1},\quad i+1<T\\ H_{i+1,i+1}&\leftarrow H_{i+1,i+1}-H_{i+1,i}H_{i,i+1}=H_{i+1,i+1}-PH_{i,i+1},\quad i+1<T\\ s_{i+1}&\leftarrow s_{i+1}-H_{i+1,i}s_{i}=s_{i+1}-Ps_{i},\quad i+1<T.\end{split}$ (29)
such that the Newton step is updated for $i=T-2,\dots,1$
$s_{i}\leftarrow s_{i}-H_{i,i+1}s_{i+1}.$ (30)
Note that we efficiently solve the equations in eq. 29 as
$\begin{split}H_{ii}y_{i}&=\left(H_{ii+1}|s_{i}\right),\quad i=1,\dots,T-2,\\ H_{T-1}y_{T-1}&=s_{T-1},\\ \end{split}$ (31)
such that only one linear system of equations has to be solved in each iteration. The Newton method does not have global convergence. In Algorithm 3 we write the Sparse Newton algorithm in pseudo-code switching between the Newton method and gradient descent with line-search. Note that the above is easily extended to the Finslerian case, where $G$ denotes the fundamental tensor and depends on both $x$ and $u$.
Algorithm 3 Sparse Newton Method
1:Input: $\mathrm{tol}$, $T$
2:Output: Geodesic estimate $x_{0:T}$
3:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$ for $t=0.,\dots,T$ and $j\leftarrow 0$
4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
5:     Compute $\left\{s_{t}^{\mathrm{grad}},s_{t}^{\mathrm{newton}}\right\}_{t=1,\dots,T-2}$ using eq. 29 and eq. 30.
6:     if $\ \textbf{then}\sum_{t=1}^{T-1}s_{t}^{\mathrm{grad}}\left(s_{t}^{\mathrm{newton}}\right)^{\top}\geq 0$
7:         $s_{t}=s_{t}^{\mathrm{newton}}$ for $t=1,\dots,T-1$
8:
9:     else
10:         $s_{t}=s_{t}^{\mathrm{grad}}$ for $t=1,\dots,T-1$
11:     end if
12:     Perform line-search for $\alpha^{*}=\operatorname*{arg\,min}_{0\leq\alpha\leq 1}E\left(x_{1:(T-1)}+\alpha s_{1:(T-1)}\right)$ with
$x_{t}^{(j+1)}\leftarrow x_{t}^{(j)}+\alpha^{*}s_{t},\quad t=1,\dots,T-1.$
13:     $j\leftarrow j+1$
14:end while
15:return $x_{t}^{(j)}$ for $t=0,\dots,T-1$
Note that the sparse Newton algorithm in Algorithm 3 has complexity $\mathcal{O}\left(Td^{3}\right)$ completely similar to GEORCE.
Since the Hessian of the energy is not necessarily positive definite, the Newton scheme can be numerically unstable. To avoid this, we represent in Algorithm 4 a regularized Newton method by adding suitable scaling in the diagonal.
Algorithm 4 Sparse Regularized Newton Method
1:Input: $\mathrm{tol}$, $T$, $\lambda$, $\kappa$
2:
3:Output: Geodesic estimate $x_{0:T}$
4:
5:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$ for $t=0.,\dots,T$ and $j\leftarrow 0$
6:
7:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
8:     $H_{ii}\leftarrow\lambda I+\partial^{2}_{x_{i}x_{i}}E(x_{0:T})$
9:
10:     $H_{ij}\leftarrow\partial^{2}_{x_{i}x_{j}}E(x_{0:T})$
11:
12:     Compute $\left\{s_{t}^{\mathrm{grad}},s_{t}^{\mathrm{newton}}\right\}_{t=1,\dots,T-2}$ using eq. 29 and eq. 30.
13:
14:     if $\ \textbf{then}\sum_{t=1}^{T-1}s_{t}^{\mathrm{grad}}\left(s_{t}^{\mathrm{newton}}\right)^{\top}\geq 0$
15:
16:         $s_{t}=s_{t}^{\mathrm{newton}}$ for $t=1,\dots,T-1$
17:
18:         $\lambda\leftarrow\kappa\lambda$
19:     else
20:         $s_{t}=s_{t}^{\mathrm{grad}}$ for $t=1,\dots,T-1$
21:
22:         $\lambda\leftarrow\nicefrac{{\lambda}}{{\kappa}}$
23:     end if
24:     Perform line-search for $\alpha^{*}=\operatorname*{arg\,min}_{0\leq\alpha\leq 1}E\left(x_{1:(T-1)}+\alpha s_{1:(T-1)}\right)$ with
$x_{t}^{(j+1)}\leftarrow x_{t}^{(j)}+\alpha^{*}s_{t},\quad t=1,\dots,T-1.$
25:     $j\leftarrow j+1$
26:end while
27:return $x_{t}^{(j)}$ for $t=0,\dots,T-1$
Note that the Newton method for the discretized energy was also proposed in [58], which also formulates the Hessian as a block matrix similar to eq. 27. However, they do not derive an adaptive scheme to solve the linear equation in the Newton step or propose the regularized version in Algorithm 4.
Appendix C Manifolds
Table 4 contains a description of the Riemannian manifolds shown in the paper.
Manifold Description Parameters Local Coordinates
$\mathbb{S}^{n}$ The n-sphere, $\{x\in\mathbb{R}^{n+1}\;|\;||x||_{2}=1\}$ Stereographic coordinates
$E(n)$ The n-Ellipsoid, $\{x\in\mathbb{R}^{n+1}\;|\;||p\odot x||_{2}=1\}$ Half-axes: $p=\left(0.5,\dots,1\right)$ equally distributed points. Stereographic coordinates
$\mathbb{T}^{2}$ Torus Major radius $R=3.0$ and minor radius $r=1.0$ Parameterized by $\left(\theta,\phi\right)$ with $x=(R+r\cos\theta)\cos\phi,y=(R+r\cos\theta)\sin\phi,z=r\sin\theta$.
$\mathbb{H}^{2}$ 2 dimensional Hyperbolic Space embedded into Minkowski pace $\mathbb{R}^{3}$ Parameterized by $\left(\alpha,\beta\right)$ with $x=\cosh\alpha,y=\sinh\alpha\cos\beta,z=\sinh\alpha\sin\beta$.
$P(n)$ Paraboloid of dimension $n$ Parameterized by standard coordinates by $\left(x^{1},\dots,x^{n},\sum_{i=1}^{n}x_{i}^{2}\right)$
$\mathcal{P}(n)$ Symmetric positive definite matrices of size $n^{2}$ Embedded into $\mathbb{R}^{n^{2}}$ by the mapping $f(x)=l(x)l(x)^{T}$, where $l:\mathbb{R}^{n(n+1)/2}\rightarrow\mathbb{R}^{n\times n}$ maps $x$ into a lower triangle matrix consisting of the elements in $x$.
’Egg-tray’ A parameterized surface in shape of an egg tray - Parameterized by $(x,y)$ with $x=x,y=y,z=2\cos x\cos y$.
Gaussian Distribution Parameters of the Gaussian distribution equipped with the Fischer-Rao metric - Parametrized by $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}$.
Fréchet Distribution Parameters of the Fréchet distribution equipped with the Fischer-Rao metric - Parametrized by $(\beta,\lambda)\in\mathbb{R}_{+}\times\mathbb{R}_{+}$.
Cauchy Distribution Parameters of the Cauchy distribution equipped with the Fischer-Rao metric - Parametrized by $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}$.
Pareto Distribution Parameters of the Pareto distribution equipped with the Fischer-Rao metric - Parametrized by $(\theta,\alpha)\in\mathbb{R}_{+}\times\mathbb{R}_{+}$.
VAE MNIST Variational-Autoencoder equipped with the pull-back metric for MNIST-data - Parametrized by the encoded latent space.
VAE SVHN Variational-Autoencoder equipped with the pull-back metric for SVHN-data - Parametrized by the encoded latent space.
VAE CelebA Variational-Autoencoder equipped with the pull-back metric for CelebA-data - Parametrized by the encoded latent space.
Table 4: Description of the manifolds used in the paper.
In the Finsler cases we consider a Riemannian manifold with a background metric, $G$, equipped with a force-field.
$f(x)=\frac{\sin x\odot\cos x}{(\cos x)^{\top}G(x)\cos x},$
where $x$ denotes the local coordinates of the manifold. In this case the corresponding metric is a Randers metric, which is a special case of Riemannian metrics, of the form [51]
$\begin{split}F&=\sqrt{a_{ij}\dot{r}^{i}\dot{r}^{j}}+b_{i}\dot{r}^{i}\\ a_{ij}&=g_{ij}\lambda+f_{i}f_{j}\lambda^{2}\\ b_{i}&=-f_{i}\lambda\\ f_{i}&=g_{ij}f^{j}\\ \lambda^{-1}&=v_{0}^{2}-g_{ij}f^{i}f^{j},\end{split}$ (32)
where $g_{ij}$ denotes the elements of the Riemannian background metric, $f^{j}$ denotes the $j$th element of the force field, while $v_{0}$ denotes the initial velocity Riemannian length of an object moving with constant velocity $v$ on the surface. For the manifold to be a proper Finsler manifold it is required that $||f||_{g}<v_{0}$.
Appendix D Variational autoencoders
D.1 Architecture
We train a Variational-Autoencoer (VAE) [52], where we use the normal distribution as a variational family to approximate the data distribution. We train a VAE for each of the following datasets
  • MNIST data [54]: The data consists of $28\times 28$ images of handwritten digits between 0 and 9. The data consists of $60,000$ training images and $10,000$ testing images.
  • SVHN data [55]: The data consists of $32\times 32\times 3$ images of house numbers obtained through Google street view. The data consists of $73,257$ digits for training and $26,032$ digits for testing. We use approximately $80\%$ of the training data for training similar to [12].
  • CelebA data [53]: The data consists of $202,599$ images of famous people of size $64\times 64\times 3$, where we use approximately $80\%$ for training similar to [12].
The architectures for the VAE for MNIST data, SVHN data and CelebA data are shown in Table 5, Table 6 and Table 7, respectively. The architecture is similar to [12] with some modifications. All models are trained for $50,000$ epochs on a GPU described in Appendix E.2.
Encoder
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=1,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
Latent Parameters
$\mu$: $\mathrm{Linear}(\mathrm{out\_feature}=8,\mathrm{bias}=\mathrm{True})$, $\mathrm{Identity}$
$\sigma$: $\mathrm{Linear}(\mathrm{out\_feature}=8,\mathrm{bias}=\mathrm{True})$, $\mathrm{Sigmoid}$
Decoder
$\mathrm{Linear}\left(\mathrm{out\_feature}=50,\mathrm{bias}=\mathrm{True}\right)$
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
$\mathrm{Linear}\left(\mathrm{out\_feature}=784,\mathrm{bias}=\mathrm{True}\right)$
Table 5: Architecture for VAE for the MNIST-dataset
Encoder
$\mathrm{Conv}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
Latent Parameters
$\mu$: $\mathrm{Linear}(\mathrm{out\_feature}=32,\mathrm{bias}=\mathrm{True})$, $\mathrm{Identity}$
$\sigma$: $\mathrm{Linear}(\mathrm{out\_feature}=32,\mathrm{bias}=\mathrm{True})$, $\mathrm{Sigmoid}$
Decoder
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
Table 6: Architecture for VAE for the SVHN-dataset
Encoder
$\mathrm{Conv}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
$\mathrm{Conv}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
$\mathrm{GELU}$
Latent Parameters
$\mu$: $\mathrm{Linear}(\mathrm{out\_feature}=32,\mathrm{bias}=\mathrm{True})$, $\mathrm{Identity}$
$\sigma$: $\mathrm{Linear}(\mathrm{out\_feature}=32,\mathrm{bias}=\mathrm{True})$, $\mathrm{Sigmoid}$
Decoder
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
$\mathrm{GELU}$
[…]
Table 7: Architecture for VAE for the CelebA-dataset
D.2 Additional experiments
In Fig. 10 and Fig. 11 we show the geodesics between two reconstructed images using the learned VAE for the MNIST VAE and SVHN VAE, respectively, where we show the results using GEORCE, the initialization and alternative methods.
Figure 10: The figure shows the results of using GEORCE and alternative methods for constructing geodesics for a Variational-Autoencoder [52] (VAE) for the MNIST dataset [54] All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional in eq. 5 is less than $10^{-3}$, or if the number of iterations exceeds $100$. We see that GEORCE uses less iterations and obtains the smallest length compared to the other algorithms. The details of the manifolds and experiments can be found in the supplementary material.
Figure 11: The figure shows the results of using GEORCE and alternative methods for constructing geodesics for a Variational-Autoencoder [52] (VAE) for the SVHN dataset [55]. All algorithms are terminated if the $\ell^{2}$-norm of the gradient of the discretized energy functional in eq. 5 is less than $10^{-3}$, or if the number of iterations exceeds $100$. We see that GEORCE uses less iterations and obtains the smallest length compared to the other algorithms. The details of the manifolds and experiments can be found in the supplementary material.
Appendix E Experiments
The following contains details of hyper-parameters, hardware and additional experiments. The implementation and code can be found at
https://github.com/FrederikMR/georce.
E.1 Hyper-parameters and methods
Table 8 shows the hyper parameters used for the different methods. It is the same hyper-parameters that have been used for all manifolds across dimension and grid size in the paper.
Method Description Parameters
GEORCE See algorthm in main paper Backtracking: $\rho=0.5$.
ADAM [41] Moments based gradient descent method $\gamma=0.01$ (step size), $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$
SGD [59] Stochastic gradient desent $\gamma=0.01$ (step size), momentum=0, dampening=0, weight_decay=0.
BFGS [37, 38, 39, 40] Quasi-Newton method -
CG [60] Conjugate-Gradient method -
dogleg [60] trust region algorithm -
trust_ncg [60] Newton conjugate trust-region algorithm -
trust_exact [60] nearly exact trust region algorithm -
Sparse Newton See Appendix B.2 Backtracking: $\rho=0.5$.
Sparse Regularized Newton See Appendix B.2 Backtracking: $\rho=0.5$ with $\lambda=1.0$ and $\kappa=0.5$.
Table 8: The Hyper-parameters for estimating of the geodesics. We refer to Scipy [61] for the description and implementation of BFGS, CG, dogleg, trust_ncg and trust_exact.
Table 9 shows the integration methods used for solving the shooting problem.
Method Description Parameters
RK45 [62] The explicit Runge-Kutta of order four to five -
RK23 [63] The explicit Runge-Kutta of order two to three -
DOP853 [64] The explicit Runge-Kutta of order eight -
Radau [65] Implicit Runge-Kutta method using Radau method -
BDF [66] Implicit multi-step order using backward differentiation -
LSODA [67] Method combining BDF and ADAM with stifness detection -
Table 9: The ode methods. The descriptions are from the Scipy-library [61].
All the timing estimates are computed by compiling the algorithms first, and then running the algorithm five times to compute the mean and standard deviation of the runtime. Table 10 states the start and end point of the geodesics constructed on the different manifolds in a local coordinate system as described in Table 4.
Manifold Start point End point
$\mathbb{S}^{n}$ $n$ equally spaced points between 00 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
$E(n)$ $n$ equally spaced points between 00 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
$\mathbb{T}^{2}$ $(0.0,0.0)$ $(5\pi/4,5\pi/4)$
$\mathbb{H}^{2}$ $(1.0,1.0)$ $(0.1,0.1)$
Paraboloid $(1,1)$ $(0,0.5)$
$\mathcal{P}(n)$ The identity matrix in local coordinates $n(n+1)/2$ equally spaced points between $0.5$ and $1.0$.
’Egg-tray’ $(-5.0,5.0)$ $(5.0,5.0)$
Gaussian Distribution $\left(\mu_{0},\sigma_{0}\right)=\left(-1.0,0.5\right)$ $\left(\mu_{T},\sigma_{T}\right)=\left(1.0,1.0\right)$
Fréchet Distribution $\left(\beta_{0},\lambda_{0}\right)=\left(0.5,0.5\right)$ $\left(\beta_{T},\lambda_{T}\right)=\left(1.0,1.0\right)$
Cauchy Distribution $\left(\mu_{0},\sigma_{0}\right)=\left(-1.0,0.5\right)$ $\left(\mu_{T},\sigma_{T}\right)=\left(1.0,1.0\right)$
Pareto Distribution $\left(\theta_{0},\alpha_{0}\right)=\left(0.5,0.5\right)$ $\left(\theta_{T},\alpha_{T}\right)=\left(1.0,1.0\right)$
Table 10: The start and end point for the geodesics estimated for each manifold.
E.2 Hardware
All figures have been computed on a HP computer with Intel Core i9-11950H 2.6 GHz 8C, 15.6” FHD, 720P CAM, 32 GB (2$\times$16GB) DDR4 3200 So-Dimm, Nvidia Quadro TI2004GB Descrete Graphics, 1TB PCle NVMe SSD, backlit Keyboard, 150W PSU, 8cell, W11Home 64 Advanced, 4YR Onsite NBD.
The runtime experiments on a CPU have been run on a CPU model XeonE5_2660v3 with $4$ nodes for at most $24$ hours with a maximum memory of $10$ GB. .
The runtime estimates on a GPU and training of the VAE have been computed on a GPU for at most 24 hours with a maximum memory of $10$ GB. The $GPU$ consists of $4$ nodes on a Tesla V100.
E.3 Additional experiments
In the following we show run time estimates computed similar to the description in the main paper with other methods than BFGS and ADAM. It should be noted that a gradient method with exact line search has also been tested against GEORCE, but shows similar computational behavior as ADAM, but often with higher computational times and therefore not included in the following. We also show the results on a GPU for CUDA compatible methods.
Note that for the Finsler manifold for the ellipsoid of dimension $50$ with a grid size of $T=50$ points, none of the methods converge including GEORCE. This is due to the strong variation by the force field causing numerical instability. However, when the grid size is increased to $T=100$, GEORCE converges to the best solution across all algorithms indicating that it requires a finer grid for this example.
E.4 Run-Time Estimates on VAE Manifolds
We show in Table 11 the run-time estimates on learned manifolds using a VAE.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
VAE MNIST $10.3075$ $1407.5743\pm 6.4430$ $10.3312$ $645.5201\pm 2.6648$ $\boldsymbol{10.3074}$ $\boldsymbol{67.6149}\pm\boldsymbol{0.0304}$
VAE SVHN -- -- $19.9855$ $13325.2285\pm 123.4978$ $\boldsymbol{19.9346}$ $\boldsymbol{882.4120}\pm\boldsymbol{1.8839}$
VAE CelebA -- -- -- -- $\boldsymbol{42.2666}$ $\boldsymbol{2222.5259}\pm\boldsymbol{8.2988}$
Table 11: The table shows the length of the estimated geodesics for the BFGS-algorithm, ADAM and GEORCE on a CPU. The methods were terminated if the $\ell^{2}$-norm of the gradient was less than $10^{-3}$ or $1,000$ iterations have been reached. When the computational time was longer than $24$ hour, the value is set to --.
E.4.1 Additional Run-time estimates on a GPU
We show in Table 12 and Table 13 the runtime of GPU compatible methods for different Riemannian manifolds, and in Table 14 and Table 15 the runtime of GPU compatible methods for different Finsler manifolds.
ADAM (T=100) Sparse Newton (T=100) GEORCE (T=100)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2755$ $0.2729\pm 0.0016$ $2.2742$ $0.1417\pm 0.0041$ $\boldsymbol{2.2741}$ $\boldsymbol{0.0163}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{3}$ $2.4376$ $0.2890\pm 0.0005$ $\boldsymbol{2.4342}$ $0.2188\pm 0.0039$ $2.4342$ $\boldsymbol{0.0735}\pm\boldsymbol{0.0018}$
$\mathbb{S}^{5}$ $2.4226$ $0.3153\pm 0.0027$ $\boldsymbol{2.3755}$ $0.8657\pm 0.0015$ $2.3755$ $\boldsymbol{0.0494}\pm\boldsymbol{0.0014}$
$\mathbb{S}^{10}$ $2.1031$ $0.3052\pm 0.0079$ $1.9653$ $117.4589\pm 0.0140$ $\boldsymbol{1.9523}$ $\boldsymbol{0.0351}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{20}$ $1.6824$ $0.3401\pm 0.0117$ $1.7069$ $146.3024\pm 0.0191$ $\boldsymbol{1.4808}$ $\boldsymbol{0.0483}\pm\boldsymbol{0.0020}$
$\mathbb{S}^{50}$ $1.1456$ $0.5339\pm 0.0014$ $1.0359$ $310.2160\pm 0.0200$ $\boldsymbol{0.9749}$ $\boldsymbol{0.0587}\pm\boldsymbol{0.0023}$
$\mathbb{S}^{100}$ $0.8475$ $1.2935\pm 0.0014$ $0.7522$ $6.6464\pm 0.0016$ $\boldsymbol{0.4188}$ $\boldsymbol{0.0717}\pm\boldsymbol{0.0022}$
$\mathbb{S}^{250}$ $0.6251$ $3.8386\pm 0.0018$ $0.4958$ $98.6502\pm 0.0147$ $\boldsymbol{0.3192}$ $\boldsymbol{0.3296}\pm\boldsymbol{0.0005}$
$\mathbb{S}^{500}$ $0.5714$ $8.2031\pm 0.0030$ $0.3787$ $899.4549\pm 0.0188$ $\boldsymbol{0.2198}$ $\boldsymbol{1.2925}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{1000}$ $0.5270$ $19.2234\pm 0.0712$- -- $\boldsymbol{0.1573}$ $\boldsymbol{6.8579}\pm\boldsymbol{0.0023}$
$1.9595$ $0.2735\pm 0.0074$ $\boldsymbol{1.9595}$ $0.2177\pm 0.0075$ $1.9595$ $\boldsymbol{0.0138}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(3\right)$ $2.0523$ $0.2839\pm 0.0056$ $\boldsymbol{2.0498}$ $0.3705\pm 0.0063$ $2.0501$ $\boldsymbol{0.0167}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(5\right)$ $1.9594$ $0.2948\pm 0.0116$- -- $\boldsymbol{1.8410}$ $\boldsymbol{0.0827}\pm\boldsymbol{0.0036}$
$\mathrm{E}\left(10\right)$ $1.6421$ $0.3097\pm 0.0133$ $\boldsymbol{1.4247}$ $116.6572\pm 0.0186$ $1.5139$ $\boldsymbol{0.0547}\pm\boldsymbol{0.0019}$
$\mathrm{E}\left(20\right)$ $1.3052$ $0.3875\pm 0.0058$ $1.4583$ $148.6781\pm 0.0307$ $\boldsymbol{0.6514}$ $\boldsymbol{0.1449}\pm\boldsymbol{0.0054}$
$\mathrm{E}\left(50\right)$ $0.9245$ $0.5704\pm 0.0013$ $0.8849$ $316.9808\pm 0.0179$ $\boldsymbol{0.4642}$ $\boldsymbol{0.0370}\pm\boldsymbol{0.0014}$
$\mathrm{E}\left(100\right)$ $0.6873$ $1.3063\pm 0.0020$ $0.5922$ $5.2868\pm 0.0007$ $\boldsymbol{0.3249}$ $\boldsymbol{0.0524}\pm\boldsymbol{0.0002}$
$\mathrm{E}\left(250\right)$ $0.4974$ $4.2084\pm 0.0029$ $0.3917$ $84.6331\pm 0.0439$ $\boldsymbol{0.2180}$ $\boldsymbol{0.2325}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(500\right)$ $0.4559$ $7.4772\pm 0.0024$ $0.3039$ $827.7163\pm 0.1003$ $\boldsymbol{0.1468}$ $\boldsymbol{1.2897}\pm\boldsymbol{0.0004}$
$\mathrm{E}\left(1000\right)$ $0.4238$ $19.8033\pm 0.0970$- -- $\boldsymbol{0.1129}$ $\boldsymbol{5.1862}\pm\boldsymbol{0.0071}$
$\mathbb{T}^{2}$ $9.9853$ $0.2890\pm 0.0016$ $9.9851$ $0.3353\pm 0.0126$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0549}\pm\boldsymbol{0.0010}$
$\mathbb{H}^{2}$ $1.2500$ $0.3201\pm 0.0004$ $1.2498$ $0.2125\pm 0.0128$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0089}\pm\boldsymbol{0.0003}$
$\mathcal{P}\left(2\right)$ $1.0786$ $0.3801\pm 0.0047$ $\boldsymbol{1.0669}$ $0.2419\pm 0.0069$ $1.0669$ $\boldsymbol{0.0227}\pm\boldsymbol{0.0010}$
$\mathcal{P}\left(3\right)$ $2.3006$ $0.4115\pm 0.0009$ $2.3810$ $106.2814\pm 0.0135$ $\boldsymbol{2.2639}$ $\boldsymbol{0.0343}\pm\boldsymbol{0.0004}$
Gaussian Distribution $\boldsymbol{2.5951}$ $0.1809\pm 0.0010$ $2.5952$ $0.1108\pm 0.0055$ $2.5952$ $\boldsymbol{0.0123}\pm\boldsymbol{0.0002}$
Fréchet Distribution $\boldsymbol{1.1538}$ $0.1778\pm 0.0029$ $1.1539$ $0.0813\pm 0.0044$ $1.1539$ $\boldsymbol{0.0055}\pm\boldsymbol{0.0003}$
Cauchy Distribution $1.6452$ $0.1818\pm 0.0007$ $1.6453$ $0.1136\pm 0.0067$ $\boldsymbol{1.6452}$ $\boldsymbol{0.0137}\pm\boldsymbol{0.0001}$
Pareto Distribution $0.8314$ $0.1779\pm 0.0026$ $\boldsymbol{0.8297}$ $0.0758\pm 0.0036$ $0.8297$ $\boldsymbol{0.0043}\pm\boldsymbol{0.0002}$
Table 12: The table shows the length of the estimated geodesics for different methods on a GPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
ADAM (T=50) Sparse Newton (T=50) GEORCE (T=50)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2479$ $0.2683\pm 0.0082$ $2.2478$ $0.1094\pm 0.0047$ $\boldsymbol{2.2477}$ $\boldsymbol{0.0187}\pm\boldsymbol{0.0002}$
$\mathbb{S}^{3}$ $2.4069$ $0.2593\pm 0.0110$ $\boldsymbol{2.4064}$ $0.1370\pm 0.0082$ $2.4064$ $\boldsymbol{0.0866}\pm\boldsymbol{0.0031}$
$\mathbb{S}^{5}$ $2.3494$ $0.2907\pm 0.0103$ $2.3480$ $0.3153\pm 0.0063$ $\boldsymbol{2.3480}$ $\boldsymbol{0.0622}\pm\boldsymbol{0.0019}$
$\mathbb{S}^{10}$ $1.9470$ $0.3026\pm 0.0014$ $1.9274$ $0.8966\pm 0.0034$ $\boldsymbol{1.9272}$ $\boldsymbol{0.0540}\pm\boldsymbol{0.0016}$
$\mathbb{S}^{20}$ $1.5036$ $0.3349\pm 0.0069$ $1.4610$ $28.7634\pm 0.0016$ $\boldsymbol{1.4566}$ $\boldsymbol{0.0670}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{50}$ $1.0320$ $0.4713\pm 0.0025$ $1.0060$ $39.8257\pm 0.0010$ $\boldsymbol{0.5279}$ $\boldsymbol{0.0438}\pm\boldsymbol{0.0002}$
$\mathbb{S}^{100}$ $0.7642$ $0.8173\pm 0.0009$ $0.6922$ $4.2461\pm 0.0007$ $\boldsymbol{0.3932}$ $\boldsymbol{0.0704}\pm\boldsymbol{0.0027}$
$\mathbb{S}^{250}$ $0.5136$ $3.2362\pm 0.0009$ $0.4589$ $60.6258\pm 0.0179$ $\boldsymbol{0.2748}$ $\boldsymbol{0.1525}\pm\boldsymbol{0.0002}$
$\mathbb{S}^{500}$ $0.4253$ $7.7012\pm 0.0038$ $0.3278$ $706.9603\pm 0.0293$ $\boldsymbol{0.1858}$ $\boldsymbol{0.7340}\pm\boldsymbol{0.0006}$
$\mathbb{S}^{1000}$ $0.3790$ $20.5850\pm 0.0098$- -- $\boldsymbol{0.1626}$ $\boldsymbol{3.1229}\pm\boldsymbol{0.0007}$
$\mathrm{E}\left(2\right)$ $1.9331$ $0.2723\pm 0.0013$ $\boldsymbol{1.9330}$ $0.1365\pm 0.0083$ $1.9331$ $\boldsymbol{0.0155}\pm\boldsymbol{0.0003}$
$\mathrm{E}\left(3\right)$ $2.0243$ $0.2852\pm 0.0017$ $\boldsymbol{2.0233}$ $0.1828\pm 0.0074$ $2.0235$ $\boldsymbol{0.0305}\pm\boldsymbol{0.0012}$
$\mathrm{E}\left(5\right)$ $1.8262$ $0.2996\pm 0.0079$ $\boldsymbol{1.8047}$ $0.6859\pm 0.0047$ $1.8054$ $\boldsymbol{0.1172}\pm\boldsymbol{0.0050}$
$\mathrm{E}\left(10\right)$ $1.5193$ $0.3094\pm 0.0086$ $\boldsymbol{0.9472}$ $57.6302\pm 0.0107$ $1.4788$ $\boldsymbol{0.0855}\pm\boldsymbol{0.0036}$
$\mathrm{E}\left(20\right)$ $1.2028$ $0.3379\pm 0.0049$ $\boldsymbol{0.7603}$ $48.7781\pm 0.0029$ $1.1366$ $\boldsymbol{0.0727}\pm\boldsymbol{0.0033}$
$\mathrm{E}\left(50\right)$ $0.8382$ $0.4596\pm 0.0092$ $0.6157$ $153.9548\pm 0.0165$ $\boldsymbol{0.4417}$ $\boldsymbol{0.0284}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(100\right)$ $0.6175$ $0.8219\pm 0.0036$ $0.5822$ $3.9507\pm 0.0014$ $\boldsymbol{0.3160}$ $\boldsymbol{0.0373}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(250\right)$ $0.4216$ $2.4940\pm 0.0028$ $0.3882$ $48.5547\pm 0.0120$ $\boldsymbol{0.2039}$ $\boldsymbol{0.1519}\pm\boldsymbol{0.0003}$
$\mathrm{E}\left(500\right)$ $0.3454$ $5.5967\pm 0.0145$- -- $\boldsymbol{0.1434}$ $\boldsymbol{0.7326}\pm\boldsymbol{0.0005}$
$\mathrm{E}\left(1000\right)$ $0.3056$ $14.1006\pm 0.0361$- -- $\boldsymbol{0.0996}$ $\boldsymbol{4.1512}\pm\boldsymbol{0.0043}$
$\mathbb{T}^{2}$ $9.8952$ $0.2821\pm 0.0029$ $\boldsymbol{9.8943}$ $38.0816\pm 0.0018$ $9.8955$ $\boldsymbol{0.0531}\pm\boldsymbol{0.0013}$
$\mathbb{H}^{2}$ $1.2435$ $0.3072\pm 0.0003$ $\boldsymbol{1.2434}$ $0.1172\pm 0.0048$ $1.2434$ $\boldsymbol{0.0082}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(2\right)$ $1.0704$ $0.3571\pm 0.0072$ $\boldsymbol{1.0567}$ $0.1231\pm 0.0051$ $1.0567$ $\boldsymbol{0.0243}\pm\boldsymbol{0.0002}$
$\mathcal{P}\left(3\right)$ $2.2400$ $0.3924\pm 0.0057$ $2.3328$ $57.1278\pm 0.0125$ $\boldsymbol{2.2304}$ $\boldsymbol{0.0198}\pm\boldsymbol{0.0001}$
Gaussian Distribution $2.5779$ $0.1785\pm 0.0003$ $\boldsymbol{2.5778}$ $0.0617\pm 0.0027$ $2.5778$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
Fréchet Distribution $1.1471$ $0.1696\pm 0.0004$ $\boldsymbol{1.1457}$ $0.0442\pm 0.0016$ $1.1457$ $\boldsymbol{0.0063}\pm\boldsymbol{0.0002}$
Cauchy Distribution $1.6341$ $0.0230\pm 0.0002$ $1.6340$ $0.0810\pm 0.0037$ $\boldsymbol{1.6340}$ $\boldsymbol{0.0135}\pm\boldsymbol{0.0003}$
Pareto Distribution $0.8235$ $0.0357\pm 0.0003$ $0.8235$ $0.0439\pm 0.0014$ $\boldsymbol{0.8234}$ $\boldsymbol{0.0052}\pm\boldsymbol{0.0003}$
Table 13: The table shows the length of the estimated geodesics for different methods on a GPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
ADAM (T=100) Sparse Newton (T=100) GEORCE (T=100)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.6530$ $0.7706\pm 0.0018$ $\boldsymbol{1.6525}$ $0.2512\pm 0.0089$ $1.6525$ $\boldsymbol{0.0372}\pm\boldsymbol{0.0004}$
$\mathbb{S}^{3}$ $1.7451$ $0.7362\pm 0.0036$- -- $\boldsymbol{1.7432}$ $\boldsymbol{0.0431}\pm\boldsymbol{0.0005}$
$\mathbb{S}^{5}$ $1.7371$ $0.9292\pm 0.0038$- -- $\boldsymbol{1.7029}$ $\boldsymbol{0.1036}\pm\boldsymbol{0.0048}$
$\mathbb{S}^{10}$ $1.5174$ $1.1358\pm 0.0018$ $1.4783$ $208.1472\pm 0.0176$ $\boldsymbol{1.4109}$ $\boldsymbol{0.1000}\pm\boldsymbol{0.0033}$
$\mathbb{S}^{20}$ $1.2009$ $1.5876\pm 0.0028$- -- $\boldsymbol{1.0760}$ $\boldsymbol{0.1509}\pm\boldsymbol{0.0058}$
$\mathbb{S}^{50}$ $0.8130$ $2.0381\pm 0.0026$- -- $\boldsymbol{0.7283}$ $\boldsymbol{0.1373}\pm\boldsymbol{0.0032}$
$\mathbb{S}^{100}$ $0.6146$ $2.7712\pm 0.0032$ $0.6210$ $972.0195\pm 0.0307$ $\boldsymbol{0.2553}$ $\boldsymbol{0.1106}\pm\boldsymbol{0.0028}$
$\mathbb{S}^{250}$ $0.5196$ $3.9803\pm 0.0019$- -- $\boldsymbol{0.1636}$ $\boldsymbol{0.5069}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{500}$ $0.4849$ $6.6164\pm 0.0023$- -- $\boldsymbol{0.1417}$ $\boldsymbol{1.6591}\pm\boldsymbol{0.0004}$
$\mathbb{S}^{1000}$ $0.3835$ $25.4638\pm 0.1448$- -- $\boldsymbol{0.1134}$ $\boldsymbol{7.9991}\pm\boldsymbol{0.0074}$
$\mathrm{E}\left(2\right)$ $1.3634$ $0.7758\pm 0.0036$ $\boldsymbol{1.3616}$ $0.3735\pm 0.0010$ $1.3618$ $\boldsymbol{0.0282}\pm\boldsymbol{0.0009}$
$\mathrm{E}\left(3\right)$ $1.4085$ $0.7541\pm 0.0032$- -- $\boldsymbol{1.4023}$ $\boldsymbol{0.1273}\pm\boldsymbol{0.0047}$
$\mathrm{E}\left(5\right)$ $1.3305$ $1.0252\pm 0.0017$ $1.5359$ $170.9661\pm 0.0296$ $\boldsymbol{1.2956}$ $\boldsymbol{0.1135}\pm\boldsymbol{0.0039}$
$\mathrm{E}\left(10\right)$ $1.1235$ $1.1254\pm 0.0045$- -- $\boldsymbol{1.0804}$ $\boldsymbol{0.0825}\pm\boldsymbol{0.0031}$
$\mathrm{E}\left(20\right)$ $0.8987$ $1.6039\pm 0.0035$- -- $\boldsymbol{0.8121}$ $\boldsymbol{0.1282}\pm\boldsymbol{0.0054}$
$\mathrm{E}\left(50\right)$ $0.6332$ $2.0424\pm 0.0034$- -- $\boldsymbol{0.2845}$ $\boldsymbol{0.0770}\pm\boldsymbol{0.0004}$
$\mathrm{E}\left(100\right)$ $0.4953$ $2.6704\pm 0.0031$ $0.4577$ $470.9805\pm 0.1448$ $\boldsymbol{0.2018}$ $\boldsymbol{0.1127}\pm\boldsymbol{0.0023}$
$\mathrm{E}\left(250\right)$ $0.4179$ $4.0843\pm 0.0009$- -- $\boldsymbol{0.1287}$ $\boldsymbol{0.5239}\pm\boldsymbol{0.0047}$
$\mathrm{E}\left(500\right)$ $0.4139$ $6.9391\pm 0.0073$ $0.2580$ $1435.3881\pm 0.0832$ $\boldsymbol{0.0968}$ $\boldsymbol{1.7109}\pm\boldsymbol{0.0010}$
$\mathrm{E}\left(1000\right)$ $0.3442$ $25.9660\pm 0.1038$- -- $\boldsymbol{0.0798}$ $\boldsymbol{7.9988}\pm\boldsymbol{0.0079}$
$\mathbb{T}^{2}$ $7.4054$ $0.8017\pm 0.0045$ $9.1903$ $152.0036\pm 0.0158$ $\boldsymbol{6.8067}$ $\boldsymbol{0.3145}\pm\boldsymbol{0.0050}$
$\mathbb{H}^{2}$ $1.3341$ $0.7949\pm 0.0072$ $1.3338$ $0.4562\pm 0.0004$ $\boldsymbol{1.3337}$ $\boldsymbol{0.0256}\pm\boldsymbol{0.0006}$
$\mathcal{P}\left(2\right)$ $0.7334$ $0.9344\pm 0.0013$ $\boldsymbol{0.7291}$ $0.4309\pm 0.0014$ $0.7292$ $\boldsymbol{0.0286}\pm\boldsymbol{0.0005}$
$\mathcal{P}\left(3\right)$ $1.2987$ $1.0812\pm 0.0006$ $1.2856$ $0.6885\pm 0.0009$ $\boldsymbol{1.2856}$ $\boldsymbol{0.0365}\pm\boldsymbol{0.0005}$
Gaussian Distribution $1.6879$ $0.4962\pm 0.0034$ $1.6878$ $0.4148\pm 0.0055$ $\boldsymbol{1.6878}$ $\boldsymbol{0.0255}\pm\boldsymbol{0.0004}$
Fréchet Distribution $0.5406$ $0.2307\pm 0.0069$ $\boldsymbol{0.5405}$ $0.2060\pm 0.0042$ $0.5405$ $\boldsymbol{0.0088}\pm\boldsymbol{0.0003}$
Cauchy Distribution $1.1140$ $0.4994\pm 0.0017$ $\boldsymbol{1.1133}$ $0.5167\pm 0.0042$ $1.1133$ $\boldsymbol{0.0367}\pm\boldsymbol{0.0017}$
Pareto Distribution $0.3885$ $0.1014\pm 0.0038$ $0.3884$ $0.1042\pm 0.0031$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0096}\pm\boldsymbol{0.0004}$
Table 14: The table shows the length of the estimated geodesics for different methods on a GPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
ADAM (T=50) Sparse Newton (T=50) GEORCE (T=50)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $\boldsymbol{1.6383}$ $0.7833\pm 0.0056$ $1.6385$ $0.2064\pm 0.0073$ $1.6385$ $\boldsymbol{0.0457}\pm\boldsymbol{0.0004}$
$\mathbb{S}^{3}$ $1.7300$ $0.7587\pm 0.0049$- -- $\boldsymbol{1.7297}$ $\boldsymbol{0.0499}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{5}$ $1.6875$ $0.9909\pm 0.0022$- -- $\boldsymbol{1.6868}$ $\boldsymbol{0.1432}\pm\boldsymbol{0.0050}$
$\mathbb{S}^{10}$ $1.4013$ $1.1472\pm 0.0024$- -- $\boldsymbol{1.3908}$ $\boldsymbol{0.1747}\pm\boldsymbol{0.0067}$
$\mathbb{S}^{20}$ $1.0829$ $1.5583\pm 0.0016$- -- $\boldsymbol{1.0560}$ $\boldsymbol{0.2663}\pm\boldsymbol{0.0005}$
$\mathbb{S}^{50}$ $0.7427$ $1.4442\pm 0.0038$- -- $\boldsymbol{0.3508}$ $\boldsymbol{0.0594}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{100}$ $0.5511$ $2.5558\pm 0.0020$- -- $\boldsymbol{0.2626}$ $\boldsymbol{0.0850}\pm\boldsymbol{0.0024}$
$\mathbb{S}^{250}$ $0.4094$ $3.8872\pm 0.0020$- -- $\boldsymbol{0.1686}$ $\boldsymbol{0.2660}\pm\boldsymbol{0.0042}$
$\mathbb{S}^{500}$ $0.3531$ $7.5214\pm 0.0022$ $0.2861$ $1388.8278\pm 0.0965$ $\boldsymbol{0.1193}$ $\boldsymbol{1.1879}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{1000}$ $0.3269$ $13.1209\pm 0.0379$- -- $\boldsymbol{0.0828}$ $\boldsymbol{6.0238}\pm\boldsymbol{0.0039}$
$\mathrm{E}\left(2\right)$ $1.3460$ $0.1327\pm 0.0060$ $\boldsymbol{1.3460}$ $0.2847\pm 0.0043$ $1.3461$ $\boldsymbol{0.0332}\pm\boldsymbol{0.0011}$
$\mathrm{E}\left(3\right)$ $1.3870$ $0.7621\pm 0.0034$- -- $\boldsymbol{1.3826}$ $\boldsymbol{0.2156}\pm\boldsymbol{0.0006}$
$\mathrm{E}\left(5\right)$ $1.2780$ $0.9718\pm 0.0022$- -- $\boldsymbol{1.2757}$ $\boldsymbol{0.1654}\pm\boldsymbol{0.0050}$
$\mathrm{E}\left(10\right)$ $1.0665$ $1.1072\pm 0.0025$- -- $\boldsymbol{1.0126}$ $\boldsymbol{0.2215}\pm\boldsymbol{0.0062}$
$\mathrm{E}\left(20\right)$ $0.8405$ $1.5702\pm 0.0031$- -- $\boldsymbol{0.7946}$ $\boldsymbol{0.1937}\pm\boldsymbol{0.0061}$
$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5797}$ $1.4336\pm 0.0029$- -- -- --
$\mathrm{E}\left(100\right)$ $0.4334$ $2.0750\pm 0.0026$- -- $\boldsymbol{0.2046}$ $\boldsymbol{0.0830}\pm\boldsymbol{0.0004}$
$\mathrm{E}\left(250\right)$ $0.3234$ $2.8610\pm 0.0021$- -- $\boldsymbol{0.1365}$ $\boldsymbol{0.2644}\pm\boldsymbol{0.0029}$
$\mathrm{E}\left(500\right)$ $0.2862$ $5.1735\pm 0.0122$ $0.2428$ $905.0696\pm 0.0365$ $\boldsymbol{0.0940}$ $\boldsymbol{0.9444}\pm\boldsymbol{0.0011}$
$\mathrm{E}\left(1000\right)$ $0.2723$ $10.4951\pm 0.0347$- -- $\boldsymbol{0.0893}$ $\boldsymbol{4.4688}\pm\boldsymbol{0.0060}$
$\boldsymbol{7.2683}$ $\boldsymbol{0.7530}\pm\boldsymbol{0.0015}$ $9.0475$ $78.7132\pm 0.0222$ $9.0475$ $13.7393\pm 0.0093$
$\mathbb{H}^{2}$ $1.3328$ $0.7454\pm 0.0052$ $1.3316$ $0.2303\pm 0.0058$ $\boldsymbol{1.3316}$ $\boldsymbol{0.0317}\pm\boldsymbol{0.0003}$
$\mathcal{P}\left(2\right)$ $0.7249$ $0.5927\pm 0.0008$ $\boldsymbol{0.7249}$ $0.2189\pm 0.0051$ $0.7249$ $\boldsymbol{0.0331}\pm\boldsymbol{0.0011}$
$\mathcal{P}\left(3\right)$ $1.2710$ $1.0184\pm 0.0008$ $1.2707$ $0.4325\pm 0.0020$ $\boldsymbol{1.2707}$ $\boldsymbol{0.0434}\pm\boldsymbol{0.0006}$
Gaussian Distribution $\boldsymbol{1.6796}$ $0.4337\pm 0.0111$ $1.6806$ $0.2081\pm 0.0047$ $1.6806$ $\boldsymbol{0.0282}\pm\boldsymbol{0.0017}$
Fréchet Distribution $0.5370$ $0.0829\pm 0.0032$ $0.5370$ $0.1059\pm 0.0042$ $\boldsymbol{0.5370}$ $\boldsymbol{0.0111}\pm\boldsymbol{0.0003}$
Cauchy Distribution $1.1085$ $0.0626\pm 0.0010$ $\boldsymbol{1.1084}$ $0.2270\pm 0.0081$ $1.1084$ $\boldsymbol{0.0410}\pm\boldsymbol{0.0015}$
Pareto Distribution $0.3857$ $0.0752\pm 0.0025$ $\boldsymbol{0.3856}$ $0.1054\pm 0.0028$ $0.3857$ $\boldsymbol{0.0082}\pm\boldsymbol{0.0003}$
Table 15: The table shows the length of the estimated geodesics for different methods on a GPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
E.4.2 Additional Run-time estimates on a CPU
Methods on a Riemannian manifold
We show in Table 16, Table 17, Table 18, Table 19, Table 20, Table 21, Table 22 and Table 23 runtime estimates for different methods on a CPU for the different Riemannian manifolds.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2742$ $57.4164\pm 0.0616$ $2.2772$ $0.2335\pm 0.0005$ $\boldsymbol{2.2741}$ $\boldsymbol{0.0078}\pm\boldsymbol{0.0000}$
$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $67.4657\pm 0.4874$ $2.4377$ $0.3475\pm 0.0010$ $2.4342$ $\boldsymbol{0.0635}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{5}$ $2.3757$ $77.6975\pm 0.0741$ $2.4222$ $1.9894\pm 0.0188$ $\boldsymbol{2.3755}$ $\boldsymbol{0.2219}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{10}$ $1.9529$ $150.0621\pm 0.8919$ $2.1016$ $4.4208\pm 0.0308$ $\boldsymbol{1.9523}$ $\boldsymbol{0.3184}\pm\boldsymbol{0.0002}$
$\mathbb{S}^{20}$ $1.4810$ $1068.8645\pm 17.2001$ $1.6839$ $8.7097\pm 0.0352$ $\boldsymbol{1.4808}$ $\boldsymbol{0.6732}\pm\boldsymbol{0.0204}$
$\mathbb{S}^{50}$ $1.0464$ $7350.6001\pm 244.3737$ $1.1463$ $61.8926\pm 0.3034$ $\boldsymbol{0.9749}$ $\boldsymbol{4.1016}\pm\boldsymbol{0.0201}$
$\mathbb{S}^{100}$- -- $0.8466$ $277.4217\pm 0.2280$ $\boldsymbol{0.4187}$ $\boldsymbol{10.9208}\pm\boldsymbol{0.0071}$
$\mathbb{S}^{250}$- -- $0.6244$ $1424.9031\pm 22.1592$ $\boldsymbol{0.3192}$ $\boldsymbol{78.2937}\pm\boldsymbol{0.0991}$
$\mathbb{S}^{500}$- -- $0.5717$ $2242.4697\pm 17.9769$ $\boldsymbol{0.2198}$ $\boldsymbol{224.2761}\pm\boldsymbol{0.4894}$
$\mathbb{S}^{1000}$- -- $0.5279$ $4329.6621\pm 38.0163$ $\boldsymbol{0.1573}$ $\boldsymbol{893.4682}\pm\boldsymbol{7.3130}$
$\mathrm{E}\left(2\right)$ $1.9595$ $56.9650\pm 0.4106$ $1.9596$ $0.2328\pm 0.0001$ $\boldsymbol{1.9595}$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0499}$ $94.1769\pm 0.7044$ $2.0525$ $0.3392\pm 0.0024$ $2.0501$ $\boldsymbol{0.0119}\pm\boldsymbol{0.0000}$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8403}$ $194.5403\pm 1.3293$ $1.9521$ $2.2227\pm 0.0186$ $1.8410$ $\boldsymbol{0.4220}\pm\boldsymbol{0.0038}$
$\mathrm{E}\left(10\right)$ $\boldsymbol{1.5114}$ $427.6278\pm 4.1044$ $1.6384$ $4.5697\pm 0.0541$ $1.5139$ $\boldsymbol{0.4566}\pm\boldsymbol{0.0016}$
$\mathrm{E}\left(20\right)$ $1.1795$ $1669.2692\pm 10.2877$ $1.3049$ $9.4335\pm 0.0507$ $\boldsymbol{0.6548}$ $\boldsymbol{1.9786}\pm\boldsymbol{0.0031}$
$\mathrm{E}\left(50\right)$ $0.8338$ $10685.6133\pm 321.6447$ $0.9244$ $73.5682\pm 0.3306$ $\boldsymbol{0.4642}$ $\boldsymbol{2.4208}\pm\boldsymbol{0.0222}$
$\mathrm{E}\left(100\right)$- -- $0.6872$ $289.4196\pm 2.5224$ $\boldsymbol{0.3249}$ $\boldsymbol{7.7128}\pm\boldsymbol{0.0714}$
$\mathrm{E}\left(250\right)$- -- $0.4978$ $1620.6428\pm 16.9546$ $\boldsymbol{0.2178}$ $\boldsymbol{56.8527}\pm\boldsymbol{0.1526}$
$\mathrm{E}\left(500\right)$- -- $0.4559$ $2135.6899\pm 30.5699$ $\boldsymbol{0.1467}$ $\boldsymbol{220.1656}\pm\boldsymbol{0.9902}$
$\mathrm{E}\left(1000\right)$- -- $0.4232$ $4398.9453\pm 25.8738$ $\boldsymbol{0.1131}$ $\boldsymbol{641.7199}\pm\boldsymbol{20.9901}$
$\mathbb{T}^{2}$ $9.9851$ $92.7400\pm 3.1267$ $9.9855$ $0.1060\pm 0.0002$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0182}\pm\boldsymbol{0.0003}$
$\mathbb{H}^{2}$ $1.2499$ $75.0079\pm 0.1710$ $1.2499$ $0.0988\pm 0.0001$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0025}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0669}$ $88.2739\pm 0.4152$ $1.0879$ $0.2611\pm 0.0004$ $1.0669$ $\boldsymbol{0.0141}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2639}$ $168.5082\pm 0.3039$ $2.2896$ $2.3780\pm 0.0046$ $2.2639$ $\boldsymbol{0.1442}\pm\boldsymbol{0.0002}$
Gaussian Distribution $\boldsymbol{2.5952}$ $47.6288\pm 0.1359$ $2.5956$ $0.0167\pm 0.0001$ $2.5952$ $\boldsymbol{0.0015}\pm\boldsymbol{0.0000}$
Fréchet Distribution $\boldsymbol{1.1538}$ $51.3063\pm 0.0566$ $1.1542$ $0.0185\pm 0.0001$ $1.1539$ $\boldsymbol{0.0009}\pm\boldsymbol{0.0000}$
Cauchy Distribution $1.6454$ $36.5031\pm 0.0437$ $1.6466$ $0.0177\pm 0.0001$ $\boldsymbol{1.6452}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0000}$
Pareto Distribution $0.8297$ $39.0465\pm 0.0471$ $0.8336$ $0.0168\pm 0.0000$ $\boldsymbol{0.8297}$ $\boldsymbol{0.0007}\pm\boldsymbol{0.0000}$
Table 16: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
BFGS (T=50) ADAM (T=50) GEORCE (T=50)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2478$ $34.0090\pm 0.7480$ $2.2479$ $0.0996\pm 0.0000$ $\boldsymbol{2.2477}$ $\boldsymbol{0.0048}\pm\boldsymbol{0.0000}$
$\mathbb{S}^{3}$ $\boldsymbol{2.4064}$ $33.7983\pm 0.2472$ $2.4069$ $0.1720\pm 0.0001$ $\boldsymbol{2.4064}$ $\boldsymbol{0.0408}\pm\boldsymbol{0.0000}$
$\mathbb{S}^{5}$ $2.3480$ $42.4323\pm 0.4287$ $2.3489$ $1.0908\pm 0.0111$ $\boldsymbol{2.3480}$ $\boldsymbol{0.1484}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{10}$ $1.9288$ $62.2979\pm 0.5799$ $1.9469$ $2.2366\pm 0.0018$ $\boldsymbol{1.9272}$ $\boldsymbol{0.2462}\pm\boldsymbol{0.0015}$
$\mathbb{S}^{20}$ $1.4614$ $161.7671\pm 0.3882$ $1.5037$ $4.3590\pm 0.0024$ $\boldsymbol{1.4566}$ $\boldsymbol{0.5572}\pm\boldsymbol{0.0009}$
$\mathbb{S}^{50}$ $0.9934$ $1028.2557\pm 1.4276$ $1.0320$ $34.7336\pm 0.1530$ $\boldsymbol{0.5279}$ $\boldsymbol{1.6104}\pm\boldsymbol{0.0016}$
$\mathbb{S}^{100}$ $0.7500$ $4076.6101\pm 188.1693$ $0.7642$ $142.7405\pm 0.5381$ $\boldsymbol{0.3932}$ $\boldsymbol{6.4462}\pm\boldsymbol{0.0468}$
$\mathbb{S}^{250}$- -- $0.5136$ $1163.6172\pm 5.2656$ $\boldsymbol{0.2748}$ $\boldsymbol{25.8527}\pm\boldsymbol{0.0436}$
$\mathbb{S}^{500}$- -- $0.4253$ $2272.3826\pm 44.3587$ $\boldsymbol{0.1858}$ $\boldsymbol{117.6824}\pm\boldsymbol{0.2510}$
$\mathbb{S}^{1000}$- -- $0.3790$ $4507.5815\pm 11.6850$ $\boldsymbol{0.1626}$ $\boldsymbol{333.4426}\pm\boldsymbol{3.6572}$
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.9330}$ $38.1118\pm 0.0767$ $1.9337$ $0.0986\pm 0.0001$ $1.9331$ $\boldsymbol{0.0038}\pm\boldsymbol{0.0000}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0233}$ $41.8059\pm 0.0366$ $2.0237$ $0.1594\pm 0.0056$ $2.0235$ $\boldsymbol{0.0110}\pm\boldsymbol{0.0000}$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8045}$ $140.1273\pm 1.0396$ $1.8262$ $1.0991\pm 0.0024$ $1.8054$ $\boldsymbol{0.3319}\pm\boldsymbol{0.0007}$
$\mathrm{E}\left(10\right)$ $\boldsymbol{1.4760}$ $187.4378\pm 1.0853$ $1.5193$ $2.2661\pm 0.0032$ $1.4788$ $\boldsymbol{0.3963}\pm\boldsymbol{0.0019}$
$\mathrm{E}\left(20\right)$ $1.1489$ $329.3681\pm 2.5049$ $1.2029$ $3.8156\pm 0.0185$ $\boldsymbol{1.1366}$ $\boldsymbol{0.4997}\pm\boldsymbol{0.0015}$
$\mathrm{E}\left(50\right)$ $0.7913$ $1755.3668\pm 3.7051$ $0.8382$ $34.8794\pm 0.1429$ $\boldsymbol{0.4417}$ $\boldsymbol{1.0645}\pm\boldsymbol{0.0041}$
$\mathrm{E}\left(100\right)$ $0.5923$ $6260.6895\pm 30.3624$ $0.6175$ $144.6669\pm 1.0968$ $\boldsymbol{0.3160}$ $\boldsymbol{3.4531}\pm\boldsymbol{0.0102}$
$\mathrm{E}\left(250\right)$- -- $0.4216$ $871.3954\pm 0.6870$ $\boldsymbol{0.2039}$ $\boldsymbol{25.3168}\pm\boldsymbol{0.0372}$
$\mathrm{E}\left(500\right)$- -- $0.3454$ $1695.4600\pm 10.5103$ $\boldsymbol{0.1434}$ $\boldsymbol{114.6093}\pm\boldsymbol{0.4502}$
$\mathrm{E}\left(1000\right)$- -- $0.3056$ $3026.5793\pm 37.3366$ $\boldsymbol{0.0996}$ $\boldsymbol{429.9206}\pm\boldsymbol{0.2607}$
$\mathbb{T}^{2}$ $9.8955$ $59.2719\pm 0.3871$ $\boldsymbol{9.8950}$ $0.0539\pm 0.0000$ $9.8955$ $\boldsymbol{0.0116}\pm\boldsymbol{0.0001}$
$\mathbb{H}^{2}$ $1.2434$ $37.3124\pm 0.0360$ $\boldsymbol{1.2434}$ $0.0439\pm 0.0002$ $1.2434$ $\boldsymbol{0.0014}\pm\boldsymbol{0.0000}$
$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0567}$ $43.7359\pm 0.0490$ $1.0610$ $0.1343\pm 0.0015$ $1.0567$ $\boldsymbol{0.0081}\pm\boldsymbol{0.0002}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2304}$ $98.5211\pm 0.6390$ $2.2307$ $1.1962\pm 0.0194$ $2.2304$ $\boldsymbol{0.0500}\pm\boldsymbol{0.0001}$
Gaussian Distribution $2.5778$ $23.5395\pm 0.0607$ $\boldsymbol{2.5769}$ $0.0084\pm 0.0001$ $2.5778$ $\boldsymbol{0.0011}\pm\boldsymbol{0.0000}$
Fréchet Distribution $\boldsymbol{1.1457}$ $26.5810\pm 0.1445$ $1.1457$ $0.0097\pm 0.0001$ $1.1457$ $\boldsymbol{0.0006}\pm\boldsymbol{0.0000}$
Cauchy Distribution $1.6341$ $20.1153\pm 0.0478$ $1.6341$ $\boldsymbol{0.0012}\pm\boldsymbol{0.0000}$ $\boldsymbol{1.6340}$ $0.0012\pm 0.0000$
Pareto Distribution $0.8235$ $22.4667\pm 0.0646$ $0.8235$ $0.0088\pm 0.0000$ $\boldsymbol{0.8234}$ $\boldsymbol{0.0004}\pm\boldsymbol{0.0000}$
Table 17: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
SGD (T=100) CG (T=100) dogleg (T=100)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.4063$ $\boldsymbol{0.2238}\pm\boldsymbol{0.0075}$ $2.2744$ $193.4884\pm 2.7955$ $\boldsymbol{2.2742}$ $3.8305\pm 0.0143$
$\mathbb{S}^{3}$ $2.7102$ $\boldsymbol{0.3350}\pm\boldsymbol{0.0007}$ $2.4346$ $248.6363\pm 2.7041$ $\boldsymbol{2.4342}$ $5.6319\pm 0.0156$
$\mathbb{S}^{5}$ $3.0827$ $2.1406\pm 0.0095$ $\boldsymbol{2.3810}$ $217.8721\pm 4.5710$ $3.1069$ $\boldsymbol{1.7988}\pm\boldsymbol{0.0217}$
$\mathbb{S}^{10}$ $3.4578$ $\boldsymbol{4.6462}\pm\boldsymbol{0.0732}$ $\boldsymbol{1.9747}$ $178.0182\pm 0.0418$ $3.5092$ $6.9953\pm 0.0104$
$\mathbb{S}^{20}$ $3.5726$ $\boldsymbol{8.8376}\pm\boldsymbol{0.0048}$ $\boldsymbol{1.4961}$ $205.9335\pm 1.9411$ $3.6632$ $41.3160\pm 0.0809$
$\mathbb{S}^{50}$ $3.2831$ $\boldsymbol{72.7622}\pm\boldsymbol{0.0207}$ $\boldsymbol{1.2180}$ $81.5957\pm 0.0459$- --
$\mathbb{S}^{100}$ $2.8231$ $298.8796\pm 0.0082$ $\boldsymbol{1.0540}$ $\boldsymbol{109.8878}\pm\boldsymbol{0.4428}$- --
$\mathbb{S}^{250}$ $2.0949$ $2608.4075\pm 40.8368$ $\boldsymbol{0.8580}$ $\boldsymbol{391.9100}\pm\boldsymbol{0.9132}$- --
$\mathbb{S}^{500}$ $1.5779$ $10754.5576\pm 300.1625$ $\boldsymbol{1.0325}$ $\boldsymbol{709.8522}\pm\boldsymbol{4.6273}$- --
$\mathbb{S}^{1000}$- -- $\boldsymbol{1.1545}$ $\boldsymbol{68.2710}\pm\boldsymbol{0.7146}$- --
$\mathrm{E}\left(2\right)$ $1.9693$ $\boldsymbol{0.2259}\pm\boldsymbol{0.0018}$ $1.9598$ $159.7566\pm 1.6439$ $\boldsymbol{1.9595}$ $3.9798\pm 0.0834$
$\mathrm{E}\left(3\right)$ $2.0769$ $\boldsymbol{0.3266}\pm\boldsymbol{0.0029}$ $\boldsymbol{2.0504}$ $269.9406\pm 1.5354$ $2.0795$ $1.2249\pm 0.0080$
$\mathrm{E}\left(5\right)$ $2.2219$ $2.2046\pm 0.0154$ $\boldsymbol{1.8503}$ $516.0936\pm 11.2331$ $2.2269$ $\boldsymbol{1.7201}\pm\boldsymbol{0.0152}$
$\mathrm{E}\left(10\right)$ $2.4040$ $\boldsymbol{4.5520}\pm\boldsymbol{0.0117}$ $\boldsymbol{1.5794}$ $286.3407\pm 2.5997$ $2.4153$ $7.0164\pm 0.0183$
$\mathrm{E}\left(20\right)$ $2.4961$ $\boldsymbol{8.8872}\pm\boldsymbol{0.0029}$ $\boldsymbol{1.1925}$ $305.8136\pm 2.2537$ $2.5175$ $40.8913\pm 0.1881$
$\mathrm{E}\left(50\right)$ $2.3587$ $\boldsymbol{73.9033}\pm\boldsymbol{0.2967}$ $\boldsymbol{0.8949}$ $181.0294\pm 2.8838$- --
$\mathrm{E}\left(100\right)$ $2.0627$ $292.5031\pm 1.1563$ $\boldsymbol{0.8718}$ $\boldsymbol{177.2222}\pm\boldsymbol{0.6547}$- --
$\mathrm{E}\left(250\right)$ $1.5510$ $2521.6353\pm 22.3619$ $\boldsymbol{0.8962}$ $\boldsymbol{341.6886}\pm\boldsymbol{1.5827}$- --
$\mathrm{E}\left(500\right)$ $1.1756$ $10801.0918\pm 301.4005$ $\boldsymbol{0.7106}$ $\boldsymbol{1545.6230}\pm\boldsymbol{10.8368}$- --
$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.8633}$ $\boldsymbol{71.7434}\pm\boldsymbol{0.1687}$- --
$\mathbb{T}^{2}$ $11.3048$ $\boldsymbol{0.1001}\pm\boldsymbol{0.0000}$ $\boldsymbol{9.9854}$ $609.3215\pm 9.3287$ $11.6903$ $1.2105\pm 0.0095$
$\mathbb{H}^{2}$ $1.3308$ $\boldsymbol{0.0839}\pm\boldsymbol{0.0000}$ $1.2532$ $549.9518\pm 3.2561$ $\boldsymbol{1.2498}$ $4.1077\pm 0.0109$
$\mathcal{P}\left(2\right)$ $1.1512$ $\boldsymbol{0.2608}\pm\boldsymbol{0.0021}$ $1.0676$ $420.3352\pm 1.1349$ $\boldsymbol{1.0669}$ $4.5709\pm 0.0106$
$\mathcal{P}\left(3\right)$ $2.3892$ $\boldsymbol{2.1047}\pm\boldsymbol{0.0039}$ $\boldsymbol{2.2648}$ $570.4799\pm 5.5919$ $2.4047$ $2.3496\pm 0.0109$
Gaussian Distribution $2.8543$ $\boldsymbol{0.0115}\pm\boldsymbol{0.0001}$ $2.5955$ $374.0846\pm 7.3067$ $\boldsymbol{2.5952}$ $3.5420\pm 0.0317$
Fréchet Distribution $1.1823$ $\boldsymbol{0.0134}\pm\boldsymbol{0.0000}$ $1.1547$ $321.4810\pm 1.0877$ $\boldsymbol{1.1539}$ $2.7293\pm 0.0160$
Cauchy Distribution $1.9752$ $\boldsymbol{0.0119}\pm\boldsymbol{0.0001}$ $1.6461$ $192.8583\pm 1.3344$ $\boldsymbol{1.6452}$ $3.7594\pm 0.0544$
Pareto Distribution $0.8500$ $\boldsymbol{0.0111}\pm\boldsymbol{0.0001}$ $0.8299$ $87.2695\pm 0.1665$ $\boldsymbol{0.8297}$ $2.2831\pm 0.0390$
Table 18: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
SGD (T=50) CG (T=50) dogleg (T=50)
Riemannian Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.3607$ $\boldsymbol{0.0963}\pm\boldsymbol{0.0000}$ $2.2478$ $112.6833\pm 0.0636$ $\boldsymbol{2.2478}$ $3.6126\pm 0.0092$
$\mathbb{S}^{3}$ $2.6515$ $\boldsymbol{0.1450}\pm\boldsymbol{0.0005}$ $2.4065$ $119.3687\pm 0.3268$ $\boldsymbol{2.4064}$ $4.7996\pm 0.0220$
$\mathbb{S}^{5}$ $2.9958$ $\boldsymbol{1.0907}\pm\boldsymbol{0.0047}$ $\boldsymbol{2.3483}$ $132.6204\pm 0.7172$ $3.0849$ $1.1818\pm 0.0089$
$\mathbb{S}^{10}$ $3.3093$ $2.2309\pm 0.0017$ $\boldsymbol{1.9287}$ $131.0700\pm 2.7469$ $3.4892$ $\boldsymbol{2.2054}\pm\boldsymbol{0.0121}$
$\mathbb{S}^{20}$ $3.3585$ $\boldsymbol{4.3963}\pm\boldsymbol{0.0043}$ $\boldsymbol{1.4644}$ $197.9373\pm 3.1216$ $3.6466$ $10.9397\pm 0.0310$
$\mathbb{S}^{50}$ $3.0604$ $\boldsymbol{33.8548}\pm\boldsymbol{0.0592}$ $\boldsymbol{1.0403}$ $66.8326\pm 0.4691$- --
$\mathbb{S}^{100}$ $2.6628$ $142.5886\pm 0.5246$ $\boldsymbol{0.8374}$ $\boldsymbol{71.4308}\pm\boldsymbol{0.1749}$- --
$\mathbb{S}^{250}$ $2.0317$ $1351.8007\pm 6.9661$ $\boldsymbol{0.7906}$ $\boldsymbol{151.2937}\pm\boldsymbol{0.2754}$- --
$\mathbb{S}^{500}$ $1.5557$ $5792.4668\pm 19.3864$ $\boldsymbol{0.9376}$ $\boldsymbol{245.6238}\pm\boldsymbol{1.3281}$- --
$\mathbb{S}^{1000}$- -- $\boldsymbol{0.8765}$ $\boldsymbol{587.6923}\pm\boldsymbol{4.5177}$- --
$\mathrm{E}\left(2\right)$ $1.9493$ $\boldsymbol{0.0959}\pm\boldsymbol{0.0001}$ $1.9331$ $91.2580\pm 0.7684$ $\boldsymbol{1.9330}$ $3.6286\pm 0.0089$
$\mathrm{E}\left(3\right)$ $2.0568$ $\boldsymbol{0.1565}\pm\boldsymbol{0.0004}$ $\boldsymbol{2.0234}$ $188.5732\pm 1.4986$ $2.0661$ $1.0738\pm 0.0245$
$\mathrm{E}\left(5\right)$ $2.1983$ $\boldsymbol{1.0906}\pm\boldsymbol{0.0026}$ $\boldsymbol{1.8099}$ $319.5569\pm 2.8321$ $2.2150$ $1.1382\pm 0.0093$
$\mathrm{E}\left(10\right)$ $2.3647$ $2.2644\pm 0.0033$ $\boldsymbol{1.4779}$ $364.8791\pm 3.5348$ $2.4045$ $\boldsymbol{2.2572}\pm\boldsymbol{0.0141}$
$\mathrm{E}\left(20\right)$ $2.4309$ $\boldsymbol{4.5347}\pm\boldsymbol{0.0025}$ $\boldsymbol{1.1536}$ $256.8932\pm 2.5821$ $2.5079$ $11.0630\pm 0.0303$
$\mathrm{E}\left(50\right)$ $2.2709$ $\boldsymbol{29.4273}\pm\boldsymbol{0.1562}$ $\boldsymbol{0.8494}$ $99.0591\pm 0.7024$- --
$\mathrm{E}\left(100\right)$ $1.9914$ $123.1176\pm 0.7069$ $\boldsymbol{0.6421}$ $\boldsymbol{104.5576}\pm\boldsymbol{0.4612}$- --
$\mathrm{E}\left(250\right)$ $1.5207$ $1334.2307\pm 3.4387$ $\boldsymbol{0.6019}$ $\boldsymbol{199.9802}\pm\boldsymbol{0.7304}$- --
$\mathrm{E}\left(500\right)$ $1.1643$ $5472.2515\pm 84.0089$ $\boldsymbol{0.6644}$ $\boldsymbol{534.2835}\pm\boldsymbol{2.8548}$- --
$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.7192}$ $\boldsymbol{640.3465}\pm\boldsymbol{1.6748}$- --
$\mathbb{T}^{2}$ $10.5379$ $\boldsymbol{0.0498}\pm\boldsymbol{0.0001}$ $\boldsymbol{9.8956}$ $239.9891\pm 0.9619$ $11.5938$ $1.0782\pm 0.0071$
$\mathbb{H}^{2}$ $1.3210$ $\boldsymbol{0.0425}\pm\boldsymbol{0.0000}$ $1.2434$ $367.4767\pm 2.7398$ $\boldsymbol{1.2434}$ $3.6867\pm 0.0105$
$\mathcal{P}\left(2\right)$ $1.1238$ $\boldsymbol{0.1225}\pm\boldsymbol{0.0009}$ $1.0567$ $222.3136\pm 1.0460$ $\boldsymbol{1.0567}$ $3.6711\pm 0.0168$
$\mathcal{P}\left(3\right)$ $2.3231$ $\boldsymbol{1.2252}\pm\boldsymbol{0.0004}$ $\boldsymbol{2.2306}$ $364.5162\pm 0.9199$ $2.3721$ $1.3057\pm 0.0108$
Gaussian Distribution $2.7366$ $\boldsymbol{0.0055}\pm\boldsymbol{0.0001}$ $2.5780$ $130.6218\pm 0.8519$ $\boldsymbol{2.5778}$ $3.4534\pm 0.0148$
Fréchet Distribution $1.1685$ $\boldsymbol{0.0066}\pm\boldsymbol{0.0001}$ $1.1458$ $169.4907\pm 0.5148$ $\boldsymbol{1.1457}$ $2.5564\pm 0.0123$
Cauchy Distribution $1.9013$ $\boldsymbol{0.0057}\pm\boldsymbol{0.0000}$ $1.6341$ $109.2575\pm 0.1122$ $\boldsymbol{1.6340}$ $3.8239\pm 0.0166$
Pareto Distribution $0.8399$ $\boldsymbol{0.0056}\pm\boldsymbol{0.0000}$ $0.8235$ $64.6855\pm 0.0367$ $\boldsymbol{0.8235}$ $2.1945\pm 0.0138$
Table 19: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
trust-ncg (T=100) trust-exact (T=100)
Riemannian Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2742$ $237.8402\pm 3.6964$ $\boldsymbol{2.2742}$ $\boldsymbol{3.8066}\pm\boldsymbol{0.0204}$
$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $340.5377\pm 1.5571$ $2.4342$ $\boldsymbol{6.8560}\pm\boldsymbol{0.0218}$
$\mathbb{S}^{5}$ $\boldsymbol{2.3754}$ $833.7228\pm 9.9481$ $2.3754$ $\boldsymbol{12.4160}\pm\boldsymbol{0.0325}$
$\mathbb{S}^{10}$ $\boldsymbol{1.9521}$ $9097.0312\pm 53.0355$ $1.9523$ $\boldsymbol{65.4820}\pm\boldsymbol{0.4559}$
$\mathbb{S}^{20}$- -- $\boldsymbol{1.4816}$ $\boldsymbol{424.0417}\pm\boldsymbol{1.6674}$
$\mathbb{S}^{50}$- -- -- --
$\mathbb{S}^{100}$- -- -- --
$\mathbb{S}^{250}$- -- -- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $1.9595$ $265.0164\pm 3.4124$ $\boldsymbol{1.9595}$ $\boldsymbol{3.8934}\pm\boldsymbol{0.0312}$
$\mathrm{E}\left(3\right)$ $2.0499$ $726.9606\pm 6.6789$ $\boldsymbol{2.0498}$ $\boldsymbol{4.7604}\pm\boldsymbol{0.1015}$
$\mathrm{E}\left(5\right)$ $1.8394$ $3843.9355\pm 39.4764$ $\boldsymbol{1.8393}$ $\boldsymbol{16.1207}\pm\boldsymbol{0.2033}$
$\mathrm{E}\left(10\right)$- -- $\boldsymbol{1.5091}$ $\boldsymbol{73.9657}\pm\boldsymbol{0.1229}$
$\mathrm{E}\left(20\right)$- -- $\boldsymbol{1.1718}$ $\boldsymbol{398.0183}\pm\boldsymbol{1.7858}$
$\mathrm{E}\left(50\right)$- -- -- --
$\mathrm{E}\left(100\right)$- -- -- --
$\mathrm{E}\left(250\right)$- -- -- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $9.9851$ $374.0822\pm 5.9408$ $\boldsymbol{9.9851}$ $\boldsymbol{6.7989}\pm\boldsymbol{0.0304}$
$\mathbb{H}^{2}$ $\boldsymbol{1.2498}$ $627.2646\pm 7.3623$ $1.2498$ $\boldsymbol{5.0876}\pm\boldsymbol{0.0134}$
$\mathcal{P}\left(2\right)$ $1.0669$ $594.2135\pm 5.1381$ $\boldsymbol{1.0669}$ $\boldsymbol{4.0894}\pm\boldsymbol{0.0212}$
$\mathcal{P}\left(3\right)$ $2.2639$ $2158.6843\pm 25.1696$ $\boldsymbol{2.2639}$ $\boldsymbol{13.2320}\pm\boldsymbol{0.0803}$
Gaussian Distribution $\boldsymbol{2.5952}$ $202.1364\pm 4.9068$ $2.5952$ $\boldsymbol{3.0117}\pm\boldsymbol{0.0225}$
Fréchet Distribution $1.1539$ $286.7279\pm 9.4552$ $\boldsymbol{1.1539}$ $\boldsymbol{2.6002}\pm\boldsymbol{0.0147}$
Cauchy Distribution $\boldsymbol{1.6452}$ $201.4917\pm 0.1496$ $1.6452$ $\boldsymbol{3.9414}\pm\boldsymbol{0.0098}$
Pareto Distribution $\boldsymbol{0.8297}$ $233.9446\pm 0.2961$ $0.8297$ $\boldsymbol{2.3762}\pm\boldsymbol{0.0172}$
Table 20: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
trust-ncg (T=50) trust-exact (T=50)
Riemannian Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2478$ $115.7323\pm 1.0615$ $\boldsymbol{2.2478}$ $\boldsymbol{3.6180}\pm\boldsymbol{0.0205}$
$\mathbb{S}^{3}$ $2.4064$ $138.8976\pm 0.8824$ $\boldsymbol{2.4064}$ $\boldsymbol{4.6065}\pm\boldsymbol{0.0202}$
$\mathbb{S}^{5}$ $2.3480$ $272.6404\pm 1.2856$ $\boldsymbol{2.3480}$ $\boldsymbol{7.6100}\pm\boldsymbol{0.0293}$
$\mathbb{S}^{10}$ $\boldsymbol{1.9274}$ $1127.8496\pm 3.9055$ $1.9275$ $\boldsymbol{19.4117}\pm\boldsymbol{0.0371}$
$\mathbb{S}^{20}$ $\boldsymbol{1.4608}$ $8353.8154\pm 153.6381$ $1.4610$ $\boldsymbol{110.3418}\pm\boldsymbol{0.4293}$
$\mathbb{S}^{50}$- -- -- --
$\mathbb{S}^{100}$- -- -- --
$\mathbb{S}^{250}$- -- -- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $1.9330$ $101.3354\pm 0.2792$ $\boldsymbol{1.9330}$ $\boldsymbol{3.8779}\pm\boldsymbol{0.0182}$
$\mathrm{E}\left(3\right)$ $2.0233$ $256.4745\pm 4.0010$ $\boldsymbol{2.0233}$ $\boldsymbol{4.6605}\pm\boldsymbol{0.0146}$
$\mathrm{E}\left(5\right)$ $1.8046$ $1220.0194\pm 7.2318$ $\boldsymbol{1.8046}$ $\boldsymbol{10.3930}\pm\boldsymbol{0.0328}$
$\mathrm{E}\left(10\right)$ $1.4753$ $3603.1384\pm 81.3956$ $\boldsymbol{1.4751}$ $\boldsymbol{23.5777}\pm\boldsymbol{0.0401}$
$\mathrm{E}\left(20\right)$- -- $\boldsymbol{1.1434}$ $\boldsymbol{109.4113}\pm\boldsymbol{0.0144}$
$\mathrm{E}\left(50\right)$- -- -- --
$\mathrm{E}\left(100\right)$- -- -- --
$\mathrm{E}\left(250\right)$- -- -- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{9.8955}$ $189.0475\pm 0.1325$ $\boldsymbol{9.8955}$ $\boldsymbol{7.5805}\pm\boldsymbol{0.0152}$
$\mathbb{H}^{2}$ $1.2434$ $200.3231\pm 0.3463$ $\boldsymbol{1.2434}$ $\boldsymbol{3.6404}\pm\boldsymbol{0.0076}$
$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0567}$ $153.1577\pm 0.7454$ $1.0567$ $\boldsymbol{3.6134}\pm\boldsymbol{0.0305}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2304}$ $612.8415\pm 8.6455$ $2.2304$ $\boldsymbol{6.2574}\pm\boldsymbol{0.0139}$
Gaussian Distribution $\boldsymbol{2.5778}$ $122.4479\pm 0.1045$ $2.5778$ $\boldsymbol{2.8004}\pm\boldsymbol{0.0155}$
Fréchet Distribution $\boldsymbol{1.1457}$ $215.6763\pm 0.4161$ $1.1457$ $\boldsymbol{2.5323}\pm\boldsymbol{0.0123}$
Cauchy Distribution $1.6340$ $113.7271\pm 0.0720$ $\boldsymbol{1.6340}$ $\boldsymbol{3.6122}\pm\boldsymbol{0.0207}$
Pareto Distribution $0.8235$ $80.3471\pm 0.8322$ $\boldsymbol{0.8235}$ $\boldsymbol{2.2065}\pm\boldsymbol{0.0180}$
Table 21: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
Sparse Newton (T=100) Sparse Regularized Newton (T=100)
Riemannian Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $\boldsymbol{2.2742}$ $\boldsymbol{0.0180}\pm\boldsymbol{0.0002}$ $2.2742$ $0.1017\pm 0.0017$
$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $\boldsymbol{0.0837}\pm\boldsymbol{0.0157}$ $2.4342$ $0.3089\pm 0.0428$
$\mathbb{S}^{5}$ $\boldsymbol{2.3755}$ $\boldsymbol{0.6723}\pm\boldsymbol{0.0409}$ $2.3756$ $1.2203\pm 0.0887$
$\mathbb{S}^{10}$ $\boldsymbol{1.9449}$ $256.4748\pm 20.0998$ $1.9531$ $\boldsymbol{4.3786}\pm\boldsymbol{0.3614}$
$\mathbb{S}^{20}$ $1.7069$ $1130.3478\pm 127.1948$ $\boldsymbol{1.4943}$ $\boldsymbol{20.3206}\pm\boldsymbol{1.5336}$
$\mathbb{S}^{50}$ $1.0815$ $16216.8467\pm 2136.6802$ $\boldsymbol{1.0106}$ $\boldsymbol{330.6598}\pm\boldsymbol{26.1208}$
$\mathbb{S}^{100}$ $\boldsymbol{0.7522}$ $\boldsymbol{1041.2721}\pm\boldsymbol{89.4723}$ $0.7781$ $2863.6448\pm 233.2495$
$\mathbb{S}^{250}$ $\boldsymbol{0.4958}$ $\boldsymbol{17070.7520}\pm\boldsymbol{2024.9161}$- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.9595}$ $\boldsymbol{0.0254}\pm\boldsymbol{0.0002}$ $1.9596$ $0.0915\pm 0.0028$
$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0498}$ $\boldsymbol{0.1217}\pm\boldsymbol{0.0219}$ $2.0499$ $0.3029\pm 0.0288$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8392}$ $\boldsymbol{0.6774}\pm\boldsymbol{0.0721}$ $1.8397$ $1.3500\pm 0.1241$
$\mathrm{E}\left(10\right)$ $1.5247$ $300.2892\pm 23.3273$ $\boldsymbol{1.5114}$ $\boldsymbol{4.4472}\pm\boldsymbol{0.3697}$
$\mathrm{E}\left(20\right)$ $1.4595$ $1101.5878\pm 115.0272$ $\boldsymbol{1.1961}$ $\boldsymbol{17.6133}\pm\boldsymbol{1.2503}$
$\mathrm{E}\left(50\right)$ $0.8825$ $15145.3818\pm 1401.9972$ $\boldsymbol{0.8279}$ $\boldsymbol{300.5634}\pm\boldsymbol{20.9595}$
$\mathrm{E}\left(100\right)$ $\boldsymbol{0.5922}$ $\boldsymbol{839.4538}\pm\boldsymbol{68.5559}$ $0.6402$ $2509.0547\pm 213.2646$
$\mathrm{E}\left(250\right)$ $\boldsymbol{0.3919}$ $\boldsymbol{16334.5381}\pm\boldsymbol{1370.7971}$- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0308}\pm\boldsymbol{0.0002}$ $9.9851$ $0.0725\pm 0.0014$
$\mathbb{H}^{2}$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0225}\pm\boldsymbol{0.0004}$ $1.2498$ $0.0948\pm 0.0036$
$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0669}$ $\boldsymbol{0.0692}\pm\boldsymbol{0.0182}$ $1.0669$ $0.2395\pm 0.0334$
$\mathcal{P}\left(3\right)$ $2.3954$ $81.4573\pm 5.9683$ $\boldsymbol{2.2639}$ $\boldsymbol{0.9140}\pm\boldsymbol{0.0747}$
Gaussian Distribution $\boldsymbol{2.5952}$ $\boldsymbol{0.0123}\pm\boldsymbol{0.0001}$ $2.5952$ $0.0472\pm 0.0007$
Fréchet Distribution $\boldsymbol{1.1539}$ $\boldsymbol{0.0085}\pm\boldsymbol{0.0001}$ $1.1539$ $0.0500\pm 0.0004$
Cauchy Distribution $\boldsymbol{1.6453}$ $\boldsymbol{0.0120}\pm\boldsymbol{0.0001}$ $1.6453$ $0.0519\pm 0.0002$
Pareto Distribution $\boldsymbol{0.8297}$ $\boldsymbol{0.0087}\pm\boldsymbol{0.0001}$ $0.8297$ $0.0444\pm 0.0005$
Table 22: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
Sparse Newton (T=50) Sparse Regularized Newton (T=50)
Riemannian Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $2.2478$ $\boldsymbol{0.0145}\pm\boldsymbol{0.0001}$ $\boldsymbol{2.2478}$ $0.0500\pm 0.0001$
$\mathbb{S}^{3}$ $2.4064$ $\boldsymbol{0.0581}\pm\boldsymbol{0.0177}$ $\boldsymbol{2.4064}$ $0.1629\pm 0.0247$
$\mathbb{S}^{5}$ $2.3480$ $\boldsymbol{0.2479}\pm\boldsymbol{0.0150}$ $\boldsymbol{2.3480}$ $0.5728\pm 0.0373$
$\mathbb{S}^{10}$ $\boldsymbol{1.9274}$ $2.2659\pm 0.1698$ $1.9277$ $\boldsymbol{1.9943}\pm\boldsymbol{0.1543}$
$\mathbb{S}^{20}$ $\boldsymbol{1.4610}$ $230.5914\pm 17.8627$ $1.4629$ $\boldsymbol{9.6611}\pm\boldsymbol{0.7412}$
$\mathbb{S}^{50}$ $0.9948$ $7394.6870\pm 845.2023$ $\boldsymbol{0.9720}$ $\boldsymbol{128.6652}\pm\boldsymbol{12.2889}$
$\mathbb{S}^{100}$ $\boldsymbol{0.6922}$ $\boldsymbol{712.1899}\pm\boldsymbol{54.3496}$ $0.7163$ $1410.9242\pm 104.6961$
$\mathbb{S}^{250}$ $\boldsymbol{0.4581}$ $\boldsymbol{13080.0830}\pm\boldsymbol{1387.7233}$- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.9330}$ $\boldsymbol{0.0207}\pm\boldsymbol{0.0002}$ $1.9330$ $0.0467\pm 0.0002$
$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0233}$ $\boldsymbol{0.0713}\pm\boldsymbol{0.0186}$ $2.0233$ $0.1612\pm 0.0257$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8046}$ $\boldsymbol{0.5449}\pm\boldsymbol{0.0668}$ $1.8050$ $0.7055\pm 0.0764$
$\mathrm{E}\left(10\right)$ $\boldsymbol{0.9473}$ $133.1282\pm 9.1672$ $1.4766$ $\boldsymbol{2.2189}\pm\boldsymbol{0.2136}$
$\mathrm{E}\left(20\right)$ $1.2745$ $591.8414\pm 35.1453$ $\boldsymbol{1.1484}$ $\boldsymbol{9.3243}\pm\boldsymbol{1.2287}$
$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5985}$ $7950.0571\pm 865.0385$ $0.7965$ $\boldsymbol{146.5869}\pm\boldsymbol{7.8365}$
$\mathrm{E}\left(100\right)$ $\boldsymbol{0.5707}$ $\boldsymbol{775.0126}\pm\boldsymbol{53.9842}$ $0.5863$ $1405.4775\pm 100.2025$
$\mathrm{E}\left(250\right)$ $\boldsymbol{0.3882}$ $\boldsymbol{9355.2041}\pm\boldsymbol{978.4170}$- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{9.8943}$ $3.4085\pm 0.2846$ $9.8955$ $\boldsymbol{0.0322}\pm\boldsymbol{0.0004}$
$\mathbb{H}^{2}$ $\boldsymbol{1.2434}$ $\boldsymbol{0.0115}\pm\boldsymbol{0.0003}$ $1.2435$ $0.0441\pm 0.0015$
$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0567}$ $\boldsymbol{0.0351}\pm\boldsymbol{0.0159}$ $1.0567$ $0.1206\pm 0.0212$
$\mathcal{P}\left(3\right)$ $2.3328$ $42.5019\pm 3.4439$ $\boldsymbol{2.2304}$ $\boldsymbol{0.3846}\pm\boldsymbol{0.0256}$
Gaussian Distribution $\boldsymbol{2.5778}$ $\boldsymbol{0.0060}\pm\boldsymbol{0.0001}$ $2.5778$ $0.0199\pm 0.0003$
Fréchet Distribution $\boldsymbol{1.1457}$ $\boldsymbol{0.0045}\pm\boldsymbol{0.0001}$ $1.1457$ $0.0217\pm 0.0002$
Cauchy Distribution $\boldsymbol{1.6340}$ $\boldsymbol{0.0079}\pm\boldsymbol{0.0002}$ $1.6340$ $0.0223\pm 0.0004$
Pareto Distribution $\boldsymbol{0.8235}$ $\boldsymbol{0.0042}\pm\boldsymbol{0.0001}$ $0.8235$ $0.0212\pm 0.0003$
Table 23: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
Methods on a Finsler manifold
We show in Table 24, Table 25, Table 26, Table 27, Table 28, Table 29, Table 30 and Table 31 runtime estimates for different methods on a CPU for different Finsler manifolds.
BFGS (T=100) ADAM (T=100) GEORCE (T=100)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.6525$ $244.9129\pm 0.0680$ $1.6536$ $0.4967\pm 0.0007$ $\boldsymbol{1.6525}$ $\boldsymbol{0.0194}\pm\boldsymbol{0.0009}$
$\mathbb{S}^{3}$ $1.7433$ $277.8015\pm 0.5819$ $1.7456$ $0.7639\pm 0.0031$ $\boldsymbol{1.7432}$ $\boldsymbol{0.0362}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{5}$ $1.7032$ $435.6987\pm 4.6022$ $1.7377$ $4.6711\pm 0.0047$ $\boldsymbol{1.7029}$ $\boldsymbol{0.4556}\pm\boldsymbol{0.0021}$
$\mathbb{S}^{10}$ $1.4151$ $812.2809\pm 13.5968$ $1.5174$ $9.7974\pm 0.0052$ $\boldsymbol{1.4109}$ $\boldsymbol{0.7473}\pm\boldsymbol{0.0029}$
$\mathbb{S}^{20}$ $1.0951$ $1446.0956\pm 4.8038$ $1.2018$ $18.3489\pm 0.0747$ $\boldsymbol{1.0760}$ $\boldsymbol{1.2856}\pm\boldsymbol{0.0016}$
$\mathbb{S}^{50}$ $0.8216$ $5992.7261\pm 56.5973$ $0.8128$ $146.4556\pm 0.9341$ $\boldsymbol{0.7285}$ $\boldsymbol{8.4758}\pm\boldsymbol{0.0234}$
$\mathbb{S}^{100}$- -- $0.6187$ $498.9705\pm 0.4032$ $\boldsymbol{0.2570}$ $\boldsymbol{15.3786}\pm\boldsymbol{0.0092}$
$\mathbb{S}^{250}$- -- $0.5195$ $1397.6732\pm 0.7793$ $\boldsymbol{0.1622}$ $\boldsymbol{132.5889}\pm\boldsymbol{0.1106}$
$\mathbb{S}^{500}$- -- $0.4850$ $1954.3231\pm 14.1107$ $\boldsymbol{0.1404}$ $\boldsymbol{364.5326}\pm\boldsymbol{1.8222}$
$\mathbb{S}^{1000}$- -- $0.3843$ $5982.0884\pm 204.1833$ $\boldsymbol{0.1140}$ $\boldsymbol{1338.9869}\pm\boldsymbol{2.3954}$
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3616}$ $230.0809\pm 1.3967$ $1.3621$ $0.5310\pm 0.0025$ $1.3618$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.4001}$ $416.6281\pm 5.5354$ $1.4087$ $0.8183\pm 0.0026$ $1.4023$ $\boldsymbol{0.1081}\pm\boldsymbol{0.0001}$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2951}$ $638.7185\pm 2.7818$ $1.3265$ $4.3802\pm 0.0039$ $1.2956$ $\boldsymbol{0.4554}\pm\boldsymbol{0.0014}$
$\mathrm{E}\left(10\right)$ $\boldsymbol{1.0733}$ $741.9274\pm 12.3382$ $1.1229$ $9.7899\pm 0.0117$ $1.0804$ $\boldsymbol{0.5815}\pm\boldsymbol{0.0022}$
$\mathrm{E}\left(20\right)$ $0.8510$ $1820.3137\pm 4.0734$ $0.8980$ $19.4614\pm 0.0039$ $\boldsymbol{0.8121}$ $\boldsymbol{1.0671}\pm\boldsymbol{0.0044}$
$\mathrm{E}\left(50\right)$ $0.5981$ $11191.9424\pm 597.7002$ $0.6325$ $146.6409\pm 0.1668$ $\boldsymbol{0.2842}$ $\boldsymbol{4.4377}\pm\boldsymbol{0.0451}$
$\mathrm{E}\left(100\right)$- -- $0.4936$ $491.8223\pm 0.2346$ $\boldsymbol{0.2009}$ $\boldsymbol{14.5367}\pm\boldsymbol{0.0092}$
$\mathrm{E}\left(250\right)$- -- $0.4159$ $1475.5049\pm 4.1430$ $\boldsymbol{0.1287}$ $\boldsymbol{125.4187}\pm\boldsymbol{0.0888}$
$\mathrm{E}\left(500\right)$- -- $0.4142$ $1984.7845\pm 10.5689$ $\boldsymbol{0.0967}$ $\boldsymbol{373.2816}\pm\boldsymbol{0.4227}$
$\mathrm{E}\left(1000\right)$- -- $0.3444$ $5929.4438\pm 167.1873$ $\boldsymbol{0.0799}$ $\boldsymbol{1211.8812}\pm\boldsymbol{0.2511}$
$\mathbb{T}^{2}$ $7.3890$ $449.4798\pm 10.3282$ $7.3585$ $0.3294\pm 0.0018$ $\boldsymbol{6.8067}$ $\boldsymbol{0.1126}\pm\boldsymbol{0.0002}$
$\mathbb{H}^{2}$ $1.3340$ $384.3149\pm 0.3044$ $1.3360$ $0.2744\pm 0.0024$ $\boldsymbol{1.3337}$ $\boldsymbol{0.0094}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7291}$ $414.8355\pm 3.0837$ $0.7338$ $0.6368\pm 0.0026$ $0.7292$ $\boldsymbol{0.0222}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $536.6575\pm 0.9323$ $1.2980$ $5.2824\pm 0.0540$ $1.2856$ $\boldsymbol{0.2094}\pm\boldsymbol{0.0004}$
Gaussian Distribution $1.6878$ $163.4399\pm 0.1859$ $1.6936$ $0.1145\pm 0.0001$ $\boldsymbol{1.6878}$ $\boldsymbol{0.0047}\pm\boldsymbol{0.0001}$
Fréchet Distribution $0.5408$ $123.8418\pm 0.0499$ $0.5406$ $0.0585\pm 0.0001$ $\boldsymbol{0.5405}$ $\boldsymbol{0.0021}\pm\boldsymbol{0.0001}$
Cauchy Distribution $1.1134$ $182.3980\pm 2.2559$ $\boldsymbol{1.1133}$ $0.1105\pm 0.0007$ $1.1133$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0002}$
Pareto Distribution $0.3884$ $84.8186\pm 0.0832$ $0.3885$ $0.0235\pm 0.0007$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0001}$
Table 24: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
BFGS (T=50) ADAM (T=50) GEORCE (T=50)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.6385$ $143.8810\pm 0.2377$ $\boldsymbol{1.6385}$ $0.2635\pm 0.0001$ $1.6385$ $\boldsymbol{0.0134}\pm\boldsymbol{0.0004}$
$\mathbb{S}^{3}$ $1.7298$ $146.9142\pm 1.0720$ $1.7298$ $0.3666\pm 0.0157$ $\boldsymbol{1.7297}$ $\boldsymbol{0.0245}\pm\boldsymbol{0.0001}$
$\mathbb{S}^{5}$ $1.6869$ $268.8883\pm 1.4754$ $1.6878$ $2.2747\pm 0.0047$ $\boldsymbol{1.6868}$ $\boldsymbol{0.3351}\pm\boldsymbol{0.0003}$
$\mathbb{S}^{10}$ $1.3919$ $452.0089\pm 2.1364$ $1.4013$ $4.5715\pm 0.0023$ $\boldsymbol{1.3908}$ $\boldsymbol{0.6533}\pm\boldsymbol{0.0180}$
$\mathbb{S}^{20}$ $1.0636$ $726.7038\pm 5.3951$ $1.0830$ $8.3452\pm 0.0040$ $\boldsymbol{1.0560}$ $\boldsymbol{1.4084}\pm\boldsymbol{0.0021}$
$\mathbb{S}^{50}$ $0.7320$ $1616.3099\pm 8.8835$ $0.7427$ $71.9378\pm 0.0786$ $\boldsymbol{0.3492}$ $\boldsymbol{2.2147}\pm\boldsymbol{0.0047}$
$\mathbb{S}^{100}$ $0.5832$ $4253.1743\pm 13.8589$ $0.5511$ $291.3856\pm 2.1209$ $\boldsymbol{0.2639}$ $\boldsymbol{7.4223}\pm\boldsymbol{0.0056}$
$\mathbb{S}^{250}$- -- $0.4094$ $1241.9021\pm 7.3069$ $\boldsymbol{0.1686}$ $\boldsymbol{52.1050}\pm\boldsymbol{0.0283}$
$\mathbb{S}^{500}$- -- $0.3531$ $2204.0300\pm 1.4159$ $\boldsymbol{0.1193}$ $\boldsymbol{239.8835}\pm\boldsymbol{0.2815}$
$\mathbb{S}^{1000}$- -- $0.3269$ $3347.6938\pm 44.0626$ $\boldsymbol{0.0828}$ $\boldsymbol{958.4282}\pm\boldsymbol{7.2917}$
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3459}$ $146.3793\pm 0.7961$ $1.3460$ $0.0436\pm 0.0001$ $1.3461$ $\boldsymbol{0.0084}\pm\boldsymbol{0.0002}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3826}$ $220.9877\pm 2.8045$ $1.3870$ $0.3788\pm 0.0023$ $1.3826$ $\boldsymbol{0.0972}\pm\boldsymbol{0.0002}$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2756}$ $333.3406\pm 0.9750$ $1.2780$ $2.2999\pm 0.0031$ $1.2757$ $\boldsymbol{0.3920}\pm\boldsymbol{0.0009}$
$\mathrm{E}\left(10\right)$ $\boldsymbol{0.5551}$ $1225.0781\pm 1.1283$ $1.0666$ $4.8530\pm 0.0046$ $1.0126$ $\boldsymbol{0.8426}\pm\boldsymbol{0.0013}$
$\mathrm{E}\left(20\right)$ $\boldsymbol{0.7340}$ $1834.0602\pm 20.3805$ $0.8404$ $8.9581\pm 0.0200$ $0.7946$ $\boldsymbol{0.9224}\pm\boldsymbol{0.0015}$
$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5735}$ $1238.1737\pm 3.1700$ $0.5797$ $67.3310\pm 0.0869$- --
$\mathrm{E}\left(100\right)$ $0.4369$ $6365.4741\pm 34.9823$ $0.4342$ $221.8947\pm 1.4868$ $\boldsymbol{0.2042}$ $\boldsymbol{6.5394}\pm\boldsymbol{0.0437}$
$\mathrm{E}\left(250\right)$- -- $0.3234$ $935.4907\pm 5.1887$ $\boldsymbol{0.1365}$ $\boldsymbol{51.8779}\pm\boldsymbol{0.0096}$
$\mathrm{E}\left(500\right)$- -- $0.2862$ $1512.1372\pm 0.9319$ $\boldsymbol{0.0940}$ $\boldsymbol{174.0314}\pm\boldsymbol{0.2612}$
$\mathrm{E}\left(1000\right)$- -- $0.2723$ $2488.8240\pm 38.0558$ $\boldsymbol{0.0893}$ $\boldsymbol{681.7435}\pm\boldsymbol{0.2447}$
$\mathbb{T}^{2}$ $\boldsymbol{7.0791}$ $598.7396\pm 0.0929$ $7.1796$ $\boldsymbol{0.1417}\pm\boldsymbol{0.0002}$ $9.0475$ $2.1109\pm 0.0035$
$\mathbb{H}^{2}$ $1.3317$ $189.9143\pm 0.0787$ $\boldsymbol{1.3316}$ $0.1403\pm 0.0026$ $1.3316$ $\boldsymbol{0.0070}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(2\right)$ $0.7249$ $224.4045\pm 0.4114$ $\boldsymbol{0.7249}$ $0.2194\pm 0.0025$ $0.7249$ $\boldsymbol{0.0153}\pm\boldsymbol{0.0001}$
$\mathcal{P}\left(3\right)$ $1.2707$ $306.2382\pm 0.1092$ $1.2735$ $2.7082\pm 0.0075$ $\boldsymbol{1.2707}$ $\boldsymbol{0.1276}\pm\boldsymbol{0.0014}$
Gaussian Distribution $1.6806$ $88.1351\pm 0.1027$ $1.6811$ $0.0569\pm 0.0001$ $\boldsymbol{1.6806}$ $\boldsymbol{0.0033}\pm\boldsymbol{0.0001}$
Fréchet Distribution $0.5371$ $71.0994\pm 0.0709$ $0.5370$ $0.0095\pm 0.0000$ $\boldsymbol{0.5370}$ $\boldsymbol{0.0015}\pm\boldsymbol{0.0001}$
Cauchy Distribution $1.1084$ $94.0708\pm 0.0792$ $1.1085$ $0.0075\pm 0.0001$ $\boldsymbol{1.1084}$ $\boldsymbol{0.0051}\pm\boldsymbol{0.0001}$
Pareto Distribution $0.3857$ $46.7253\pm 0.2142$ $0.3857$ $0.0084\pm 0.0000$ $\boldsymbol{0.3857}$ $\boldsymbol{0.0011}\pm\boldsymbol{0.0001}$
Table 25: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
SGD (T=100) CG (T=100) dogleg (T=100)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.7321$ $\boldsymbol{0.5226}\pm\boldsymbol{0.0033}$ $1.6536$ $864.7791\pm 2.1507$ $\boldsymbol{1.6525}$ $14.6421\pm 0.0605$
$\mathbb{S}^{3}$ $1.9179$ $\boldsymbol{0.8067}\pm\boldsymbol{0.0005}$ $\boldsymbol{1.7453}$ $976.1121\pm 13.5118$ $1.9228$ $4.1058\pm 0.0123$
$\mathbb{S}^{5}$ $2.1622$ $\boldsymbol{4.3635}\pm\boldsymbol{0.0015}$ $\boldsymbol{1.7099}$ $1326.8875\pm 10.6284$ $2.1703$ $5.8463\pm 0.0436$
$\mathbb{S}^{10}$ $2.4132$ $\boldsymbol{8.9293}\pm\boldsymbol{0.0051}$ $\boldsymbol{1.4427}$ $801.4553\pm 1.2792$ $2.4298$ $18.1452\pm 0.0392$
$\mathbb{S}^{20}$ $2.4912$ $\boldsymbol{19.4868}\pm\boldsymbol{0.0023}$ $\boldsymbol{1.2646}$ $397.4008\pm 2.7750$ $2.5207$ $98.6398\pm 0.1329$
$\mathbb{S}^{50}$ $2.2856$ $\boldsymbol{155.1010}\pm\boldsymbol{0.2023}$ $\boldsymbol{0.9295}$ $274.3937\pm 7.9883$- --
$\mathbb{S}^{100}$ $1.9542$ $619.2194\pm 0.4311$ $\boldsymbol{1.0756}$ $\boldsymbol{221.9957}\pm\boldsymbol{0.5803}$- --
$\mathbb{S}^{250}$ $1.4363$ $5475.3242\pm 1.4705$ $\boldsymbol{0.8813}$ $\boldsymbol{304.5134}\pm\boldsymbol{0.0238}$- --
$\mathbb{S}^{500}$- -- $\boldsymbol{1.0790}$ $\boldsymbol{31.3454}\pm\boldsymbol{0.0388}$- --
$\mathbb{S}^{1000}$- -- $\boldsymbol{0.7859}$ $\boldsymbol{137.1001}\pm\boldsymbol{0.4771}$- --
$\mathrm{E}\left(2\right)$ $1.3705$ $\boldsymbol{0.5219}\pm\boldsymbol{0.0024}$ $1.3620$ $401.8943\pm 0.7424$ $\boldsymbol{1.3616}$ $14.9190\pm 0.0640$
$\mathrm{E}\left(3\right)$ $1.4388$ $\boldsymbol{0.8083}\pm\boldsymbol{0.0019}$ $\boldsymbol{1.4147}$ $621.6700\pm 3.4375$ $1.4397$ $4.6115\pm 0.0556$
$\mathrm{E}\left(5\right)$ $1.5358$ $\boldsymbol{4.6840}\pm\boldsymbol{0.0044}$ $\boldsymbol{1.3355}$ $1523.9905\pm 19.2195$ $1.5373$ $5.7719\pm 0.0192$
$\mathrm{E}\left(10\right)$ $1.6639$ $\boldsymbol{9.7218}\pm\boldsymbol{0.0077}$ $\boldsymbol{1.0329}$ $2475.9639\pm 93.5344$ $1.6672$ $18.9040\pm 0.6761$
$\mathrm{E}\left(20\right)$ $1.7336$ $\boldsymbol{19.6573}\pm\boldsymbol{0.0115}$ $\boldsymbol{0.9341}$ $552.4011\pm 0.4176$ $1.7403$ $95.1917\pm 0.7952$
$\mathrm{E}\left(50\right)$ $1.6453$ $\boldsymbol{145.3531}\pm\boldsymbol{0.3967}$ $\boldsymbol{0.7045}$ $573.5270\pm 6.1649$- --
$\mathrm{E}\left(100\right)$ $1.4396$ $616.7589\pm 0.2897$ $\boldsymbol{0.7510}$ $\boldsymbol{394.1170}\pm\boldsymbol{2.5401}$- --
$\mathrm{E}\left(250\right)$ $1.0800$ $4937.1353\pm 67.9600$ $\boldsymbol{0.6266}$ $\boldsymbol{784.1846}\pm\boldsymbol{4.4711}$- --
$\mathrm{E}\left(500\right)$- -- $\boldsymbol{0.8183}$ $\boldsymbol{31.6893}\pm\boldsymbol{0.0422}$- --
$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.5997}$ $\boldsymbol{152.4971}\pm\boldsymbol{0.5062}$- --
$\mathbb{T}^{2}$- -- $\boldsymbol{8.5402}$ $198.5802\pm 0.2435$ $9.1903$ $4.4851\pm 0.0162$
$\mathbb{H}^{2}$ $1.4881$ $\boldsymbol{0.2459}\pm\boldsymbol{0.0024}$ $1.3515$ $2582.8269\pm 190.4313$ $\boldsymbol{1.3338}$ $24.6057\pm 0.0985$
$\mathcal{P}\left(2\right)$ $0.7860$ $\boldsymbol{0.6287}\pm\boldsymbol{0.0027}$ $0.7321$ $1333.9152\pm 50.5504$ $\boldsymbol{0.7291}$ $17.3661\pm 0.0366$
$\mathcal{P}\left(3\right)$ $1.3773$ $\boldsymbol{5.3062}\pm\boldsymbol{0.0015}$ $1.2912$ $920.9217\pm 15.9820$ $\boldsymbol{1.2856}$ $35.0991\pm 0.0589$
Gaussian Distribution $1.9173$ $\boldsymbol{0.1056}\pm\boldsymbol{0.0002}$ $1.6881$ $820.4708\pm 1.3159$ $\boldsymbol{1.6878}$ $15.1765\pm 0.0966$
Fréchet Distribution $0.5499$ $\boldsymbol{0.1143}\pm\boldsymbol{0.0009}$ $0.5414$ $302.1487\pm 0.2635$ $\boldsymbol{0.5405}$ $9.8568\pm 0.0634$
Cauchy Distribution $1.4099$ $\boldsymbol{0.1025}\pm\boldsymbol{0.0010}$ $1.1138$ $821.1230\pm 0.8392$ $\boldsymbol{1.1133}$ $17.1469\pm 0.0529$
Pareto Distribution $0.3951$ $\boldsymbol{0.1124}\pm\boldsymbol{0.0002}$ $0.3899$ $206.6691\pm 0.1273$ $\boldsymbol{0.3884}$ $5.9249\pm 0.0319$
Table 26: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
SGD (T=50) CG (T=50) dogleg (T=50)
Finsler Manifold Length Runtime Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.7109$ $\boldsymbol{0.2588}\pm\boldsymbol{0.0008}$ $1.6385$ $439.9231\pm 4.5034$ $\boldsymbol{1.6385}$ $14.8097\pm 0.0251$
$\mathbb{S}^{3}$ $1.8925$ $\boldsymbol{0.3882}\pm\boldsymbol{0.0126}$ $\boldsymbol{1.7298}$ $496.1117\pm 4.7323$ $1.9113$ $3.6687\pm 0.0188$
$\mathbb{S}^{5}$ $2.1279$ $\boldsymbol{2.2943}\pm\boldsymbol{0.0029}$ $\boldsymbol{1.6874}$ $658.4536\pm 4.7574$ $2.1588$ $4.0045\pm 0.0993$
$\mathbb{S}^{10}$ $2.3586$ $\boldsymbol{4.8071}\pm\boldsymbol{0.0036}$ $\boldsymbol{1.3911}$ $896.1763\pm 6.7043$ $2.4194$ $6.7420\pm 0.0149$
$\mathbb{S}^{20}$ $2.4105$ $\boldsymbol{9.5867}\pm\boldsymbol{0.0165}$ $\boldsymbol{1.1016}$ $622.5931\pm 6.0175$ $2.5121$ $25.6215\pm 0.0523$
$\mathbb{S}^{50}$ $2.1971$ $\boldsymbol{71.2637}\pm\boldsymbol{0.0976}$ $\boldsymbol{0.9319}$ $177.8937\pm 1.2493$- --
$\mathbb{S}^{100}$ $1.8917$ $295.5522\pm 1.0715$ $\boldsymbol{0.6486}$ $\boldsymbol{219.8582}\pm\boldsymbol{1.3192}$- --
$\mathbb{S}^{250}$ $1.4136$ $2760.2507\pm 17.1501$ $\boldsymbol{0.5816}$ $\boldsymbol{333.5332}\pm\boldsymbol{1.1282}$- --
$\mathbb{S}^{500}$ $1.0691$ $11508.7256\pm 47.0186$ $\boldsymbol{0.6111}$ $\boldsymbol{510.0441}\pm\boldsymbol{1.7312}$- --
$\mathbb{S}^{1000}$- -- $\boldsymbol{0.6762}$ $\boldsymbol{565.0285}\pm\boldsymbol{2.4234}$- --
$\mathrm{E}\left(2\right)$ $1.3601$ $\boldsymbol{0.2615}\pm\boldsymbol{0.0002}$ $1.3461$ $422.3503\pm 12.1072$ $\boldsymbol{1.3460}$ $20.7184\pm 0.0617$
$\mathrm{E}\left(3\right)$ $1.4291$ $\boldsymbol{0.4036}\pm\boldsymbol{0.0006}$ $\boldsymbol{1.3897}$ $1200.0380\pm 33.6871$ $1.4325$ $4.2812\pm 0.0319$
$\mathrm{E}\left(5\right)$ $1.5254$ $\boldsymbol{2.1031}\pm\boldsymbol{0.0024}$ $\boldsymbol{1.2803}$ $993.6235\pm 15.9966$ $1.5310$ $4.2148\pm 0.0296$
$\mathrm{E}\left(10\right)$ $1.6487$ $\boldsymbol{4.7998}\pm\boldsymbol{0.0110}$ $\boldsymbol{0.5483}$ $2210.6582\pm 37.5507$ $1.6615$ $6.6250\pm 0.0132$
$\mathrm{E}\left(20\right)$ $1.7096$ $\boldsymbol{9.5709}\pm\boldsymbol{0.0027}$ $\boldsymbol{0.8485}$ $564.2457\pm 0.9527$ $1.7353$ $23.9862\pm 0.0349$
$\mathrm{E}\left(50\right)$ $1.6121$ $\boldsymbol{73.0136}\pm\boldsymbol{0.0216}$ $\boldsymbol{0.6313}$ $344.0060\pm 1.0995$- --
$\mathrm{E}\left(100\right)$ $1.4125$ $\boldsymbol{313.0101}\pm\boldsymbol{0.3825}$ $\boldsymbol{0.4981}$ $331.0117\pm 1.3895$- --
$\mathrm{E}\left(250\right)$ $1.0688$ $2706.8379\pm 8.4708$ $\boldsymbol{0.5508}$ $\boldsymbol{378.6086}\pm\boldsymbol{0.2978}$- --
$\mathrm{E}\left(500\right)$ $0.8132$ $11307.9814\pm 281.9052$ $\boldsymbol{0.5677}$ $\boldsymbol{751.1521}\pm\boldsymbol{1.0706}$- --
$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.5987}$ $\boldsymbol{77.4691}\pm\boldsymbol{0.2770}$- --
$\mathbb{T}^{2}$ $inf$ $\boldsymbol{0.0490}\pm\boldsymbol{0.0001}$ $\boldsymbol{8.1529}$ $167.4743\pm 2.8798$ $9.0475$ $4.3053\pm 0.0156$
$\mathbb{H}^{2}$ $1.4850$ $\boldsymbol{0.1246}\pm\boldsymbol{0.0040}$ $1.3320$ $2460.6926\pm 57.6646$ $\boldsymbol{1.3316}$ $19.7572\pm 0.0371$
$\mathcal{P}\left(2\right)$ $0.7739$ $\boldsymbol{0.3208}\pm\boldsymbol{0.0030}$ $0.7250$ $712.9343\pm 11.1547$ $\boldsymbol{0.7249}$ $14.5680\pm 0.0587$
$\mathcal{P}\left(3\right)$ $1.3514$ $\boldsymbol{2.6554}\pm\boldsymbol{0.0101}$ $1.2709$ $788.9128\pm 8.9808$ $\boldsymbol{1.2707}$ $22.6905\pm 0.0425$
Gaussian Distribution $1.8463$ $\boldsymbol{0.0519}\pm\boldsymbol{0.0001}$ $1.6807$ $283.3971\pm 6.7184$ $\boldsymbol{1.6806}$ $16.2727\pm 0.1111$
Fréchet Distribution $0.5455$ $\boldsymbol{0.0596}\pm\boldsymbol{0.0001}$ $0.5372$ $253.2346\pm 0.2384$ $\boldsymbol{0.5370}$ $9.3157\pm 0.0269$
Cauchy Distribution $1.3641$ $\boldsymbol{0.0522}\pm\boldsymbol{0.0001}$ $1.1088$ $360.8286\pm 0.4404$ $\boldsymbol{1.1084}$ $16.8439\pm 0.0398$
Pareto Distribution $0.3919$ $\boldsymbol{0.0555}\pm\boldsymbol{0.0001}$ $0.3857$ $142.1006\pm 0.0492$ $\boldsymbol{0.3856}$ $7.9442\pm 0.0303$
Table 27: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
trust-ncg (T=100) trust-exact (T=100)
Finsler Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.6525$ $1261.5378\pm 7.7825$ $\boldsymbol{1.6525}$ $\boldsymbol{14.7265}\pm\boldsymbol{0.0327}$
$\mathbb{S}^{3}$ $1.7432$ $1994.5762\pm 20.6694$ $\boldsymbol{1.7432}$ $\boldsymbol{20.4180}\pm\boldsymbol{0.0359}$
$\mathbb{S}^{5}$ $\boldsymbol{1.7027}$ $5603.5640\pm 76.5314$ $1.7027$ $\boldsymbol{38.7189}\pm\boldsymbol{0.0167}$
$\mathbb{S}^{10}$- -- $\boldsymbol{1.4108}$ $\boldsymbol{184.4648}\pm\boldsymbol{0.3001}$
$\mathbb{S}^{20}$- -- $\boldsymbol{1.0769}$ $\boldsymbol{1087.9535}\pm\boldsymbol{2.3390}$
$\mathbb{S}^{50}$- -- -- --
$\mathbb{S}^{100}$- -- -- --
$\mathbb{S}^{250}$- -- -- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $1.3616$ $1019.1379\pm 8.5157$ $\boldsymbol{1.3616}$ $\boldsymbol{15.4908}\pm\boldsymbol{0.0374}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3997}$ $4148.8574\pm 2.3006$ $1.3997$ $\boldsymbol{31.8261}\pm\boldsymbol{0.1322}$
$\mathrm{E}\left(5\right)$ $1.2950$ $5965.7002\pm 44.3871$ $\boldsymbol{1.2950}$ $\boldsymbol{40.3707}\pm\boldsymbol{0.0405}$
$\mathrm{E}\left(10\right)$- -- $\boldsymbol{1.0318}$ $\boldsymbol{249.5034}\pm\boldsymbol{0.2415}$
$\mathrm{E}\left(20\right)$- -- $\boldsymbol{0.8109}$ $\boldsymbol{1033.4261}\pm\boldsymbol{3.7266}$
$\mathrm{E}\left(50\right)$- -- -- --
$\mathrm{E}\left(100\right)$- -- -- --
$\mathrm{E}\left(250\right)$- -- -- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{8.9015}$ $348.9266\pm 0.5435$ $9.1722$ $\boldsymbol{56.1951}\pm\boldsymbol{0.0567}$
$\mathbb{H}^{2}$ $\boldsymbol{1.3338}$ $3365.4121\pm 15.3287$ $1.3338$ $\boldsymbol{24.6266}\pm\boldsymbol{0.1178}$
$\mathcal{P}\left(2\right)$ $0.7291$ $2314.0154\pm 22.6535$ $\boldsymbol{0.7291}$ $\boldsymbol{16.6480}\pm\boldsymbol{0.0684}$
$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $3887.0437\pm 81.9548$ $1.2856$ $\boldsymbol{42.8468}\pm\boldsymbol{0.1088}$
Gaussian Distribution $\boldsymbol{1.6878}$ $1176.3495\pm 4.4250$ $1.6878$ $\boldsymbol{13.9002}\pm\boldsymbol{0.0735}$
Fréchet Distribution $0.5405$ $1173.9244\pm 4.9721$ $\boldsymbol{0.5405}$ $\boldsymbol{9.5119}\pm\boldsymbol{0.0199}$
Cauchy Distribution $\boldsymbol{1.1133}$ $1425.5028\pm 8.3454$ $1.1133$ $\boldsymbol{14.6313}\pm\boldsymbol{0.0416}$
Pareto Distribution $\boldsymbol{0.3884}$ $406.4951\pm 4.4740$ $0.3884$ $\boldsymbol{5.9478}\pm\boldsymbol{0.0531}$
Table 28: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
trust-ncg (T=50) trust-exact (T=50)
Finsler Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $1.6385$ $461.6090\pm 2.4283$ $\boldsymbol{1.6385}$ $\boldsymbol{14.9718}\pm\boldsymbol{0.0477}$
$\mathbb{S}^{3}$ $1.7297$ $738.8096\pm 14.1991$ $\boldsymbol{1.7297}$ $\boldsymbol{18.2119}\pm\boldsymbol{0.0850}$
$\mathbb{S}^{5}$ $1.6868$ $1845.5012\pm 28.1314$ $\boldsymbol{1.6868}$ $\boldsymbol{27.8877}\pm\boldsymbol{0.0334}$
$\mathbb{S}^{10}$ $\boldsymbol{1.3881}$ $11723.0781\pm 267.5437$ $1.3881$ $\boldsymbol{72.3786}\pm\boldsymbol{0.5046}$
$\mathbb{S}^{20}$- -- $\boldsymbol{1.0591}$ $\boldsymbol{321.4109}\pm\boldsymbol{0.3724}$
$\mathbb{S}^{50}$- -- -- --
$\mathbb{S}^{100}$- -- -- --
$\mathbb{S}^{250}$- -- -- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $1.3460$ $760.6122\pm 4.1086$ $\boldsymbol{1.3460}$ $\boldsymbol{18.9955}\pm\boldsymbol{0.0502}$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3825}$ $1558.1522\pm 1.8558$ $1.3825$ $\boldsymbol{25.1379}\pm\boldsymbol{0.0384}$
$\mathrm{E}\left(5\right)$ $1.2756$ $1658.7789\pm 4.1942$ $\boldsymbol{1.2755}$ $\boldsymbol{31.1768}\pm\boldsymbol{0.0425}$
$\mathrm{E}\left(10\right)$ $1.0112$ $15954.7256\pm 487.2155$ $\boldsymbol{1.0109}$ $\boldsymbol{104.7452}\pm\boldsymbol{1.3478}$
$\mathrm{E}\left(20\right)$- -- $\boldsymbol{0.8000}$ $\boldsymbol{371.1157}\pm\boldsymbol{2.4557}$
$\mathrm{E}\left(50\right)$- -- -- --
$\mathrm{E}\left(100\right)$- -- -- --
$\mathrm{E}\left(250\right)$- -- -- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{8.4108}$ $\boldsymbol{364.3664}\pm\boldsymbol{2.8448}$- --
$\mathbb{H}^{2}$ $1.3316$ $1348.8287\pm 29.8787$ $\boldsymbol{1.3316}$ $\boldsymbol{19.5780}\pm\boldsymbol{0.0558}$
$\mathcal{P}\left(2\right)$ $0.7249$ $702.5833\pm 1.9384$ $\boldsymbol{0.7249}$ $\boldsymbol{14.2773}\pm\boldsymbol{0.0611}$
$\mathcal{P}\left(3\right)$ $1.2707$ $1289.7513\pm 26.6135$ $\boldsymbol{1.2707}$ $\boldsymbol{20.2860}\pm\boldsymbol{0.0433}$
Gaussian Distribution $\boldsymbol{1.6806}$ $586.1818\pm 0.5433$ $1.6806$ $\boldsymbol{13.2321}\pm\boldsymbol{0.0358}$
Fréchet Distribution $\boldsymbol{0.5370}$ $458.0856\pm 0.2181$ $0.5370$ $\boldsymbol{9.8644}\pm\boldsymbol{0.0257}$
Cauchy Distribution $\boldsymbol{1.1084}$ $584.2502\pm 1.5708$ $1.1084$ $\boldsymbol{13.5337}\pm\boldsymbol{0.0222}$
Pareto Distribution $\boldsymbol{0.3856}$ $366.4953\pm 0.2211$ $\boldsymbol{0.3856}$ $\boldsymbol{8.6113}\pm\boldsymbol{0.0271}$
Table 29: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
Sparse Newton (T=100) Sparse Regularized Newton (T=100)
Finsler Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $\boldsymbol{1.6525}$ $\boldsymbol{0.0632}\pm\boldsymbol{0.0009}$ $1.6525$ $0.4062\pm 0.0273$
$\mathbb{S}^{3}$ $\boldsymbol{1.7432}$ $\boldsymbol{0.3018}\pm\boldsymbol{0.0366}$ $1.7433$ $1.0369\pm 0.0880$
$\mathbb{S}^{5}$ $\boldsymbol{1.7027}$ $11.3808\pm 0.8228$ $1.7028$ $\boldsymbol{3.5064}\pm\boldsymbol{0.3096}$
$\mathbb{S}^{10}$ $\boldsymbol{1.4092}$ $121.2282\pm 12.4696$ $1.4166$ $\boldsymbol{11.3672}\pm\boldsymbol{0.8836}$
$\mathbb{S}^{20}$ $1.3030$ $2840.8333\pm 230.3178$ $\boldsymbol{1.0865}$ $\boldsymbol{52.2612}\pm\boldsymbol{5.1479}$
$\mathbb{S}^{50}$- -- $\boldsymbol{0.7556}$ $\boldsymbol{809.2408}\pm\boldsymbol{68.4436}$
$\mathbb{S}^{100}$ $0.6141$ $16876.4219\pm 1314.8346$ $\boldsymbol{0.5968}$ $\boldsymbol{6984.5508}\pm\boldsymbol{566.5103}$
$\mathbb{S}^{250}$- -- -- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3616}$ $\boldsymbol{0.0916}\pm\boldsymbol{0.0029}$ $1.3619$ $0.3423\pm 0.0174$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3997}$ $\boldsymbol{0.5216}\pm\boldsymbol{0.0511}$ $1.3997$ $1.3200\pm 0.1149$
$\mathrm{E}\left(5\right)$ $1.5359$ $195.1668\pm 21.6446$ $\boldsymbol{1.2952}$ $\boldsymbol{3.7467}\pm\boldsymbol{0.3126}$
$\mathrm{E}\left(10\right)$ $1.2272$ $560.2833\pm 45.3723$ $\boldsymbol{1.0807}$ $\boldsymbol{10.0806}\pm\boldsymbol{0.7869}$
$\mathrm{E}\left(20\right)$ $1.0210$ $2939.9907\pm 244.8212$ $\boldsymbol{0.8555}$ $\boldsymbol{52.9693}\pm\boldsymbol{5.7635}$
$\mathrm{E}\left(50\right)$- -- $\boldsymbol{0.6211}$ $\boldsymbol{733.2034}\pm\boldsymbol{60.1911}$
$\mathrm{E}\left(100\right)$- -- $\boldsymbol{0.5132}$ $\boldsymbol{6657.5815}\pm\boldsymbol{538.6103}$
$\mathrm{E}\left(250\right)$- -- -- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{9.1903}$ $\boldsymbol{27.4443}\pm\boldsymbol{2.4951}$ $\boldsymbol{9.1903}$ $27.5325\pm 2.2826$
$\mathbb{H}^{2}$ $\boldsymbol{1.3338}$ $\boldsymbol{0.0816}\pm\boldsymbol{0.0019}$ $1.3340$ $0.3529\pm 0.0227$
$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7291}$ $\boldsymbol{0.1944}\pm\boldsymbol{0.0248}$ $0.7291$ $0.7200\pm 0.0868$
$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $\boldsymbol{0.8253}\pm\boldsymbol{0.0772}$ $1.2857$ $2.7333\pm 0.2234$
Gaussian Distribution $\boldsymbol{1.6878}$ $\boldsymbol{0.0636}\pm\boldsymbol{0.0006}$ $1.6878$ $0.2146\pm 0.0122$
Fréchet Distribution $\boldsymbol{0.5405}$ $\boldsymbol{0.0368}\pm\boldsymbol{0.0001}$ $0.5411$ $0.1835\pm 0.0107$
Cauchy Distribution $\boldsymbol{1.1133}$ $\boldsymbol{0.0720}\pm\boldsymbol{0.0015}$ $1.1133$ $0.2079\pm 0.0126$
Pareto Distribution $\boldsymbol{0.3884}$ $\boldsymbol{0.0159}\pm\boldsymbol{0.0002}$ $0.3886$ $0.1791\pm 0.0097$
Table 30: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.
Sparse Newton (T=50) Sparse Regularized Newton (T=50)
Finsler Manifold Length Runtime Length Runtime
$\mathbb{S}^{2}$ $\boldsymbol{1.6385}$ $\boldsymbol{0.0504}\pm\boldsymbol{0.0005}$ $1.6385$ $0.1821\pm 0.0097$
$\mathbb{S}^{3}$ $\boldsymbol{1.7297}$ $\boldsymbol{0.1977}\pm\boldsymbol{0.0270}$ $1.7297$ $0.4563\pm 0.0481$
$\mathbb{S}^{5}$ $1.6868$ $\boldsymbol{0.8981}\pm\boldsymbol{0.0788}$ $\boldsymbol{1.6868}$ $1.7423\pm 0.1372$
$\mathbb{S}^{10}$ $\boldsymbol{1.3354}$ $332.2007\pm 26.8370$ $1.3881$ $\boldsymbol{6.2534}\pm\boldsymbol{0.5135}$
$\mathbb{S}^{20}$ $\boldsymbol{0.8588}$ $1562.1842\pm 125.8776$ $1.0579$ $\boldsymbol{30.9976}\pm\boldsymbol{2.3280}$
$\mathbb{S}^{50}$- -- $\boldsymbol{0.7149}$ $\boldsymbol{364.6181}\pm\boldsymbol{32.4896}$
$\mathbb{S}^{100}$- -- $\boldsymbol{0.5364}$ $\boldsymbol{3463.5500}\pm\boldsymbol{274.5852}$
$\mathbb{S}^{250}$- -- -- --
$\mathbb{S}^{500}$- -- -- --
$\mathbb{S}^{1000}$- -- -- --
$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3460}$ $\boldsymbol{0.0658}\pm\boldsymbol{0.0015}$ $1.3461$ $0.1592\pm 0.0057$
$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3825}$ $\boldsymbol{0.2757}\pm\boldsymbol{0.0323}$ $1.3825$ $0.6126\pm 0.0610$
$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2756}$ $\boldsymbol{0.8444}\pm\boldsymbol{0.0827}$ $1.2756$ $1.6410\pm 0.1521$
$\mathrm{E}\left(10\right)$ $\boldsymbol{1.0106}$ $\boldsymbol{276.2837}\pm\boldsymbol{26.7637}$ $1.0575$ $284.1870\pm 27.1057$
$\mathrm{E}\left(20\right)$ $0.8571$ $1488.3325\pm 147.3422$ $\boldsymbol{0.7945}$ $\boldsymbol{30.3454}\pm\boldsymbol{1.4186}$
$\mathrm{E}\left(50\right)$- -- $\boldsymbol{0.5718}$ $\boldsymbol{384.6765}\pm\boldsymbol{30.0164}$
$\mathrm{E}\left(100\right)$- -- $\boldsymbol{0.4397}$ $\boldsymbol{3289.7231}\pm\boldsymbol{270.1520}$
$\mathrm{E}\left(250\right)$- -- -- --
$\mathrm{E}\left(500\right)$- -- -- --
$\mathrm{E}\left(1000\right)$- -- -- --
$\mathbb{T}^{2}$ $\boldsymbol{9.0475}$ $\boldsymbol{14.0754}\pm\boldsymbol{1.0615}$ $\boldsymbol{9.0475}$ $14.1199\pm 1.0787$
$\mathbb{H}^{2}$ $\boldsymbol{1.3316}$ $\boldsymbol{0.0443}\pm\boldsymbol{0.0002}$ $1.3317$ $0.1778\pm 0.0101$
$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7249}$ $\boldsymbol{0.1060}\pm\boldsymbol{0.0219}$ $0.7249$ $0.3272\pm 0.0354$
$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2707}$ $\boldsymbol{0.4780}\pm\boldsymbol{0.0503}$ $1.2707$ $1.1371\pm 0.1027$
Gaussian Distribution $\boldsymbol{1.6806}$ $\boldsymbol{0.0348}\pm\boldsymbol{0.0003}$ $1.6807$ $0.0914\pm 0.0039$
Fréchet Distribution $\boldsymbol{0.5370}$ $\boldsymbol{0.0194}\pm\boldsymbol{0.0004}$ $0.5371$ $0.0880\pm 0.0009$
Cauchy Distribution $\boldsymbol{1.1084}$ $\boldsymbol{0.0411}\pm\boldsymbol{0.0002}$ $1.1085$ $0.0975\pm 0.0039$
Pareto Distribution $\boldsymbol{0.3856}$ $\boldsymbol{0.0174}\pm\boldsymbol{0.0004}$ $0.3857$ $0.0884\pm 0.0048$
Table 31: The table shows the length of the estimated geodesics for different methods on a CPU. The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached. When the computational time was longer than $24$ hours, the value is set to --.

### v2

@@
 However, these approximations tend to be either numerically unstable, have slow convergence, or scale poorly with manifold dimension and number of grid points.
-We introduce a new algorithm called GEORCE that computes geodesics via a transformation into a discrete control problem.
+We introduce a new algorithm called GEORCE that computes geodesics in a local chart via a transformation into a discrete control problem.
 We show that GEORCE has global convergence and quadratic local convergence.
@@
 The lengths of geodesic curves inform us about manifold distances.
-Likewise, the exponential and logarithm maps, which generalize addition and subtraction, are rooted in geodesics [22].
+Likewise, the exponential and logarithm maps, which generalize vector addition and subtraction, are rooted in geodesics [22, Chapter 9].
 Unfortunately, the solution to the boundary-value problem for geodesics are only available in closed form for the simplest manifolds, and numerical approximations are therefore required in practice.
-Such approximations are typically achieved by minimizing the energy functional over a suitable set of curves connecting the given boundary points and parametrized by $\gamma:[0,1]\rightarrow\mathcal{M}$ [23]
+Such approximations are typically achieved by minimizing the energy functional over a suitable set of curves connecting the given boundary points and parametrized by $\gamma:[0,1]\rightarrow U$ [23, Page 194]
 $\mathcal{E}(\gamma)=\frac{1}{2}\int_{0}^{1}\dot{\gamma}(t)^{\top}G\left(\gamma(t)\right)\dot{\gamma}(t)\,\mathrm{d}t,$ (1)
-where $G$ denotes the metric matrix function of the manifold in local coordinates.
+where $G$ denotes the metric matrix function of the manifold in local coordinates in a chart $\phi:U\rightarrow\mathcal{M}$, where $U$ is an open subset of $\mathbb{R}^{d}$.
 This energy is commonly minimized with general-purpose gradient-based optimization methods [24, 11, 12] or, in particular when considering the initial value problem and using shooting methods, by solving the system of ordinary differential equations that arise by applying the Euler-Lagrange equations [25, 26].
@@
 Also note, that in the case of the torus the initial curve determines the direction around the hole so that the obtained geodesics follow this direction and are therefore clearly not the globally shortest geodesics between the given start and end point.
-Experimental details and benchmark data are in Appendix E.
+Experimental details and benchmark data are in Appendix F.
 In this paper, we introduce the GEORCE algorithm (GEodesic Optimization algoRithm using Control tEchniques) that computes numerical solutions to the geodesic boundary value problem via an optimal control strategy.
@@
 The metric, $g:T_{x}\mathcal{M}\times T_{x}\mathcal{M}\rightarrow\mathbb{R}$, defines an inner product that smoothly varies over the tangent spaces of $\mathcal{M}$ such that in a local chart $g(v,w)=v^{\top}G(x)w$, where $G$ is the metric matrix function evaluated in $x\in\mathcal{M}$ for $v,w\in T_{x}\mathcal{M}$.
-The tangent space, $T_{x}\mathcal{M}$, consists of tangents to all curves at $x\in\mathcal{M}$ [23].
-A geodesic can then be defined using the inner product in each tangent space. The length of a curve, $\gamma:U\subseteq[0,1]\rightarrow\mathcal{M}$ parameterized by $t\in[0,1]$ is given by
+The tangent space, $T_{x}\mathcal{M}$, consists of tangents to all curves at $x\in\mathcal{M}$ [23, Page 7 Definition 2.6].
+A geodesic can then be defined using the inner product in each tangent space. The length of a curve, $\gamma:[0,1]\rightarrow U$ parameterized by $t\in[0,1]$ is given by
 $\mathcal{L}(\gamma)=\int_{0}^{1}\sqrt{\dot{\gamma}(t)^{\top}G(\gamma(t))\dot{\gamma}(t)}\,\mathrm{d}t,$ (2)
-with the corresponding energy given by eq.
-1.
-A geodesic is a smooth curve that locally minimizes the length between two points $x,y\in\mathcal{M}$ and gives rise to the Riemannian distance, $\mathrm{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}_{+}$, which is defined as the length of the corresponding shortest geodesic between the given two points on the manifold [23].
-It can be shown that if a curve $\gamma$ locally minimizes the energy (1), then the curve has constant velocity and locally minimizes the length [30].
-For general Riemannian manifolds, existence and uniqueness are only guaranteed in a star-shaped neighborhood around any point $x\in\mathcal{M}$ [23].
-All points, where the geodesics from a point $x\in\mathcal{M}$ are no longer length minimizing, are denoted the cut locus of $x$ [23].
-In this paper, we assume that the manifold is geodesically complete such that at least one length-minimizing geodesic exists between any two points $x,y\in\mathcal{M}$.
-A Riemannian manifold can be equipped with a unique metrically compatible connection, $\nabla$, known as the Levi-Civita Connection [23], where uniqueness and existence are guaranteed by completeness of the metric [23].
+where $\phi:U\rightarrow\mathcal{M}$ is a local chart for an open set $U\subset\mathbb{R}^{d}$.
+A Riemannian manifold can be equipped with a unique metrically compatible connection, $\nabla$, known as the Levi-Civita Connection [23, Page 55 Theorem 3.6], where uniqueness and existence are guaranteed by completeness of the metric [23, Page 63 Lemma 2.3].
+Note that curves are geodesics if and only if they are critical points of the energy functional, i.e., the first variation of the energy functional is zero for all variations with fixed endpoints assuming that the end points are not conjugate points [23, Page 196 Proposition 2.5].
+This also holds in the Finslerian case [29, Page 81].
 An equivalent way to define geodesics is to define these as curves, $\gamma$, with zero acceleration using the Levi-Civita connection as the covariant derivative, i.e., $\nabla_{\dot{\gamma}(t)}\dot{\gamma}(t)=0$.
-In a local chart, this gives rise to the following (usually non-linear) ode-system [23]
+In a local chart, this gives rise to the following (usually non-linear) ode-system [23, Page 62]
 $\frac{\mathrm{d}^{2}\gamma^{k}}{\mathrm{d}t^{2}}+\Gamma_{ij}^{k}\frac{\mathrm{d}\gamma^{i}}{\mathrm{d}t}\frac{\mathrm{d}\gamma^{j}}{\mathrm{d}t}=0,\quad\gamma(0)=a,\gamma(1)=b,$ (3)
@@
 $\dot{\gamma}(0)=v$.
-Within the cut locus of $x\in\mathcal{M}$ the exponential map, $\mathrm{Exp}_{x}$, is a diffeomorphism and its inverse is the logarithmic map, $\mathrm{Log}_{x}:\mathcal{M}\rightarrow T_{x}\mathcal{M}$.
+Within the injectivity radius of $x$ the exponential map, $\mathrm{Exp}_{x}$, is a diffeomorphism and its inverse is the logarithmic map, $\mathrm{Log}_{x}:\mathcal{M}\rightarrow T_{x}\mathcal{M}$.
 Examples of Riemannian manifolds include information geometric spaces, i.e., statistical manifolds, where “points” in the manifold correspond to statistical distributions.
@@
 1 can be discretized and minimized using standard optimization methods.
-Consider fixed start and end points $x_{0}=a\in\mathcal{M}$ and $x_{T}=b\in\mathcal{M}$, and let $x_{0:T}:=\{x_{t}\}_{t=0}^{T}$ denote a discretized version of the candidate curve $\gamma$ in $T+1$ points.
-Minimizing the energy functional can be re-written as a discrete constrained optimization problem, i.e., subdivision of $[0,1]$ into $T$ subintervals, in the chosen local coordinate system [25, 12]
+Consider fixed start and end points $x_{0}=a\in U$ and $x_{T}=b\in U$, and let $x_{0:T}:=\{x_{t}\}_{t=0}^{T}$ denote a discretized version of the candidate curve $\gamma$ in $T+1$ points in a local chart $U$.
+Minimizing the energy functional can be re-written as a discrete constrained optimization problem, i.e., subdivision of $[0,1]$ into $T$ subintervals, in the chosen local chart [11, 12]
 $\begin{split}E(x_{0:T})=\quad\min_{x_{0:T}}\quad&\sum_{t=0}^{T-1}(x_{t+1}-x_{t})^{\top}G(x_{t})(x_{t+1}-x_{t})\\ \text{s.t.}\quad&x_{0}=a,x_{T}=b,\end{split}$ (5)
@@
 This is not restrictive since $\gamma$ can be approximated into a curve for any given precision $T$ or converted to a smooth function, e.g., using splines [24].
+Note that by approximating the functional energy by forward differences, the computational error is linear in $\nicefrac{{1}}{{T}}$, that is, in the form $\mathcal{O}\left(\nicefrac{{1}}{{T}}\right)$.
+Figure 2: The estimates of bvp-solvers with different integration methods compared to GEORCE on $\mathbb{S}^{n}$ for $n=2,3,5,10$.
+The bvp-solvers minimize the squared error to the boundary condition using BFGS [32, 33, 34, 35].
+All methods are terminated if they take more than 24 hours or use more than 10 GB of memory on a CPU.
+Details on the experiment is found in Appendix F.1.
+Similarly to the energy functional in Eq.
+1, it is assumed that a stationary point to the discretized energy functional in Eq.
+5 is either a local or global minimum point.
+Thus, the number of points, T, in Eq.
+5 needs to be sufficiently high to achieve this.
+If this is not fulfilled, then there may exist saddle points for the discretized energy functional.
 The optimization problem in Eq.
 5 can be solved using standard solvers in the variables $x_{1:(T-1)}$.
-Gradient descent converges to a local minimum (global convergence), but will have local linear convergence, while a Newton method or quasi-Newton method will have local quadratic and super-linear convergence, respectively, but will involve solving a linear equation system and will, in general, not have global convergence [33].
+Gradient descent converges to a local minimum (global convergence), but will have local linear convergence, while a Newton method or quasi-Newton method will have local quadratic and super-linear convergence, respectively, but will involve solving a linear equation system and will, in general, not have global convergence [36, Chapter 8].
 Limitations of existing numerical solvers.
@@
 However, these methods often require access to higher-order information about the metric tensor, e.g in the form of Christoffel symbols.
-Figure 2: The estimates of bvp-solvers with different integration methods compared to GEORCE on $\mathbb{S}^{n}$ for $n=2,3,5,10$.
-The bvp-solvers minimize the squared error to the boundary condition using BFGS [37, 38, 39, 40].
-All methods are terminated if they take more than 24 hours or use more than 10 GB of memory on a CPU.
-Details on the experiment is found in Appendix E.1.
 Directly minimizing the energy in Eq.
-5 can, e.g., be done using off-the-self solvers like ADAM [41] or BFGS [37, 38, 39, 40].
+5 can, e.g., be done using off-the-shelf solvers like ADAM [40] or BFGS [32, 33, 34, 35].
 However, these methods tend to have slow convergence and scale poorly in the manifold dimension and number of grid points (see Sec.
@@
 3 GEORCE
-To develop an efficient primal-dual algorithm for computing geodesics, we will first reformulate the discretized energy functional (5) as a discrete control problem.
+To develop an efficient primal-dual algorithm for computing geodesics, we will first reformulate the discretized energy functional (5) as a discrete control problem in a local chart.
 We then derive the necessary conditions for a minimum for the control problem using techniques from optimal control to decompose the discrete control problem into strictly convex subproblems; from this, we deduce an iterative scheme to compute geodesics.
 Finally, we prove that our proposed algorithm exhibits global convergence and local quadratic convergence.
-As a starting point, we reformulate the problem of finding a minimum of the discrete energy functional (5) to a control problem
+We will assume the following for our algorithm
+  • We assume access to a local chart of the manifold and assume that the entire candidate curve of the geodesic can be represented within the same chart.
+  • All computations are performed within the local coordinate chart $U$, which is defined based on the induced metric of the Riemannian manifold.
+  • For computational efficiency, we will leverage the Euclidean properties of the local chart (i.e., straight-line interpolation within $U$).
+  • We will assume that the end points are not conjugate points, and that the discretization is sufficiently fine such that a critical point of the discretized energy functional in Eq.
+5 is either a local or global minimum point.
+3.1 Deriviaton of GEORCE
+As a starting point, we reformulate the problem of finding a minimum of the discrete energy functional (5) to a control problem in a local chart
 $\begin{split}\min_{(x_{1:T-1},u_{0:T-1})}\quad&\left\{\sum_{t=0}^{T-1}u_{t}^{\top}G(x_{t})u_{t}\right\}\\ \text{s.t.}\quad&x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a,x_{T}=b.\end{split}$ (6)
@@
 Proof.
-See Appendix A.2. ∎
+The Hamiltonian of Eq. 6 is
+$H_{t}(x_{t},u_{t},\mu_{t})=u_{t}^{\top}G(x_{t})u_{t}+\mu_{t}^{\top}(x_{t}+u_{t}).$
+The Hamiltonian $H_{t}(x_{t},u_{t},\mu_{t})$ is strictly convex in $u_{t}$, since $G(x_{t})$ is positive definite.
+The time-discrete version of Pontryagins maximum principle [47] then yields the following when applied to our control problem (6)
+$\begin{split}&2G(x_{t})u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &{\left.\kern-1.2pt\nabla_{y}\left[u_{t}^{\top}G(y)u_{t}\right]\vphantom{\big|}\right|_{y=x_{t}}}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1.\\ &x_{0}=a,x_{T}=b.\end{split}$ (8)
 Proposition 1 decomposes the minimization problem in Eq.
@@
 6 and are sufficient in the case where the objective function in Eq.
-6 is convex in both $x_{t}$ and $u_{t}$.
-However, this is generally not the case.
+6 is convex in both $x_{t}$ and $u_{t}$ [48, Page 244].
+However, this is generally not the case, but due to the assumption applying a sufficiently fine discretization the solution to Eq.
+7 will be a local or global minimum point.
 The problem with the necessary conditions in Proposition 1 is that the equations cannot be solved in closed form for a general metric matrix function, $G$.
@@
 Fixing these in iteration $i$, the system of equations in Eq. 7 reduces to
-$\begin{split}&2G_{t}u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &\nu_{t}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1,\\ &\sum_{t=0}^{T-1}u_{t}=b-a,\\ \end{split}$ (9)
+[…]
 where $\nu_{1:(T-1)}$ and $G_{0:(T-1)}$ are fixed (9).
@@
 Proof.
-See Appendix A.3. ∎
+From the update formula $\nu_{t}+\mu_{t}=\mu_{t-1}$ for $t=1,\dots,T-1$ we deduce that $\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}$.
+From the update scheme $2G_{t}u_{t}+\mu_{t}=0$ we deduce that $u_{t}=-\frac{1}{2}G_{t}^{-1}\mu_{t}$, since $G_{t}$ is positive definite and hence has a well-defined inverse.
+Noting that $\sum_{t=0}^{T-1}u_{t}=b-a$ we can re-write the iterative formulas as
+$\begin{split}&\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j},\quad t=0,\dots,T-1,\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &-\frac{1}{2}\sum_{t=0}^{T-1}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t+1}^{T-1}\nu_{j}\right)=b-a.\end{split}$
+The last equation can be used to solve for $\mu_{T-1}$ utilizing that the sum of (inverse) positive definite matrices is positive definite and has a well-defined inverse
+$\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),$
+which proves the result. ∎
 Algorithm 1 GEORCE for Riemannian manifolds
@@
 3:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$, $u_{t}^{(0)}\leftarrow\frac{b-a}{T}$ for $t=0.,\dots,T$ and $i\leftarrow 0$
-4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
+4:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
 5:  $G_{t}\leftarrow G\left(x_{t}^{(i)}\right)$ for $t=0,\dots,T-1$
@@
 7:  $\mu_{T-1}\leftarrow\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right)$
-8:     $u_{t}\leftarrow-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right)$ for $t=0,\dots,T-1$
-9:     $x_{t+1}\leftarrow x_{t}+u_{t}$ for $t=0,\dots,T-1$
-10:     Using line search find $\alpha^{*}$ for the following optimization problem with the discrete energy functional $E$
-$\begin{split}\alpha^{*}=\operatorname*{arg\,min}_{\alpha}\quad&E\left(x_{0:T}\right)\quad\text{(exact line search)}\\ \text{s.t.}\quad&x_{t+1}=x_{t}+\alpha u_{t}+(1-\alpha)u_{t}^{(i)},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
-11:     Set $u_{t}^{(i+1)}\leftarrow\alpha^{*}u_{t}+(1-\alpha^{*})u_{t}^{(i)}$ for $t=0,\dots,T-1$
-12:     Set $x_{t+1}^{(i+1)}\leftarrow x_{t}^{(i+1)}+u_{t}^{(i+1)}$ for $t=0,\dots,T-1$
-13:     $i\leftarrow i+1$
-14:end while
-15:return $x_{t}$ for $t=0,\dots,T-1$
-Proposition 2 immediately leads to an iterative updating of $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ to estimate the geodesic, which we denote the GEORCE-algorithm.
-GEORCE is described in pseudo-code in Algorithm 1.
+8:  $\tilde{u}_{t}\leftarrow-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right)$ for $t=0,\dots,T-1$
+9:  Using line search find $\alpha^{*}$ for the following optimization problem with the discrete energy functional $E$
+$\begin{split}\alpha^{*}=\operatorname*{arg\,min}_{\alpha}\quad&E\left(\tilde{x}_{0:T}\right)\quad\text{(exact line search)}\\ \text{s.t.}\quad&\tilde{x}_{t+1}=\tilde{x}_{t}+\alpha\tilde{u}_{t}+(1-\alpha)u_{t}^{(i)},\quad t=0,\dots,T-1,\\ &\tilde{x}_{0}=a.\end{split}$
+10:  Set $u_{t}^{(i+1)}\leftarrow\alpha^{*}\tilde{u}_{t}+(1-\alpha^{*})u_{t}^{(i)}$ for $t=0,\dots,T-1$
+11:  Set $x_{t+1}^{(i+1)}\leftarrow x_{t}^{(i+1)}+u_{t}^{(i+1)}$ for $t=0,\dots,T-1$
+12:  $i\leftarrow i+1$
+13:end while
+14:return $x_{t}$ for $t=0,\dots,T-1$
+[…]
+As we show below, GEORCE exhibits global convergence, which means that the algorithm will converge to a local minimum.
+If there are multiple local minima, then GEORCE will not necessarily find the global minimum.
+To investigate whether GEORCE has found the global minimum, one can apply GEORCE for different initial curves to compute locally-length minimizing curves and choose the curve with the shortest length as the global minimum.
+In Appendix C, we illustrate the effect of different choices of initialization curves and how it affects the solution using GEORCE.
 Line search and stopping criteria
@@
 We instead use soft line search, which merely ensures that the objective function decreases in each iteration [36].
-We implement this using backtracking, where an initial step size is multiplied by a decay rate, $10<\rho<$ 1, until some condition is met, e.g., the Wolfe condition [48], the Armijo condition [49], or the curvature condition [48].
+We implement this using backtracking, where an initial step size is multiplied by a decay rate, $0<\rho<1$, until some condition is met, e.g., the Wolfe condition [49], the Armijo condition [50], or the curvature condition [^2] Note that the term “curvature condition” is rather misleading in a Riemannian context as it refers to conditions on the derivative of the objective function and not the Riemannian curvature.
+[49].
 In practice, we find that the Armijo condition [50] with $\rho=0.5$ is sufficient for most manifolds, although the decay rate can be fine-tuned to increase the performance of GEORCE.
@@
 In all empirical comparisons, identical stopping criteria are applied to all optimization algorithms.
-Convergence results
-In the following, we show that GEORCE is guaranteed to converge to a local minimum, i.e., it exhibits global convergence.
-Proposition 3 (Global convergence).
-Let $E^{(i)}$ be the value of the discretized energy functional for the solution after iteration $i$ (with line search) in GEORCE.
-If the starting point $\left(x^{(0)},u^{(0)}\right)$ is feasible in the sense that the geodesic candidate has the correct start and end points, then the series $\left\{E^{(i)}\right\}_{i>0}$ will converge to a (local) minimum.
-Proof.
-See Appendix A.5 for the proof. ∎
-In theory, GEORCE could jump between two local minima with the same objective value unless the local minimum is strongly unique, as described in Appendix A.1, and seen in conjugate points [23].
-Proposition 3 ensures that GEORCE converges to a local minimum (global convergence) similar to gradient descent.
-In contrast, gradient descent has only linear local convergence, but we show in the following that under sufficient regularity of the discretized energy functional (see Appendix A.1), GEORCE has quadratic local convergence.
-Figure 3: The figure shows the $\ell^{2}$-norm of the gradient of eq.
-5 for each iteration for GEORCE and alternative methods applied to construct geodesics for $100$ iterations for four different manifolds corresponding to Fig.
-1.
-We see that the $\ell^{2}$-norm of the gradient of eq.
-5 for GEORCE converges considerably faster than for alternative algorithms.
-Proposition 4 (Local quadratic convergence).
-Under sufficient regularity of the discretized energy functional (see Appendix A.1) it holds that if the discretized energy functional has a strongly unique (local) minimum point $z^{*}$ (see Appendix A.1) and locally the step-size $\alpha^{*}=1$, i.e., no line search, then GEORCE has locally quadratic convergence, i.e.,
-$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*})\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
-where $z^{(i)}=(x^{(i)},u^{(i)})$ is the solution from GEORCE in iteration $i$, and $B_{\epsilon}\left(z^{*}\right)$ is the ball with radius $\epsilon>0$ and center $z^{*}$.
-Proof.
-See Appendix A.6 for the proof and Appendix A.1 for the assumptions. ∎
-Fig.
-3 demonstrates the convergence by plotting the $\ell^{2}$-norm of the gradient of the discretized energy functional in each iteration for four different manifolds.
-Note that by formulating the geodesic boundary problem as a control problem, we estimate the discretized tangent vectors along the curve $\{u_{t}\}_{t=0}^{T}$.
-Modulo scaling $u_{0}$ will in the limit of a sufficiently fine grid correspond to the logarithmic map, $\mathrm{Log}_{a}b$.
 Algorithm Global convergence Local convergence Gradient Hessian Complexity
@@
 Table 1: Convergence, complexity, and use of higher order derivatives for different optimization algorithms in each iteration, where $T$ is the number of grid points and $d$ is the manifold dimension.
+A checkmark indicates that the algorithm exhibits the property, while a cross indicates that it is not the case.
 Comparison with other optimization methods
@@
 However, the sparse Newton step requires second-order derivatives, unlike GEORCE, which only requires first-order derivatives of the discretized energy.
+Note that in Table 1, we assume that the iterations for the different methods remain well-defined in the chosen chart.
+3.2 Convergence results
+In this section, we show that GEORCE exhibits global convergence, i.e., converges to a local minimum, and that GEORCE exhibits local quadratic convergence under certain regularity of the discretized energy functional.
+To simplify the notation, we let $\langle\cdot,\cdot\rangle$, $\left\lVert\cdot\right\rVert$ and $\nabla$ denote the Euclidean inner product, norm and gradient in $\mathbb{R}^{d}$, respectively.
+Global convergence
+We will prove that GEORCE converges to a (local) minimum.
+To prove this, we will use the fact that the discretized energy functional is smooth and hence locally Lipschitz.
+This implies that the first order Taylor approximation of the discretized energy functional can be written as
+$\Delta E=\langle\nabla E(z_{0}),\Delta z\rangle+\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert,$ (12)
+Thus, the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{O}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally.
+To prove global convergence for the algorithm GEORCE we first show existence and uniqueness by proving two lemmas.
+Secondly, we prove that in each iteration the discretized energy functional is decreasing, unless a (local) minimum has been found.
+For notation set $x^{(i)}=(a,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b)$ and $u^{(i)}=(u_{0}^{(i)},u_{1}^{(i)},\dots,u_{T-2}^{(i)},u_{T-1}^{(i)})$ as the solution for iteration $i$ such that $(x^{(i+1)},u^{(i+1)})$ is the solution using the update scheme in proposition 2.
+Lemma 3.
+Assume that $\left(x^{(i)},u^{(i)}\right)$ is a feasible solution, then
+  1. 1.
+    There exists a unique solution $\left(x^{(i+1)},u^{(i+1)}\right)$ to the system of equations in Eq. 11 based on $\left(x^{(i)},u^{(i)}\right)$.
+  2. 2.
+    All linear combinations $(1-\alpha)\left(x^{(i)},u^{(i)}\right)+\alpha\left(x^{(i+1)},u^{(i+1)}\right)$ for $0<\alpha\leq 1$ are feasible solutions.
+Proof.
+Since the matrices, $\left\{G\left(x_{t}^{(i)}\right)\right\}_{t=0}^{T}$ are positive definite, then the matrices are regular with unique inverse, and so are the sum of the inverse matrices.
+This means that there exists an unique solution to the update scheme in proposition 2 proving 1.
+Initially, assume that $\left(x^{(i+1)},u^{(i+1)}\right)$ is well-defined in the local chart.
+Since the solution $\left(x^{(i+1)},u^{(i+1)}\right)$ is also a feasible solution for the update scheme in proposition 2, then the linear combination of two feasible solutions will also be a feasible solution, i.e.
+$\sum_{t=0}^{T-1}u_{t}^{(j)}=(b-a),\quad j=1,\dots,(i+1).$
+The new solution based on the linear combination gives
+$\begin{split}&(1-\alpha)\sum_{t=0}^{T-1}u_{t}^{(i)}+\alpha\sum_{t=0}^{T-1}u_{t}^{(i+1)}=(1-\alpha)(b-a)+\alpha(b-a)=(b-a)\\ &(1-\alpha)x^{(i)}+\alpha x^{(i+1)}=(1-\alpha)\left(\alpha,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b\right)+\alpha\left(a,x_{1}^{(i+1)},\dots,x_{T-1}^{(i+1)},b\right)\\ &=\left(a,(1-\alpha)x_{1}^{(i)}+\alpha x_{1}^{(i+1)},\dots,(1-\alpha)x_{T-1}^{(i)}+\alpha x_{T-1}^{(i+1)},b\right),\end{split}$
+where for any $t=0,\dots,T-1$
+$\begin{split}(1-\alpha)x_{t}^{(i)}+\alpha x_{t}^{(i+1)}&=(1-\alpha)\left(a+\sum_{j=0}^{t-1}u_{j}^{(i)}\right)+\alpha\left(a+\sum_{j=0}^{t-1}u_{j}^{(i+1)}\right)\\ &=a+\sum_{j=0}^{t-1}\left((1-\alpha)u_{j}^{(i)}+\alpha u_{j}^{(i+1)}\right).\end{split}$
+This shows that the linear combination in the state is feasible as it produces feasible state variables in terms of start and end point.
+Furthermore, each state variable is feasible, if they were determined from the linear combination of the control vectors, which proves that the linear combinations of the state and control variables are also feasible solutions.
+If $\left(x^{(i+1)},u^{(i+1)}\right)$ is not well-defined in the local chart, then the line-search in GEORCE will be able to determine a point belonging to the manifold and that will be on the line between $\left(x^{(i+1)},u^{(i+1)}\right)$ and $\left(x^{(i)},u^{(i)}\right)$ as the local chart is defined on an open set and $x^{(i)}$ is well-defined in the local chart.
+The new solution is then a feasible solution following the argument above.
+∎
+Lemma 4.
+Let $\left\{x_{t}^{(i)},u_{t}^{(i)}\right\}_{t=0}^{T}$ denote the feasible solution after iteration $i$ in GEORCE.
+If $\left\{x_{t}^{(i)},u_{t}^{(i)}\right\}_{t=0}^{T}$ is not a (local) minimum point, then the feasible solution from iteration $\left\{x_{t}^{(i+1)},u_{t}^{(i+1)}\right\}_{t=0}^{T}$ will decrease the objective function in the sense that there exists an $\eta>0$ such that for all $\alpha$ with $0<\alpha\leq\eta\leq 1$, then
+$E\left(x^{(i)}+\alpha\left(x^{(i+1)}-x^{(i)}\right),u^{(i)}+\alpha\left(u^{(i+1)}-u^{(i)}\right)\right)<E\left(x^{(i)},u^{(i)}\right),$
+where $x^{(i)}=\left(a,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b\right)$ and $u^{(i)}=\left(u_{0}^{(i)},u_{1}^{(i)},\dots,u_{T-2}^{(i)},u_{T-1}^{(i)}\right)$.
+Proof.
+$E(x,u)$ is a smooth function. The first order Taylor approximation of the energy function $E(x,u)$ in Eq. 5 in $(x,u)$ is
+$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{x_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta x_{t}\right\rangle+\mathcal{O}(\Delta x_{t})\left\lVert\Delta x_{t}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right),\end{split}$
+where $\Delta u_{t}:=u_{t}^{(i+1)}-u_{t}^{(i)}$ and $\Delta x_{t}:=x_{t}^{(i+1)}-x_{t}^{(i)}$. By the optimally conditions in Eq. 7 we have that
+$\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E(x,u)\vphantom{\big|}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{t-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=-\mu_{t},\quad t=0,\dots,T-1.\end{split}$ (13)
+Inserting the optimality conditions in the Taylor expansion we have that
+$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\langle\mu_{t-1}-\mu_{t},\Delta x_{t}\rangle+\mathcal{O}(\Delta x_{t})\left\lVert\Delta x_{t}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big|}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$
+Since $\Delta x_{t}=\sum_{j=0}^{t-1}\Delta u_{j}$, then
+$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\langle\mu_{t-1}-\mu_{t},\sum_{j=0}^{t-1}\Delta u_{j}\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big|}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$
+Re-arranging the terms we get
+$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-2}\left\langle\mu_{t},\sum_{j=0}^{t}\Delta u_{j}\right\rangle\\ &-\sum_{t=1}^{T-1}\left(\left\langle\mu_{t},\sum_{j=0}^{t-1}\Delta u_{j}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert u_{t}\right\rVert\right)\\ &=\langle\mu_{0},\Delta u_{0}\rangle+\sum_{t=1}^{T-2}\langle\mu_{t},\Delta u_{t}\rangle-\left\langle\mu_{T-1},\sum_{j=0}^{T-1}\Delta u_{j}\right\rangle+\langle\mu_{T-1},\Delta u_{T-1}\rangle\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\end{split}$
+Since $\Delta u_{j}=u_{j}^{(i+1)}-u_{j}^{(i)}$ and $\Delta x_{j}=x_{j}^{(i+1)}-x_{j}^{(i)}$ for $j=0,\dots,T$, and since the solution is feasible we have that $\sum_{j=0}^{T-1}u_{j}^{(i)}=b-a$ for any $i$, which implies that $\sum_{j=0}^{T-1}\Delta u_{j}=0$.
+This reduces the above expression to
+$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left(\langle\mu_{t},\Delta u_{t}\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\\ &=\sum_{t=0}^{T-1}\left\langle\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big|}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\sum_{t=0}^{T-1}\mathcal{O}\left(\Delta u_{t}\right)\left\lVert u_{t}\right\rVert.\end{split}$
+By Eq. 13 we have
+$\begin{split}\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=-{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}\\ &+{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big|}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}\\ &=-2G\left(x_{t}^{(i)}\right)\left(u_{t}^{(i+1)}-u_{t}^{(i)}\right),\end{split}$ (14)
+which implies that
+$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$ (15)
+Since $G\left(x_{t}^{(i)}\right)$ is positive definite, the first summation is negative assuming that at least one non-zero vector $\{\Delta u_{t}\}_{t=0}^{T-1}$, which is the case, since we have assumed that $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ is not a (local) minimum.
+Now scale all vectors, $\left\{\Delta u_{t}\right\}_{t=0}^{T-1}$ with a scalar $0<\alpha<1$.
+Note that the term, $2G\left(x_{t}^{(i)}\right)\Delta u_{t}$, is unaffected by this, since it equals $\mu_{t}+\nabla_{u_{t}}E(x,u)\left(x_{t}^{(i)},u_{t}^{(i)}\right)$, which is independent of $\alpha$.
+The first order Taylor order approximation is
+$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\alpha\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\alpha\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\alpha\Delta u_{j}\right\rVert+\mathcal{O}\left(\alpha\Delta u_{t}\right)\left\lVert\alpha\Delta u_{t}\right\rVert\right).\end{split}$
+In the limit it now follows that there exists an $\eta>0$ such that $\frac{\Delta E(x,u)}{\alpha}<0$ for some $\alpha$ with $0<\alpha\leq\eta\leq 1$, which implies that $\Delta E(x,u)<0$. ∎
+Proposition 5.
+Let $E^{(i)}$ be the value of the discretized energy functional for the solution after iteration $i$ (with line search) in GEORCE.
+If the starting point $\left(x^{(0)},u^{(0)}\right)$ is feasible, then the series $\left\{E^{(i)}\right\}_{i>0}$ will converge to a (local) minimum.
+Proof.
+If the starting point $\left(x^{(0)},u^{(0)}\right)$ in GEORCE is feasible, then the points $\left(x^{(i)},u^{(i)}\right)$ after each iteration (with line-search) in GEORCE will also be feasible solutions by lemma 3.
+[…]
+It is not possible for GEORCE to jump between two local minimum points.
+To see this, assume that there are two local minimum points that have the same discretized energy, and assume that GEORCE is jumping between the two points, i.e., there is convergence to a local minimum in terms of the discretized energy functional according to Proposition 5, but not to a single local minimum point.
+In the GEORCE iteration, the line search will try to improve the discretized energy functional by finding a point closer to one of the two minimum points.
+Either this leads to a new point with improved discretized energy, or GEORCE stops concluding that a local minimum point has been found since no improvements in discretized energy were possible by line-search in accordance with Lemma 4.
+But this contradicts the assumption that GEORCE jumps between two minimum points with the same discretized energy.
+Thus, GEORCE converges to only one local minimum point.
+Local convergence
+We will show that GEORCE under certain regularity of the discretized energy functional exhibits local convergence in the sense that
+$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$ (16)
+for some $c>0$. We will assume the following regularity of the discretized energy functional
+Assumption 6 (Local convergence).
+We assume the following regarding the discretized energy functional.
+  • We assume that the discretized energy functional $E(z)$ is locally strictly convex in the (local) minimum point $z^{*}=\left(x^{*},u^{*}\right)$ in the sense that
+    $\exists\epsilon>0:\,\forall z\in B_{\epsilon}\left(z^{*}\right),z\neq z^{*}:\,\forall\alpha]0,1[:\,E\left(\left(1-\alpha\right)z+\alpha z^{*}\right)<\left(1-\alpha\right)E(z)+\alpha E\left(z^{*}\right),$
+    where $B_{\epsilon}=\left\{z\,|\,\left\lVert z-z^{*}\right\rVert<\epsilon\right\}$.
+  • Assume that the discretized energy functional $E(z)$ is smooth, and consider the first order Taylor approximation of the discretized energy functional
+    $\Delta E=\langle\nabla E(z_{0}),\Delta z\rangle+\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert,$
+    Then, the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{O}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally.
+Figure 3: Illustration of the convergence of GEORCE and baseline algorithms.
+We plot $\log\left\lVert z^{(i)}-z^{*}\right\rVert$ over $\log\left\lVert z^{(i+1)}-z^{*}\right\rVert$ for each iteration $i$ for the different methods.
+We use the convergence point of GEORCE as $z^{*}$.
+If the norms are 0, we set the value to 20-20.
+All initial curves are set to the same random deformation of a straight line similar to Appendix C.
+Intuitively, it is expected that GEORCE exhibits local quadratic convergence.
+GEORCE solves iteratively quadratic minimization problems by fixing the metric matrix function.
+Since GEORCE exhibits global convergence, the metric matrix function in each iteration of GEORCE will converge to the metric matrix function at the local minimum.
+Close to the local minimum, the metric matrix function will be approximately constant, since it is smooth and hence continuous.
+At that point, GEORCE is simply solving a quadratic optimization problem, which is solved in one iteration similar to the Newton method, i.e,.
+has quadratic convergence.
+It is therefore expected that when GEORCE is sufficiently close to the minimum a large “dip” will occur.
+This is especially clear for the Euclidean space, where $G=I$ is independent of $x$.
+Therefore, GEORCE will find the exact solution after one iteration, as it simply solves a quadratic minimization problem.
+We illustrate the convergence of GEORCE and the baseline algorithms for $\mathbb{S}^{2}$ and $\mathbb{R}^{2}$ in Fig.
+3.
+We first prove the following result, before proving quadratic local convergence.
+Lemma 7.
+Assume $f(x)$ is a smooth function. Let $z^{*}$ be a (local) minimum point for $f$, and assume that $f$ is locally strictly convex in $z^{*}$. Then
+$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}(z^{*})\setminus\{z^{*}\}:\quad\langle\nabla f(z),z^{*}-z\rangle<0.$
+Proof.
+Since $f$ is locally strictly convex in $z^{*}$, it follows that
+$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}\left(z^{*}\right),z\neq z^{*}:\quad\forall\alpha\in]0,1[:\quad f\left(\left(1-\alpha\right)z+\alpha z^{*}\right)<\left(1-\alpha\right)f(z)+\alpha f\left(z^{*}\right),$
+which implies that
+$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}\left(z^{*}\right):\quad\forall\alpha\in]0,1[:\quad\frac{f\left(z+\alpha\left(z^{*}-z\right)\right)-f(z)}{\alpha}<f\left(z^{*}\right)-f(z).$
+As $f$ is locally strictly convex in the minimum point $z^{*}$
+$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}\left(z^{*}\right),z\neq z^{*}:\quad f\left(z^{*}\right)-f(z)<0,$
+which follows from the fact that if there exists a $z^{\prime}\in B_{\epsilon}\left(z^{*}\right)$ with $f\left(z^{\prime}\right)=f\left(z^{*}\right)$, then since $f$ is locally strictly convex in $z^{*}$, then
+$\forall\alpha\in]0,1[:\quad(1-\alpha)f\left(z^{\prime}\right)+\alpha f\left(z^{*}\right)=f\left(z^{*}\right)>f\left(\left(1-\alpha\right)z+\alpha z^{*}\right).$
+However, this contradicts that $z^{\prime}$ and $z^{*}$ are local minimum points. Since $f$ is a smooth function, then the first order Taylor approximation gives
+$\frac{f\left(z+\alpha\left(z^{*}-z\right)\right)-f(z)}{\alpha}=\left\langle\nabla f(z),z^{*}-z\right\rangle+\mathcal{O}\left(\alpha\left(z^{*}-z\right)\right)\left\lVert z^{*}-z\right\rVert<f\left(z^{*}\right)-f(z)<0.$
+In the limit, then $\lim_{\alpha\searrow 0}\mathcal{O}\left(\alpha\left(z^{*}-z\right)\right)=0$, which implies that $\left\langle\nabla f(z),z^{*}-z\right\rangle<0$. ∎
+Proposition 8.
+If the discretized energy functional is locally strictly convex in the (local) minimum point $z^{*}$ and locally $\alpha^{*}=1$, i.e.
+no line search, then GEORCE has locally quadratic convergence, i.e.
+$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
+where $z^{(i)}=(x^{(i)},u^{(i)})$ is the solution from GEORCE at iteration $i$.
+Proof.
+Assume that $z^{*}$ is a (local) minimum point.
+Since GEORCE has global convergence to a (local) minimum point by theorem 5, we will assume that the algorithm has convergence point $E(z^{*})$.
+Consider iteration $i$ with $z^{(i)}=(x^{(i)},u^{(i)})$ from GEORCE, and assume that all following iterations in GEORCE are in the ball $B_{\epsilon_{1}}(z^{*})$, and fulfill the locally strictly convex assumption in 6.
+Note that as the series $\left\{z^{(i)}\right\}$ converges to $z^{*}$, then $\left\{z^{(i)}\right\}_{i\geq k}\subset B_{\epsilon_{1}}(z^{*})$ for some $k>0$.
+It follows as $E(z)$ is locally Lipschitz continuous that
+$E\left(z^{(i+1)}-z^{(*)}\right)-E\left(z^{*}\right)=\mathcal{O}\left(\left\lVert z^{(i+1)}-z^{*}\right\rVert\right),$
+since $z^{(i+1)}\in B_{\epsilon}\left(z^{*}\right)$. The left hand term above can be rearranged into
+$E\left(z^{(i+1)}\right)-E\left(z^{*}\right)=\left(E\left(z^{(i+1)}\right)-E\left(z^{(i)}\right)\right)+\left(E\left(z^{(i)}\right)-E\left(z^{*}\right)\right).$
+By Eq. 15 in the proof of lemma 4 and utilizing $\alpha^{*}=1$, it follows that
+$\begin{split}&E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\\ &=\left(E\left(z^{(i+1)}\right)-E\left(z^{(i)}\right)\right)+\left(E\left(z^{(i)}\right)-E\left(z^{*}\right)\right)\\ &=\sum_{t=0}^{T-1}\left(\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\\ &-\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\left(u_{t}^{*}-u_{t}^{(i)}\right)\right\rangle\\ &-\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert-\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\\ &=\sum_{t=0}^{T-1}\left(\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}-\left(u_{t}^{*}-u_{t}^{(i)}\right)\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert+\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\right)\\ &=\sum_{t=0}^{T-1}\left(\left\langle 2G\left(x_{t}^{(i)}\right)\Delta u_{t},u_{t}^{*}-u_{t}^{(i+1)}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert+\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\right).\end{split}$
+Observe now that the term $2G\left(x_{t}^{(i)}\right)\Delta u_{t}$ is the gradient for $z_{t}^{(i)}$.
+Define $\epsilon_{2}$ such that the $\mathcal{O}$-terms are sufficiently close to zero for all $z\in B_{\epsilon_{2}}(z^{*})$.
+Thus, for $\epsilon=\min\left\{\epsilon_{1},\epsilon_{2}\right\}$ we have from lemma 7
+$\sum_{t=0}^{T-1}\left\langle 2G\left(x_{t}^{(i)}\right)\Delta u_{t},u_{t}^{*}-u_{t}^{(i+1)}\right\rangle<0,$
+since $E$ is assumed locally strictly convex in the local minimum point $z^{*}$ by assumptions 6.
+Combining the inequalities and utilizing that $E(z)$ is locally Lipschitz continuous by assumptions 6, then
+$\mathcal{O}\left(\left\lVert z^{(i+1)}-z^{*}\right\rVert\right)=E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\leq\mathcal{O}\left(\left\lVert z^{(i)}-z^{*}\right\rVert^{2}\right),$
+which implies that there exists an $\epsilon>0$ such that
+$\forall z^{(i)}\in B_{\epsilon}\left(z^{*}\right):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
+for $c>0$. ∎
+The assumption in the above proof that the line search parameter, $\alpha^{*}=1$ locally at the minimum point is based on that locally the approximation of the discretized energy functional becomes close to the true function.
+Figure 4: The figure shows the $\ell^{2}$-norm of the gradient of Eq.
+5 for each iteration for GEORCE and alternative methods applied to construct geodesics for $100$ iterations for four different manifolds corresponding to Fig.
+1.
+We see that the $\ell^{2}$-norm of the gradient of Eq.
+5 for GEORCE converges considerably faster than for alternative algorithms.
+Fig.
+4 shows the convergence by plotting the $\ell^{2}$-norm of the gradient of the discretized energy functional in each iteration for four different manifolds.
+Note that by formulating the geodesic boundary problem as a control problem, we estimate the discretized tangent vectors along the curve $\{u_{t}\}_{t=0}^{T}$.
+Modulo scaling $u_{0}$ will in the limit of a sufficiently fine grid correspond to the logarithmic map, $\mathrm{Log}_{a}b$.
 4 Extension to Finslerian geometry
 A Finsler geometry can be seen as a generalization of Riemannian geometry, where the tangent spaces are not equipped with an quadratic form metric but with a Finsler metric instead.
-A Finsler metric, $F:\mathcal{M}\times T\mathcal{M}\rightarrow\mathbb{R}_{+}$, where $\mathcal{M}$ is a differentiable manifold, is a Minkowski norm in the sense that $F$ is $C^{\infty}$ on $T\mathcal{M}\setminus\{0\}$, $F(x,cv)=cF(x,v)$ for all $v\in T\mathcal{M}$, $x\in\mathcal{M}$ and $c>0$ and for any $v\in T\mathcal{M}$, then $\frac{\partial^{2}F}{\partial v^{i}\partial v^{j}}(x,v)$ is positive definite [50].
-Geodesics on a Finsler manifold can be found equivalently to the Riemannian case by minimizing the energy functional, $\int_{0}^{1}F^{2}\left(\gamma(t),\dot{\gamma}(t)\right)\,\mathrm{d}t$, on a suitable set of curves connecting the two points thereby approximating a geodesic $\gamma$, [50].
-GEORCE is similarly applicable to Finsler geometry.
-First, we note that the energy functional can be written as [50]
+A Finsler metric, $F:\mathcal{M}\times T\mathcal{M}\rightarrow\mathbb{R}_{+}$, where $\mathcal{M}$ is a differentiable manifold, is a Minkowski norm in the sense that $F$ is $C^{\infty}$ on $T\mathcal{M}\setminus\{0\}$, $F(x,cv)=cF(x,v)$ for all $v\in T\mathcal{M}$, $x\in\mathcal{M}$ and $c>0$ and for any $v\in T\mathcal{M}$, then $\frac{\partial^{2}F}{\partial v^{i}\partial v^{j}}(x,v)$ is positive definite [29, Chapter 2].
+Figure 5: The red circles shows the indicatrix field, $\{v\in T_{x}\mathcal{M}\,|\,F(x,v)=1\}$, in the chart of the sphere used in the Riemannian case in Fig.
+1 and its corresponding center is marked with black dot.
+The left figure is the Riemannian case, while the right figure is the Finslerian case corresponding to a force field making the indicatrix-field become non-centered ellipsis for the geodesics (blue and green), where the arrow indicates the direction of the curve.
+Geodesics for a Finsler manifold can be found equivalently to the Riemannian case by minimizing the energy functional, $\int_{0}^{1}F^{2}\left(\gamma(t),\dot{\gamma}(t)\right)\,\mathrm{d}t$, on a suitable set of curves connecting the two points thereby approximating a geodesic $\gamma$, [29, Chapter 3].
+In this way, the indicatrix field for Finslerian manifolds consists of non-centered ellipses unlike the Riemannian case, meaning that the distance is not symmetric for Finslerian manifolds.
+We illustrate this in Fig.
+5 for the local chart for the sphere used in Fig.
+1, where we shift the indicatrix fields by a wind-field on the form $(0,0.75)$.
+GEORCE is similarly applicable to Finsler geometry. First, we note that the energy functional can be written as [29, Page 20]
 $\frac{1}{2}\int_{0}^{1}F^{2}\left(\gamma(t),\dot{\gamma}(t)\right)\,\mathrm{d}t=\int_{0}^{1}\dot{\gamma}(t)^{\top}G\left(\gamma(t),\dot{\gamma}(t)\right)\dot{\gamma}(t)\,\mathrm{d}t,$ (17)
 where $G(x,v)=\frac{1}{2}\frac{\partial^{2}F^{2}}{\partial v^{i}\partial v^{j}}\left(x,v\right)$ denotes the fundamental tensor in local coordinates [^3] In literature, the fundamental tensor is often denoted with a lower case $g$.
-To have the notation consistent with the Riemannian case, we opt to use an upper case $G$ instead., which is positive definite [50].
+To have the notation consistent with the Riemannian case, we opt to use an upper case $G$ instead., which is positive definite [29, Page 19].
 Unlike the Riemannian case, the fundamental tensor, $G$, is now a function of both the position and velocity, which implies the update formulas in Proposition 2.
@@
 Using the modified update scheme in Eq.
-14, we estimate geodesics for Finsler manifolds similar to Algorithm 1.
+20, we estimate geodesics for Finsler manifolds similar to Algorithm 1.
 For completeness, we show the algorithm in pseudo-code for the Finsler case in Appendix B.1.
@@
 5 Experiments
-Figure 4: The first two figures display the length and runtime of geodesics for $\mathbb{S}^{n}$ estimated using different algorithms from Table 2 and other methods in Appendix E.3.
-The latter row shows the length and runtime, respectively for the $n$-Ellipsoid with half axes of $n$ equally spaced points between $0.5$ and $1.0$.
-The optimization solvers are described in Appendix E.1.
+Figure 6: The first two figures display the length and runtime of geodesics for $\mathbb{S}^{n}$ estimated using different algorithms from Table 2.
+The latter row shows the length and runtime, respectively, for the $n$-Ellipsoid with half axes of $n$ equally spaced points between $0.5$ and $1.0$.
+The optimization solvers are described in Appendix F.1.
 BFGS (T=100) ADAM (T=100) GEORCE (T=100)
@@
 $\mathbb{S}^{50}$ $1.0464$ $7350.6001\pm 244.3737$ $1.1463$ $61.8926\pm 0.3034$ $\boldsymbol{0.9749}$ $\boldsymbol{4.1016}\pm\boldsymbol{0.0201}$
-$\mathbb{S}^{100}$- -- $0.8466$ $277.4217\pm 0.2280$ $\boldsymbol{0.4187}$ $\boldsymbol{10.9208}\pm\boldsymbol{0.0071}$
-$\mathbb{S}^{250}$- -- $0.6244$ $1424.9031\pm 22.1592$ $\boldsymbol{0.3192}$ $\boldsymbol{78.2937}\pm\boldsymbol{0.0991}$
-$\mathbb{S}^{500}$- -- $0.5717$ $2242.4697\pm 17.9769$ $\boldsymbol{0.2198}$ $\boldsymbol{224.2761}\pm\boldsymbol{0.4894}$
-$\mathbb{S}^{1000}$- -- $0.5279$ $4329.6621\pm 38.0163$ $\boldsymbol{0.1573}$ $\boldsymbol{893.4682}\pm\boldsymbol{7.3130}$
+$\mathbb{S}^{100}$ - - $0.8466$ $277.4217\pm 0.2280$ $\boldsymbol{0.4187}$ $\boldsymbol{10.9208}\pm\boldsymbol{0.0071}$
+$\mathbb{S}^{250}$ - - $0.6244$ $1424.9031\pm 22.1592$ $\boldsymbol{0.3192}$ $\boldsymbol{78.2937}\pm\boldsymbol{0.0991}$
+$\mathbb{S}^{500}$ - - $0.5717$ $2242.4697\pm 17.9769$ $\boldsymbol{0.2198}$ $\boldsymbol{224.2761}\pm\boldsymbol{0.4894}$
+$\mathbb{S}^{1000}$ - - $0.5279$ $4329.6621\pm 38.0163$ $\boldsymbol{0.1573}$ $\boldsymbol{893.4682}\pm\boldsymbol{7.3130}$
 $\mathrm{E}\left(2\right)$ $1.9595$ $56.9650\pm 0.4106$ $1.9596$ $0.2328\pm 0.0001$ $\boldsymbol{1.9595}$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0001}$
@@
 $\mathrm{E}\left(50\right)$ $0.8338$ $10685.6133\pm 321.6447$ $0.9244$ $73.5682\pm 0.3306$ $\boldsymbol{0.4642}$ $\boldsymbol{2.4208}\pm\boldsymbol{0.0222}$
-$\mathrm{E}\left(100\right)$- -- $0.6872$ $289.4196\pm 2.5224$ $\boldsymbol{0.3249}$ $\boldsymbol{7.7128}\pm\boldsymbol{0.0714}$
-$\mathrm{E}\left(250\right)$- -- $0.4978$ $1620.6428\pm 16.9546$ $\boldsymbol{0.2178}$ $\boldsymbol{56.8527}\pm\boldsymbol{0.1526}$
-$\mathrm{E}\left(500\right)$- -- $0.4559$ $2135.6899\pm 30.5699$ $\boldsymbol{0.1467}$ $\boldsymbol{220.1656}\pm\boldsymbol{0.9902}$
-$\mathrm{E}\left(1000\right)$- -- $0.4232$ $4398.9453\pm 25.8738$ $\boldsymbol{0.1131}$ $\boldsymbol{641.7199}\pm\boldsymbol{20.9901}$
-$\mathbb{T}^{2}$ $9.9851$ $92.7400\pm 3.1267$ $9.9855$ $0.1060\pm 0.0002$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0182}\pm\boldsymbol{0.0003}$
+$\mathrm{E}\left(100\right)$ - - $0.6872$ $289.4196\pm 2.5224$ $\boldsymbol{0.3249}$ $\boldsymbol{7.7128}\pm\boldsymbol{0.0714}$
+$\mathrm{E}\left(250\right)$ - - $0.4978$ $1620.6428\pm 16.9546$ $\boldsymbol{0.2178}$ $\boldsymbol{56.8527}\pm\boldsymbol{0.1526}$
+$\mathrm{E}\left(500\right)$ - - $0.4559$ $2135.6899\pm 30.5699$ $\boldsymbol{0.1467}$ $\boldsymbol{220.1656}\pm\boldsymbol{0.9902}$
+$\mathrm{E}\left(1000\right)$ - - $0.4232$ $4398.9453\pm 25.8738$ $\boldsymbol{0.1131}$ $\boldsymbol{641.7199}\pm\boldsymbol{20.9901}$
+$\mathbb{T}^{2}$ $9.9851$ $2.7400\pm 3.1267 9.9$8559.9855 $0.1060\pm 0.0002$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0182}\pm\boldsymbol{0.0003}$
 $\mathbb{H}^{2}$ $1.2499$ $75.0079\pm 0.1710$ $1.2499$ $0.0988\pm 0.0001$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0025}\pm\boldsymbol{0.0001}$
 $\mathcal{P}\left(2\right)$ $\boldsymbol{1.0669}$ $88.2739\pm 0.4152$ $1.0879$ $0.2611\pm 0.0004$ $1.0669$ $\boldsymbol{0.0141}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2639}$ $168.5082\pm 0.3039$ $2.2896$ $2.3780\pm 0.0046$ $2.2639$ $\boldsymbol{0.1442}\pm\boldsymbol{0.0002}$
+$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2639}$ $168.5082\pm 0.3039$ $2.2896$ $3780\pm 0.0046 2.2$6392.2639 $\boldsymbol{0.1442}\pm\boldsymbol{0.0002}$
 Gaussian Distribution $\boldsymbol{2.5952}$ $47.6288\pm 0.1359$ $2.5956$ $0.0167\pm 0.0001$ $2.5952$ $\boldsymbol{0.0015}\pm\boldsymbol{0.0000}$
@@
 $\mathrm{E}(n)$ denotes an Ellipsoid of dimension $n$ with half axes of $n$ equally spaced points between $0.5$ and $1.0$, while $\mathcal{P}(n)$ denotes the space of $n\times n$ symmetric positive definite matrices.
-When the computational time was longer than $24$ hours, the value is set to --.
+When the computational time was longer than $24$ hours, the value is set to -.
+ADAM (T=100) Sparse Newton (T=100) GEORCE (T=100)
+Riemannian Manifold Length Runtime Length Runtime Length Runtime
+$\mathbb{S}^{2}$ $2.2755$ $0.2729\pm 0.0016$ $2.2742$ $0.1417\pm 0.0041$ $\boldsymbol{2.2741}$ $\boldsymbol{0.0163}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{3}$ $2.4376$ $0.2890\pm 0.0005$ $\boldsymbol{2.4342}$ $0.2188\pm 0.0039$ $2.4342$ $\boldsymbol{0.0735}\pm\boldsymbol{0.0018}$
+$\mathbb{S}^{5}$ $2.4226$ $0.3153\pm 0.0027$ $\boldsymbol{2.3755}$ $0.8657\pm 0.0015$ $2.3755$ $\boldsymbol{0.0494}\pm\boldsymbol{0.0014}$
+$\mathbb{S}^{10}$ $2.1031$ $0.3052\pm 0.0079$ $1.9653$ $117.4589\pm 0.0140$ $\boldsymbol{1.9523}$ $\boldsymbol{0.0351}\pm\boldsymbol{0.0001}$
+$\mathbb{S}^{20}$ $1.6824$ $0.3401\pm 0.0117$ $1.7069$ $146.3024\pm 0.0191$ $\boldsymbol{1.4808}$ $\boldsymbol{0.0483}\pm\boldsymbol{0.0020}$
+$\mathbb{S}^{50}$ $1.1456$ $0.5339\pm 0.0014$ $1.0359$ $310.2160\pm 0.0200$ $\boldsymbol{0.9749}$ $\boldsymbol{0.0587}\pm\boldsymbol{0.0023}$
+$\mathbb{S}^{100}$ $0.8475$ $1.2935\pm 0.0014$ $0.7522$ $6.6464\pm 0.0016$ $\boldsymbol{0.4188}$ $\boldsymbol{0.0717}\pm\boldsymbol{0.0022}$
+$\mathbb{S}^{250}$ $0.6251$ $3.8386\pm 0.0018$ $0.4958$ $98.6502\pm 0.0147$ $\boldsymbol{0.3192}$ $\boldsymbol{0.3296}\pm\boldsymbol{0.0005}$
+$\mathbb{S}^{500}$ $0.5714$ $8.2031\pm 0.0030$ $0.3787$ $899.4549\pm 0.0188$ $\boldsymbol{0.2198}$ $\boldsymbol{1.2925}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{1000}$ $0.5270$ $19.2234\pm 0.0712$ - - $\boldsymbol{0.1573}$ $\boldsymbol{6.8579}\pm\boldsymbol{0.0023}$
+$1.9595$ $0.2735\pm 0.0074$ $\boldsymbol{1.9595}$ $0.2177\pm 0.0075$ $1.9595$ $\boldsymbol{0.0138}\pm\boldsymbol{0.0001}$
+$\mathrm{E}\left(3\right)$ $2.0523$ $0.2839\pm 0.0056$ $\boldsymbol{2.0498}$ $0.3705\pm 0.0063$ $2.0501$ $\boldsymbol{0.0167}\pm\boldsymbol{0.0001}$
+$\mathrm{E}\left(5\right)$ $1.9594$ $0.2948\pm 0.0116$ - - $\boldsymbol{1.8410}$ $\boldsymbol{0.0827}\pm\boldsymbol{0.0036}$
+$\mathrm{E}\left(10\right)$ $1.6421$ $0.3097\pm 0.0133$ $\boldsymbol{1.4247}$ $116.6572\pm 0.0186$ $1.5139$ $\boldsymbol{0.0547}\pm\boldsymbol{0.0019}$
+$\mathrm{E}\left(20\right)$ $1.3052$ $0.3875\pm 0.0058$ $1.4583$ $148.6781\pm 0.0307$ $\boldsymbol{0.6514}$ $\boldsymbol{0.1449}\pm\boldsymbol{0.0054}$
+$\mathrm{E}\left(50\right)$ $0.9245$ $0.5704\pm 0.0013$ $0.8849$ $316.9808\pm 0.0179$ $\boldsymbol{0.4642}$ $\boldsymbol{0.0370}\pm\boldsymbol{0.0014}$
+$\mathrm{E}\left(100\right)$ $0.6873$ $1.3063\pm 0.0020$ $0.5922$ $5.2868\pm 0.0007$ $\boldsymbol{0.3249}$ $\boldsymbol{0.0524}\pm\boldsymbol{0.0002}$
+$\mathrm{E}\left(250\right)$ $0.4974$ $4.2084\pm 0.0029$ $0.3917$ $84.6331\pm 0.0439$ $\boldsymbol{0.2180}$ $\boldsymbol{0.2325}\pm\boldsymbol{0.0001}$
+$\mathrm{E}\left(500\right)$ $0.4559$ $7.4772\pm 0.0024$ $0.3039$ $827.7163\pm 0.1003$ $\boldsymbol{0.1468}$ $\boldsymbol{1.2897}\pm\boldsymbol{0.0004}$
+$\mathrm{E}\left(1000\right)$ $0.4238$ $19.8033\pm 0.0970$ - - $\boldsymbol{0.1129}$ $\boldsymbol{5.1862}\pm\boldsymbol{0.0071}$
+$\mathbb{T}^{2}$ $9.9853$ $0.2890\pm 0.0016$ $9.9851$ $0.3353\pm 0.0126$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0549}\pm\boldsymbol{0.0010}$
+$\mathbb{H}^{2}$ $1.2500$ $0.3201\pm 0.0004$ $1.2498$ $0.2125\pm 0.0128$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0089}\pm\boldsymbol{0.0003}$
+$\mathcal{P}\left(2\right)$ $1.0786$ $0.3801\pm 0.0047$ $\boldsymbol{1.0669}$ $0.2419\pm 0.0069$ $1.0669$ $\boldsymbol{0.0227}\pm\boldsymbol{0.0010}$
+$\mathcal{P}\left(3\right)$ $2.3006$ $0.4115\pm 0.0009$ $2.3810$ $106.2814\pm 0.0135$ $\boldsymbol{2.2639}$ $\boldsymbol{0.0343}\pm\boldsymbol{0.0004}$
+Gaussian Distribution $\boldsymbol{2.5951}$ $0.1809\pm 0.0010$ $2.5952$ $0.1108\pm 0.0055$ $2.5952$ $\boldsymbol{0.0123}\pm\boldsymbol{0.0002}$
+Fréchet Distribution $\boldsymbol{1.1538}$ $0.1778\pm 0.0029$ $1.1539$ $0.0813\pm 0.0044$ $1.1539$ $\boldsymbol{0.0055}\pm\boldsymbol{0.0003}$
+Cauchy Distribution $1.6452$ $0.1818\pm 0.0007$ $1.6453$ $0.1136\pm 0.0067$ $\boldsymbol{1.6452}$ $\boldsymbol{0.0137}\pm\boldsymbol{0.0001}$
+Pareto Distribution $0.8314$ $0.1779\pm 0.0026$ $\boldsymbol{0.8297}$ $0.0758\pm 0.0036$ $0.8297$ $\boldsymbol{0.0043}\pm\boldsymbol{0.0002}$
+Table 3: The table shows the length of the estimated geodesics for different methods on a GPU.
+The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
+When the computational time was longer than $24$ hours, the value is set to -.
+ADAM (T=50) Sparse Newton (T=50) GEORCE (T=50)
+Riemannian Manifold Length Runtime Length Runtime Length Runtime
+$\mathbb{S}^{2}$ $2.2479$ $0.2683\pm 0.0082$ $2.2478$ $0.1094\pm 0.0047$ $\boldsymbol{2.2477}$ $\boldsymbol{0.0187}\pm\boldsymbol{0.0002}$
+$\mathbb{S}^{3}$ $2.4069$ $0.2593\pm 0.0110$ $\boldsymbol{2.4064}$ $0.1370\pm 0.0082$ $2.4064$ $\boldsymbol{0.0866}\pm\boldsymbol{0.0031}$
+$\mathbb{S}^{5}$ $2.3494$ $0.2907\pm 0.0103$ $2.3480$ $0.3153\pm 0.0063$ $\boldsymbol{2.3480}$ $\boldsymbol{0.0622}\pm\boldsymbol{0.0019}$
+$\mathbb{S}^{10}$ $1.9470$ $0.3026\pm 0.0014$ $1.9274$ $0.8966\pm 0.0034$ $\boldsymbol{1.9272}$ $\boldsymbol{0.0540}\pm\boldsymbol{0.0016}$
+$\mathbb{S}^{20}$ $1.5036$ $0.3349\pm 0.0069$ $1.4610$ $28.7634\pm 0.0016$ $\boldsymbol{1.4566}$ $\boldsymbol{0.0670}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{50}$ $1.0320$ $0.4713\pm 0.0025$ $1.0060$ $39.8257\pm 0.0010$ $\boldsymbol{0.5279}$ $\boldsymbol{0.0438}\pm\boldsymbol{0.0002}$
+$\mathbb{S}^{100}$ $0.7642$ $0.8173\pm 0.0009$ $0.6922$ $4.2461\pm 0.0007$ $\boldsymbol{0.3932}$ $\boldsymbol{0.0704}\pm\boldsymbol{0.0027}$
+$\mathbb{S}^{250}$ $0.5136$ $3.2362\pm 0.0009$ $0.4589$ $60.6258\pm 0.0179$ $\boldsymbol{0.2748}$ $\boldsymbol{0.1525}\pm\boldsymbol{0.0002}$
+$\mathbb{S}^{500}$ $0.4253$ $7.7012\pm 0.0038$ $0.3278$ $706.9603\pm 0.0293$ $\boldsymbol{0.1858}$ $\boldsymbol{0.7340}\pm\boldsymbol{0.0006}$
+$\mathbb{S}^{1000}$ $0.3790$ $20.5850\pm 0.0098$ - - $\boldsymbol{0.1626}$ $\boldsymbol{3.1229}\pm\boldsymbol{0.0007}$
+$\mathrm{E}\left(2\right)$ $1.9331$ $0.2723\pm 0.0013$ $\boldsymbol{1.9330}$ $0.1365\pm 0.0083$ $1.9331$ $\boldsymbol{0.0155}\pm\boldsymbol{0.0003}$
+$\mathrm{E}\left(3\right)$ $2.0243$ $0.2852\pm 0.0017$ $\boldsymbol{2.0233}$ $0.1828\pm 0.0074$ $2.0235$ $\boldsymbol{0.0305}\pm\boldsymbol{0.0012}$
+$\mathrm{E}\left(5\right)$ $1.8262$ $0.2996\pm 0.0079$ $\boldsymbol{1.8047}$ $0.6859\pm 0.0047$ $1.8054$ $\boldsymbol{0.1172}\pm\boldsymbol{0.0050}$
+$\mathrm{E}\left(10\right)$ $1.5193$ $0.3094\pm 0.0086$ $\boldsymbol{0.9472}$ $57.6302\pm 0.0107$ $1.4788$ $\boldsymbol{0.0855}\pm\boldsymbol{0.0036}$
+$\mathrm{E}\left(20\right)$ $1.2028$ $0.3379\pm 0.0049$ $\boldsymbol{0.7603}$ $48.7781\pm 0.0029$ $1.1366$ $\boldsymbol{0.0727}\pm\boldsymbol{0.0033}$
+$\mathrm{E}\left(50\right)$ $0.8382$ $0.4596\pm 0.0092$ $0.6157$ $153.9548\pm 0.0165$ $\boldsymbol{0.4417}$ $\boldsymbol{0.0284}\pm\boldsymbol{0.0001}$
+$\mathrm{E}\left(100\right)$ $0.6175$ $0.8219\pm 0.0036$ $0.5822$ $3.9507\pm 0.0014$ $\boldsymbol{0.3160}$ $\boldsymbol{0.0373}\pm\boldsymbol{0.0001}$
+$\mathrm{E}\left(250\right)$ $0.4216$ $2.4940\pm 0.0028$ $0.3882$ $48.5547\pm 0.0120$ $\boldsymbol{0.2039}$ $\boldsymbol{0.1519}\pm\boldsymbol{0.0003}$
+$\mathrm{E}\left(500\right)$ $0.3454$ $5.5967\pm 0.0145$ - - $\boldsymbol{0.1434}$ $\boldsymbol{0.7326}\pm\boldsymbol{0.0005}$
+$\mathrm{E}\left(1000\right)$ $0.3056$ $14.1006\pm 0.0361$ - - $\boldsymbol{0.0996}$ $\boldsymbol{4.1512}\pm\boldsymbol{0.0043}$
+$\mathbb{T}^{2}$ $9.8952$ $0.2821\pm 0.0029$ $\boldsymbol{9.8943}$ $38.0816\pm 0.0018$ $9.8955$ $\boldsymbol{0.0531}\pm\boldsymbol{0.0013}$
+$\mathbb{H}^{2}$ $1.2435$ $0.3072\pm 0.0003$ $\boldsymbol{1.2434}$ $0.1172\pm 0.0048$ $1.2434$ $\boldsymbol{0.0082}\pm\boldsymbol{0.0001}$
+$\mathcal{P}\left(2\right)$ $1.0704$ $0.3571\pm 0.0072$ $\boldsymbol{1.0567}$ $0.1231\pm 0.0051$ $1.0567$ $\boldsymbol{0.0243}\pm\boldsymbol{0.0002}$
+$\mathcal{P}\left(3\right)$ $2.2400$ $0.3924\pm 0.0057$ $2.3328$ $57.1278\pm 0.0125$ $\boldsymbol{2.2304}$ $\boldsymbol{0.0198}\pm\boldsymbol{0.0001}$
+Gaussian Distribution $2.5779$ $0.1785\pm 0.0003$ $\boldsymbol{2.5778}$ $0.0617\pm 0.0027$ $2.5778$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
+Fréchet Distribution $1.1471$ $0.1696\pm 0.0004$ $\boldsymbol{1.1457}$ $0.0442\pm 0.0016$ $1.1457$ $\boldsymbol{0.0063}\pm\boldsymbol{0.0002}$
+Cauchy Distribution $1.6341$ $0.0230\pm 0.0002$ $1.6340$ $0.0810\pm 0.0037$ $\boldsymbol{1.6340}$ $\boldsymbol{0.0135}\pm\boldsymbol{0.0003}$
+Pareto Distribution $0.8235$ $0.0357\pm 0.0003$ $0.8235$ $0.0439\pm 0.0014$ $\boldsymbol{0.8234}$ $\boldsymbol{0.0052}\pm\boldsymbol{0.0003}$
+Table 4: The table shows the length of the estimated geodesics for different methods on a GPU.
+The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
+When the computational time was longer than $24$ hours, the value is set to -.
 Riemannian manifolds
@@
 The BFGS estimate, ADAM estimate and the initial curve overlap completely.
-To compare GEORCE with other optimization algorithms, we show the runtime and length of geodesics computed for different manifolds with varying dimensions in Table 2.
+To compare GEORCE with other optimization algorithms, we show the runtime and length of the geodesics computed for different manifolds with varying dimensions in Table 2 on a CPU.
+In Table 3 and Table 4, we show the corresponding results on a GPU for compatible methods for $T=100$ and $T=50$, respectively.
 Here we compare GEORCE with BFGS [32, 33, 34, 35] and ADAM [40].
@@
 5 is less than $10^{-4}$, or if the number of iterations exceeds $1,000$.
-A full description of the manifolds and experimental details can be found in Appendix C and Appendix E, respectively.
+A full description of the manifolds and experimental details can be found in Appendix D and Appendix F, respectively.
 We show the estimated geodesics with $T=100$ grid points and summarize the results for the corresponding runtime and length of the computed geodesics for the $n$-sphere and $n$-ellipsoid in Fig.
-4.
-In Appendix E.3, we show similar run-time experiments for varying grid points as well as other optimization algorithms.
-We also show the corresponding results on a GPU with other CUDA compatible algorithms.
-In general, we see that GEORCE is considerably faster with lower length compared to alternative methods.
-To test GEORCE on manifolds with multiple length minimizing geodesics, we consider the geodesic between $(-5,-5)$ and $(5,5)$ on an “egg-tray” parameterized by $(x,y,2\cos x\cos y)$ with $T=1,000$ grid points.
+6.
+In general, we see that GEORCE is considerably faster with a shorter length compared to alternative methods.
+To test GEORCE on manifolds with multiple length minimizing geodesics, we consider the geodesic between (5,5)(-5,-5) and $(5,5)$ on an “egg-tray” parameterized by $(x,y,2\cos x\cos y)$ with $T=1,000$ grid points.
 We show the estimates in Fig.
-5, where we see that GEORCE obtains the a signficiantly smaller length, where there is also a clear symmetric to solution to GEORCE along the other side of the “egg-tops”.
+7, where we see that GEORCE obtains a significantly smaller length, where there is also a clear symmetry to solution to GEORCE along the other side of the “egg-tops”.
 Figure 8: The connecting geodesic between four different distributions equipped with the Fisher-Rao metric using GEORCE.
 From left to right the figures show the connecting geodesic (blue) between two fixed distributions (black) for the Gaussian distribution, Fréchet distribution, Cauchy distribution and the Pareto distribution.
-Figure 7: The geodesics computed under the Hessian of the log-partition function of a Gaussian distribution. The background color illustrates the log-partition function.
 To illustrate the application of GEORCE to more abstract manifolds, we compute geodesics on the statistical manifolds, i.e., between distributions [17, 18].
@@
 The computed geodesics using alternative methods are summarized in Table 2, where we see that GEORCE obtains significantly shorter length much faster, also in high-dimension, compared to alternative methods.
+Figure 9: The geodesics computed under the Hessian of the log-partition function of a Gaussian distribution. The background color illustrates the log-partition function.
 We further illustrate the application of GEORCE to information geometry by considering potential functions encountered in Bregman geometry [18].
@@
 Fig.
-7 shows geodesics computed with both GEORCE and alternative methods using at most $100$ iterations.
+9 shows geodesics computed with both GEORCE and alternative methods that use at most $100$ iterations.
 We see that GEORCE uses less iterations and obtains the smallest length compared to the alternative methods.
@@
 $\mathbb{S}^{50}$ $0.8216$ $5992.7261\pm 56.5973$ $0.8128$ $146.4556\pm 0.9341$ $\boldsymbol{0.7285}$ $\boldsymbol{8.4758}\pm\boldsymbol{0.0234}$
-$\mathbb{S}^{100}$- -- $0.6187$ $498.9705\pm 0.4032$ $\boldsymbol{0.2570}$ $\boldsymbol{15.3786}\pm\boldsymbol{0.0092}$
-$\mathbb{S}^{250}$- -- $0.5195$ $1397.6732\pm 0.7793$ $\boldsymbol{0.1622}$ $\boldsymbol{132.5889}\pm\boldsymbol{0.1106}$
-$\mathbb{S}^{500}$- -- $0.4850$ $1954.3231\pm 14.1107$ $\boldsymbol{0.1404}$ $\boldsymbol{364.5326}\pm\boldsymbol{1.8222}$
-$\mathbb{S}^{1000}$- -- $0.3843$ $5982.0884\pm 204.1833$ $\boldsymbol{0.1140}$ $\boldsymbol{1338.9869}\pm\boldsymbol{2.3954}$
+$\mathbb{S}^{100}$ - - $0.6187$ $498.9705\pm 0.4032$ $\boldsymbol{0.2570}$ $\boldsymbol{15.3786}\pm\boldsymbol{0.0092}$
+$\mathbb{S}^{250}$ - - $0.5195$ $1397.6732\pm 0.7793$ $\boldsymbol{0.1622}$ $\boldsymbol{132.5889}\pm\boldsymbol{0.1106}$
+$\mathbb{S}^{500}$ - - $0.4850$ $1954.3231\pm 14.1107$ $\boldsymbol{0.1404}$ $\boldsymbol{364.5326}\pm\boldsymbol{1.8222}$
+$\mathbb{S}^{1000}$ - - $0.3843$ $5982.0884\pm 204.1833$ $\boldsymbol{0.1140}$ $\boldsymbol{1338.9869}\pm\boldsymbol{2.3954}$
 $\mathrm{E}\left(2\right)$ $\boldsymbol{1.3616}$ $230.0809\pm 1.3967$ $1.3621$ $0.5310\pm 0.0025$ $1.3618$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
@@
 $\mathrm{E}\left(50\right)$ $0.5981$ $11191.9424\pm 597.7002$ $0.6325$ $146.6409\pm 0.1668$ $\boldsymbol{0.2842}$ $\boldsymbol{4.4377}\pm\boldsymbol{0.0451}$
-$\mathrm{E}\left(100\right)$- -- $0.4936$ $491.8223\pm 0.2346$ $\boldsymbol{0.2009}$ $\boldsymbol{14.5367}\pm\boldsymbol{0.0092}$
-$\mathrm{E}\left(250\right)$- -- $0.4159$ $1475.5049\pm 4.1430$ $\boldsymbol{0.1287}$ $\boldsymbol{125.4187}\pm\boldsymbol{0.0888}$
-$\mathrm{E}\left(500\right)$- -- $0.4142$ $1984.7845\pm 10.5689$ $\boldsymbol{0.0967}$ $\boldsymbol{373.2816}\pm\boldsymbol{0.4227}$
-$\mathrm{E}\left(1000\right)$- -- $0.3444$ $5929.4438\pm 167.1873$ $\boldsymbol{0.0799}$ $\boldsymbol{1211.8812}\pm\boldsymbol{0.2511}$
+$\mathrm{E}\left(100\right)$ - - $0.4936$ $491.8223\pm 0.2346$ $\boldsymbol{0.2009}$ $\boldsymbol{14.5367}\pm\boldsymbol{0.0092}$
+$\mathrm{E}\left(250\right)$ - - $0.4159$ $1475.5049\pm 4.1430$ $\boldsymbol{0.1287}$ $\boldsymbol{125.4187}\pm\boldsymbol{0.0888}$
+$\mathrm{E}\left(500\right)$ - - $0.4142$ $1984.7845\pm 10.5689$ $\boldsymbol{0.0967}$ $\boldsymbol{373.2816}\pm\boldsymbol{0.4227}$
+$\mathrm{E}\left(1000\right)$ - - $0.3444$ $5929.4438\pm 167.1873$ $\boldsymbol{0.0799}$ $\boldsymbol{1211.8812}\pm\boldsymbol{0.2511}$
 $\mathbb{T}^{2}$ $7.3890$ $449.4798\pm 10.3282$ $7.3585$ $0.3294\pm 0.0018$ $\boldsymbol{6.8067}$ $\boldsymbol{0.1126}\pm\boldsymbol{0.0002}$
@@
 Fréchet Distribution $0.5408$ $123.8418\pm 0.0499$ $0.5406$ $0.0585\pm 0.0001$ $\boldsymbol{0.5405}$ $\boldsymbol{0.0021}\pm\boldsymbol{0.0001}$
-Cauchy Distribution $1.1134$ $182.3980\pm 2.2559$ $\boldsymbol{1.1133}$ $0.1105\pm 0.0007$ $1.1133$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0002}$
+[…]
 Pareto Distribution $0.3884$ $84.8186\pm 0.0832$ $0.3885$ $0.0235\pm 0.0007$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0001}$
@@
 $\mathrm{E}(n)$ denotes an Ellipsoid of dimension $n$ with half axes of $n$ equally spaced points between $0.5$ and $1.0$, while $\mathcal{P}(n)$ denotes the space of $n\times n$ symmetric positive definite matrices.
-When the computational time was longer than $24$ hours, the value is set to --.
+When the computational time was longer than $24$ hours, the value is set to -.
+ADAM (T=100) Sparse Newton (T=100) GEORCE (T=100)
+Finsler Manifold Length Runtime Length Runtime Length Runtime
+$\mathbb{S}^{2}$ $1.6530$ $0.7706\pm 0.0018$ $\boldsymbol{1.6525}$ $0.2512\pm 0.0089$ $1.6525$ $\boldsymbol{0.0372}\pm\boldsymbol{0.0004}$
+$\mathbb{S}^{3}$ $1.7451$ $0.7362\pm 0.0036$ - - $\boldsymbol{1.7432}$ $\boldsymbol{0.0431}\pm\boldsymbol{0.0005}$
+$\mathbb{S}^{5}$ $1.7371$ $0.9292\pm 0.0038$ - - $\boldsymbol{1.7029}$ $\boldsymbol{0.1036}\pm\boldsymbol{0.0048}$
+$\mathbb{S}^{10}$ $1.5174$ $1.1358\pm 0.0018$ $1.4783$ $208.1472\pm 0.0176$ $\boldsymbol{1.4109}$ $\boldsymbol{0.1000}\pm\boldsymbol{0.0033}$
+$\mathbb{S}^{20}$ $1.2009$ $1.5876\pm 0.0028$ - - $\boldsymbol{1.0760}$ $\boldsymbol{0.1509}\pm\boldsymbol{0.0058}$
+$\mathbb{S}^{50}$ $0.8130$ $2.0381\pm 0.0026$ - - $\boldsymbol{0.7283}$ $\boldsymbol{0.1373}\pm\boldsymbol{0.0032}$
+$\mathbb{S}^{100}$ $0.6146$ $2.7712\pm 0.0032$ $0.6210$ $972.0195\pm 0.0307$ $\boldsymbol{0.2553}$ $\boldsymbol{0.1106}\pm\boldsymbol{0.0028}$
+$\mathbb{S}^{250}$ $0.5196$ $3.9803\pm 0.0019$ - - $\boldsymbol{0.1636}$ $\boldsymbol{0.5069}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{500}$ $0.4849$ $6.6164\pm 0.0023$ - - $\boldsymbol{0.1417}$ $\boldsymbol{1.6591}\pm\boldsymbol{0.0004}$
+$\mathbb{S}^{1000}$ $0.3835$ $25.4638\pm 0.1448$ - - $\boldsymbol{0.1134}$ $\boldsymbol{7.9991}\pm\boldsymbol{0.0074}$
+$\mathrm{E}\left(2\right)$ $1.3634$ $0.7758\pm 0.0036$ $\boldsymbol{1.3616}$ $0.3735\pm 0.0010$ $1.3618$ $\boldsymbol{0.0282}\pm\boldsymbol{0.0009}$
+$\mathrm{E}\left(3\right)$ $1.4085$ $0.7541\pm 0.0032$ - - $\boldsymbol{1.4023}$ $\boldsymbol{0.1273}\pm\boldsymbol{0.0047}$
+$\mathrm{E}\left(5\right)$ $1.3305$ $1.0252\pm 0.0017$ $1.5359$ $170.9661\pm 0.0296$ $\boldsymbol{1.2956}$ $\boldsymbol{0.1135}\pm\boldsymbol{0.0039}$
+$\mathrm{E}\left(10\right)$ $1.1235$ $1.1254\pm 0.0045$ - - $\boldsymbol{1.0804}$ $\boldsymbol{0.0825}\pm\boldsymbol{0.0031}$
+$\mathrm{E}\left(20\right)$ $0.8987$ $1.6039\pm 0.0035$ - - $\boldsymbol{0.8121}$ $\boldsymbol{0.1282}\pm\boldsymbol{0.0054}$
+$\mathrm{E}\left(50\right)$ $0.6332$ $2.0424\pm 0.0034$ - - $\boldsymbol{0.2845}$ $\boldsymbol{0.0770}\pm\boldsymbol{0.0004}$
+$\mathrm{E}\left(100\right)$ $0.4953$ $2.6704\pm 0.0031$ $0.4577$ $470.9805\pm 0.1448$ $\boldsymbol{0.2018}$ $\boldsymbol{0.1127}\pm\boldsymbol{0.0023}$
+$\mathrm{E}\left(250\right)$ $0.4179$ $4.0843\pm 0.0009$ - - $\boldsymbol{0.1287}$ $\boldsymbol{0.5239}\pm\boldsymbol{0.0047}$
+$\mathrm{E}\left(500\right)$ $0.4139$ $6.9391\pm 0.0073$ $0.2580$ $1435.3881\pm 0.0832$ $\boldsymbol{0.0968}$ $\boldsymbol{1.7109}\pm\boldsymbol{0.0010}$
+$\mathrm{E}\left(1000\right)$ $0.3442$ $25.9660\pm 0.1038$ - - $\boldsymbol{0.0798}$ $\boldsymbol{7.9988}\pm\boldsymbol{0.0079}$
+$\mathbb{T}^{2}$ $7.4054$ $0.8017\pm 0.0045$ $9.1903$ $152.0036\pm 0.0158$ $\boldsymbol{6.8067}$ $\boldsymbol{0.3145}\pm\boldsymbol{0.0050}$
+$\mathbb{H}^{2}$ $1.3341$ $0.7949\pm 0.0072$ $1.3338$ $0.4562\pm 0.0004$ $\boldsymbol{1.3337}$ $\boldsymbol{0.0256}\pm\boldsymbol{0.0006}$
+$\mathcal{P}\left(2\right)$ $0.7334$ $0.9344\pm 0.0013$ $\boldsymbol{0.7291}$ $0.4309\pm 0.0014$ $0.7292$ $\boldsymbol{0.0286}\pm\boldsymbol{0.0005}$
+$\mathcal{P}\left(3\right)$ $1.2987$ $1.0812\pm 0.0006$ $1.2856$ $0.6885\pm 0.0009$ $\boldsymbol{1.2856}$ $\boldsymbol{0.0365}\pm\boldsymbol{0.0005}$
+Gaussian Distribution $1.6879$ $0.4962\pm 0.0034$ $1.6878$ $0.4148\pm 0.0055$ $\boldsymbol{1.6878}$ $\boldsymbol{0.0255}\pm\boldsymbol{0.0004}$
+Fréchet Distribution $0.5406$ $0.2307\pm 0.0069$ $\boldsymbol{0.5405}$ $0.2060\pm 0.0042$ $0.5405$ $\boldsymbol{0.0088}\pm\boldsymbol{0.0003}$
+Cauchy Distribution $1.1140$ $0.4994\pm 0.0017$ $\boldsymbol{1.1133}$ $0.5167\pm 0.0042$ $1.1133$ $\boldsymbol{0.0367}\pm\boldsymbol{0.0017}$
+Pareto Distribution $0.3885$ $0.1014\pm 0.0038$ $0.3884$ $0.1042\pm 0.0031$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0096}\pm\boldsymbol{0.0004}$
+Table 6: The table shows the length of the estimated geodesics for different methods on a GPU.
+The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
+When the computational time was longer than $24$ hours, the value is set to -.
+ADAM (T=50) Sparse Newton (T=50) GEORCE (T=50)
+Finsler Manifold Length Runtime Length Runtime Length Runtime
+$\mathbb{S}^{2}$ $\boldsymbol{1.6383}$ $0.7833\pm 0.0056$ $1.6385$ $0.2064\pm 0.0073$ $1.6385$ $\boldsymbol{0.0457}\pm\boldsymbol{0.0004}$
+$\mathbb{S}^{3}$ $1.7300$ $0.7587\pm 0.0049$ - - $\boldsymbol{1.7297}$ $\boldsymbol{0.0499}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{5}$ $1.6875$ $0.9909\pm 0.0022$ - - $\boldsymbol{1.6868}$ $\boldsymbol{0.1432}\pm\boldsymbol{0.0050}$
+$\mathbb{S}^{10}$ $1.4013$ $1.1472\pm 0.0024$ - - $\boldsymbol{1.3908}$ $\boldsymbol{0.1747}\pm\boldsymbol{0.0067}$
+$\mathbb{S}^{20}$ $1.0829$ $1.5583\pm 0.0016$ - - $\boldsymbol{1.0560}$ $\boldsymbol{0.2663}\pm\boldsymbol{0.0005}$
+$\mathbb{S}^{50}$ $0.7427$ $1.4442\pm 0.0038$ - - $\boldsymbol{0.3508}$ $\boldsymbol{0.0594}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{100}$ $0.5511$ $2.5558\pm 0.0020$ - - $\boldsymbol{0.2626}$ $\boldsymbol{0.0850}\pm\boldsymbol{0.0024}$
+$\mathbb{S}^{250}$ $0.4094$ $3.8872\pm 0.0020$ - - $\boldsymbol{0.1686}$ $\boldsymbol{0.2660}\pm\boldsymbol{0.0042}$
+$\mathbb{S}^{500}$ $0.3531$ $7.5214\pm 0.0022$ $0.2861$ $1388.8278\pm 0.0965$ $\boldsymbol{0.1193}$ $\boldsymbol{1.1879}\pm\boldsymbol{0.0003}$
+$\mathbb{S}^{1000}$ $0.3269$ $13.1209\pm 0.0379$ - - $\boldsymbol{0.0828}$ $\boldsymbol{6.0238}\pm\boldsymbol{0.0039}$
+$\mathrm{E}\left(2\right)$ $1.3460$ $0.1327\pm 0.0060$ $\boldsymbol{1.3460}$ $0.2847\pm 0.0043$ $1.3461$ $\boldsymbol{0.0332}\pm\boldsymbol{0.0011}$
+$\mathrm{E}\left(3\right)$ $1.3870$ $0.7621\pm 0.0034$ - - $\boldsymbol{1.3826}$ $\boldsymbol{0.2156}\pm\boldsymbol{0.0006}$
+$\mathrm{E}\left(5\right)$ $1.2780$ $0.9718\pm 0.0022$ - - $\boldsymbol{1.2757}$ $\boldsymbol{0.1654}\pm\boldsymbol{0.0050}$
+$\mathrm{E}\left(10\right)$ $1.0665$ $1.1072\pm 0.0025$ - - $\boldsymbol{1.0126}$ $\boldsymbol{0.2215}\pm\boldsymbol{0.0062}$
+$\mathrm{E}\left(20\right)$ $0.8405$ $1.5702\pm 0.0031$ - - $\boldsymbol{0.7946}$ $\boldsymbol{0.1937}\pm\boldsymbol{0.0061}$
+$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5797}$ $1.4336\pm 0.0029$ - - - -
+$\mathrm{E}\left(100\right)$ $0.4334$ $2.0750\pm 0.0026$ - - $\boldsymbol{0.2046}$ $\boldsymbol{0.0830}\pm\boldsymbol{0.0004}$
+$\mathrm{E}\left(250\right)$ $0.3234$ $2.8610\pm 0.0021$ - - $\boldsymbol{0.1365}$ $\boldsymbol{0.2644}\pm\boldsymbol{0.0029}$
+$\mathrm{E}\left(500\right)$ $0.2862$ $5.1735\pm 0.0122$ $0.2428$ $905.0696\pm 0.0365$ $\boldsymbol{0.0940}$ $\boldsymbol{0.9444}\pm\boldsymbol{0.0011}$
+$\mathrm{E}\left(1000\right)$ $0.2723$ $10.4951\pm 0.0347$ - - $\boldsymbol{0.0893}$ $\boldsymbol{4.4688}\pm\boldsymbol{0.0060}$
+$\boldsymbol{7.2683}$ $\boldsymbol{0.7530}\pm\boldsymbol{0.0015}$ $9.0475$ $78.7132\pm 0.0222$ $9.0475$ $13.7393\pm 0.0093$
+$\mathbb{H}^{2}$ $1.3328$ $0.7454\pm 0.0052$ $1.3316$ $0.2303\pm 0.0058$ $\boldsymbol{1.3316}$ $\boldsymbol{0.0317}\pm\boldsymbol{0.0003}$
+$\mathcal{P}\left(2\right)$ $0.7249$ $0.5927\pm 0.0008$ $\boldsymbol{0.7249}$ $0.2189\pm 0.0051$ $0.7249$ $\boldsymbol{0.0331}\pm\boldsymbol{0.0011}$
+$\mathcal{P}\left(3\right)$ $1.2710$ $1.0184\pm 0.0008$ $1.2707$ $0.4325\pm 0.0020$ $\boldsymbol{1.2707}$ $\boldsymbol{0.0434}\pm\boldsymbol{0.0006}$
+Gaussian Distribution $\boldsymbol{1.6796}$ $0.4337\pm 0.0111$ $1.6806$ $0.2081\pm 0.0047$ $1.6806$ $\boldsymbol{0.0282}\pm\boldsymbol{0.0017}$
+Fréchet Distribution $0.5370$ $0.0829\pm 0.0032$ $0.5370$ $0.1059\pm 0.0042$ $\boldsymbol{0.5370}$ $\boldsymbol{0.0111}\pm\boldsymbol{0.0003}$
+Cauchy Distribution $1.1085$ $0.0626\pm 0.0010$ $\boldsymbol{1.1084}$ $0.2270\pm 0.0081$ $1.1084$ $\boldsymbol{0.0410}\pm\boldsymbol{0.0015}$
+Pareto Distribution $0.3857$ $0.0752\pm 0.0025$ $\boldsymbol{0.3856}$ $0.1054\pm 0.0028$ $0.3857$ $\boldsymbol{0.0082}\pm\boldsymbol{0.0003}$
+Table 7: The table shows the length of the estimated geodesics for different methods on a GPU.
+The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
+When the computational time was longer than $24$ hours, the value is set to -.
 Finsler manifolds
-To illustrate our method for different Finsler manifolds and dimensions, we consider a microswimmer on a Riemannian manifold $m$ that moves with constant self-propulsion velocity $v$ with Riemannian length, $v_{0}:=||v||_{g}$ [51] and we assume that this movement is affected by a generic force field, $f$, of the form
+[…]
 $f(x)=\frac{\sin x\odot\cos x}{(\cos x)^{\top}G(x)\cos x},$ (21)
 where $\cos x$ and $\sin x$ denote the element-wise function for each entry of the local coordinates $x$, while $G$ denotes the corresponding Riemannian background metric such that the force field has $G$-length less than or equal to 1.
-If $||f(x)||_{g}<v_{0}$, then the optimal travel time of the microswimmer corresponds to a geodesic for a Finsler manifold of the form [51]
+If $f(x)||_{g}<v_{0}$, then the optimal travel time of the microswimmer corresponds to a geodesic for a Finsler manifold of the form [51]
 $F=\sqrt{a_{ij}\dot{r}^{i}\dot{r}^{j}}+b_{i}\dot{r}^{i},$
 where $a_{ij}=g_{ij}\lambda+f_{i}f_{j}\lambda^{2}$, $b_{i}=-f_{i}\lambda$, $b_{i}=-f_{i}\lambda$, $f_{i}=g_{ij}f^{j}$ and $\lambda^{-1}=v_{0}^{2}-g_{ij}f^{i}f^{j}$ [51], while $r(t)$ denotes the position of the microswimmer at time $t$.
-We refer to [51] and Appendix C for details.
+We refer to [51] and Appendix D for details.
 We consider the same background metrics in Table 2 all equipped with the generic force field in Eq.
@@
 10 shows the estimated geodesics for four different Riemannian manifolds equipped with the generic force field in Eq.
-15.
+21.
 We show the estimated geodesics for $T=100$ grid points in Table 5.
-In Appendix E.3 we show similar runtime experiments for varying grid points as well as other optimization algorithms.
-We see that GEORCE is considerably faster and more accurate than the alternatives.
+In Table 6 and Table 7 we show the corresponding results on a GPU for compatible methods for $T=100$ and $T=50$.
 Applications to generative models
@@
 For latent space generative models, the decoder can be seen as a smooth $d$ dimensional immersion, $f:\mathcal{Z}\rightarrow\mathcal{X}$, which defines a Riemannian geometry with the pull-back metric $G(z)=J_{f}^{\top}(z)J_{f}(z)$, where $J_{f}$ denotes the Jacobian of $z$ [11, 12].
-We train a VAE for the CelebA dataset [53], which consists of $202{,}599$ facial images of size $64\times 64\times 3$, where the architecture is described in Appendix D.1.
+We train a VAE for the CelebA dataset [53], which consists of 202,599202{,}599 facial images of size $64\times 64\times 3$, where the architecture is described in Appendix E.1.
 Fig.
@@
 We see that GEORCE obtains a smaller length than comparable methods for $T=100$ grid points.
-We refer to Appendix D.2 for runtime results for the VAE, where we see that GEORCE is significantly faster than alternative methods.
+We refer to Appendix E.2 for runtime results for the VAE, where we see that GEORCE is significantly faster than alternative methods.
 Note that the tolerance for convergence is $10^{-3}$, since it was observed that it was not possible for any method to obtain an $\ell^{2}$-norm of the gradient less than $10^{-4}$.
 This is most likely due to potential instabilities for the learned manifold.
-We summarize the results in Table 11 in Appendix E.3, where we also consider the geodesic with a VAE trained on MNIST data [54] and SVHN data [55].
-For the exact architecture and further experiments of the VAE, we refer to Appendix D.1 and Appendix D.2.
+We summarize the results in Table 8 on a CPU.
 Figure 11: Geodesics for a variational autoencoder [52] trained on the CelebA dataset [53] using GEORCE and alternative methods.
@@
 The first-axis shows the grid number for the image sequence.
-The details of the manifolds and experiments can be found Appendix D.1
+The details of the manifolds and experiments can be found Appendix E.1
+BFGS (T=100) ADAM (T=100) GEORCE (T=100)
+Riemannian Manifold Length Runtime Length Runtime Length Runtime
+VAE MNIST $10.3075$ $1407.5743\pm 6.4430$ $10.3312$ $645.5201\pm 2.6648$ $\boldsymbol{10.3074}$ $\boldsymbol{67.6149}\pm\boldsymbol{0.0304}$
+VAE SVHN - - $19.9855$ $13325.2285\pm 123.4978$ $\boldsymbol{19.9346}$ $\boldsymbol{882.4120}\pm\boldsymbol{1.8839}$
+VAE CelebA - - - - $\boldsymbol{42.2666}$ $\boldsymbol{2222.5259}\pm\boldsymbol{8.2988}$
+Table 8: The table shows the length of the estimated geodesics for the BFGS-algorithm, ADAM and GEORCE on a CPU.
+The methods were terminated if the $\ell^{2}$-norm of the gradient was less than $10^{-3}$ or $1,000$ iterations have been reached.
+When the computational time was longer than $24$ hour, the value is set to -.
 6 Conclusion
@@
   • Funding: Research grant 42062 from VILLUM FONDEN; funding from the European Research Council (ERC) under the European Union’s Horizon research and innovation programme (grant agreement 101125993); funding from the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606).
-  • Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use): NA
+  • Conflict of interest/Competing interests: NA
   • Ethics approval and consent to participate: NA
@@
 Appendix A Proofs and derivations
-A.1 Assumptions
-In the following, we will prove local quadratic convergence under the following assumptions.
-Assumption 5.
-We assume the following regarding the discretized energy functional.
-  • A (local) minimum point $z^{*}=(x^{*},u^{*})$ to the discretized energy functional is defined as strongly unique minimum point if and only if
-    $\exists\epsilon>0:\quad\exists K>0:\quad\forall z\in B_{\epsilon}(z^{*}):\quad E(z)-E(z^{*})\geq K\left\lVert z-z^{*}\right\rVert,$
-    where $B_{\epsilon}(z^{*})=\left\{z|\left\lVert z-z^{*}\right\rVert<\epsilon\right\}$.
-  • We assume that the discretized energy functional $E(z)$ is locally Lipschitz.
-  • The first order Taylor approximation of the discretized energy functional is
-    $\Delta E=\langle\nabla E(z_{0}),\Delta z\rangle+\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert,$
-    i.e., we assume that $\mathcal{O}\left(\Delta z\right)$ is independent of $z_{0}$, and the term $\mathcal{O}\left(\Delta z\right)\left\lVert\Delta z\right\rVert$ can be re-written as $\mathcal{\tilde{O}}\left(\left\lVert\Delta z\right\rVert^{2}\right)$ locally (which follows from the Lipschitz condition).
-Further for global convergence we will assume Lipschitz continuity of the discretized energy functional.
-Note that since the metric varies smoothly on the manifold, then the discretized energy functional is also smooth and hence Lipschitz.
-Thus, all Finsler and Riemannian manifolds satisfies the above conditions, but an even larger class of energy functionals that does not necessarily stem from a Riemannian or Finsler metric will satisfy the requirements.
-A.2 Proof of necessary conditions
-Proposition 6.
-The necessary conditions for a minimum in Eq. 6 is
-$\begin{split}&2G(x_{t})u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &{\left.\kern-1.2pt\nabla_{y}\left[u_{t}^{\top}G(y)u_{t}\right]\vphantom{\big{|}}\right|_{y=x_{t}}}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1.\\ &x_{0}=a,x_{T}=b.\end{split}$ (16)
-where $\mu_{t}\in\mathbb{R}^{d}$ for $t=0,\dots,T-1$.
-Proof.
-The Hamiltonian of Eq. 6 is
-$H_{t}(x_{t},u_{t},\mu_{t})=u_{t}^{\top}G(x_{t})u_{t}+\mu_{t}^{\top}(x_{t}+u_{t}).$ (17)
-The Hamiltonian $H_{t}(x_{t},u_{t},\mu_{t})$ is strictly convex in $u_{t}$, since $G(x_{t})$ is positive definite.
-The time-discrete version of Pontryagins maximum principle [56] then yields the following when applied to our control problem (6)
-$\begin{split}&2G(x_{t})u_{t}+\mu_{t}=0,\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &{\left.\kern-1.2pt\nabla_{y}\left[u_{t}^{\top}G(y)u_{t}\right]\vphantom{\big{|}}\right|_{y=x_{t}}}+\mu_{t}=\mu_{t-1},\quad t=1,\dots,T-1.\\ &x_{0}=a,x_{T}=b.\end{split}$ (18)
-A.3 Proof of update scheme for Riemannian geometry
-Proposition 7.
-The update scheme in eq. 7 and eq. 9 for $u_{t},\mu_{t}$ and $x_{t}$ can in closed-form be written as
-$\begin{split}&\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &x_{t+1}=x_{t}+u_{t},\quad t=0,\dots,T-1,\\ &x_{0}=a.\end{split}$
-Proof.
-From the update formula $\nu_{t}+\mu_{t}=\mu_{t-1}$ for $t=1,\dots,T-1$ we deduce that $\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}$.
-From the update scheme $02G_{t}u_{t}+\mu_{t}=$ 0 we deduce that $u_{t}=-\frac{1}{2}G_{t}^{-1}\mu_{t}$, since $G_{t}$ is positive definite and hence has a well-defined inverse.
-Noting that $\sum_{t=0}^{T-1}u_{t}=b-a$ we can re-write the iterative formulas as
-$\begin{split}&\mu_{t}=\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j},\quad t=0,\dots,T-1,\\ &u_{t}=-\frac{1}{2}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t}^{T-1}\nu_{j}\right),\quad t=0,\dots,T-1,\\ &-\frac{1}{2}\sum_{t=0}^{T-1}G_{t}^{-1}\left(\mu_{T-1}+\sum_{j>t+1}^{T-1}\nu_{j}\right)=b-a.\end{split}$
-The last equation can be used to solve for $\mu_{T-1}$ utilizing that the sum of (inverse) positive definite matrices is positive definite and has a well-defined inverse
-$\mu_{T-1}=\left(\sum_{t=0}^{T-1}G_{t}^{-1}\right)^{-1}\left(2(a-b)-\sum_{t=0}^{T-1}G_{t}^{-1}\sum_{t>j}^{T-1}\nu_{j}\right),$
-which proves the result. ∎
-A.4 Proof of update scheme for Finslerian geometry
+A.1 Proof of update scheme for Finslerian geometry
 Consider the following control problem in the Finslerian case.
@@
 where it is exploited that the fundamental tensor is positive definite.
-A.5 Proof of global convergence
-To prove global convergence for the algorithm GEORCE we first show existence and uniqueness by proving two lemmas.
-Secondly, we prove that in each iteration the discretized energy functional is decreasing, unless a (local) minimum has been found.
-We will assume Lipschitz continuity of the metric tensor as described in 5.
-[…]
-Lemma 8.
-Assume that $(x^{(i)},u^{(i)})$ is a feasible solution, then
-  1. 1.
-    There exists a unique solution $\left(x^{(i+1)},u^{(i+1)}\right)$ to the system of equations in eq. 10 based on $\left(x^{(i)},u^{(i)}\right)$.
-  2. 2.
-    All linear combinations $(1-\alpha)\left(x^{(i)},u^{(i)}\right)+\alpha\left(x^{(i+1)},u^{(i+1)}\right)$ for $0<\alpha\leq 1$ are feasible solutions.
-Proof.
-Since the matrices, $\left\{G\left(x_{t}^{(i)}\right)\right\}_{t=0}^{T}$ are positive definite, then the matrices are regular with unique inverse, and so are the sum of the inverse matrices.
-This means that there exists an unique solution to the update scheme in proposition 2 proving 1.
-Since the solution $\left(x^{(i+1)},u^{(i+1)}\right)$ is also a feasible solution for the update scheme in proposition 2, then the linear combination of two feasible solutions will also be a feasible solution, i.e.
-$\sum_{t=0}^{T-1}u_{t}^{(j)}=(b-a),\quad j=1,\dots,(i+1).$
-The new solution based on the linear combination gives
-$\begin{split}&(1-\alpha)\sum_{t=0}^{T-1}u_{t}^{(i)}+\alpha\sum_{t=0}^{T-1}u_{t}^{(i+1)}=(1-\alpha)(b-a)+\alpha(b-a)=(b-a)\\ &(1-\alpha)x^{(i)}+\alpha x^{(i+1)}=(1-\alpha)\left(\alpha,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b\right)+\alpha\left(a,x_{1}^{(i+1)},\dots,x_{T-1}^{(i+1)},b\right)\\ &=\left(a,(1-\alpha)x_{1}^{(i)}+\alpha x_{1}^{(i+1)},\dots,(1-\alpha)x_{T-1}^{(i)}+\alpha x_{T-1}^{(i+1)},b\right),\end{split}$
-where for any $t=0,\dots,T-1$
-$\begin{split}(1-\alpha)x_{t}^{(i)}+\alpha x_{t}^{(i+1)}&=(1-\alpha)\left(a+\sum_{j=0}^{t-1}u_{j}^{(i)}\right)+\alpha\left(a+\sum_{j=0}^{t-1}u_{j}^{(i+1)}\right)\\ &=a+\sum_{j=0}^{t-1}\left((1-\alpha)u_{j}^{(i)}+\alpha u_{j}^{(i+1)}\right).\end{split}$
-This shows that the linear combination in the state is feasible as it produces feasible state variables in terms of start and end point.
-Furthermore, each state variable is feasible, if they were determined from the linear combination of the control vectors, which proves that the linear combinations of the state and control variables are also feasible solutions.
-∎
-Lemma 9.
-Let $\left\{x_{t}^{(i)},u_{t}^{(i)}\right\}_{t=0}^{T}$ denote the feasible solution after iteration $i$ in GEORCE.
-If $\left\{x_{t}^{(i)},u_{t}^{(i)}\right\}_{t=0}^{T}$ is not a (local) minimum point, then the feasible solution from iteration $\left\{x_{t}^{(i+1)},u_{t}^{(i+1)}\right\}_{t=0}^{T}$ will decrease the objective function in the sense that there exists an $\eta>0$ such that for all $\alpha$ with $0<\alpha\leq\eta\leq 1$, then
-$E\left(x^{(i)}+\alpha\left(x^{(i+1)}-x^{(i)}\right),u^{(i)}+\alpha\left(u^{(i+1)}-u^{(i)}\right)\right)<E\left(x^{(i)},u^{(i)}\right),$
-where $x^{(i)}=\left(a,x_{1}^{(i)},\dots,x_{T-1}^{(i)},b\right)$ and $u^{(i)}=\left(u_{0}^{(i)},u_{1}^{(i)},\dots,u_{T-2}^{(i)},u_{T-1}^{(i)}\right)$.
-Proof.
-$E(x,u)$ is a smooth function. The first order Taylor approximation of the energy function $E(x,u)$ in eq. 5 in $(x,u)$ is
-$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{x_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta x_{t}\right\rangle+\mathcal{O}(\Delta x_{t})\left\lVert\Delta x_{t}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right),\end{split}$
-where $\Delta u_{t}:=u_{t}^{(i+1)}-u_{t}^{(i)}$ and $\Delta x_{t}:=x_{t}^{(i+1)}-x_{t}^{(i)}$. By the optimally conditions in eq. 7 we have that
-$\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{t-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=-\mu_{t},\quad t=0,\dots,T-1.\end{split}$ (22)
-Inserting the optimality conditions in the Taylor expansion we have that
-$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\langle\mu_{t-1}-\mu_{t},\Delta x_{t}\rangle+\mathcal{O}(\Delta x_{t})\left\lVert\Delta x_{t}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$
-Since $\Delta x_{t}=\sum_{j=0}^{t-1}\Delta u_{j}$, then
-$\begin{split}\Delta E(x,u)&=\sum_{t=1}^{T-1}\left(\langle\mu_{t-1}-\mu_{t},\sum_{j=0}^{t-1}\Delta u_{j}\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$
-Re-arranging the terms we get
-$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-2}\left\langle\mu_{t},\sum_{j=0}^{t}\Delta u_{j}\right\rangle\\ &-\sum_{t=1}^{T-1}\left(\left\langle\mu_{t},\sum_{j=0}^{t-1}\Delta u_{j}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert u_{t}\right\rVert\right)\\ &=\langle\mu_{0},\Delta u_{0}\rangle+\sum_{t=1}^{T-2}\langle\mu_{t},\Delta u_{t}\rangle-\left\langle\mu_{T-1},\sum_{j=0}^{T-1}\Delta u_{j}\right\rangle+\langle\mu_{T-1},\Delta u_{T-1}\rangle\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\end{split}$
-Since $\Delta u_{j}=u_{j}^{(i+1)}-u_{j}^{(i)}$ and $\Delta x_{j}=x_{j}^{(i+1)}-x_{j}^{(i)}$ for $j=0,\dots,T$, and since the solution is feasible we have that $\sum_{j=0}^{T-1}u_{j}^{(i)}=b-a$ for any $i$, which implies that $\sum_{j=0}^{T-1}\Delta u_{j}=0$.
-This reduces the above expression to
-$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left(\langle\mu_{t},\Delta u_{t}\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\left\langle{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\\ &=\sum_{t=0}^{T-1}\left\langle\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}},\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\sum_{t=0}^{T-1}\mathcal{O}\left(\Delta u_{t}\right)\left\lVert u_{t}\right\rVert.\end{split}$
-By eq. 22 we have
-$\begin{split}\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=-{\left.\kern-1.2pt\nabla_{u_{t}}E(x_{t},u_{t})\vphantom{\big{|}}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}\\ &+{\left.\kern-1.2pt\nabla_{u_{t}}E(x,u)\vphantom{\big{|}}\right|_{(x,u)=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}\\ &=-2G\left(x_{t}^{(i)}\right)\left(u_{t}^{(i+1)}-u_{t}^{(i)}\right),\end{split}$ (23)
-which implies that
-$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right).\end{split}$ (24)
-Since $G\left(x_{t}^{(i)}\right)$ is positive definite, the first summation is negative assuming that at least one non-zero vector $\{\Delta u_{t}\}_{t=0}^{T-1}$, which is the case, since we have assumed that $\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ is not a (local) minimum.
-Now scale all vectors, $\left\{\Delta u_{t}\right\}_{t=0}^{T-1}$ with a scalar $10<\alpha<$ 1.
-Note that the term, $-2G\left(x_{t}^{(i)}\right)\Delta u_{t}$, is unaffected by this, since it equals $\mu_{t}+\nabla_{u_{t}}E(x,u)\left(x_{t}^{(i)},u_{t}^{(i)}\right)$, which is independent of $\alpha$.
-The first order Taylor order approximation is
-$\begin{split}\Delta E(x,u)&=\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\alpha\Delta u_{t}\right\rangle\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\alpha\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\alpha\Delta u_{j}\right\rVert+\mathcal{O}\left(\alpha\Delta u_{t}\right)\left\lVert\alpha\Delta u_{t}\right\rVert\right).\end{split}$
-In the limit it now follows that there exists an $\eta>0$ such that $\frac{\Delta E(x,u)}{\alpha}<0$ for some $\alpha$ with $0<\alpha\leq\eta\leq 1$, which implies that $\Delta E(x,u)<0$. ∎
-Theorem 10.
-Let $E^{(i)}$ be the value of the discretized energy functional for the solution after iteration $i$ (with line search) in GEORCE.
-If the starting point $\left(x^{(0)},u^{(0)}\right)$ is feasible, then the series $\left\{E^{(i)}\right\}_{i>0}$ will converge to a (local) minimum.
-Proof.
-If the starting point $\left(x^{(0)},u^{(0)}\right)$ in GEORCE is feasible, then the points $\left(x^{(i)},u^{(i)}\right)$ after each iteration (with line-search) in GEORCE will also be feasible solutions by lemma 8.
-The discretized energy functional is a positive function, i.e.
-a lower bound is at least 0.
-Assume that the series $\left\{E^{(i)}\right\}_{i>0}$ is not converging to a (local) minimum.
-From lemma 9 it follows that $\left\{E^{(i)}\right\}_{i>0}$ is decreasing for increasing $i$, and since $E^{(i)}$ has a lower bound, the series $\left\{E^{(i)}\right\}_{i}$ can not be non-converging.
-Assume now that the convergence value for $\left\{E^{(i)}\right\}_{i}$, denoted $\hat{E}$, is not a (local) minimum, and denote the convergence point $(\hat{x},\hat{u})$.
-According to lemma 9 GEORCE will from the solution $(\hat{x},\hat{u})$ in the following iteration produce a new point $(\tilde{x},\tilde{u})$ such that $E(\tilde{x},\tilde{u})<\hat{E}$, which is contradiction as the series $\left\{E^{(i)}\right\}_{i}$ is assumed to be converging to $\hat{E}$.
-The series, $\left\{E^{(i)}\right\}_{i}$, will therefore converge, and the convergence point is a (local) minimum.
-∎
-Under the assumption of a strongly unique minimum point in Appendix A.1 GEORCE will converge to a local minimum by 3. Thus, it is not possible that GEORCE will ”jump” between local minimas.
-A.6 Proof of local quadratic convergence
-In this section we show that GEORCE has locally quadratic convergence to a (local) minimum under certain assumptions satisfied by the discretized energy functional.
-We will assume the assumptions in 5 hold.
-We first prove the following result, before proving quadratic local convergence.
-Lemma 11.
-Assume $f(x)$ is a smooth function. Let $z^{*}$ be a strongly unique (local) minimum point for $f$. Then
-$\exists\epsilon>0:\quad\forall z\in B_{\epsilon}(z^{*})\setminus\{z^{*}\}:\quad\langle\nabla f(z),z^{*}-z\rangle<0.$
-Proof.
-Since $f(x)$ is a smooth function, then the first order Taylor approximation for $f(x)$ exists. Consider a point $z\in B_{\epsilon}(z^{*})\setminus\{z^{*}\}$, then
-$f(z^{*})-f(z)=\langle\nabla f(z),z^{*}-z\rangle+\mathcal{O}\left(z^{*}-z\right)\left\lVert z^{*}-z\right\rVert.$
-Since $z^{*}$ is a strongly unique (local) minimum point, then
-$f(z^{*})-f(z)\leq-K\left\lVert z^{*}-z\right\rVert<0,\quad K>0,$
-which implies that
-$\frac{f(z^{*})-f(z)}{\left\lVert z^{*}-z\right\rVert}=\left\langle\nabla f(z),\frac{z^{*}-z}{\left\lVert z^{*}-z\right\rVert}\right\rangle+\mathcal{O}\left(z^{*}-z\right)\leq-K.$
-Since $\lim_{z\rightarrow z^{*}}\mathcal{O}(z^{*}-z)=0$, then there exists $\epsilon>0$ such that
-$\left\langle\nabla f(z),\frac{z^{*}-z}{\left\lVert z^{*}-z\right\rVert}\right\rangle<0,$
-which implies that $\langle\nabla f(z),z^{*}-z\rangle<0$ for some $\epsilon>0$ for all $z\in B_{\epsilon}\left(z^{*}\right)\setminus\left\{z^{*}\right\}$. ∎
-Proposition 12.
-If the discretized energy functional has a strongly unique (local) minimum point $z^{*}$ and locally $\alpha^{*}=1$, i.e. no line search, then GEORCE has locally quadratic convergence, i.e.
-$\exists\epsilon>0:\quad\exists c>0:\quad\forall z^{(i)}\in B_{\epsilon}(z^{*}):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
-where $z^{(i)}=(x^{(i)},u^{(i)})$ is the solution from GEORCE at iteration $i$.
-Proof.
-Assume that $z^{*}$ is a (local) minimum point.
-Since GEORCE has global convergence to a (local) minimum point by theorem 3, we will assume that the algorithm has convergence point $E(z^{*})$.
-Consider iteration $i$ with $z^{(i)}=(x^{(i)},u^{(i)})$ from GEORCE, and assume that all following iterations in GEORCE is in the ball $B_{\epsilon_{1}}(z^{*})$, and fulfills the strongly unique assumption in 5.
-Note that as the series $\left\{z^{(i)}\right\}$ converge to $z^{*}$, then $\left\{z^{(i)}\right\}_{i\geq k}\subset B_{\epsilon_{1}}(z^{*})$ for some $k>0$.
-It then follows from the assumption of strongly unique minimizing point that for some $K>0$
-$E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\geq K\left\lVert z^{(i+1)}-z^{*}\right\rVert,$
-since $z^{(i+1)}\in B_{\epsilon_{1}}(z^{*})$. The left hand term above can be rearranged into
-$E\left(z^{(i+1)}\right)-E\left(z^{*}\right)=\left(E\left(z^{(i+1)}\right)-E\left(z^{(i)}\right)\right)+\left(E\left(z^{(i)}\right)-E\left(z^{*}\right)\right).$
-By eq. 24 in the proof of lemma 9 and utilizing $\alpha^{*}=1$, it follows that
-$\begin{split}&E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\\ &=\left(E\left(z^{(i+1)}\right)-E\left(z^{(i)}\right)\right)+\left(E\left(z^{(i)}\right)-E\left(z^{*}\right)\right)\\ &=\sum_{t=0}^{T-1}\left(\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert+\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\right)\\ &-\sum_{t=0}^{T-1}\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\left(u_{t}^{*}-u_{t}^{(i)}\right)\right\rangle\\ &-\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert-\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\\ &=\sum_{t=0}^{T-1}\left(\left\langle-2G\left(x_{t}^{(i)}\right)\Delta u_{t},\Delta u_{t}-\left(u_{t}^{*}-u_{t}^{(i)}\right)\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}(\Delta u_{t})\left\lVert\Delta u_{t}\right\rVert+\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\right)\\ &=\sum_{t=0}^{T-1}\left(\left\langle 2G\left(x_{t}^{(i)}\right)\Delta u_{t},u_{t}^{*}-u_{t}^{(i+1)}\right\rangle+\mathcal{O}\left(\sum_{j=0}^{t-1}\Delta u_{j}\right)\left\lVert\sum_{j=0}^{t-1}\Delta u_{j}\right\rVert\right)\\ &+\sum_{t=0}^{T-1}\mathcal{O}\left(\Delta u_{t}\right)\left\lVert\Delta u_{t}\right\rVert\\ &+\sum_{t=0}^{T-1}\left(\mathcal{O}\left(\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right)\left\lVert\sum_{j=0}^{t-1}\left(u_{j}^{*}-u_{j}^{(i)}\right)\right\rVert+\mathcal{O}\left(u_{t}^{*}-u_{t}^{(i)}\right)\left\lVert u_{t}^{*}-u_{t}^{(i)}\right\rVert\right).\end{split}$
-Observe now that the term $2G\left(x_{t}^{(i)}\right)\Delta u_{t}$ is the gradient for $z_{t}^{(i)}$.
-Define $\epsilon_{2}$ such that the $\mathcal{O}$-terms are sufficiently close to zero for all $z\in B_{\epsilon_{2}}(z^{*})$.
-Thus, for $\epsilon=\min\left\{\epsilon_{1},\epsilon_{2}\right\}$ we have from lemma 11
-$\sum_{t=0}^{T-1}\left\langle 2G\left(x_{t}^{(i)}\right)\Delta u_{t},u_{t}^{*}-u_{t}^{(i+1)}\right\rangle<0,$
-since $z^{*}$ is assumed to be a strongly unique minimum point by assumptions 5. Combining the inequalities and utilizing that $E(z)$ is locally Lipschitz continuous by assumptions 5, then
-$K\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq E\left(z^{(i+1)}\right)-E\left(z^{*}\right)\leq\mathcal{O}\left(\left\lVert z^{i}-z^{*}\right\rVert^{2}\right),$
-which implies that there exists an $\epsilon>0$ such that
-$\forall z^{(i)}\in B_{\epsilon}\left(z^{*}\right):\quad\left\lVert z^{(i+1)}-z^{*}\right\rVert\leq c\left\lVert z^{(i)}-z^{*}\right\rVert^{2},$
-for $c>0$. ∎
-The assumption in the above proof that the line search parameter, $\alpha^{*}=1$ locally at the minimum point is based on that locally the approximation of the discretized energy functional becomes close to the true function.
-A.7 Extension to Finslerian manifolds
+A.2 Extension to Finslerian manifolds
 The proof for global convergence and local quadratic convergence hold similar for GEORCE in the Finsler case.
@@
 $\begin{split}{\left.\kern-1.2pt\nabla_{x_{t}}E_{F}(x_{t},u_{t0})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=\mu_{T-1}-\mu_{t},\quad t=1,\dots,T-1,\\ {\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i+1)}\right)}}&=2G\left(x_{t}^{(i)},u_{t}^{(i)}\right)u_{t}^{(i+1)}+\zeta_{t}=-\mu_{t},\\ &\quad t=0,\dots,T-1\end{split}$ (25)
-The proof of global and local quadratic convergence are completely similar for the Riemannian and Finsler case except that eq. 23 in the Finsler case becomes
+The proof of global and local quadratic convergence is completely similar for the Riemannian and Finsler case except that Eq. 14 in the Finsler case becomes
 $\begin{split}\mu_{t}+{\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i)},u_{t}^{(i)}\right)}}&=-{\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i+1)},u_{t}^{(i+1)}\right)}}\\ &+{\left.\kern-1.2pt\nabla_{u_{t}}E_{F}(x_{t},u_{t})\vphantom{\big|}\right|_{(x_{t},u_{t})=\left(x_{t}^{(i+1)},u_{t}^{(i)}\right)}}\\ &=-\left(G\left(x_{t}^{(i)},u_{t}^{(i)}\right)u_{t}^{(i+1)}+\zeta_{t}\right)\\ &+\left(2G\left(x_{t}^{(i)},u_{t}^{(i)}\right)+\zeta_{t}\right)\\ &=-2G\left(x_{t}^{(i)},u_{t}^{(i)}\right)\left(u_{t}^{(i+1)}-u_{t}^{(i)}\right),\end{split}$ (26)
@@
 3:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$, $u_{t}^{(0)}\leftarrow\frac{b-a}{T}$ for $t=0.,\dots,T$ and $i\leftarrow 0$
-4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
+4:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
 5:  $G_{t}\leftarrow G\left(x_{t}^{(i)},u_{t}^{(i)}\right)$ for $t=0,\dots,T-1$
@@
 Note that we efficiently solve the equations in Eq. 29 as
-$\begin{split}H_{ii}y_{i}&=\left(H_{ii+1}|s_{i}\right),\quad i=1,\dots,T-2,\\ H_{T-1}y_{T-1}&=s_{T-1},\\ \end{split}$ (31)
+[…]
 such that only one linear system of equations has to be solved in each iteration.
@@
 3:Set $x_{t}^{(0)}\leftarrow a+\frac{b-a}{T}t$ for $t=0.,\dots,T$ and $j\leftarrow 0$
-4:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
+4:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
 5:  Compute $\left\{s_{t}^{\mathrm{grad}},s_{t}^{\mathrm{newton}}\right\}_{t=1,\dots,T-2}$ using Eq. 29 and Eq. 30.
@@
 6:
-7:while $\ \textbf{do}\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big{|}}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$
+7:while $\left\lVert{\left.\kern-1.2pt\nabla_{y}E(y)\vphantom{\big|}\right|_{y=x_{t}^{(i)}}}\right\rVert_{2}>\mathrm{tol}$ do
 8:  $H_{ii}\leftarrow\lambda I+\partial^{2}_{x_{i}x_{i}}E(x_{0:T})$
@@
 However, they do not derive an adaptive scheme to solve the linear equation in the Newton step or propose the regularized version in Algorithm 4.
-Appendix C Manifolds
+Appendix C Choice of initialization curve
+Figure 12: The first row shows the estimated curves using GEORCE for the egg-tray, i.e.
+$f(x,y)=\left(x,y,2\cos x\cos y\right)$, while the second row shows the estimated curves for the sphere for different initializations of the initial curve.
+In this section, we show the effect of the choice of initialization curve in GEORCE.
+In Algorithm 1 and the experiments, the initial curve is a straight line between the start and end point.
+However, the initial curve is not restricted to this and can in principle be any curve assuming that it is defined in the local chart.
+To illustrate the effect of the initial curve, we will consider the egg-tray with the parametrization $f(x,y)=\left(x,y,2\cos x\cos y\right)$, and the sphere for two antipodal points on the equator.
+For the egg-tray we consider setting the initial curve randomly as
+$x_{t}^{(0)}=t\frac{b-a}{T}+a+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}\left(0,I\right),$
+while we consider for the sphere the random initial curve.
+$x_{t}^{(0)}=t\frac{b-a}{T}+a+\boldsymbol{1}\mathrm{Sign}\left(\epsilon\right),\quad\epsilon\sim\mathcal{N}\left(0,1\right),$
+For each random initial curve, we estimate the energy-minimizing curve using GEORCE and show the result in Fig.
+12 for $100$ random initializations.
+The first row shows the results for the egg-tray, where the rightmost figure is the distribution of length for the computed curves.
+We display the corresponding curves in the same color as in the histogram.
+In general, we see that for a small set of curves GEORCE has found a local minimum, where the curves move on the inside of the egg-tray rather than on the outside of the egg-tray, which appears to be the global minimum.
+This illustrates how GEORCE is guaranteed to converge to a local minimum and not a global minimum, and that this can be checked by expecting different initialization curves.
+For the sphere, there are infinitely many length-minimizing curves, and since we add the sign of a standard normal distributed variable for the random curves, we see that GEORCE tends to move in the direction of the closest local minimum.
+Appendix D Manifolds
 Table 9 contains a description of the Riemannian manifolds shown in the paper.
 Manifold Description Parameters Local Coordinates
-$\mathbb{S}^{n}$ The n-sphere, $\{x\in\mathbb{R}^{n+1}\;|\;||x||_{2}=1\}$ Stereographic coordinates
-$E(n)$ The n-Ellipsoid, $\{x\in\mathbb{R}^{n+1}\;|\;||p\odot x||_{2}=1\}$ Half-axes: $p=\left(0.5,\dots,1\right)$ equally distributed points. Stereographic coordinates
+[…]
+[…]
 $\mathbb{T}^{2}$ Torus Major radius $R=3.0$ and minor radius $r=1.0$ Parameterized by $\left(\theta,\phi\right)$ with $x=(R+r\cos\theta)\cos\phi,y=(R+r\cos\theta)\sin\phi,z=r\sin\theta$.
-$\mathbb{H}^{2}$ 2 dimensional Hyperbolic Space embedded into Minkowski pace $\mathbb{R}^{3}$ Parameterized by $\left(\alpha,\beta\right)$ with $x=\cosh\alpha,y=\sinh\alpha\cos\beta,z=\sinh\alpha\sin\beta$.
-$P(n)$ Paraboloid of dimension $n$ Parameterized by standard coordinates by $\left(x^{1},\dots,x^{n},\sum_{i=1}^{n}x_{i}^{2}\right)$
-$\mathcal{P}(n)$ Symmetric positive definite matrices of size $n^{2}$ Embedded into $\mathbb{R}^{n^{2}}$ by the mapping $f(x)=l(x)l(x)^{T}$, where $l:\mathbb{R}^{n(n+1)/2}\rightarrow\mathbb{R}^{n\times n}$ maps $x$ into a lower triangle matrix consisting of the elements in $x$.
+$\mathbb{H}^{2}$ 2 dimensional Hyperbolic Space embedded into Minkowski pace $\mathbb{R}^{3}$ - Parameterized by $\left(\alpha,\beta\right)$ with $x=\cosh\alpha,y=\sinh\alpha\cos\beta,z=\sinh\alpha\sin\beta$.
+$P(n)$ Paraboloid of dimension $n$ - Parameterized by standard coordinates by $\left(x^{1},\dots,x^{n},\sum_{i=1}^{n}x_{i}^{2}\right)$
+$\mathcal{P}(n)$ Symmetric positive definite matrices of size $n^{2}$ - Embedded into $\mathbb{R}^{n^{2}}$ by the mapping $f(x)=l(x)l(x)^{T}$, where $l:\mathbb{R}^{n(n+1)/2}\rightarrow\mathbb{R}^{n\times n}$ maps $x$ into a lower triangle matrix consisting of the elements in $x$.
 ’Egg-tray’ A parameterized surface in shape of an egg tray - Parameterized by $(x,y)$ with $x=x,y=y,z=2\cos x\cos y$.
@@
 where $g_{ij}$ denotes the elements of the Riemannian background metric, $f^{j}$ denotes the $j$th element of the force field, while $v_{0}$ denotes the initial velocity Riemannian length of an object moving with constant velocity $v$ on the surface.
-For the manifold to be a proper Finsler manifold it is required that $||f||_{g}<v_{0}$.
-Appendix D Variational autoencoders
-D.1 Architecture
-We train a Variational-Autoencoer (VAE) [52], where we use the normal distribution as a variational family to approximate the data distribution. We train a VAE for each of the following datasets
+For the manifold to be a proper Finsler manifold it is required that $f||_{g}<v_{0}$.
+Appendix E Variational autoencoders
+E.1 Architecture
+We train a Variational-Autoencoder (VAE) [52], where we use the normal distribution as a variational family to approximate the data distribution. We train a VAE for each of the following datasets
   • MNIST data [56]: The data consists of $28\times 28$ images of handwritten digits between 0 and 9. The data consists of $60,000$ training images and $10,000$ testing images.
@@
 The architecture is similar to [12] with some modifications.
-All models are trained for $50,000$ epochs on a GPU described in Appendix E.2.
+All models are trained for $50,000$ epochs on a GPU described in Appendix F.2.
 Encoder
@@
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=1,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=16,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=1,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
@@
 Decoder
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=3,\mathrm{kernel\_shape}=2\times 2,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 Table 11: Architecture for VAE for the SVHN-dataset
@@
 Decoder
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=64,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=32,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=2,\mathrm{bias}=\mathrm{False}\right)$
 $\mathrm{GELU}$
-[…]
+$\mathrm{Conv2dTransposed}\left(\mathrm{output\_channels}=3,\mathrm{kernel\_shape}=4\times 4,\mathrm{stride}=4,\mathrm{bias}=\mathrm{False}\right)$
 Table 12: Architecture for VAE for the CelebA-dataset
-D.2 Additional experiments
+E.2 Additional experiments
 In Fig.
@@
 The details of the manifolds and experiments can be found in the supplementary material.
-Appendix E Experiments
+Appendix F Experiments
 The following contains details of hyper-parameters, hardware and additional experiments. The implementation and code can be found at
 https://github.com/FrederikMR/georce.
-E.1 Hyper-parameters and methods
+F.1 Hyper-parameters and methods
 Table 13 shows the hyper parameters used for the different methods. It is the same hyper-parameters that have been used for all manifolds across dimension and grid size in the paper.
@@
 Manifold Start point End point
-$\mathbb{S}^{n}$ $n$ equally spaced points between 00 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
-$E(n)$ $n$ equally spaced points between 00 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
+$\mathbb{S}^{n}$ $n$ equally spaced points between 0 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
+$E(n)$ $n$ equally spaced points between 0 and $1$ (end point is excluded) $\left(0.5,0.5,\dots,0.5\right)$
 $\mathbb{T}^{2}$ $(0.0,0.0)$ $(5\pi/4,5\pi/4)$
@@
 $\mathcal{P}(n)$ The identity matrix in local coordinates $n(n+1)/2$ equally spaced points between $0.5$ and $1.0$.
-’Egg-tray’ $(-5.0,5.0)$ $(5.0,5.0)$
+’Egg-tray’ (5.0,5.0)(-5.0,5.0) $(5.0,5.0)$
 Gaussian Distribution $\left(\mu_{0},\sigma_{0}\right)=\left(-1.0,0.5\right)$ $\left(\mu_{T},\sigma_{T}\right)=\left(1.0,1.0\right)$
@@
 Table 15: The start and end point for the geodesics estimated for each manifold.
-E.2 Hardware
-All figures have been computed on a HP computer with Intel Core i9-11950H 2.6 GHz 8C, 15.6” FHD, 720P CAM, 32 GB (2$\times$16GB) DDR4 3200 So-Dimm, Nvidia Quadro TI2004GB Descrete Graphics, 1TB PCle NVMe SSD, backlit Keyboard, 150W PSU, 8cell, W11Home 64 Advanced, 4YR Onsite NBD.
-The runtime experiments on a CPU have been run on a CPU model XeonE5_2660v3 with $4$ nodes for at most $24$ hours with a maximum memory of $10$ GB. .
-The runtime estimates on a GPU and training of the VAE have been computed on a GPU for at most 24 hours with a maximum memory of $10$ GB. The $GPU$ consists of $4$ nodes on a Tesla V100.
-E.3 Additional experiments
-In the following we show run time estimates computed similar to the description in the main paper with other methods than BFGS and ADAM.
-It should be noted that a gradient method with exact line search has also been tested against GEORCE, but shows similar computational behavior as ADAM, but often with higher computational times and therefore not included in the following.
-We also show the results on a GPU for CUDA compatible methods.
-Note that for the Finsler manifold for the ellipsoid of dimension $50$ with a grid size of $T=50$ points, none of the methods converge including GEORCE.
-This is due to the strong variation by the force field causing numerical instability.
-However, when the grid size is increased to $T=100$, GEORCE converges to the best solution across all algorithms indicating that it requires a finer grid for this example.
-E.4 Run-Time Estimates on VAE Manifolds
-We show in Table 11 the run-time estimates on learned manifolds using a VAE.
-BFGS (T=100) ADAM (T=100) GEORCE (T=100)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-VAE MNIST $10.3075$ $1407.5743\pm 6.4430$ $10.3312$ $645.5201\pm 2.6648$ $\boldsymbol{10.3074}$ $\boldsymbol{67.6149}\pm\boldsymbol{0.0304}$
-VAE SVHN -- -- $19.9855$ $13325.2285\pm 123.4978$ $\boldsymbol{19.9346}$ $\boldsymbol{882.4120}\pm\boldsymbol{1.8839}$
-VAE CelebA -- -- -- -- $\boldsymbol{42.2666}$ $\boldsymbol{2222.5259}\pm\boldsymbol{8.2988}$
-Table 11: The table shows the length of the estimated geodesics for the BFGS-algorithm, ADAM and GEORCE on a CPU.
-The methods were terminated if the $\ell^{2}$-norm of the gradient was less than $10^{-3}$ or $1,000$ iterations have been reached.
-When the computational time was longer than $24$ hour, the value is set to --.
-E.4.1 Additional Run-time estimates on a GPU
-We show in Table 12 and Table 13 the runtime of GPU compatible methods for different Riemannian manifolds, and in Table 14 and Table 15 the runtime of GPU compatible methods for different Finsler manifolds.
-ADAM (T=100) Sparse Newton (T=100) GEORCE (T=100)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2755$ $0.2729\pm 0.0016$ $2.2742$ $0.1417\pm 0.0041$ $\boldsymbol{2.2741}$ $\boldsymbol{0.0163}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{3}$ $2.4376$ $0.2890\pm 0.0005$ $\boldsymbol{2.4342}$ $0.2188\pm 0.0039$ $2.4342$ $\boldsymbol{0.0735}\pm\boldsymbol{0.0018}$
-$\mathbb{S}^{5}$ $2.4226$ $0.3153\pm 0.0027$ $\boldsymbol{2.3755}$ $0.8657\pm 0.0015$ $2.3755$ $\boldsymbol{0.0494}\pm\boldsymbol{0.0014}$
-$\mathbb{S}^{10}$ $2.1031$ $0.3052\pm 0.0079$ $1.9653$ $117.4589\pm 0.0140$ $\boldsymbol{1.9523}$ $\boldsymbol{0.0351}\pm\boldsymbol{0.0001}$
-$\mathbb{S}^{20}$ $1.6824$ $0.3401\pm 0.0117$ $1.7069$ $146.3024\pm 0.0191$ $\boldsymbol{1.4808}$ $\boldsymbol{0.0483}\pm\boldsymbol{0.0020}$
-$\mathbb{S}^{50}$ $1.1456$ $0.5339\pm 0.0014$ $1.0359$ $310.2160\pm 0.0200$ $\boldsymbol{0.9749}$ $\boldsymbol{0.0587}\pm\boldsymbol{0.0023}$
-$\mathbb{S}^{100}$ $0.8475$ $1.2935\pm 0.0014$ $0.7522$ $6.6464\pm 0.0016$ $\boldsymbol{0.4188}$ $\boldsymbol{0.0717}\pm\boldsymbol{0.0022}$
-$\mathbb{S}^{250}$ $0.6251$ $3.8386\pm 0.0018$ $0.4958$ $98.6502\pm 0.0147$ $\boldsymbol{0.3192}$ $\boldsymbol{0.3296}\pm\boldsymbol{0.0005}$
-$\mathbb{S}^{500}$ $0.5714$ $8.2031\pm 0.0030$ $0.3787$ $899.4549\pm 0.0188$ $\boldsymbol{0.2198}$ $\boldsymbol{1.2925}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{1000}$ $0.5270$ $19.2234\pm 0.0712$- -- $\boldsymbol{0.1573}$ $\boldsymbol{6.8579}\pm\boldsymbol{0.0023}$
-$1.9595$ $0.2735\pm 0.0074$ $\boldsymbol{1.9595}$ $0.2177\pm 0.0075$ $1.9595$ $\boldsymbol{0.0138}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(3\right)$ $2.0523$ $0.2839\pm 0.0056$ $\boldsymbol{2.0498}$ $0.3705\pm 0.0063$ $2.0501$ $\boldsymbol{0.0167}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(5\right)$ $1.9594$ $0.2948\pm 0.0116$- -- $\boldsymbol{1.8410}$ $\boldsymbol{0.0827}\pm\boldsymbol{0.0036}$
-$\mathrm{E}\left(10\right)$ $1.6421$ $0.3097\pm 0.0133$ $\boldsymbol{1.4247}$ $116.6572\pm 0.0186$ $1.5139$ $\boldsymbol{0.0547}\pm\boldsymbol{0.0019}$
-$\mathrm{E}\left(20\right)$ $1.3052$ $0.3875\pm 0.0058$ $1.4583$ $148.6781\pm 0.0307$ $\boldsymbol{0.6514}$ $\boldsymbol{0.1449}\pm\boldsymbol{0.0054}$
-$\mathrm{E}\left(50\right)$ $0.9245$ $0.5704\pm 0.0013$ $0.8849$ $316.9808\pm 0.0179$ $\boldsymbol{0.4642}$ $\boldsymbol{0.0370}\pm\boldsymbol{0.0014}$
-$\mathrm{E}\left(100\right)$ $0.6873$ $1.3063\pm 0.0020$ $0.5922$ $5.2868\pm 0.0007$ $\boldsymbol{0.3249}$ $\boldsymbol{0.0524}\pm\boldsymbol{0.0002}$
-$\mathrm{E}\left(250\right)$ $0.4974$ $4.2084\pm 0.0029$ $0.3917$ $84.6331\pm 0.0439$ $\boldsymbol{0.2180}$ $\boldsymbol{0.2325}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(500\right)$ $0.4559$ $7.4772\pm 0.0024$ $0.3039$ $827.7163\pm 0.1003$ $\boldsymbol{0.1468}$ $\boldsymbol{1.2897}\pm\boldsymbol{0.0004}$
-$\mathrm{E}\left(1000\right)$ $0.4238$ $19.8033\pm 0.0970$- -- $\boldsymbol{0.1129}$ $\boldsymbol{5.1862}\pm\boldsymbol{0.0071}$
-$\mathbb{T}^{2}$ $9.9853$ $0.2890\pm 0.0016$ $9.9851$ $0.3353\pm 0.0126$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0549}\pm\boldsymbol{0.0010}$
-$\mathbb{H}^{2}$ $1.2500$ $0.3201\pm 0.0004$ $1.2498$ $0.2125\pm 0.0128$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0089}\pm\boldsymbol{0.0003}$
-$\mathcal{P}\left(2\right)$ $1.0786$ $0.3801\pm 0.0047$ $\boldsymbol{1.0669}$ $0.2419\pm 0.0069$ $1.0669$ $\boldsymbol{0.0227}\pm\boldsymbol{0.0010}$
-$\mathcal{P}\left(3\right)$ $2.3006$ $0.4115\pm 0.0009$ $2.3810$ $106.2814\pm 0.0135$ $\boldsymbol{2.2639}$ $\boldsymbol{0.0343}\pm\boldsymbol{0.0004}$
-Gaussian Distribution $\boldsymbol{2.5951}$ $0.1809\pm 0.0010$ $2.5952$ $0.1108\pm 0.0055$ $2.5952$ $\boldsymbol{0.0123}\pm\boldsymbol{0.0002}$
-Fréchet Distribution $\boldsymbol{1.1538}$ $0.1778\pm 0.0029$ $1.1539$ $0.0813\pm 0.0044$ $1.1539$ $\boldsymbol{0.0055}\pm\boldsymbol{0.0003}$
-Cauchy Distribution $1.6452$ $0.1818\pm 0.0007$ $1.6453$ $0.1136\pm 0.0067$ $\boldsymbol{1.6452}$ $\boldsymbol{0.0137}\pm\boldsymbol{0.0001}$
-Pareto Distribution $0.8314$ $0.1779\pm 0.0026$ $\boldsymbol{0.8297}$ $0.0758\pm 0.0036$ $0.8297$ $\boldsymbol{0.0043}\pm\boldsymbol{0.0002}$
-Table 12: The table shows the length of the estimated geodesics for different methods on a GPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-ADAM (T=50) Sparse Newton (T=50) GEORCE (T=50)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2479$ $0.2683\pm 0.0082$ $2.2478$ $0.1094\pm 0.0047$ $\boldsymbol{2.2477}$ $\boldsymbol{0.0187}\pm\boldsymbol{0.0002}$
-$\mathbb{S}^{3}$ $2.4069$ $0.2593\pm 0.0110$ $\boldsymbol{2.4064}$ $0.1370\pm 0.0082$ $2.4064$ $\boldsymbol{0.0866}\pm\boldsymbol{0.0031}$
-$\mathbb{S}^{5}$ $2.3494$ $0.2907\pm 0.0103$ $2.3480$ $0.3153\pm 0.0063$ $\boldsymbol{2.3480}$ $\boldsymbol{0.0622}\pm\boldsymbol{0.0019}$
-$\mathbb{S}^{10}$ $1.9470$ $0.3026\pm 0.0014$ $1.9274$ $0.8966\pm 0.0034$ $\boldsymbol{1.9272}$ $\boldsymbol{0.0540}\pm\boldsymbol{0.0016}$
-$\mathbb{S}^{20}$ $1.5036$ $0.3349\pm 0.0069$ $1.4610$ $28.7634\pm 0.0016$ $\boldsymbol{1.4566}$ $\boldsymbol{0.0670}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{50}$ $1.0320$ $0.4713\pm 0.0025$ $1.0060$ $39.8257\pm 0.0010$ $\boldsymbol{0.5279}$ $\boldsymbol{0.0438}\pm\boldsymbol{0.0002}$
-$\mathbb{S}^{100}$ $0.7642$ $0.8173\pm 0.0009$ $0.6922$ $4.2461\pm 0.0007$ $\boldsymbol{0.3932}$ $\boldsymbol{0.0704}\pm\boldsymbol{0.0027}$
-$\mathbb{S}^{250}$ $0.5136$ $3.2362\pm 0.0009$ $0.4589$ $60.6258\pm 0.0179$ $\boldsymbol{0.2748}$ $\boldsymbol{0.1525}\pm\boldsymbol{0.0002}$
-$\mathbb{S}^{500}$ $0.4253$ $7.7012\pm 0.0038$ $0.3278$ $706.9603\pm 0.0293$ $\boldsymbol{0.1858}$ $\boldsymbol{0.7340}\pm\boldsymbol{0.0006}$
-$\mathbb{S}^{1000}$ $0.3790$ $20.5850\pm 0.0098$- -- $\boldsymbol{0.1626}$ $\boldsymbol{3.1229}\pm\boldsymbol{0.0007}$
-$\mathrm{E}\left(2\right)$ $1.9331$ $0.2723\pm 0.0013$ $\boldsymbol{1.9330}$ $0.1365\pm 0.0083$ $1.9331$ $\boldsymbol{0.0155}\pm\boldsymbol{0.0003}$
-$\mathrm{E}\left(3\right)$ $2.0243$ $0.2852\pm 0.0017$ $\boldsymbol{2.0233}$ $0.1828\pm 0.0074$ $2.0235$ $\boldsymbol{0.0305}\pm\boldsymbol{0.0012}$
-$\mathrm{E}\left(5\right)$ $1.8262$ $0.2996\pm 0.0079$ $\boldsymbol{1.8047}$ $0.6859\pm 0.0047$ $1.8054$ $\boldsymbol{0.1172}\pm\boldsymbol{0.0050}$
-$\mathrm{E}\left(10\right)$ $1.5193$ $0.3094\pm 0.0086$ $\boldsymbol{0.9472}$ $57.6302\pm 0.0107$ $1.4788$ $\boldsymbol{0.0855}\pm\boldsymbol{0.0036}$
-$\mathrm{E}\left(20\right)$ $1.2028$ $0.3379\pm 0.0049$ $\boldsymbol{0.7603}$ $48.7781\pm 0.0029$ $1.1366$ $\boldsymbol{0.0727}\pm\boldsymbol{0.0033}$
-$\mathrm{E}\left(50\right)$ $0.8382$ $0.4596\pm 0.0092$ $0.6157$ $153.9548\pm 0.0165$ $\boldsymbol{0.4417}$ $\boldsymbol{0.0284}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(100\right)$ $0.6175$ $0.8219\pm 0.0036$ $0.5822$ $3.9507\pm 0.0014$ $\boldsymbol{0.3160}$ $\boldsymbol{0.0373}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(250\right)$ $0.4216$ $2.4940\pm 0.0028$ $0.3882$ $48.5547\pm 0.0120$ $\boldsymbol{0.2039}$ $\boldsymbol{0.1519}\pm\boldsymbol{0.0003}$
-$\mathrm{E}\left(500\right)$ $0.3454$ $5.5967\pm 0.0145$- -- $\boldsymbol{0.1434}$ $\boldsymbol{0.7326}\pm\boldsymbol{0.0005}$
-$\mathrm{E}\left(1000\right)$ $0.3056$ $14.1006\pm 0.0361$- -- $\boldsymbol{0.0996}$ $\boldsymbol{4.1512}\pm\boldsymbol{0.0043}$
-$\mathbb{T}^{2}$ $9.8952$ $0.2821\pm 0.0029$ $\boldsymbol{9.8943}$ $38.0816\pm 0.0018$ $9.8955$ $\boldsymbol{0.0531}\pm\boldsymbol{0.0013}$
-$\mathbb{H}^{2}$ $1.2435$ $0.3072\pm 0.0003$ $\boldsymbol{1.2434}$ $0.1172\pm 0.0048$ $1.2434$ $\boldsymbol{0.0082}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(2\right)$ $1.0704$ $0.3571\pm 0.0072$ $\boldsymbol{1.0567}$ $0.1231\pm 0.0051$ $1.0567$ $\boldsymbol{0.0243}\pm\boldsymbol{0.0002}$
-$\mathcal{P}\left(3\right)$ $2.2400$ $0.3924\pm 0.0057$ $2.3328$ $57.1278\pm 0.0125$ $\boldsymbol{2.2304}$ $\boldsymbol{0.0198}\pm\boldsymbol{0.0001}$
-Gaussian Distribution $2.5779$ $0.1785\pm 0.0003$ $\boldsymbol{2.5778}$ $0.0617\pm 0.0027$ $2.5778$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
-Fréchet Distribution $1.1471$ $0.1696\pm 0.0004$ $\boldsymbol{1.1457}$ $0.0442\pm 0.0016$ $1.1457$ $\boldsymbol{0.0063}\pm\boldsymbol{0.0002}$
-Cauchy Distribution $1.6341$ $0.0230\pm 0.0002$ $1.6340$ $0.0810\pm 0.0037$ $\boldsymbol{1.6340}$ $\boldsymbol{0.0135}\pm\boldsymbol{0.0003}$
-Pareto Distribution $0.8235$ $0.0357\pm 0.0003$ $0.8235$ $0.0439\pm 0.0014$ $\boldsymbol{0.8234}$ $\boldsymbol{0.0052}\pm\boldsymbol{0.0003}$
-Table 13: The table shows the length of the estimated geodesics for different methods on a GPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-ADAM (T=100) Sparse Newton (T=100) GEORCE (T=100)
-Finsler Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.6530$ $0.7706\pm 0.0018$ $\boldsymbol{1.6525}$ $0.2512\pm 0.0089$ $1.6525$ $\boldsymbol{0.0372}\pm\boldsymbol{0.0004}$
-$\mathbb{S}^{3}$ $1.7451$ $0.7362\pm 0.0036$- -- $\boldsymbol{1.7432}$ $\boldsymbol{0.0431}\pm\boldsymbol{0.0005}$
-$\mathbb{S}^{5}$ $1.7371$ $0.9292\pm 0.0038$- -- $\boldsymbol{1.7029}$ $\boldsymbol{0.1036}\pm\boldsymbol{0.0048}$
-$\mathbb{S}^{10}$ $1.5174$ $1.1358\pm 0.0018$ $1.4783$ $208.1472\pm 0.0176$ $\boldsymbol{1.4109}$ $\boldsymbol{0.1000}\pm\boldsymbol{0.0033}$
-$\mathbb{S}^{20}$ $1.2009$ $1.5876\pm 0.0028$- -- $\boldsymbol{1.0760}$ $\boldsymbol{0.1509}\pm\boldsymbol{0.0058}$
-$\mathbb{S}^{50}$ $0.8130$ $2.0381\pm 0.0026$- -- $\boldsymbol{0.7283}$ $\boldsymbol{0.1373}\pm\boldsymbol{0.0032}$
-$\mathbb{S}^{100}$ $0.6146$ $2.7712\pm 0.0032$ $0.6210$ $972.0195\pm 0.0307$ $\boldsymbol{0.2553}$ $\boldsymbol{0.1106}\pm\boldsymbol{0.0028}$
-$\mathbb{S}^{250}$ $0.5196$ $3.9803\pm 0.0019$- -- $\boldsymbol{0.1636}$ $\boldsymbol{0.5069}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{500}$ $0.4849$ $6.6164\pm 0.0023$- -- $\boldsymbol{0.1417}$ $\boldsymbol{1.6591}\pm\boldsymbol{0.0004}$
-$\mathbb{S}^{1000}$ $0.3835$ $25.4638\pm 0.1448$- -- $\boldsymbol{0.1134}$ $\boldsymbol{7.9991}\pm\boldsymbol{0.0074}$
-$\mathrm{E}\left(2\right)$ $1.3634$ $0.7758\pm 0.0036$ $\boldsymbol{1.3616}$ $0.3735\pm 0.0010$ $1.3618$ $\boldsymbol{0.0282}\pm\boldsymbol{0.0009}$
-$\mathrm{E}\left(3\right)$ $1.4085$ $0.7541\pm 0.0032$- -- $\boldsymbol{1.4023}$ $\boldsymbol{0.1273}\pm\boldsymbol{0.0047}$
-$\mathrm{E}\left(5\right)$ $1.3305$ $1.0252\pm 0.0017$ $1.5359$ $170.9661\pm 0.0296$ $\boldsymbol{1.2956}$ $\boldsymbol{0.1135}\pm\boldsymbol{0.0039}$
-$\mathrm{E}\left(10\right)$ $1.1235$ $1.1254\pm 0.0045$- -- $\boldsymbol{1.0804}$ $\boldsymbol{0.0825}\pm\boldsymbol{0.0031}$
-$\mathrm{E}\left(20\right)$ $0.8987$ $1.6039\pm 0.0035$- -- $\boldsymbol{0.8121}$ $\boldsymbol{0.1282}\pm\boldsymbol{0.0054}$
-$\mathrm{E}\left(50\right)$ $0.6332$ $2.0424\pm 0.0034$- -- $\boldsymbol{0.2845}$ $\boldsymbol{0.0770}\pm\boldsymbol{0.0004}$
-$\mathrm{E}\left(100\right)$ $0.4953$ $2.6704\pm 0.0031$ $0.4577$ $470.9805\pm 0.1448$ $\boldsymbol{0.2018}$ $\boldsymbol{0.1127}\pm\boldsymbol{0.0023}$
-$\mathrm{E}\left(250\right)$ $0.4179$ $4.0843\pm 0.0009$- -- $\boldsymbol{0.1287}$ $\boldsymbol{0.5239}\pm\boldsymbol{0.0047}$
-$\mathrm{E}\left(500\right)$ $0.4139$ $6.9391\pm 0.0073$ $0.2580$ $1435.3881\pm 0.0832$ $\boldsymbol{0.0968}$ $\boldsymbol{1.7109}\pm\boldsymbol{0.0010}$
-$\mathrm{E}\left(1000\right)$ $0.3442$ $25.9660\pm 0.1038$- -- $\boldsymbol{0.0798}$ $\boldsymbol{7.9988}\pm\boldsymbol{0.0079}$
-$\mathbb{T}^{2}$ $7.4054$ $0.8017\pm 0.0045$ $9.1903$ $152.0036\pm 0.0158$ $\boldsymbol{6.8067}$ $\boldsymbol{0.3145}\pm\boldsymbol{0.0050}$
-$\mathbb{H}^{2}$ $1.3341$ $0.7949\pm 0.0072$ $1.3338$ $0.4562\pm 0.0004$ $\boldsymbol{1.3337}$ $\boldsymbol{0.0256}\pm\boldsymbol{0.0006}$
-$\mathcal{P}\left(2\right)$ $0.7334$ $0.9344\pm 0.0013$ $\boldsymbol{0.7291}$ $0.4309\pm 0.0014$ $0.7292$ $\boldsymbol{0.0286}\pm\boldsymbol{0.0005}$
-$\mathcal{P}\left(3\right)$ $1.2987$ $1.0812\pm 0.0006$ $1.2856$ $0.6885\pm 0.0009$ $\boldsymbol{1.2856}$ $\boldsymbol{0.0365}\pm\boldsymbol{0.0005}$
-Gaussian Distribution $1.6879$ $0.4962\pm 0.0034$ $1.6878$ $0.4148\pm 0.0055$ $\boldsymbol{1.6878}$ $\boldsymbol{0.0255}\pm\boldsymbol{0.0004}$
-Fréchet Distribution $0.5406$ $0.2307\pm 0.0069$ $\boldsymbol{0.5405}$ $0.2060\pm 0.0042$ $0.5405$ $\boldsymbol{0.0088}\pm\boldsymbol{0.0003}$
-Cauchy Distribution $1.1140$ $0.4994\pm 0.0017$ $\boldsymbol{1.1133}$ $0.5167\pm 0.0042$ $1.1133$ $\boldsymbol{0.0367}\pm\boldsymbol{0.0017}$
-Pareto Distribution $0.3885$ $0.1014\pm 0.0038$ $0.3884$ $0.1042\pm 0.0031$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0096}\pm\boldsymbol{0.0004}$
-Table 14: The table shows the length of the estimated geodesics for different methods on a GPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-ADAM (T=50) Sparse Newton (T=50) GEORCE (T=50)
-Finsler Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $\boldsymbol{1.6383}$ $0.7833\pm 0.0056$ $1.6385$ $0.2064\pm 0.0073$ $1.6385$ $\boldsymbol{0.0457}\pm\boldsymbol{0.0004}$
-$\mathbb{S}^{3}$ $1.7300$ $0.7587\pm 0.0049$- -- $\boldsymbol{1.7297}$ $\boldsymbol{0.0499}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{5}$ $1.6875$ $0.9909\pm 0.0022$- -- $\boldsymbol{1.6868}$ $\boldsymbol{0.1432}\pm\boldsymbol{0.0050}$
-$\mathbb{S}^{10}$ $1.4013$ $1.1472\pm 0.0024$- -- $\boldsymbol{1.3908}$ $\boldsymbol{0.1747}\pm\boldsymbol{0.0067}$
-$\mathbb{S}^{20}$ $1.0829$ $1.5583\pm 0.0016$- -- $\boldsymbol{1.0560}$ $\boldsymbol{0.2663}\pm\boldsymbol{0.0005}$
-$\mathbb{S}^{50}$ $0.7427$ $1.4442\pm 0.0038$- -- $\boldsymbol{0.3508}$ $\boldsymbol{0.0594}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{100}$ $0.5511$ $2.5558\pm 0.0020$- -- $\boldsymbol{0.2626}$ $\boldsymbol{0.0850}\pm\boldsymbol{0.0024}$
-$\mathbb{S}^{250}$ $0.4094$ $3.8872\pm 0.0020$- -- $\boldsymbol{0.1686}$ $\boldsymbol{0.2660}\pm\boldsymbol{0.0042}$
-$\mathbb{S}^{500}$ $0.3531$ $7.5214\pm 0.0022$ $0.2861$ $1388.8278\pm 0.0965$ $\boldsymbol{0.1193}$ $\boldsymbol{1.1879}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{1000}$ $0.3269$ $13.1209\pm 0.0379$- -- $\boldsymbol{0.0828}$ $\boldsymbol{6.0238}\pm\boldsymbol{0.0039}$
-$\mathrm{E}\left(2\right)$ $1.3460$ $0.1327\pm 0.0060$ $\boldsymbol{1.3460}$ $0.2847\pm 0.0043$ $1.3461$ $\boldsymbol{0.0332}\pm\boldsymbol{0.0011}$
-$\mathrm{E}\left(3\right)$ $1.3870$ $0.7621\pm 0.0034$- -- $\boldsymbol{1.3826}$ $\boldsymbol{0.2156}\pm\boldsymbol{0.0006}$
-$\mathrm{E}\left(5\right)$ $1.2780$ $0.9718\pm 0.0022$- -- $\boldsymbol{1.2757}$ $\boldsymbol{0.1654}\pm\boldsymbol{0.0050}$
-$\mathrm{E}\left(10\right)$ $1.0665$ $1.1072\pm 0.0025$- -- $\boldsymbol{1.0126}$ $\boldsymbol{0.2215}\pm\boldsymbol{0.0062}$
-$\mathrm{E}\left(20\right)$ $0.8405$ $1.5702\pm 0.0031$- -- $\boldsymbol{0.7946}$ $\boldsymbol{0.1937}\pm\boldsymbol{0.0061}$
-$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5797}$ $1.4336\pm 0.0029$- -- -- --
-$\mathrm{E}\left(100\right)$ $0.4334$ $2.0750\pm 0.0026$- -- $\boldsymbol{0.2046}$ $\boldsymbol{0.0830}\pm\boldsymbol{0.0004}$
-$\mathrm{E}\left(250\right)$ $0.3234$ $2.8610\pm 0.0021$- -- $\boldsymbol{0.1365}$ $\boldsymbol{0.2644}\pm\boldsymbol{0.0029}$
-$\mathrm{E}\left(500\right)$ $0.2862$ $5.1735\pm 0.0122$ $0.2428$ $905.0696\pm 0.0365$ $\boldsymbol{0.0940}$ $\boldsymbol{0.9444}\pm\boldsymbol{0.0011}$
-$\mathrm{E}\left(1000\right)$ $0.2723$ $10.4951\pm 0.0347$- -- $\boldsymbol{0.0893}$ $\boldsymbol{4.4688}\pm\boldsymbol{0.0060}$
-$\boldsymbol{7.2683}$ $\boldsymbol{0.7530}\pm\boldsymbol{0.0015}$ $9.0475$ $78.7132\pm 0.0222$ $9.0475$ $13.7393\pm 0.0093$
-$\mathbb{H}^{2}$ $1.3328$ $0.7454\pm 0.0052$ $1.3316$ $0.2303\pm 0.0058$ $\boldsymbol{1.3316}$ $\boldsymbol{0.0317}\pm\boldsymbol{0.0003}$
-$\mathcal{P}\left(2\right)$ $0.7249$ $0.5927\pm 0.0008$ $\boldsymbol{0.7249}$ $0.2189\pm 0.0051$ $0.7249$ $\boldsymbol{0.0331}\pm\boldsymbol{0.0011}$
-$\mathcal{P}\left(3\right)$ $1.2710$ $1.0184\pm 0.0008$ $1.2707$ $0.4325\pm 0.0020$ $\boldsymbol{1.2707}$ $\boldsymbol{0.0434}\pm\boldsymbol{0.0006}$
-Gaussian Distribution $\boldsymbol{1.6796}$ $0.4337\pm 0.0111$ $1.6806$ $0.2081\pm 0.0047$ $1.6806$ $\boldsymbol{0.0282}\pm\boldsymbol{0.0017}$
-Fréchet Distribution $0.5370$ $0.0829\pm 0.0032$ $0.5370$ $0.1059\pm 0.0042$ $\boldsymbol{0.5370}$ $\boldsymbol{0.0111}\pm\boldsymbol{0.0003}$
-Cauchy Distribution $1.1085$ $0.0626\pm 0.0010$ $\boldsymbol{1.1084}$ $0.2270\pm 0.0081$ $1.1084$ $\boldsymbol{0.0410}\pm\boldsymbol{0.0015}$
-Pareto Distribution $0.3857$ $0.0752\pm 0.0025$ $\boldsymbol{0.3856}$ $0.1054\pm 0.0028$ $0.3857$ $\boldsymbol{0.0082}\pm\boldsymbol{0.0003}$
-Table 15: The table shows the length of the estimated geodesics for different methods on a GPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-E.4.2 Additional Run-time estimates on a CPU
-Methods on a Riemannian manifold
-We show in Table 16, Table 17, Table 18, Table 19, Table 20, Table 21, Table 22 and Table 23 runtime estimates for different methods on a CPU for the different Riemannian manifolds.
-BFGS (T=100) ADAM (T=100) GEORCE (T=100)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2742$ $57.4164\pm 0.0616$ $2.2772$ $0.2335\pm 0.0005$ $\boldsymbol{2.2741}$ $\boldsymbol{0.0078}\pm\boldsymbol{0.0000}$
-$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $67.4657\pm 0.4874$ $2.4377$ $0.3475\pm 0.0010$ $2.4342$ $\boldsymbol{0.0635}\pm\boldsymbol{0.0001}$
-$\mathbb{S}^{5}$ $2.3757$ $77.6975\pm 0.0741$ $2.4222$ $1.9894\pm 0.0188$ $\boldsymbol{2.3755}$ $\boldsymbol{0.2219}\pm\boldsymbol{0.0001}$
-$\mathbb{S}^{10}$ $1.9529$ $150.0621\pm 0.8919$ $2.1016$ $4.4208\pm 0.0308$ $\boldsymbol{1.9523}$ $\boldsymbol{0.3184}\pm\boldsymbol{0.0002}$
-$\mathbb{S}^{20}$ $1.4810$ $1068.8645\pm 17.2001$ $1.6839$ $8.7097\pm 0.0352$ $\boldsymbol{1.4808}$ $\boldsymbol{0.6732}\pm\boldsymbol{0.0204}$
-$\mathbb{S}^{50}$ $1.0464$ $7350.6001\pm 244.3737$ $1.1463$ $61.8926\pm 0.3034$ $\boldsymbol{0.9749}$ $\boldsymbol{4.1016}\pm\boldsymbol{0.0201}$
-$\mathbb{S}^{100}$- -- $0.8466$ $277.4217\pm 0.2280$ $\boldsymbol{0.4187}$ $\boldsymbol{10.9208}\pm\boldsymbol{0.0071}$
-$\mathbb{S}^{250}$- -- $0.6244$ $1424.9031\pm 22.1592$ $\boldsymbol{0.3192}$ $\boldsymbol{78.2937}\pm\boldsymbol{0.0991}$
-$\mathbb{S}^{500}$- -- $0.5717$ $2242.4697\pm 17.9769$ $\boldsymbol{0.2198}$ $\boldsymbol{224.2761}\pm\boldsymbol{0.4894}$
-$\mathbb{S}^{1000}$- -- $0.5279$ $4329.6621\pm 38.0163$ $\boldsymbol{0.1573}$ $\boldsymbol{893.4682}\pm\boldsymbol{7.3130}$
-$\mathrm{E}\left(2\right)$ $1.9595$ $56.9650\pm 0.4106$ $1.9596$ $0.2328\pm 0.0001$ $\boldsymbol{1.9595}$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0499}$ $94.1769\pm 0.7044$ $2.0525$ $0.3392\pm 0.0024$ $2.0501$ $\boldsymbol{0.0119}\pm\boldsymbol{0.0000}$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8403}$ $194.5403\pm 1.3293$ $1.9521$ $2.2227\pm 0.0186$ $1.8410$ $\boldsymbol{0.4220}\pm\boldsymbol{0.0038}$
-$\mathrm{E}\left(10\right)$ $\boldsymbol{1.5114}$ $427.6278\pm 4.1044$ $1.6384$ $4.5697\pm 0.0541$ $1.5139$ $\boldsymbol{0.4566}\pm\boldsymbol{0.0016}$
-$\mathrm{E}\left(20\right)$ $1.1795$ $1669.2692\pm 10.2877$ $1.3049$ $9.4335\pm 0.0507$ $\boldsymbol{0.6548}$ $\boldsymbol{1.9786}\pm\boldsymbol{0.0031}$
-$\mathrm{E}\left(50\right)$ $0.8338$ $10685.6133\pm 321.6447$ $0.9244$ $73.5682\pm 0.3306$ $\boldsymbol{0.4642}$ $\boldsymbol{2.4208}\pm\boldsymbol{0.0222}$
-$\mathrm{E}\left(100\right)$- -- $0.6872$ $289.4196\pm 2.5224$ $\boldsymbol{0.3249}$ $\boldsymbol{7.7128}\pm\boldsymbol{0.0714}$
-$\mathrm{E}\left(250\right)$- -- $0.4978$ $1620.6428\pm 16.9546$ $\boldsymbol{0.2178}$ $\boldsymbol{56.8527}\pm\boldsymbol{0.1526}$
-$\mathrm{E}\left(500\right)$- -- $0.4559$ $2135.6899\pm 30.5699$ $\boldsymbol{0.1467}$ $\boldsymbol{220.1656}\pm\boldsymbol{0.9902}$
-$\mathrm{E}\left(1000\right)$- -- $0.4232$ $4398.9453\pm 25.8738$ $\boldsymbol{0.1131}$ $\boldsymbol{641.7199}\pm\boldsymbol{20.9901}$
-$\mathbb{T}^{2}$ $9.9851$ $92.7400\pm 3.1267$ $9.9855$ $0.1060\pm 0.0002$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0182}\pm\boldsymbol{0.0003}$
-$\mathbb{H}^{2}$ $1.2499$ $75.0079\pm 0.1710$ $1.2499$ $0.0988\pm 0.0001$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0025}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0669}$ $88.2739\pm 0.4152$ $1.0879$ $0.2611\pm 0.0004$ $1.0669$ $\boldsymbol{0.0141}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2639}$ $168.5082\pm 0.3039$ $2.2896$ $2.3780\pm 0.0046$ $2.2639$ $\boldsymbol{0.1442}\pm\boldsymbol{0.0002}$
-Gaussian Distribution $\boldsymbol{2.5952}$ $47.6288\pm 0.1359$ $2.5956$ $0.0167\pm 0.0001$ $2.5952$ $\boldsymbol{0.0015}\pm\boldsymbol{0.0000}$
-Fréchet Distribution $\boldsymbol{1.1538}$ $51.3063\pm 0.0566$ $1.1542$ $0.0185\pm 0.0001$ $1.1539$ $\boldsymbol{0.0009}\pm\boldsymbol{0.0000}$
-Cauchy Distribution $1.6454$ $36.5031\pm 0.0437$ $1.6466$ $0.0177\pm 0.0001$ $\boldsymbol{1.6452}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0000}$
-Pareto Distribution $0.8297$ $39.0465\pm 0.0471$ $0.8336$ $0.0168\pm 0.0000$ $\boldsymbol{0.8297}$ $\boldsymbol{0.0007}\pm\boldsymbol{0.0000}$
-Table 16: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-BFGS (T=50) ADAM (T=50) GEORCE (T=50)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2478$ $34.0090\pm 0.7480$ $2.2479$ $0.0996\pm 0.0000$ $\boldsymbol{2.2477}$ $\boldsymbol{0.0048}\pm\boldsymbol{0.0000}$
-$\mathbb{S}^{3}$ $\boldsymbol{2.4064}$ $33.7983\pm 0.2472$ $2.4069$ $0.1720\pm 0.0001$ $\boldsymbol{2.4064}$ $\boldsymbol{0.0408}\pm\boldsymbol{0.0000}$
-$\mathbb{S}^{5}$ $2.3480$ $42.4323\pm 0.4287$ $2.3489$ $1.0908\pm 0.0111$ $\boldsymbol{2.3480}$ $\boldsymbol{0.1484}\pm\boldsymbol{0.0001}$
-$\mathbb{S}^{10}$ $1.9288$ $62.2979\pm 0.5799$ $1.9469$ $2.2366\pm 0.0018$ $\boldsymbol{1.9272}$ $\boldsymbol{0.2462}\pm\boldsymbol{0.0015}$
-$\mathbb{S}^{20}$ $1.4614$ $161.7671\pm 0.3882$ $1.5037$ $4.3590\pm 0.0024$ $\boldsymbol{1.4566}$ $\boldsymbol{0.5572}\pm\boldsymbol{0.0009}$
-$\mathbb{S}^{50}$ $0.9934$ $1028.2557\pm 1.4276$ $1.0320$ $34.7336\pm 0.1530$ $\boldsymbol{0.5279}$ $\boldsymbol{1.6104}\pm\boldsymbol{0.0016}$
-$\mathbb{S}^{100}$ $0.7500$ $4076.6101\pm 188.1693$ $0.7642$ $142.7405\pm 0.5381$ $\boldsymbol{0.3932}$ $\boldsymbol{6.4462}\pm\boldsymbol{0.0468}$
-$\mathbb{S}^{250}$- -- $0.5136$ $1163.6172\pm 5.2656$ $\boldsymbol{0.2748}$ $\boldsymbol{25.8527}\pm\boldsymbol{0.0436}$
-$\mathbb{S}^{500}$- -- $0.4253$ $2272.3826\pm 44.3587$ $\boldsymbol{0.1858}$ $\boldsymbol{117.6824}\pm\boldsymbol{0.2510}$
-$\mathbb{S}^{1000}$- -- $0.3790$ $4507.5815\pm 11.6850$ $\boldsymbol{0.1626}$ $\boldsymbol{333.4426}\pm\boldsymbol{3.6572}$
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.9330}$ $38.1118\pm 0.0767$ $1.9337$ $0.0986\pm 0.0001$ $1.9331$ $\boldsymbol{0.0038}\pm\boldsymbol{0.0000}$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0233}$ $41.8059\pm 0.0366$ $2.0237$ $0.1594\pm 0.0056$ $2.0235$ $\boldsymbol{0.0110}\pm\boldsymbol{0.0000}$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8045}$ $140.1273\pm 1.0396$ $1.8262$ $1.0991\pm 0.0024$ $1.8054$ $\boldsymbol{0.3319}\pm\boldsymbol{0.0007}$
-$\mathrm{E}\left(10\right)$ $\boldsymbol{1.4760}$ $187.4378\pm 1.0853$ $1.5193$ $2.2661\pm 0.0032$ $1.4788$ $\boldsymbol{0.3963}\pm\boldsymbol{0.0019}$
-$\mathrm{E}\left(20\right)$ $1.1489$ $329.3681\pm 2.5049$ $1.2029$ $3.8156\pm 0.0185$ $\boldsymbol{1.1366}$ $\boldsymbol{0.4997}\pm\boldsymbol{0.0015}$
-$\mathrm{E}\left(50\right)$ $0.7913$ $1755.3668\pm 3.7051$ $0.8382$ $34.8794\pm 0.1429$ $\boldsymbol{0.4417}$ $\boldsymbol{1.0645}\pm\boldsymbol{0.0041}$
-$\mathrm{E}\left(100\right)$ $0.5923$ $6260.6895\pm 30.3624$ $0.6175$ $144.6669\pm 1.0968$ $\boldsymbol{0.3160}$ $\boldsymbol{3.4531}\pm\boldsymbol{0.0102}$
-$\mathrm{E}\left(250\right)$- -- $0.4216$ $871.3954\pm 0.6870$ $\boldsymbol{0.2039}$ $\boldsymbol{25.3168}\pm\boldsymbol{0.0372}$
-$\mathrm{E}\left(500\right)$- -- $0.3454$ $1695.4600\pm 10.5103$ $\boldsymbol{0.1434}$ $\boldsymbol{114.6093}\pm\boldsymbol{0.4502}$
-$\mathrm{E}\left(1000\right)$- -- $0.3056$ $3026.5793\pm 37.3366$ $\boldsymbol{0.0996}$ $\boldsymbol{429.9206}\pm\boldsymbol{0.2607}$
-$\mathbb{T}^{2}$ $9.8955$ $59.2719\pm 0.3871$ $\boldsymbol{9.8950}$ $0.0539\pm 0.0000$ $9.8955$ $\boldsymbol{0.0116}\pm\boldsymbol{0.0001}$
-$\mathbb{H}^{2}$ $1.2434$ $37.3124\pm 0.0360$ $\boldsymbol{1.2434}$ $0.0439\pm 0.0002$ $1.2434$ $\boldsymbol{0.0014}\pm\boldsymbol{0.0000}$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0567}$ $43.7359\pm 0.0490$ $1.0610$ $0.1343\pm 0.0015$ $1.0567$ $\boldsymbol{0.0081}\pm\boldsymbol{0.0002}$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2304}$ $98.5211\pm 0.6390$ $2.2307$ $1.1962\pm 0.0194$ $2.2304$ $\boldsymbol{0.0500}\pm\boldsymbol{0.0001}$
-Gaussian Distribution $2.5778$ $23.5395\pm 0.0607$ $\boldsymbol{2.5769}$ $0.0084\pm 0.0001$ $2.5778$ $\boldsymbol{0.0011}\pm\boldsymbol{0.0000}$
-Fréchet Distribution $\boldsymbol{1.1457}$ $26.5810\pm 0.1445$ $1.1457$ $0.0097\pm 0.0001$ $1.1457$ $\boldsymbol{0.0006}\pm\boldsymbol{0.0000}$
-Cauchy Distribution $1.6341$ $20.1153\pm 0.0478$ $1.6341$ $\boldsymbol{0.0012}\pm\boldsymbol{0.0000}$ $\boldsymbol{1.6340}$ $0.0012\pm 0.0000$
-Pareto Distribution $0.8235$ $22.4667\pm 0.0646$ $0.8235$ $0.0088\pm 0.0000$ $\boldsymbol{0.8234}$ $\boldsymbol{0.0004}\pm\boldsymbol{0.0000}$
-Table 17: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-SGD (T=100) CG (T=100) dogleg (T=100)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.4063$ $\boldsymbol{0.2238}\pm\boldsymbol{0.0075}$ $2.2744$ $193.4884\pm 2.7955$ $\boldsymbol{2.2742}$ $3.8305\pm 0.0143$
-$\mathbb{S}^{3}$ $2.7102$ $\boldsymbol{0.3350}\pm\boldsymbol{0.0007}$ $2.4346$ $248.6363\pm 2.7041$ $\boldsymbol{2.4342}$ $5.6319\pm 0.0156$
-$\mathbb{S}^{5}$ $3.0827$ $2.1406\pm 0.0095$ $\boldsymbol{2.3810}$ $217.8721\pm 4.5710$ $3.1069$ $\boldsymbol{1.7988}\pm\boldsymbol{0.0217}$
-$\mathbb{S}^{10}$ $3.4578$ $\boldsymbol{4.6462}\pm\boldsymbol{0.0732}$ $\boldsymbol{1.9747}$ $178.0182\pm 0.0418$ $3.5092$ $6.9953\pm 0.0104$
-$\mathbb{S}^{20}$ $3.5726$ $\boldsymbol{8.8376}\pm\boldsymbol{0.0048}$ $\boldsymbol{1.4961}$ $205.9335\pm 1.9411$ $3.6632$ $41.3160\pm 0.0809$
-$\mathbb{S}^{50}$ $3.2831$ $\boldsymbol{72.7622}\pm\boldsymbol{0.0207}$ $\boldsymbol{1.2180}$ $81.5957\pm 0.0459$- --
-$\mathbb{S}^{100}$ $2.8231$ $298.8796\pm 0.0082$ $\boldsymbol{1.0540}$ $\boldsymbol{109.8878}\pm\boldsymbol{0.4428}$- --
-$\mathbb{S}^{250}$ $2.0949$ $2608.4075\pm 40.8368$ $\boldsymbol{0.8580}$ $\boldsymbol{391.9100}\pm\boldsymbol{0.9132}$- --
-$\mathbb{S}^{500}$ $1.5779$ $10754.5576\pm 300.1625$ $\boldsymbol{1.0325}$ $\boldsymbol{709.8522}\pm\boldsymbol{4.6273}$- --
-$\mathbb{S}^{1000}$- -- $\boldsymbol{1.1545}$ $\boldsymbol{68.2710}\pm\boldsymbol{0.7146}$- --
-$\mathrm{E}\left(2\right)$ $1.9693$ $\boldsymbol{0.2259}\pm\boldsymbol{0.0018}$ $1.9598$ $159.7566\pm 1.6439$ $\boldsymbol{1.9595}$ $3.9798\pm 0.0834$
-$\mathrm{E}\left(3\right)$ $2.0769$ $\boldsymbol{0.3266}\pm\boldsymbol{0.0029}$ $\boldsymbol{2.0504}$ $269.9406\pm 1.5354$ $2.0795$ $1.2249\pm 0.0080$
-$\mathrm{E}\left(5\right)$ $2.2219$ $2.2046\pm 0.0154$ $\boldsymbol{1.8503}$ $516.0936\pm 11.2331$ $2.2269$ $\boldsymbol{1.7201}\pm\boldsymbol{0.0152}$
-$\mathrm{E}\left(10\right)$ $2.4040$ $\boldsymbol{4.5520}\pm\boldsymbol{0.0117}$ $\boldsymbol{1.5794}$ $286.3407\pm 2.5997$ $2.4153$ $7.0164\pm 0.0183$
-$\mathrm{E}\left(20\right)$ $2.4961$ $\boldsymbol{8.8872}\pm\boldsymbol{0.0029}$ $\boldsymbol{1.1925}$ $305.8136\pm 2.2537$ $2.5175$ $40.8913\pm 0.1881$
-$\mathrm{E}\left(50\right)$ $2.3587$ $\boldsymbol{73.9033}\pm\boldsymbol{0.2967}$ $\boldsymbol{0.8949}$ $181.0294\pm 2.8838$- --
-$\mathrm{E}\left(100\right)$ $2.0627$ $292.5031\pm 1.1563$ $\boldsymbol{0.8718}$ $\boldsymbol{177.2222}\pm\boldsymbol{0.6547}$- --
-$\mathrm{E}\left(250\right)$ $1.5510$ $2521.6353\pm 22.3619$ $\boldsymbol{0.8962}$ $\boldsymbol{341.6886}\pm\boldsymbol{1.5827}$- --
-$\mathrm{E}\left(500\right)$ $1.1756$ $10801.0918\pm 301.4005$ $\boldsymbol{0.7106}$ $\boldsymbol{1545.6230}\pm\boldsymbol{10.8368}$- --
-$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.8633}$ $\boldsymbol{71.7434}\pm\boldsymbol{0.1687}$- --
-$\mathbb{T}^{2}$ $11.3048$ $\boldsymbol{0.1001}\pm\boldsymbol{0.0000}$ $\boldsymbol{9.9854}$ $609.3215\pm 9.3287$ $11.6903$ $1.2105\pm 0.0095$
-$\mathbb{H}^{2}$ $1.3308$ $\boldsymbol{0.0839}\pm\boldsymbol{0.0000}$ $1.2532$ $549.9518\pm 3.2561$ $\boldsymbol{1.2498}$ $4.1077\pm 0.0109$
-$\mathcal{P}\left(2\right)$ $1.1512$ $\boldsymbol{0.2608}\pm\boldsymbol{0.0021}$ $1.0676$ $420.3352\pm 1.1349$ $\boldsymbol{1.0669}$ $4.5709\pm 0.0106$
-$\mathcal{P}\left(3\right)$ $2.3892$ $\boldsymbol{2.1047}\pm\boldsymbol{0.0039}$ $\boldsymbol{2.2648}$ $570.4799\pm 5.5919$ $2.4047$ $2.3496\pm 0.0109$
-Gaussian Distribution $2.8543$ $\boldsymbol{0.0115}\pm\boldsymbol{0.0001}$ $2.5955$ $374.0846\pm 7.3067$ $\boldsymbol{2.5952}$ $3.5420\pm 0.0317$
-Fréchet Distribution $1.1823$ $\boldsymbol{0.0134}\pm\boldsymbol{0.0000}$ $1.1547$ $321.4810\pm 1.0877$ $\boldsymbol{1.1539}$ $2.7293\pm 0.0160$
-Cauchy Distribution $1.9752$ $\boldsymbol{0.0119}\pm\boldsymbol{0.0001}$ $1.6461$ $192.8583\pm 1.3344$ $\boldsymbol{1.6452}$ $3.7594\pm 0.0544$
-Pareto Distribution $0.8500$ $\boldsymbol{0.0111}\pm\boldsymbol{0.0001}$ $0.8299$ $87.2695\pm 0.1665$ $\boldsymbol{0.8297}$ $2.2831\pm 0.0390$
-Table 18: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-SGD (T=50) CG (T=50) dogleg (T=50)
-Riemannian Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.3607$ $\boldsymbol{0.0963}\pm\boldsymbol{0.0000}$ $2.2478$ $112.6833\pm 0.0636$ $\boldsymbol{2.2478}$ $3.6126\pm 0.0092$
-$\mathbb{S}^{3}$ $2.6515$ $\boldsymbol{0.1450}\pm\boldsymbol{0.0005}$ $2.4065$ $119.3687\pm 0.3268$ $\boldsymbol{2.4064}$ $4.7996\pm 0.0220$
-$\mathbb{S}^{5}$ $2.9958$ $\boldsymbol{1.0907}\pm\boldsymbol{0.0047}$ $\boldsymbol{2.3483}$ $132.6204\pm 0.7172$ $3.0849$ $1.1818\pm 0.0089$
-$\mathbb{S}^{10}$ $3.3093$ $2.2309\pm 0.0017$ $\boldsymbol{1.9287}$ $131.0700\pm 2.7469$ $3.4892$ $\boldsymbol{2.2054}\pm\boldsymbol{0.0121}$
-$\mathbb{S}^{20}$ $3.3585$ $\boldsymbol{4.3963}\pm\boldsymbol{0.0043}$ $\boldsymbol{1.4644}$ $197.9373\pm 3.1216$ $3.6466$ $10.9397\pm 0.0310$
-$\mathbb{S}^{50}$ $3.0604$ $\boldsymbol{33.8548}\pm\boldsymbol{0.0592}$ $\boldsymbol{1.0403}$ $66.8326\pm 0.4691$- --
-$\mathbb{S}^{100}$ $2.6628$ $142.5886\pm 0.5246$ $\boldsymbol{0.8374}$ $\boldsymbol{71.4308}\pm\boldsymbol{0.1749}$- --
-$\mathbb{S}^{250}$ $2.0317$ $1351.8007\pm 6.9661$ $\boldsymbol{0.7906}$ $\boldsymbol{151.2937}\pm\boldsymbol{0.2754}$- --
-$\mathbb{S}^{500}$ $1.5557$ $5792.4668\pm 19.3864$ $\boldsymbol{0.9376}$ $\boldsymbol{245.6238}\pm\boldsymbol{1.3281}$- --
-$\mathbb{S}^{1000}$- -- $\boldsymbol{0.8765}$ $\boldsymbol{587.6923}\pm\boldsymbol{4.5177}$- --
-$\mathrm{E}\left(2\right)$ $1.9493$ $\boldsymbol{0.0959}\pm\boldsymbol{0.0001}$ $1.9331$ $91.2580\pm 0.7684$ $\boldsymbol{1.9330}$ $3.6286\pm 0.0089$
-$\mathrm{E}\left(3\right)$ $2.0568$ $\boldsymbol{0.1565}\pm\boldsymbol{0.0004}$ $\boldsymbol{2.0234}$ $188.5732\pm 1.4986$ $2.0661$ $1.0738\pm 0.0245$
-$\mathrm{E}\left(5\right)$ $2.1983$ $\boldsymbol{1.0906}\pm\boldsymbol{0.0026}$ $\boldsymbol{1.8099}$ $319.5569\pm 2.8321$ $2.2150$ $1.1382\pm 0.0093$
-$\mathrm{E}\left(10\right)$ $2.3647$ $2.2644\pm 0.0033$ $\boldsymbol{1.4779}$ $364.8791\pm 3.5348$ $2.4045$ $\boldsymbol{2.2572}\pm\boldsymbol{0.0141}$
-$\mathrm{E}\left(20\right)$ $2.4309$ $\boldsymbol{4.5347}\pm\boldsymbol{0.0025}$ $\boldsymbol{1.1536}$ $256.8932\pm 2.5821$ $2.5079$ $11.0630\pm 0.0303$
-$\mathrm{E}\left(50\right)$ $2.2709$ $\boldsymbol{29.4273}\pm\boldsymbol{0.1562}$ $\boldsymbol{0.8494}$ $99.0591\pm 0.7024$- --
-$\mathrm{E}\left(100\right)$ $1.9914$ $123.1176\pm 0.7069$ $\boldsymbol{0.6421}$ $\boldsymbol{104.5576}\pm\boldsymbol{0.4612}$- --
-$\mathrm{E}\left(250\right)$ $1.5207$ $1334.2307\pm 3.4387$ $\boldsymbol{0.6019}$ $\boldsymbol{199.9802}\pm\boldsymbol{0.7304}$- --
-$\mathrm{E}\left(500\right)$ $1.1643$ $5472.2515\pm 84.0089$ $\boldsymbol{0.6644}$ $\boldsymbol{534.2835}\pm\boldsymbol{2.8548}$- --
-$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.7192}$ $\boldsymbol{640.3465}\pm\boldsymbol{1.6748}$- --
-$\mathbb{T}^{2}$ $10.5379$ $\boldsymbol{0.0498}\pm\boldsymbol{0.0001}$ $\boldsymbol{9.8956}$ $239.9891\pm 0.9619$ $11.5938$ $1.0782\pm 0.0071$
-$\mathbb{H}^{2}$ $1.3210$ $\boldsymbol{0.0425}\pm\boldsymbol{0.0000}$ $1.2434$ $367.4767\pm 2.7398$ $\boldsymbol{1.2434}$ $3.6867\pm 0.0105$
-$\mathcal{P}\left(2\right)$ $1.1238$ $\boldsymbol{0.1225}\pm\boldsymbol{0.0009}$ $1.0567$ $222.3136\pm 1.0460$ $\boldsymbol{1.0567}$ $3.6711\pm 0.0168$
-$\mathcal{P}\left(3\right)$ $2.3231$ $\boldsymbol{1.2252}\pm\boldsymbol{0.0004}$ $\boldsymbol{2.2306}$ $364.5162\pm 0.9199$ $2.3721$ $1.3057\pm 0.0108$
-Gaussian Distribution $2.7366$ $\boldsymbol{0.0055}\pm\boldsymbol{0.0001}$ $2.5780$ $130.6218\pm 0.8519$ $\boldsymbol{2.5778}$ $3.4534\pm 0.0148$
-Fréchet Distribution $1.1685$ $\boldsymbol{0.0066}\pm\boldsymbol{0.0001}$ $1.1458$ $169.4907\pm 0.5148$ $\boldsymbol{1.1457}$ $2.5564\pm 0.0123$
-Cauchy Distribution $1.9013$ $\boldsymbol{0.0057}\pm\boldsymbol{0.0000}$ $1.6341$ $109.2575\pm 0.1122$ $\boldsymbol{1.6340}$ $3.8239\pm 0.0166$
-Pareto Distribution $0.8399$ $\boldsymbol{0.0056}\pm\boldsymbol{0.0000}$ $0.8235$ $64.6855\pm 0.0367$ $\boldsymbol{0.8235}$ $2.1945\pm 0.0138$
-Table 19: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-trust-ncg (T=100) trust-exact (T=100)
-Riemannian Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2742$ $237.8402\pm 3.6964$ $\boldsymbol{2.2742}$ $\boldsymbol{3.8066}\pm\boldsymbol{0.0204}$
-$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $340.5377\pm 1.5571$ $2.4342$ $\boldsymbol{6.8560}\pm\boldsymbol{0.0218}$
-$\mathbb{S}^{5}$ $\boldsymbol{2.3754}$ $833.7228\pm 9.9481$ $2.3754$ $\boldsymbol{12.4160}\pm\boldsymbol{0.0325}$
-$\mathbb{S}^{10}$ $\boldsymbol{1.9521}$ $9097.0312\pm 53.0355$ $1.9523$ $\boldsymbol{65.4820}\pm\boldsymbol{0.4559}$
-$\mathbb{S}^{20}$- -- $\boldsymbol{1.4816}$ $\boldsymbol{424.0417}\pm\boldsymbol{1.6674}$
-$\mathbb{S}^{50}$- -- -- --
-$\mathbb{S}^{100}$- -- -- --
-$\mathbb{S}^{250}$- -- -- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $1.9595$ $265.0164\pm 3.4124$ $\boldsymbol{1.9595}$ $\boldsymbol{3.8934}\pm\boldsymbol{0.0312}$
-$\mathrm{E}\left(3\right)$ $2.0499$ $726.9606\pm 6.6789$ $\boldsymbol{2.0498}$ $\boldsymbol{4.7604}\pm\boldsymbol{0.1015}$
-$\mathrm{E}\left(5\right)$ $1.8394$ $3843.9355\pm 39.4764$ $\boldsymbol{1.8393}$ $\boldsymbol{16.1207}\pm\boldsymbol{0.2033}$
-$\mathrm{E}\left(10\right)$- -- $\boldsymbol{1.5091}$ $\boldsymbol{73.9657}\pm\boldsymbol{0.1229}$
-$\mathrm{E}\left(20\right)$- -- $\boldsymbol{1.1718}$ $\boldsymbol{398.0183}\pm\boldsymbol{1.7858}$
-$\mathrm{E}\left(50\right)$- -- -- --
-$\mathrm{E}\left(100\right)$- -- -- --
-$\mathrm{E}\left(250\right)$- -- -- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $9.9851$ $374.0822\pm 5.9408$ $\boldsymbol{9.9851}$ $\boldsymbol{6.7989}\pm\boldsymbol{0.0304}$
-$\mathbb{H}^{2}$ $\boldsymbol{1.2498}$ $627.2646\pm 7.3623$ $1.2498$ $\boldsymbol{5.0876}\pm\boldsymbol{0.0134}$
-$\mathcal{P}\left(2\right)$ $1.0669$ $594.2135\pm 5.1381$ $\boldsymbol{1.0669}$ $\boldsymbol{4.0894}\pm\boldsymbol{0.0212}$
-$\mathcal{P}\left(3\right)$ $2.2639$ $2158.6843\pm 25.1696$ $\boldsymbol{2.2639}$ $\boldsymbol{13.2320}\pm\boldsymbol{0.0803}$
-Gaussian Distribution $\boldsymbol{2.5952}$ $202.1364\pm 4.9068$ $2.5952$ $\boldsymbol{3.0117}\pm\boldsymbol{0.0225}$
-Fréchet Distribution $1.1539$ $286.7279\pm 9.4552$ $\boldsymbol{1.1539}$ $\boldsymbol{2.6002}\pm\boldsymbol{0.0147}$
-Cauchy Distribution $\boldsymbol{1.6452}$ $201.4917\pm 0.1496$ $1.6452$ $\boldsymbol{3.9414}\pm\boldsymbol{0.0098}$
-Pareto Distribution $\boldsymbol{0.8297}$ $233.9446\pm 0.2961$ $0.8297$ $\boldsymbol{2.3762}\pm\boldsymbol{0.0172}$
-Table 20: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-trust-ncg (T=50) trust-exact (T=50)
-Riemannian Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2478$ $115.7323\pm 1.0615$ $\boldsymbol{2.2478}$ $\boldsymbol{3.6180}\pm\boldsymbol{0.0205}$
-$\mathbb{S}^{3}$ $2.4064$ $138.8976\pm 0.8824$ $\boldsymbol{2.4064}$ $\boldsymbol{4.6065}\pm\boldsymbol{0.0202}$
-$\mathbb{S}^{5}$ $2.3480$ $272.6404\pm 1.2856$ $\boldsymbol{2.3480}$ $\boldsymbol{7.6100}\pm\boldsymbol{0.0293}$
-$\mathbb{S}^{10}$ $\boldsymbol{1.9274}$ $1127.8496\pm 3.9055$ $1.9275$ $\boldsymbol{19.4117}\pm\boldsymbol{0.0371}$
-$\mathbb{S}^{20}$ $\boldsymbol{1.4608}$ $8353.8154\pm 153.6381$ $1.4610$ $\boldsymbol{110.3418}\pm\boldsymbol{0.4293}$
-$\mathbb{S}^{50}$- -- -- --
-$\mathbb{S}^{100}$- -- -- --
-$\mathbb{S}^{250}$- -- -- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $1.9330$ $101.3354\pm 0.2792$ $\boldsymbol{1.9330}$ $\boldsymbol{3.8779}\pm\boldsymbol{0.0182}$
-$\mathrm{E}\left(3\right)$ $2.0233$ $256.4745\pm 4.0010$ $\boldsymbol{2.0233}$ $\boldsymbol{4.6605}\pm\boldsymbol{0.0146}$
-$\mathrm{E}\left(5\right)$ $1.8046$ $1220.0194\pm 7.2318$ $\boldsymbol{1.8046}$ $\boldsymbol{10.3930}\pm\boldsymbol{0.0328}$
-$\mathrm{E}\left(10\right)$ $1.4753$ $3603.1384\pm 81.3956$ $\boldsymbol{1.4751}$ $\boldsymbol{23.5777}\pm\boldsymbol{0.0401}$
-$\mathrm{E}\left(20\right)$- -- $\boldsymbol{1.1434}$ $\boldsymbol{109.4113}\pm\boldsymbol{0.0144}$
-$\mathrm{E}\left(50\right)$- -- -- --
-$\mathrm{E}\left(100\right)$- -- -- --
-$\mathrm{E}\left(250\right)$- -- -- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{9.8955}$ $189.0475\pm 0.1325$ $\boldsymbol{9.8955}$ $\boldsymbol{7.5805}\pm\boldsymbol{0.0152}$
-$\mathbb{H}^{2}$ $1.2434$ $200.3231\pm 0.3463$ $\boldsymbol{1.2434}$ $\boldsymbol{3.6404}\pm\boldsymbol{0.0076}$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0567}$ $153.1577\pm 0.7454$ $1.0567$ $\boldsymbol{3.6134}\pm\boldsymbol{0.0305}$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{2.2304}$ $612.8415\pm 8.6455$ $2.2304$ $\boldsymbol{6.2574}\pm\boldsymbol{0.0139}$
-Gaussian Distribution $\boldsymbol{2.5778}$ $122.4479\pm 0.1045$ $2.5778$ $\boldsymbol{2.8004}\pm\boldsymbol{0.0155}$
-Fréchet Distribution $\boldsymbol{1.1457}$ $215.6763\pm 0.4161$ $1.1457$ $\boldsymbol{2.5323}\pm\boldsymbol{0.0123}$
-Cauchy Distribution $1.6340$ $113.7271\pm 0.0720$ $\boldsymbol{1.6340}$ $\boldsymbol{3.6122}\pm\boldsymbol{0.0207}$
-Pareto Distribution $0.8235$ $80.3471\pm 0.8322$ $\boldsymbol{0.8235}$ $\boldsymbol{2.2065}\pm\boldsymbol{0.0180}$
-Table 21: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-Sparse Newton (T=100) Sparse Regularized Newton (T=100)
-Riemannian Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $\boldsymbol{2.2742}$ $\boldsymbol{0.0180}\pm\boldsymbol{0.0002}$ $2.2742$ $0.1017\pm 0.0017$
-$\mathbb{S}^{3}$ $\boldsymbol{2.4342}$ $\boldsymbol{0.0837}\pm\boldsymbol{0.0157}$ $2.4342$ $0.3089\pm 0.0428$
-$\mathbb{S}^{5}$ $\boldsymbol{2.3755}$ $\boldsymbol{0.6723}\pm\boldsymbol{0.0409}$ $2.3756$ $1.2203\pm 0.0887$
-$\mathbb{S}^{10}$ $\boldsymbol{1.9449}$ $256.4748\pm 20.0998$ $1.9531$ $\boldsymbol{4.3786}\pm\boldsymbol{0.3614}$
-$\mathbb{S}^{20}$ $1.7069$ $1130.3478\pm 127.1948$ $\boldsymbol{1.4943}$ $\boldsymbol{20.3206}\pm\boldsymbol{1.5336}$
-$\mathbb{S}^{50}$ $1.0815$ $16216.8467\pm 2136.6802$ $\boldsymbol{1.0106}$ $\boldsymbol{330.6598}\pm\boldsymbol{26.1208}$
-$\mathbb{S}^{100}$ $\boldsymbol{0.7522}$ $\boldsymbol{1041.2721}\pm\boldsymbol{89.4723}$ $0.7781$ $2863.6448\pm 233.2495$
-$\mathbb{S}^{250}$ $\boldsymbol{0.4958}$ $\boldsymbol{17070.7520}\pm\boldsymbol{2024.9161}$- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.9595}$ $\boldsymbol{0.0254}\pm\boldsymbol{0.0002}$ $1.9596$ $0.0915\pm 0.0028$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0498}$ $\boldsymbol{0.1217}\pm\boldsymbol{0.0219}$ $2.0499$ $0.3029\pm 0.0288$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8392}$ $\boldsymbol{0.6774}\pm\boldsymbol{0.0721}$ $1.8397$ $1.3500\pm 0.1241$
-$\mathrm{E}\left(10\right)$ $1.5247$ $300.2892\pm 23.3273$ $\boldsymbol{1.5114}$ $\boldsymbol{4.4472}\pm\boldsymbol{0.3697}$
-$\mathrm{E}\left(20\right)$ $1.4595$ $1101.5878\pm 115.0272$ $\boldsymbol{1.1961}$ $\boldsymbol{17.6133}\pm\boldsymbol{1.2503}$
-$\mathrm{E}\left(50\right)$ $0.8825$ $15145.3818\pm 1401.9972$ $\boldsymbol{0.8279}$ $\boldsymbol{300.5634}\pm\boldsymbol{20.9595}$
-$\mathrm{E}\left(100\right)$ $\boldsymbol{0.5922}$ $\boldsymbol{839.4538}\pm\boldsymbol{68.5559}$ $0.6402$ $2509.0547\pm 213.2646$
-$\mathrm{E}\left(250\right)$ $\boldsymbol{0.3919}$ $\boldsymbol{16334.5381}\pm\boldsymbol{1370.7971}$- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{9.9851}$ $\boldsymbol{0.0308}\pm\boldsymbol{0.0002}$ $9.9851$ $0.0725\pm 0.0014$
-$\mathbb{H}^{2}$ $\boldsymbol{1.2498}$ $\boldsymbol{0.0225}\pm\boldsymbol{0.0004}$ $1.2498$ $0.0948\pm 0.0036$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0669}$ $\boldsymbol{0.0692}\pm\boldsymbol{0.0182}$ $1.0669$ $0.2395\pm 0.0334$
-$\mathcal{P}\left(3\right)$ $2.3954$ $81.4573\pm 5.9683$ $\boldsymbol{2.2639}$ $\boldsymbol{0.9140}\pm\boldsymbol{0.0747}$
-Gaussian Distribution $\boldsymbol{2.5952}$ $\boldsymbol{0.0123}\pm\boldsymbol{0.0001}$ $2.5952$ $0.0472\pm 0.0007$
-Fréchet Distribution $\boldsymbol{1.1539}$ $\boldsymbol{0.0085}\pm\boldsymbol{0.0001}$ $1.1539$ $0.0500\pm 0.0004$
-Cauchy Distribution $\boldsymbol{1.6453}$ $\boldsymbol{0.0120}\pm\boldsymbol{0.0001}$ $1.6453$ $0.0519\pm 0.0002$
-Pareto Distribution $\boldsymbol{0.8297}$ $\boldsymbol{0.0087}\pm\boldsymbol{0.0001}$ $0.8297$ $0.0444\pm 0.0005$
-Table 22: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-Sparse Newton (T=50) Sparse Regularized Newton (T=50)
-Riemannian Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $2.2478$ $\boldsymbol{0.0145}\pm\boldsymbol{0.0001}$ $\boldsymbol{2.2478}$ $0.0500\pm 0.0001$
-$\mathbb{S}^{3}$ $2.4064$ $\boldsymbol{0.0581}\pm\boldsymbol{0.0177}$ $\boldsymbol{2.4064}$ $0.1629\pm 0.0247$
-$\mathbb{S}^{5}$ $2.3480$ $\boldsymbol{0.2479}\pm\boldsymbol{0.0150}$ $\boldsymbol{2.3480}$ $0.5728\pm 0.0373$
-$\mathbb{S}^{10}$ $\boldsymbol{1.9274}$ $2.2659\pm 0.1698$ $1.9277$ $\boldsymbol{1.9943}\pm\boldsymbol{0.1543}$
-$\mathbb{S}^{20}$ $\boldsymbol{1.4610}$ $230.5914\pm 17.8627$ $1.4629$ $\boldsymbol{9.6611}\pm\boldsymbol{0.7412}$
-$\mathbb{S}^{50}$ $0.9948$ $7394.6870\pm 845.2023$ $\boldsymbol{0.9720}$ $\boldsymbol{128.6652}\pm\boldsymbol{12.2889}$
-$\mathbb{S}^{100}$ $\boldsymbol{0.6922}$ $\boldsymbol{712.1899}\pm\boldsymbol{54.3496}$ $0.7163$ $1410.9242\pm 104.6961$
-$\mathbb{S}^{250}$ $\boldsymbol{0.4581}$ $\boldsymbol{13080.0830}\pm\boldsymbol{1387.7233}$- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.9330}$ $\boldsymbol{0.0207}\pm\boldsymbol{0.0002}$ $1.9330$ $0.0467\pm 0.0002$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{2.0233}$ $\boldsymbol{0.0713}\pm\boldsymbol{0.0186}$ $2.0233$ $0.1612\pm 0.0257$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.8046}$ $\boldsymbol{0.5449}\pm\boldsymbol{0.0668}$ $1.8050$ $0.7055\pm 0.0764$
-$\mathrm{E}\left(10\right)$ $\boldsymbol{0.9473}$ $133.1282\pm 9.1672$ $1.4766$ $\boldsymbol{2.2189}\pm\boldsymbol{0.2136}$
-$\mathrm{E}\left(20\right)$ $1.2745$ $591.8414\pm 35.1453$ $\boldsymbol{1.1484}$ $\boldsymbol{9.3243}\pm\boldsymbol{1.2287}$
-$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5985}$ $7950.0571\pm 865.0385$ $0.7965$ $\boldsymbol{146.5869}\pm\boldsymbol{7.8365}$
-$\mathrm{E}\left(100\right)$ $\boldsymbol{0.5707}$ $\boldsymbol{775.0126}\pm\boldsymbol{53.9842}$ $0.5863$ $1405.4775\pm 100.2025$
-$\mathrm{E}\left(250\right)$ $\boldsymbol{0.3882}$ $\boldsymbol{9355.2041}\pm\boldsymbol{978.4170}$- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{9.8943}$ $3.4085\pm 0.2846$ $9.8955$ $\boldsymbol{0.0322}\pm\boldsymbol{0.0004}$
-$\mathbb{H}^{2}$ $\boldsymbol{1.2434}$ $\boldsymbol{0.0115}\pm\boldsymbol{0.0003}$ $1.2435$ $0.0441\pm 0.0015$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{1.0567}$ $\boldsymbol{0.0351}\pm\boldsymbol{0.0159}$ $1.0567$ $0.1206\pm 0.0212$
-$\mathcal{P}\left(3\right)$ $2.3328$ $42.5019\pm 3.4439$ $\boldsymbol{2.2304}$ $\boldsymbol{0.3846}\pm\boldsymbol{0.0256}$
-Gaussian Distribution $\boldsymbol{2.5778}$ $\boldsymbol{0.0060}\pm\boldsymbol{0.0001}$ $2.5778$ $0.0199\pm 0.0003$
-Fréchet Distribution $\boldsymbol{1.1457}$ $\boldsymbol{0.0045}\pm\boldsymbol{0.0001}$ $1.1457$ $0.0217\pm 0.0002$
-Cauchy Distribution $\boldsymbol{1.6340}$ $\boldsymbol{0.0079}\pm\boldsymbol{0.0002}$ $1.6340$ $0.0223\pm 0.0004$
-Pareto Distribution $\boldsymbol{0.8235}$ $\boldsymbol{0.0042}\pm\boldsymbol{0.0001}$ $0.8235$ $0.0212\pm 0.0003$
-Table 23: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-Methods on a Finsler manifold
-We show in Table 24, Table 25, Table 26, Table 27, Table 28, Table 29, Table 30 and Table 31 runtime estimates for different methods on a CPU for different Finsler manifolds.
-BFGS (T=100) ADAM (T=100) GEORCE (T=100)
-Finsler Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.6525$ $244.9129\pm 0.0680$ $1.6536$ $0.4967\pm 0.0007$ $\boldsymbol{1.6525}$ $\boldsymbol{0.0194}\pm\boldsymbol{0.0009}$
-$\mathbb{S}^{3}$ $1.7433$ $277.8015\pm 0.5819$ $1.7456$ $0.7639\pm 0.0031$ $\boldsymbol{1.7432}$ $\boldsymbol{0.0362}\pm\boldsymbol{0.0001}$
-$\mathbb{S}^{5}$ $1.7032$ $435.6987\pm 4.6022$ $1.7377$ $4.6711\pm 0.0047$ $\boldsymbol{1.7029}$ $\boldsymbol{0.4556}\pm\boldsymbol{0.0021}$
-$\mathbb{S}^{10}$ $1.4151$ $812.2809\pm 13.5968$ $1.5174$ $9.7974\pm 0.0052$ $\boldsymbol{1.4109}$ $\boldsymbol{0.7473}\pm\boldsymbol{0.0029}$
-$\mathbb{S}^{20}$ $1.0951$ $1446.0956\pm 4.8038$ $1.2018$ $18.3489\pm 0.0747$ $\boldsymbol{1.0760}$ $\boldsymbol{1.2856}\pm\boldsymbol{0.0016}$
-$\mathbb{S}^{50}$ $0.8216$ $5992.7261\pm 56.5973$ $0.8128$ $146.4556\pm 0.9341$ $\boldsymbol{0.7285}$ $\boldsymbol{8.4758}\pm\boldsymbol{0.0234}$
-$\mathbb{S}^{100}$- -- $0.6187$ $498.9705\pm 0.4032$ $\boldsymbol{0.2570}$ $\boldsymbol{15.3786}\pm\boldsymbol{0.0092}$
-$\mathbb{S}^{250}$- -- $0.5195$ $1397.6732\pm 0.7793$ $\boldsymbol{0.1622}$ $\boldsymbol{132.5889}\pm\boldsymbol{0.1106}$
-$\mathbb{S}^{500}$- -- $0.4850$ $1954.3231\pm 14.1107$ $\boldsymbol{0.1404}$ $\boldsymbol{364.5326}\pm\boldsymbol{1.8222}$
-$\mathbb{S}^{1000}$- -- $0.3843$ $5982.0884\pm 204.1833$ $\boldsymbol{0.1140}$ $\boldsymbol{1338.9869}\pm\boldsymbol{2.3954}$
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3616}$ $230.0809\pm 1.3967$ $1.3621$ $0.5310\pm 0.0025$ $1.3618$ $\boldsymbol{0.0130}\pm\boldsymbol{0.0002}$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{1.4001}$ $416.6281\pm 5.5354$ $1.4087$ $0.8183\pm 0.0026$ $1.4023$ $\boldsymbol{0.1081}\pm\boldsymbol{0.0001}$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2951}$ $638.7185\pm 2.7818$ $1.3265$ $4.3802\pm 0.0039$ $1.2956$ $\boldsymbol{0.4554}\pm\boldsymbol{0.0014}$
-$\mathrm{E}\left(10\right)$ $\boldsymbol{1.0733}$ $741.9274\pm 12.3382$ $1.1229$ $9.7899\pm 0.0117$ $1.0804$ $\boldsymbol{0.5815}\pm\boldsymbol{0.0022}$
-$\mathrm{E}\left(20\right)$ $0.8510$ $1820.3137\pm 4.0734$ $0.8980$ $19.4614\pm 0.0039$ $\boldsymbol{0.8121}$ $\boldsymbol{1.0671}\pm\boldsymbol{0.0044}$
-$\mathrm{E}\left(50\right)$ $0.5981$ $11191.9424\pm 597.7002$ $0.6325$ $146.6409\pm 0.1668$ $\boldsymbol{0.2842}$ $\boldsymbol{4.4377}\pm\boldsymbol{0.0451}$
-$\mathrm{E}\left(100\right)$- -- $0.4936$ $491.8223\pm 0.2346$ $\boldsymbol{0.2009}$ $\boldsymbol{14.5367}\pm\boldsymbol{0.0092}$
-$\mathrm{E}\left(250\right)$- -- $0.4159$ $1475.5049\pm 4.1430$ $\boldsymbol{0.1287}$ $\boldsymbol{125.4187}\pm\boldsymbol{0.0888}$
-$\mathrm{E}\left(500\right)$- -- $0.4142$ $1984.7845\pm 10.5689$ $\boldsymbol{0.0967}$ $\boldsymbol{373.2816}\pm\boldsymbol{0.4227}$
-$\mathrm{E}\left(1000\right)$- -- $0.3444$ $5929.4438\pm 167.1873$ $\boldsymbol{0.0799}$ $\boldsymbol{1211.8812}\pm\boldsymbol{0.2511}$
-$\mathbb{T}^{2}$ $7.3890$ $449.4798\pm 10.3282$ $7.3585$ $0.3294\pm 0.0018$ $\boldsymbol{6.8067}$ $\boldsymbol{0.1126}\pm\boldsymbol{0.0002}$
-$\mathbb{H}^{2}$ $1.3340$ $384.3149\pm 0.3044$ $1.3360$ $0.2744\pm 0.0024$ $\boldsymbol{1.3337}$ $\boldsymbol{0.0094}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7291}$ $414.8355\pm 3.0837$ $0.7338$ $0.6368\pm 0.0026$ $0.7292$ $\boldsymbol{0.0222}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $536.6575\pm 0.9323$ $1.2980$ $5.2824\pm 0.0540$ $1.2856$ $\boldsymbol{0.2094}\pm\boldsymbol{0.0004}$
-Gaussian Distribution $1.6878$ $163.4399\pm 0.1859$ $1.6936$ $0.1145\pm 0.0001$ $\boldsymbol{1.6878}$ $\boldsymbol{0.0047}\pm\boldsymbol{0.0001}$
-Fréchet Distribution $0.5408$ $123.8418\pm 0.0499$ $0.5406$ $0.0585\pm 0.0001$ $\boldsymbol{0.5405}$ $\boldsymbol{0.0021}\pm\boldsymbol{0.0001}$
-Cauchy Distribution $1.1134$ $182.3980\pm 2.2559$ $\boldsymbol{1.1133}$ $0.1105\pm 0.0007$ $1.1133$ $\boldsymbol{0.0067}\pm\boldsymbol{0.0002}$
-Pareto Distribution $0.3884$ $84.8186\pm 0.0832$ $0.3885$ $0.0235\pm 0.0007$ $\boldsymbol{0.3883}$ $\boldsymbol{0.0020}\pm\boldsymbol{0.0001}$
-Table 24: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-BFGS (T=50) ADAM (T=50) GEORCE (T=50)
-Finsler Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.6385$ $143.8810\pm 0.2377$ $\boldsymbol{1.6385}$ $0.2635\pm 0.0001$ $1.6385$ $\boldsymbol{0.0134}\pm\boldsymbol{0.0004}$
-$\mathbb{S}^{3}$ $1.7298$ $146.9142\pm 1.0720$ $1.7298$ $0.3666\pm 0.0157$ $\boldsymbol{1.7297}$ $\boldsymbol{0.0245}\pm\boldsymbol{0.0001}$
-$\mathbb{S}^{5}$ $1.6869$ $268.8883\pm 1.4754$ $1.6878$ $2.2747\pm 0.0047$ $\boldsymbol{1.6868}$ $\boldsymbol{0.3351}\pm\boldsymbol{0.0003}$
-$\mathbb{S}^{10}$ $1.3919$ $452.0089\pm 2.1364$ $1.4013$ $4.5715\pm 0.0023$ $\boldsymbol{1.3908}$ $\boldsymbol{0.6533}\pm\boldsymbol{0.0180}$
-$\mathbb{S}^{20}$ $1.0636$ $726.7038\pm 5.3951$ $1.0830$ $8.3452\pm 0.0040$ $\boldsymbol{1.0560}$ $\boldsymbol{1.4084}\pm\boldsymbol{0.0021}$
-$\mathbb{S}^{50}$ $0.7320$ $1616.3099\pm 8.8835$ $0.7427$ $71.9378\pm 0.0786$ $\boldsymbol{0.3492}$ $\boldsymbol{2.2147}\pm\boldsymbol{0.0047}$
-$\mathbb{S}^{100}$ $0.5832$ $4253.1743\pm 13.8589$ $0.5511$ $291.3856\pm 2.1209$ $\boldsymbol{0.2639}$ $\boldsymbol{7.4223}\pm\boldsymbol{0.0056}$
-$\mathbb{S}^{250}$- -- $0.4094$ $1241.9021\pm 7.3069$ $\boldsymbol{0.1686}$ $\boldsymbol{52.1050}\pm\boldsymbol{0.0283}$
-$\mathbb{S}^{500}$- -- $0.3531$ $2204.0300\pm 1.4159$ $\boldsymbol{0.1193}$ $\boldsymbol{239.8835}\pm\boldsymbol{0.2815}$
-$\mathbb{S}^{1000}$- -- $0.3269$ $3347.6938\pm 44.0626$ $\boldsymbol{0.0828}$ $\boldsymbol{958.4282}\pm\boldsymbol{7.2917}$
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3459}$ $146.3793\pm 0.7961$ $1.3460$ $0.0436\pm 0.0001$ $1.3461$ $\boldsymbol{0.0084}\pm\boldsymbol{0.0002}$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3826}$ $220.9877\pm 2.8045$ $1.3870$ $0.3788\pm 0.0023$ $1.3826$ $\boldsymbol{0.0972}\pm\boldsymbol{0.0002}$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2756}$ $333.3406\pm 0.9750$ $1.2780$ $2.2999\pm 0.0031$ $1.2757$ $\boldsymbol{0.3920}\pm\boldsymbol{0.0009}$
-$\mathrm{E}\left(10\right)$ $\boldsymbol{0.5551}$ $1225.0781\pm 1.1283$ $1.0666$ $4.8530\pm 0.0046$ $1.0126$ $\boldsymbol{0.8426}\pm\boldsymbol{0.0013}$
-$\mathrm{E}\left(20\right)$ $\boldsymbol{0.7340}$ $1834.0602\pm 20.3805$ $0.8404$ $8.9581\pm 0.0200$ $0.7946$ $\boldsymbol{0.9224}\pm\boldsymbol{0.0015}$
-$\mathrm{E}\left(50\right)$ $\boldsymbol{0.5735}$ $1238.1737\pm 3.1700$ $0.5797$ $67.3310\pm 0.0869$- --
-$\mathrm{E}\left(100\right)$ $0.4369$ $6365.4741\pm 34.9823$ $0.4342$ $221.8947\pm 1.4868$ $\boldsymbol{0.2042}$ $\boldsymbol{6.5394}\pm\boldsymbol{0.0437}$
-$\mathrm{E}\left(250\right)$- -- $0.3234$ $935.4907\pm 5.1887$ $\boldsymbol{0.1365}$ $\boldsymbol{51.8779}\pm\boldsymbol{0.0096}$
-$\mathrm{E}\left(500\right)$- -- $0.2862$ $1512.1372\pm 0.9319$ $\boldsymbol{0.0940}$ $\boldsymbol{174.0314}\pm\boldsymbol{0.2612}$
-$\mathrm{E}\left(1000\right)$- -- $0.2723$ $2488.8240\pm 38.0558$ $\boldsymbol{0.0893}$ $\boldsymbol{681.7435}\pm\boldsymbol{0.2447}$
-$\mathbb{T}^{2}$ $\boldsymbol{7.0791}$ $598.7396\pm 0.0929$ $7.1796$ $\boldsymbol{0.1417}\pm\boldsymbol{0.0002}$ $9.0475$ $2.1109\pm 0.0035$
-$\mathbb{H}^{2}$ $1.3317$ $189.9143\pm 0.0787$ $\boldsymbol{1.3316}$ $0.1403\pm 0.0026$ $1.3316$ $\boldsymbol{0.0070}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(2\right)$ $0.7249$ $224.4045\pm 0.4114$ $\boldsymbol{0.7249}$ $0.2194\pm 0.0025$ $0.7249$ $\boldsymbol{0.0153}\pm\boldsymbol{0.0001}$
-$\mathcal{P}\left(3\right)$ $1.2707$ $306.2382\pm 0.1092$ $1.2735$ $2.7082\pm 0.0075$ $\boldsymbol{1.2707}$ $\boldsymbol{0.1276}\pm\boldsymbol{0.0014}$
-Gaussian Distribution $1.6806$ $88.1351\pm 0.1027$ $1.6811$ $0.0569\pm 0.0001$ $\boldsymbol{1.6806}$ $\boldsymbol{0.0033}\pm\boldsymbol{0.0001}$
-Fréchet Distribution $0.5371$ $71.0994\pm 0.0709$ $0.5370$ $0.0095\pm 0.0000$ $\boldsymbol{0.5370}$ $\boldsymbol{0.0015}\pm\boldsymbol{0.0001}$
-Cauchy Distribution $1.1084$ $94.0708\pm 0.0792$ $1.1085$ $0.0075\pm 0.0001$ $\boldsymbol{1.1084}$ $\boldsymbol{0.0051}\pm\boldsymbol{0.0001}$
-Pareto Distribution $0.3857$ $46.7253\pm 0.2142$ $0.3857$ $0.0084\pm 0.0000$ $\boldsymbol{0.3857}$ $\boldsymbol{0.0011}\pm\boldsymbol{0.0001}$
-Table 25: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-SGD (T=100) CG (T=100) dogleg (T=100)
-Finsler Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.7321$ $\boldsymbol{0.5226}\pm\boldsymbol{0.0033}$ $1.6536$ $864.7791\pm 2.1507$ $\boldsymbol{1.6525}$ $14.6421\pm 0.0605$
-$\mathbb{S}^{3}$ $1.9179$ $\boldsymbol{0.8067}\pm\boldsymbol{0.0005}$ $\boldsymbol{1.7453}$ $976.1121\pm 13.5118$ $1.9228$ $4.1058\pm 0.0123$
-$\mathbb{S}^{5}$ $2.1622$ $\boldsymbol{4.3635}\pm\boldsymbol{0.0015}$ $\boldsymbol{1.7099}$ $1326.8875\pm 10.6284$ $2.1703$ $5.8463\pm 0.0436$
-$\mathbb{S}^{10}$ $2.4132$ $\boldsymbol{8.9293}\pm\boldsymbol{0.0051}$ $\boldsymbol{1.4427}$ $801.4553\pm 1.2792$ $2.4298$ $18.1452\pm 0.0392$
-$\mathbb{S}^{20}$ $2.4912$ $\boldsymbol{19.4868}\pm\boldsymbol{0.0023}$ $\boldsymbol{1.2646}$ $397.4008\pm 2.7750$ $2.5207$ $98.6398\pm 0.1329$
-$\mathbb{S}^{50}$ $2.2856$ $\boldsymbol{155.1010}\pm\boldsymbol{0.2023}$ $\boldsymbol{0.9295}$ $274.3937\pm 7.9883$- --
-$\mathbb{S}^{100}$ $1.9542$ $619.2194\pm 0.4311$ $\boldsymbol{1.0756}$ $\boldsymbol{221.9957}\pm\boldsymbol{0.5803}$- --
-$\mathbb{S}^{250}$ $1.4363$ $5475.3242\pm 1.4705$ $\boldsymbol{0.8813}$ $\boldsymbol{304.5134}\pm\boldsymbol{0.0238}$- --
-$\mathbb{S}^{500}$- -- $\boldsymbol{1.0790}$ $\boldsymbol{31.3454}\pm\boldsymbol{0.0388}$- --
-$\mathbb{S}^{1000}$- -- $\boldsymbol{0.7859}$ $\boldsymbol{137.1001}\pm\boldsymbol{0.4771}$- --
-$\mathrm{E}\left(2\right)$ $1.3705$ $\boldsymbol{0.5219}\pm\boldsymbol{0.0024}$ $1.3620$ $401.8943\pm 0.7424$ $\boldsymbol{1.3616}$ $14.9190\pm 0.0640$
-$\mathrm{E}\left(3\right)$ $1.4388$ $\boldsymbol{0.8083}\pm\boldsymbol{0.0019}$ $\boldsymbol{1.4147}$ $621.6700\pm 3.4375$ $1.4397$ $4.6115\pm 0.0556$
-$\mathrm{E}\left(5\right)$ $1.5358$ $\boldsymbol{4.6840}\pm\boldsymbol{0.0044}$ $\boldsymbol{1.3355}$ $1523.9905\pm 19.2195$ $1.5373$ $5.7719\pm 0.0192$
-$\mathrm{E}\left(10\right)$ $1.6639$ $\boldsymbol{9.7218}\pm\boldsymbol{0.0077}$ $\boldsymbol{1.0329}$ $2475.9639\pm 93.5344$ $1.6672$ $18.9040\pm 0.6761$
-$\mathrm{E}\left(20\right)$ $1.7336$ $\boldsymbol{19.6573}\pm\boldsymbol{0.0115}$ $\boldsymbol{0.9341}$ $552.4011\pm 0.4176$ $1.7403$ $95.1917\pm 0.7952$
-$\mathrm{E}\left(50\right)$ $1.6453$ $\boldsymbol{145.3531}\pm\boldsymbol{0.3967}$ $\boldsymbol{0.7045}$ $573.5270\pm 6.1649$- --
-$\mathrm{E}\left(100\right)$ $1.4396$ $616.7589\pm 0.2897$ $\boldsymbol{0.7510}$ $\boldsymbol{394.1170}\pm\boldsymbol{2.5401}$- --
-$\mathrm{E}\left(250\right)$ $1.0800$ $4937.1353\pm 67.9600$ $\boldsymbol{0.6266}$ $\boldsymbol{784.1846}\pm\boldsymbol{4.4711}$- --
-$\mathrm{E}\left(500\right)$- -- $\boldsymbol{0.8183}$ $\boldsymbol{31.6893}\pm\boldsymbol{0.0422}$- --
-$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.5997}$ $\boldsymbol{152.4971}\pm\boldsymbol{0.5062}$- --
-$\mathbb{T}^{2}$- -- $\boldsymbol{8.5402}$ $198.5802\pm 0.2435$ $9.1903$ $4.4851\pm 0.0162$
-$\mathbb{H}^{2}$ $1.4881$ $\boldsymbol{0.2459}\pm\boldsymbol{0.0024}$ $1.3515$ $2582.8269\pm 190.4313$ $\boldsymbol{1.3338}$ $24.6057\pm 0.0985$
-$\mathcal{P}\left(2\right)$ $0.7860$ $\boldsymbol{0.6287}\pm\boldsymbol{0.0027}$ $0.7321$ $1333.9152\pm 50.5504$ $\boldsymbol{0.7291}$ $17.3661\pm 0.0366$
-$\mathcal{P}\left(3\right)$ $1.3773$ $\boldsymbol{5.3062}\pm\boldsymbol{0.0015}$ $1.2912$ $920.9217\pm 15.9820$ $\boldsymbol{1.2856}$ $35.0991\pm 0.0589$
-Gaussian Distribution $1.9173$ $\boldsymbol{0.1056}\pm\boldsymbol{0.0002}$ $1.6881$ $820.4708\pm 1.3159$ $\boldsymbol{1.6878}$ $15.1765\pm 0.0966$
-Fréchet Distribution $0.5499$ $\boldsymbol{0.1143}\pm\boldsymbol{0.0009}$ $0.5414$ $302.1487\pm 0.2635$ $\boldsymbol{0.5405}$ $9.8568\pm 0.0634$
-Cauchy Distribution $1.4099$ $\boldsymbol{0.1025}\pm\boldsymbol{0.0010}$ $1.1138$ $821.1230\pm 0.8392$ $\boldsymbol{1.1133}$ $17.1469\pm 0.0529$
-Pareto Distribution $0.3951$ $\boldsymbol{0.1124}\pm\boldsymbol{0.0002}$ $0.3899$ $206.6691\pm 0.1273$ $\boldsymbol{0.3884}$ $5.9249\pm 0.0319$
-Table 26: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-SGD (T=50) CG (T=50) dogleg (T=50)
-Finsler Manifold Length Runtime Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.7109$ $\boldsymbol{0.2588}\pm\boldsymbol{0.0008}$ $1.6385$ $439.9231\pm 4.5034$ $\boldsymbol{1.6385}$ $14.8097\pm 0.0251$
-$\mathbb{S}^{3}$ $1.8925$ $\boldsymbol{0.3882}\pm\boldsymbol{0.0126}$ $\boldsymbol{1.7298}$ $496.1117\pm 4.7323$ $1.9113$ $3.6687\pm 0.0188$
-$\mathbb{S}^{5}$ $2.1279$ $\boldsymbol{2.2943}\pm\boldsymbol{0.0029}$ $\boldsymbol{1.6874}$ $658.4536\pm 4.7574$ $2.1588$ $4.0045\pm 0.0993$
-$\mathbb{S}^{10}$ $2.3586$ $\boldsymbol{4.8071}\pm\boldsymbol{0.0036}$ $\boldsymbol{1.3911}$ $896.1763\pm 6.7043$ $2.4194$ $6.7420\pm 0.0149$
-$\mathbb{S}^{20}$ $2.4105$ $\boldsymbol{9.5867}\pm\boldsymbol{0.0165}$ $\boldsymbol{1.1016}$ $622.5931\pm 6.0175$ $2.5121$ $25.6215\pm 0.0523$
-$\mathbb{S}^{50}$ $2.1971$ $\boldsymbol{71.2637}\pm\boldsymbol{0.0976}$ $\boldsymbol{0.9319}$ $177.8937\pm 1.2493$- --
-$\mathbb{S}^{100}$ $1.8917$ $295.5522\pm 1.0715$ $\boldsymbol{0.6486}$ $\boldsymbol{219.8582}\pm\boldsymbol{1.3192}$- --
-$\mathbb{S}^{250}$ $1.4136$ $2760.2507\pm 17.1501$ $\boldsymbol{0.5816}$ $\boldsymbol{333.5332}\pm\boldsymbol{1.1282}$- --
-$\mathbb{S}^{500}$ $1.0691$ $11508.7256\pm 47.0186$ $\boldsymbol{0.6111}$ $\boldsymbol{510.0441}\pm\boldsymbol{1.7312}$- --
-$\mathbb{S}^{1000}$- -- $\boldsymbol{0.6762}$ $\boldsymbol{565.0285}\pm\boldsymbol{2.4234}$- --
-$\mathrm{E}\left(2\right)$ $1.3601$ $\boldsymbol{0.2615}\pm\boldsymbol{0.0002}$ $1.3461$ $422.3503\pm 12.1072$ $\boldsymbol{1.3460}$ $20.7184\pm 0.0617$
-$\mathrm{E}\left(3\right)$ $1.4291$ $\boldsymbol{0.4036}\pm\boldsymbol{0.0006}$ $\boldsymbol{1.3897}$ $1200.0380\pm 33.6871$ $1.4325$ $4.2812\pm 0.0319$
-$\mathrm{E}\left(5\right)$ $1.5254$ $\boldsymbol{2.1031}\pm\boldsymbol{0.0024}$ $\boldsymbol{1.2803}$ $993.6235\pm 15.9966$ $1.5310$ $4.2148\pm 0.0296$
-$\mathrm{E}\left(10\right)$ $1.6487$ $\boldsymbol{4.7998}\pm\boldsymbol{0.0110}$ $\boldsymbol{0.5483}$ $2210.6582\pm 37.5507$ $1.6615$ $6.6250\pm 0.0132$
-$\mathrm{E}\left(20\right)$ $1.7096$ $\boldsymbol{9.5709}\pm\boldsymbol{0.0027}$ $\boldsymbol{0.8485}$ $564.2457\pm 0.9527$ $1.7353$ $23.9862\pm 0.0349$
-$\mathrm{E}\left(50\right)$ $1.6121$ $\boldsymbol{73.0136}\pm\boldsymbol{0.0216}$ $\boldsymbol{0.6313}$ $344.0060\pm 1.0995$- --
-$\mathrm{E}\left(100\right)$ $1.4125$ $\boldsymbol{313.0101}\pm\boldsymbol{0.3825}$ $\boldsymbol{0.4981}$ $331.0117\pm 1.3895$- --
-$\mathrm{E}\left(250\right)$ $1.0688$ $2706.8379\pm 8.4708$ $\boldsymbol{0.5508}$ $\boldsymbol{378.6086}\pm\boldsymbol{0.2978}$- --
-$\mathrm{E}\left(500\right)$ $0.8132$ $11307.9814\pm 281.9052$ $\boldsymbol{0.5677}$ $\boldsymbol{751.1521}\pm\boldsymbol{1.0706}$- --
-$\mathrm{E}\left(1000\right)$- -- $\boldsymbol{0.5987}$ $\boldsymbol{77.4691}\pm\boldsymbol{0.2770}$- --
-$\mathbb{T}^{2}$ $inf$ $\boldsymbol{0.0490}\pm\boldsymbol{0.0001}$ $\boldsymbol{8.1529}$ $167.4743\pm 2.8798$ $9.0475$ $4.3053\pm 0.0156$
-$\mathbb{H}^{2}$ $1.4850$ $\boldsymbol{0.1246}\pm\boldsymbol{0.0040}$ $1.3320$ $2460.6926\pm 57.6646$ $\boldsymbol{1.3316}$ $19.7572\pm 0.0371$
-$\mathcal{P}\left(2\right)$ $0.7739$ $\boldsymbol{0.3208}\pm\boldsymbol{0.0030}$ $0.7250$ $712.9343\pm 11.1547$ $\boldsymbol{0.7249}$ $14.5680\pm 0.0587$
-$\mathcal{P}\left(3\right)$ $1.3514$ $\boldsymbol{2.6554}\pm\boldsymbol{0.0101}$ $1.2709$ $788.9128\pm 8.9808$ $\boldsymbol{1.2707}$ $22.6905\pm 0.0425$
-Gaussian Distribution $1.8463$ $\boldsymbol{0.0519}\pm\boldsymbol{0.0001}$ $1.6807$ $283.3971\pm 6.7184$ $\boldsymbol{1.6806}$ $16.2727\pm 0.1111$
-Fréchet Distribution $0.5455$ $\boldsymbol{0.0596}\pm\boldsymbol{0.0001}$ $0.5372$ $253.2346\pm 0.2384$ $\boldsymbol{0.5370}$ $9.3157\pm 0.0269$
-Cauchy Distribution $1.3641$ $\boldsymbol{0.0522}\pm\boldsymbol{0.0001}$ $1.1088$ $360.8286\pm 0.4404$ $\boldsymbol{1.1084}$ $16.8439\pm 0.0398$
-Pareto Distribution $0.3919$ $\boldsymbol{0.0555}\pm\boldsymbol{0.0001}$ $0.3857$ $142.1006\pm 0.0492$ $\boldsymbol{0.3856}$ $7.9442\pm 0.0303$
-Table 27: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-trust-ncg (T=100) trust-exact (T=100)
-Finsler Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.6525$ $1261.5378\pm 7.7825$ $\boldsymbol{1.6525}$ $\boldsymbol{14.7265}\pm\boldsymbol{0.0327}$
-$\mathbb{S}^{3}$ $1.7432$ $1994.5762\pm 20.6694$ $\boldsymbol{1.7432}$ $\boldsymbol{20.4180}\pm\boldsymbol{0.0359}$
-$\mathbb{S}^{5}$ $\boldsymbol{1.7027}$ $5603.5640\pm 76.5314$ $1.7027$ $\boldsymbol{38.7189}\pm\boldsymbol{0.0167}$
-$\mathbb{S}^{10}$- -- $\boldsymbol{1.4108}$ $\boldsymbol{184.4648}\pm\boldsymbol{0.3001}$
-$\mathbb{S}^{20}$- -- $\boldsymbol{1.0769}$ $\boldsymbol{1087.9535}\pm\boldsymbol{2.3390}$
-$\mathbb{S}^{50}$- -- -- --
-$\mathbb{S}^{100}$- -- -- --
-$\mathbb{S}^{250}$- -- -- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $1.3616$ $1019.1379\pm 8.5157$ $\boldsymbol{1.3616}$ $\boldsymbol{15.4908}\pm\boldsymbol{0.0374}$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3997}$ $4148.8574\pm 2.3006$ $1.3997$ $\boldsymbol{31.8261}\pm\boldsymbol{0.1322}$
-$\mathrm{E}\left(5\right)$ $1.2950$ $5965.7002\pm 44.3871$ $\boldsymbol{1.2950}$ $\boldsymbol{40.3707}\pm\boldsymbol{0.0405}$
-$\mathrm{E}\left(10\right)$- -- $\boldsymbol{1.0318}$ $\boldsymbol{249.5034}\pm\boldsymbol{0.2415}$
-$\mathrm{E}\left(20\right)$- -- $\boldsymbol{0.8109}$ $\boldsymbol{1033.4261}\pm\boldsymbol{3.7266}$
-$\mathrm{E}\left(50\right)$- -- -- --
-$\mathrm{E}\left(100\right)$- -- -- --
-$\mathrm{E}\left(250\right)$- -- -- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{8.9015}$ $348.9266\pm 0.5435$ $9.1722$ $\boldsymbol{56.1951}\pm\boldsymbol{0.0567}$
-$\mathbb{H}^{2}$ $\boldsymbol{1.3338}$ $3365.4121\pm 15.3287$ $1.3338$ $\boldsymbol{24.6266}\pm\boldsymbol{0.1178}$
-$\mathcal{P}\left(2\right)$ $0.7291$ $2314.0154\pm 22.6535$ $\boldsymbol{0.7291}$ $\boldsymbol{16.6480}\pm\boldsymbol{0.0684}$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $3887.0437\pm 81.9548$ $1.2856$ $\boldsymbol{42.8468}\pm\boldsymbol{0.1088}$
-Gaussian Distribution $\boldsymbol{1.6878}$ $1176.3495\pm 4.4250$ $1.6878$ $\boldsymbol{13.9002}\pm\boldsymbol{0.0735}$
-Fréchet Distribution $0.5405$ $1173.9244\pm 4.9721$ $\boldsymbol{0.5405}$ $\boldsymbol{9.5119}\pm\boldsymbol{0.0199}$
-Cauchy Distribution $\boldsymbol{1.1133}$ $1425.5028\pm 8.3454$ $1.1133$ $\boldsymbol{14.6313}\pm\boldsymbol{0.0416}$
-Pareto Distribution $\boldsymbol{0.3884}$ $406.4951\pm 4.4740$ $0.3884$ $\boldsymbol{5.9478}\pm\boldsymbol{0.0531}$
-Table 28: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-trust-ncg (T=50) trust-exact (T=50)
-Finsler Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $1.6385$ $461.6090\pm 2.4283$ $\boldsymbol{1.6385}$ $\boldsymbol{14.9718}\pm\boldsymbol{0.0477}$
-$\mathbb{S}^{3}$ $1.7297$ $738.8096\pm 14.1991$ $\boldsymbol{1.7297}$ $\boldsymbol{18.2119}\pm\boldsymbol{0.0850}$
-$\mathbb{S}^{5}$ $1.6868$ $1845.5012\pm 28.1314$ $\boldsymbol{1.6868}$ $\boldsymbol{27.8877}\pm\boldsymbol{0.0334}$
-$\mathbb{S}^{10}$ $\boldsymbol{1.3881}$ $11723.0781\pm 267.5437$ $1.3881$ $\boldsymbol{72.3786}\pm\boldsymbol{0.5046}$
-$\mathbb{S}^{20}$- -- $\boldsymbol{1.0591}$ $\boldsymbol{321.4109}\pm\boldsymbol{0.3724}$
-$\mathbb{S}^{50}$- -- -- --
-$\mathbb{S}^{100}$- -- -- --
-$\mathbb{S}^{250}$- -- -- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $1.3460$ $760.6122\pm 4.1086$ $\boldsymbol{1.3460}$ $\boldsymbol{18.9955}\pm\boldsymbol{0.0502}$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3825}$ $1558.1522\pm 1.8558$ $1.3825$ $\boldsymbol{25.1379}\pm\boldsymbol{0.0384}$
-$\mathrm{E}\left(5\right)$ $1.2756$ $1658.7789\pm 4.1942$ $\boldsymbol{1.2755}$ $\boldsymbol{31.1768}\pm\boldsymbol{0.0425}$
-$\mathrm{E}\left(10\right)$ $1.0112$ $15954.7256\pm 487.2155$ $\boldsymbol{1.0109}$ $\boldsymbol{104.7452}\pm\boldsymbol{1.3478}$
-$\mathrm{E}\left(20\right)$- -- $\boldsymbol{0.8000}$ $\boldsymbol{371.1157}\pm\boldsymbol{2.4557}$
-$\mathrm{E}\left(50\right)$- -- -- --
-$\mathrm{E}\left(100\right)$- -- -- --
-$\mathrm{E}\left(250\right)$- -- -- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{8.4108}$ $\boldsymbol{364.3664}\pm\boldsymbol{2.8448}$- --
-$\mathbb{H}^{2}$ $1.3316$ $1348.8287\pm 29.8787$ $\boldsymbol{1.3316}$ $\boldsymbol{19.5780}\pm\boldsymbol{0.0558}$
-$\mathcal{P}\left(2\right)$ $0.7249$ $702.5833\pm 1.9384$ $\boldsymbol{0.7249}$ $\boldsymbol{14.2773}\pm\boldsymbol{0.0611}$
-$\mathcal{P}\left(3\right)$ $1.2707$ $1289.7513\pm 26.6135$ $\boldsymbol{1.2707}$ $\boldsymbol{20.2860}\pm\boldsymbol{0.0433}$
-Gaussian Distribution $\boldsymbol{1.6806}$ $586.1818\pm 0.5433$ $1.6806$ $\boldsymbol{13.2321}\pm\boldsymbol{0.0358}$
-Fréchet Distribution $\boldsymbol{0.5370}$ $458.0856\pm 0.2181$ $0.5370$ $\boldsymbol{9.8644}\pm\boldsymbol{0.0257}$
-Cauchy Distribution $\boldsymbol{1.1084}$ $584.2502\pm 1.5708$ $1.1084$ $\boldsymbol{13.5337}\pm\boldsymbol{0.0222}$
-Pareto Distribution $\boldsymbol{0.3856}$ $366.4953\pm 0.2211$ $\boldsymbol{0.3856}$ $\boldsymbol{8.6113}\pm\boldsymbol{0.0271}$
-Table 29: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-Sparse Newton (T=100) Sparse Regularized Newton (T=100)
-Finsler Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $\boldsymbol{1.6525}$ $\boldsymbol{0.0632}\pm\boldsymbol{0.0009}$ $1.6525$ $0.4062\pm 0.0273$
-$\mathbb{S}^{3}$ $\boldsymbol{1.7432}$ $\boldsymbol{0.3018}\pm\boldsymbol{0.0366}$ $1.7433$ $1.0369\pm 0.0880$
-$\mathbb{S}^{5}$ $\boldsymbol{1.7027}$ $11.3808\pm 0.8228$ $1.7028$ $\boldsymbol{3.5064}\pm\boldsymbol{0.3096}$
-$\mathbb{S}^{10}$ $\boldsymbol{1.4092}$ $121.2282\pm 12.4696$ $1.4166$ $\boldsymbol{11.3672}\pm\boldsymbol{0.8836}$
-$\mathbb{S}^{20}$ $1.3030$ $2840.8333\pm 230.3178$ $\boldsymbol{1.0865}$ $\boldsymbol{52.2612}\pm\boldsymbol{5.1479}$
-$\mathbb{S}^{50}$- -- $\boldsymbol{0.7556}$ $\boldsymbol{809.2408}\pm\boldsymbol{68.4436}$
-$\mathbb{S}^{100}$ $0.6141$ $16876.4219\pm 1314.8346$ $\boldsymbol{0.5968}$ $\boldsymbol{6984.5508}\pm\boldsymbol{566.5103}$
-$\mathbb{S}^{250}$- -- -- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3616}$ $\boldsymbol{0.0916}\pm\boldsymbol{0.0029}$ $1.3619$ $0.3423\pm 0.0174$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3997}$ $\boldsymbol{0.5216}\pm\boldsymbol{0.0511}$ $1.3997$ $1.3200\pm 0.1149$
-$\mathrm{E}\left(5\right)$ $1.5359$ $195.1668\pm 21.6446$ $\boldsymbol{1.2952}$ $\boldsymbol{3.7467}\pm\boldsymbol{0.3126}$
-$\mathrm{E}\left(10\right)$ $1.2272$ $560.2833\pm 45.3723$ $\boldsymbol{1.0807}$ $\boldsymbol{10.0806}\pm\boldsymbol{0.7869}$
-$\mathrm{E}\left(20\right)$ $1.0210$ $2939.9907\pm 244.8212$ $\boldsymbol{0.8555}$ $\boldsymbol{52.9693}\pm\boldsymbol{5.7635}$
-$\mathrm{E}\left(50\right)$- -- $\boldsymbol{0.6211}$ $\boldsymbol{733.2034}\pm\boldsymbol{60.1911}$
-$\mathrm{E}\left(100\right)$- -- $\boldsymbol{0.5132}$ $\boldsymbol{6657.5815}\pm\boldsymbol{538.6103}$
-$\mathrm{E}\left(250\right)$- -- -- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{9.1903}$ $\boldsymbol{27.4443}\pm\boldsymbol{2.4951}$ $\boldsymbol{9.1903}$ $27.5325\pm 2.2826$
-$\mathbb{H}^{2}$ $\boldsymbol{1.3338}$ $\boldsymbol{0.0816}\pm\boldsymbol{0.0019}$ $1.3340$ $0.3529\pm 0.0227$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7291}$ $\boldsymbol{0.1944}\pm\boldsymbol{0.0248}$ $0.7291$ $0.7200\pm 0.0868$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2856}$ $\boldsymbol{0.8253}\pm\boldsymbol{0.0772}$ $1.2857$ $2.7333\pm 0.2234$
-Gaussian Distribution $\boldsymbol{1.6878}$ $\boldsymbol{0.0636}\pm\boldsymbol{0.0006}$ $1.6878$ $0.2146\pm 0.0122$
-Fréchet Distribution $\boldsymbol{0.5405}$ $\boldsymbol{0.0368}\pm\boldsymbol{0.0001}$ $0.5411$ $0.1835\pm 0.0107$
-Cauchy Distribution $\boldsymbol{1.1133}$ $\boldsymbol{0.0720}\pm\boldsymbol{0.0015}$ $1.1133$ $0.2079\pm 0.0126$
-Pareto Distribution $\boldsymbol{0.3884}$ $\boldsymbol{0.0159}\pm\boldsymbol{0.0002}$ $0.3886$ $0.1791\pm 0.0097$
-Table 30: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
-Sparse Newton (T=50) Sparse Regularized Newton (T=50)
-Finsler Manifold Length Runtime Length Runtime
-$\mathbb{S}^{2}$ $\boldsymbol{1.6385}$ $\boldsymbol{0.0504}\pm\boldsymbol{0.0005}$ $1.6385$ $0.1821\pm 0.0097$
-$\mathbb{S}^{3}$ $\boldsymbol{1.7297}$ $\boldsymbol{0.1977}\pm\boldsymbol{0.0270}$ $1.7297$ $0.4563\pm 0.0481$
-$\mathbb{S}^{5}$ $1.6868$ $\boldsymbol{0.8981}\pm\boldsymbol{0.0788}$ $\boldsymbol{1.6868}$ $1.7423\pm 0.1372$
-$\mathbb{S}^{10}$ $\boldsymbol{1.3354}$ $332.2007\pm 26.8370$ $1.3881$ $\boldsymbol{6.2534}\pm\boldsymbol{0.5135}$
-$\mathbb{S}^{20}$ $\boldsymbol{0.8588}$ $1562.1842\pm 125.8776$ $1.0579$ $\boldsymbol{30.9976}\pm\boldsymbol{2.3280}$
-$\mathbb{S}^{50}$- -- $\boldsymbol{0.7149}$ $\boldsymbol{364.6181}\pm\boldsymbol{32.4896}$
-$\mathbb{S}^{100}$- -- $\boldsymbol{0.5364}$ $\boldsymbol{3463.5500}\pm\boldsymbol{274.5852}$
-$\mathbb{S}^{250}$- -- -- --
-$\mathbb{S}^{500}$- -- -- --
-$\mathbb{S}^{1000}$- -- -- --
-$\mathrm{E}\left(2\right)$ $\boldsymbol{1.3460}$ $\boldsymbol{0.0658}\pm\boldsymbol{0.0015}$ $1.3461$ $0.1592\pm 0.0057$
-$\mathrm{E}\left(3\right)$ $\boldsymbol{1.3825}$ $\boldsymbol{0.2757}\pm\boldsymbol{0.0323}$ $1.3825$ $0.6126\pm 0.0610$
-$\mathrm{E}\left(5\right)$ $\boldsymbol{1.2756}$ $\boldsymbol{0.8444}\pm\boldsymbol{0.0827}$ $1.2756$ $1.6410\pm 0.1521$
-$\mathrm{E}\left(10\right)$ $\boldsymbol{1.0106}$ $\boldsymbol{276.2837}\pm\boldsymbol{26.7637}$ $1.0575$ $284.1870\pm 27.1057$
-$\mathrm{E}\left(20\right)$ $0.8571$ $1488.3325\pm 147.3422$ $\boldsymbol{0.7945}$ $\boldsymbol{30.3454}\pm\boldsymbol{1.4186}$
-$\mathrm{E}\left(50\right)$- -- $\boldsymbol{0.5718}$ $\boldsymbol{384.6765}\pm\boldsymbol{30.0164}$
-$\mathrm{E}\left(100\right)$- -- $\boldsymbol{0.4397}$ $\boldsymbol{3289.7231}\pm\boldsymbol{270.1520}$
-$\mathrm{E}\left(250\right)$- -- -- --
-$\mathrm{E}\left(500\right)$- -- -- --
-$\mathrm{E}\left(1000\right)$- -- -- --
-$\mathbb{T}^{2}$ $\boldsymbol{9.0475}$ $\boldsymbol{14.0754}\pm\boldsymbol{1.0615}$ $\boldsymbol{9.0475}$ $14.1199\pm 1.0787$
-$\mathbb{H}^{2}$ $\boldsymbol{1.3316}$ $\boldsymbol{0.0443}\pm\boldsymbol{0.0002}$ $1.3317$ $0.1778\pm 0.0101$
-$\mathcal{P}\left(2\right)$ $\boldsymbol{0.7249}$ $\boldsymbol{0.1060}\pm\boldsymbol{0.0219}$ $0.7249$ $0.3272\pm 0.0354$
-$\mathcal{P}\left(3\right)$ $\boldsymbol{1.2707}$ $\boldsymbol{0.4780}\pm\boldsymbol{0.0503}$ $1.2707$ $1.1371\pm 0.1027$
-Gaussian Distribution $\boldsymbol{1.6806}$ $\boldsymbol{0.0348}\pm\boldsymbol{0.0003}$ $1.6807$ $0.0914\pm 0.0039$
-Fréchet Distribution $\boldsymbol{0.5370}$ $\boldsymbol{0.0194}\pm\boldsymbol{0.0004}$ $0.5371$ $0.0880\pm 0.0009$
-Cauchy Distribution $\boldsymbol{1.1084}$ $\boldsymbol{0.0411}\pm\boldsymbol{0.0002}$ $1.1085$ $0.0975\pm 0.0039$
-Pareto Distribution $\boldsymbol{0.3856}$ $\boldsymbol{0.0174}\pm\boldsymbol{0.0004}$ $0.3857$ $0.0884\pm 0.0048$
-Table 31: The table shows the length of the estimated geodesics for different methods on a CPU.
-The methods have been stopped if the norm of the gradient was less than $10^{-4}$ or if a maximum of $1,000$ iterations has been reached.
-When the computational time was longer than $24$ hours, the value is set to --.
+F.2 Hardware
+[…]
+The runtime experiments on a CPU have been run on a CPU model XeonE5_2660v3 with $1$ nodes for at most $24$ hours with a maximum memory of $10$ GB.
+The runtime estimates on a GPU and training of the VAE have been computed on a GPU for at most 24 hours with a maximum memory of $16$ GB. The $GPU$ consists of $1$ node on a Tesla V100.